A Bibliography of O’Reilly & Associates and O’Reilly Media. Inc. Publishers

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

26 October 2023 Version 3.71

Title word cross-reference

#70 [1272, 1273]. #70-059 [1272]. #70-068 [1273].

2 [955]. 2 + 2 = 5986 [1466]. 3 [1157, 1580].

*# [1230].

.Mac [1952]. .NET [1871, 22, 188, 345, 446, 508, 596, 719, 721, 726, 735, 758, 791, 1005, 1041, 1044, 1045, 1050, 1056, 1097, 1098, 1099, 1127, 1265, 1478, 1869, 1870, 1874, 1910, 1911, 1912, 1928, 2009, 2041].

’05 [466, 1542]. 08 [1551].

1 [1423]. 1.0 [1016]. 1.1 [59]. 1.2 [1592]. 1000 [1521]. 1000D [1081]. 10g [716, 715]. 10th [2122]. 11 [1394].

2 [53, 212, 272, 586, 2147, 924, 946, 1525, 1531, 1540, 2035, 2057]. 2.0 [2, 55, 206, 397, 671, 947, 1007, 1051, 1248, 1285, 1514, 1756, 1813, 2085]. 2.1 [506]. 2.2 [204]. 2000 [38, 205, 609, 615, 674, 932, 992, 1095, 1275, 1367, 1368, 1668, 1763, 1793, 1885, 1971, 2081]. 2001 [96]. 2003 [70, 71, 72, 73, 74, 282, 356, 367, 368, 794, 795, 861, 993, 1154, 1972, 2038]. 2003-2013 [1758]. 2004 [1203]. 2005 [84, 151, 760, 761, 1008, 1048, 1049, 1127, 1130, 1477, 2133, 2030, 2068]. 2006 [152, 153]. 2007 [623, 731, 732, 733, 1131, 1133, 1134, 1135, 2135, 1985, 1986, 2042]. 2008 [1478]. 2009 [154]. 2013 [1758]. 21st [630, 631, 948]. 220-601 [336].

3 [130, 228, 468, 637, 706, 819, 1087, 1193, 1356, 1509, 1532, 1716]. 3.0 [23, 24, 239, 827, 852, 1055, 1085, 1286, 1487, 1623, 1713, 1714]. 3.5 [55, 763, 1052, 1054, 1056, 1478]. 3.x [279].

4 [44, 229, 671, 1192, 1201, 1268, 1536, 1537, 1722]. 4.0 [26, 27, 724, 1174, 1272, 1273, 1449, 1919]. 4.4BSD [339, 340, 341, 342, 343, 344]. 4.4BSD-Lite [339]. 450D [1082].

5 [230, 350, 1541, 1543, 1797, 1962, 2002, 2003, 2004]. 5.0 [725, 726]. 5.5 [2]. 56K [420].

6 [231, 465, 1547, 1548, 1602, 1603]. 6.1 [1372]. 6.2 [1173]. 6th [2127].

7 [1647, 2031]. 70-058 [1271]. 70-620 [337]. 7th [2128].

8 [58, 258, 298, 1086, 1205, 1556, 1777, 1983, 2036]. 8.13 [325]. 8.5 [1577]. 802.11b [1698]. 802.11n [574]. 8th [2129].

9 [299, 1523]. 90 [934, 935]. 92 [958]. 95 [279, 982, 1508, 1737]. 97 [1017, 1018, 1019]. 98 [208, 910, 1448, 2106]. 9th [2121].

= [602, 1109, 1664].

A100 [1246]. Absolute [823, 1108, 1106]. Abstraction [1581]. abuse [83]. Academic [981]. Access [86, 793, 1754, 852, 71, 175, 176, 282, 658, 659, 1131, 1133, 1263, 1472, 1676, 1679, 1684, 1727]. accessible [1882]. Accidental [1617, 1618, 1619]. acquainted [135]. Across [1690]. ActionScript [1778, 1084, 1085, 1282, 1283, 1284, 1285, 1286, 1713, 1714]. Active [392, 911, 1095, 47, 48, 51, 52, 864, 1261]. ActiveX [1672]. Add [1677]. Add-Ins [1677]. Adding [1566, 379]. Addressing [599, 600, 601, 602, 603]. admin [1960]. Administration [171, 240, 611, 612, 613, 616, 617, 854, 856, 857, 858, 970, 1233, 1299, 1442, 1443, 2107, 14, 112, 265, 288, 347, 662, 861, 1301, 1323, 1971].

Administrator [943, 944, 1296, 1298, 2036, 122]. Administrators [609, 751, 1062, 1368, 1482, 1508, 1919, 598, 1003, 1063, 1071]. ADO [1672]. ADO.NET [756, 757, 763]. Adobe [4, 5, 53, 275, 276, 400, 1189, 1190, 1191, 1192, 1194, 1195, 1196]. Adolescent [83]. adolescents [2025, 2026]. Advanced [123, 246, 330, 459, 511, 1178, 1825, 1905]. Advancing [2095]. advantage [1682]. Adventure [1842, 5]. advertising [378]. AEs [223]. Africa [1258]. African [1258]. After [64]. Age [98, 692, 1887, 2060]. Agent [800]. Agents [936, 1733]. agile [1776, 1917, 1918]. aguide [438]. AI [195]. Air [276]. AirPort [574]. AJAX [1029, 2056, 224, 350, 531, 832, 1051, 1140, 1223, 1410, 1501, 1502, 1519, 1566, 1578, 1621, 1644, 1795, 2048]. alcohol [83]. Algorithms [802, 1088, 1254, 2062, 1456]. Alien [1197]. Alpha [1246]. alternative [2043]. amateur [933]. Amazing [720, 1474]. Amazon [119, 1306]. Ambient [1288]. America [187, 384, 1761]. American [471]. Analysis [246, 1083, 1215, 176, 313, 327, 1158, 1225, 1715]. Analytics [1254, 359]. Analyzing [953]. Android [891, 1232, 2154, 2099, 2100]. Animating [959]. Animation [309, 764, 959, 435, 1777]. Annotated [1111]. Annoyances [208, 908, 910, 912, 913, 1017, 1018, 1019, 1020, 115, 116, 164, 622, 694, 792, 878, 885, 918, 920, 921, 1263, 1646, 1881, 1978]. Annoyances.org [912]. annoying [115, 116, 164, 622, 694, 792, 878, 920, 1263, 1646, 1881, 1978]. Annual [2121, 2122, 2127, 2128, 2129]. answers [1135]. Ant [840, 1932]. anthology [57, 460, 2123]. antipatterns [922]. ANTLR [1475]. AOL [384]. Apache [204, 224, 296, 297, 584, 586, 995, 996, 997, 1645, 1833]. API [531, 1682, 1833]. Apollo [275]. App [1718, 1719, 2138, 1829]. Apple [644, 1066, 1557, 1757]. Apples [1557]. AppleScript [681, 1327, 1328, 1498]. AppleWorks [465]. appliance [2137]. appliances [2137]. Application [209, 715, 825, 1105, 2113, 2114]. Applications [273, 632, 726, 734, 755, 801, 845, 1294, 1408, 1608, 1688, 1690, 1772, 1773, 2070, 2076, 2079, 192, 226, 235, 294, 360, 362, 379, 494, 730, 791, 937, 1010, 1045, 1219, 1268, 1322, 1478, 1756, 2048, 2072, 2073, 2088]. Applied [1836]. Applying [179]. approach [178, 232]. Apps [44, 295, 1014, 1829, 1974, 1002]. Architecture [1694, 1695, 1219, 1821, 1853, 2087]. Architektur [364]. Arduino [109, 95, 109, 891, 2150]. Arduino-Projekte [109]. Argentina [785, 1363]. Argentine [785]. Art [334, 1062, 2067, 6, 2117, 128, 144, 213, 482, 493, 824, 865, 1157, 1171, 1268, 1776, 2112]. Artificial [2158]. artists [173]. ASP [446, 1560, 1561, 2045, 2046]. ASP.NET [54, 55, 445, 947, 1029, 1007, 1039, 1042, 1046, 1051, 1052, 1054, 2056]. AspectJ [1247]. Assembling [177]. assembly [865]. assessing [450]. assessment [1227, 1228]. asset [977]. Asterisk [1802, 1991, 1992, 1993]. Astronomy [1924]. Atlas [2055]. Atom [766]. Atomics [190]. attack [1609]. attacks [2109]. Audience [1484, 1198]. audiences [1882]. Audio [132, 610, 1606]. Aufgaben [399]. Auflage [552]. Augmented [43, 44]. Ausdrücke [1871]. Ausgabe [552, 550, 551]. Australia [743, 1366]. authentication [1999]. AutoCAD [932]. AutoIt [576]. Automated [1001]. Automating [903, 1057, 1299, 1680, 1832, 1058, 1661]. automation [294]. automatisieren [1661]. aux [1658]. available [961]. Avian [744]. avoid [1978]. avoiding [922]. awk [413, 1653, 1654, 1652, 69, 412, 417, 1651]. AWT [2115]. AxKit [768].

back [573]. Backbone [2076]. Backup [1003, 1569, 1571]. backups [2111]. backyards [1474]. Balancing [197, 2111]. Bandit [2062]. Bandits [101]. barefoot [2060]. Baseball [15]. Based [1671, 2104, 705, 1057, 1058, 1157, 1179]. basements [1474]. Bash [30, 1660, 399, 739, 1330, 1331, 1332]. Basic [735, 758, 1008, 1041, 1044, 1048, 1127, 1130, 1477, 2133, 1478, 2030, 1295, 398, 591, 734, 755, 833, 1265, 1677, 1682, 1907]. Basics [605, 1013, 1704]. basierte [1642]. battling [1887]. Bayesian [2112]. Bazaar [1617, 1618, 1619]. Be [123, 124, 1883, 1905]. BEA [958]. BeagleBone [1641, 1642, 2103]. Beans [473]. Beautiful [168, 2130, 1821, 125, 1860]. BEEP [1688]. before [1699]. Beginner [1340, 1346]. beginner‘s [1778, 95, 1643]. Beginning [1641, 1933]. behandelt [399, 1120, 1722]. Behind [1700, 1230]. Below [1445]. Ben [1664]. BeOS [123, 124, 1905]. Berkeley [667, 35]. Bernhard [1145, 1146]. Best [189, 2142, 295, 519, 637, 707, 742, 991, 1457, 308, 314, 456, 1135, 1458, 1459, 1628]. Bestselling [1452, 66, 69, 518, 1343, 1655]. Beta [539]. Better [1894, 1246]. Betting [2158]. Beyond [1065, 1197, 1340, 1678, 1895, 2061]. BGP [1981]. Big [699, 1757, 692]. BIND [31, 32, 33, 34, 35, 37, 1071, 1072]. binding [1220]. Bioinformatics [660, 1755, 1933, 1934]. Bipolar [2025]. Birds [744]. Bitcoin [76, 1518, 1878]. bits [487]. black [1960]. Blackberry [1125]. BLAST [962]. bleeding [1213]. blending [127]. Blockchain [1879]. Blocks [1631]. Blogging [403, 817]. Blokzijl [1142]. blue [704]. Blueprint [1878, 1879]. blunders [1959]. Bob [1145]. Book [160, 1127, 316, 438, 589, 699, 824, 875, 977, 1268, 1370, 1698, 1722, 1757, 1860, 1913, 1960, 2007, 2029, 2031, 2032, 2033, 2110, 1213]. Books [1424, 1452, 66, 69, 518, 1343, 1655]. Bookshelf [1446, 1450, 1451, 1452, 1453, 1784, 1786, 66, 69, 518, 526, 559, 780, 1343, 1655, 1787, 1817, 2053]. boot [1860, 1957]. Border [1981]. boss [709]. Bots [2150]. Bourne [399]. Bourne-Shell [399]. Box [1822, 1823, 830, 2007]. Boy [708]. BPF [246]. Braden [1145]. brain [875, 1211]. brain-friendly [1211]. Brazil [746, 816]. Brazilian [816]. Breast [1178, 1177]. Brian [1146]. Bridge [4]. British [786]. broken [1466]. browsing [1210]. BSD [960, 999]. Buch [1722]. Build [58, 1073, 1608, 2101, 294, 424, 869, 961, 1478, 1632, 2085]. Building [54, 55, 65, 117, 211, 233, 278, 577, 578, 632, 726, 772, 814, 855, 937, 1219, 1254, 1294, 1333, 1484, 1485, 1486, 1596, 1631, 1818, 1829, 1923, 1926, 2076, 2097, 2098, 2106, 2116, 1981, 334, 730, 791, 1268, 1475, 1756, 2137, 1935]. Built [513, 514]. Built-in [513]. Built-ins [514]. burned [2094]. burning [1845, 2031]. bus [281]. Business [62, 63, 303, 353, 355, 471, 472, 534, 593, 654, 672, 736, 785, 786, 816, 899, 900, 925, 994, 1148, 1258, 1260, 1357, 1358, 1359, 1360, 1361, 1362, 1363, 1364, 1365, 1366, 1371, 1596, 1696, 2064, 2080, 797, 953, 1478, 1601, 2090]. Buyer [486, 1925].

C [99, 100, 300, 351, 352, 369, 376, 440, 948, 1068, 1088, 1089, 1462, 1463, 1464, 1465, 1466, 1467, 1575, 1576, 1586, 1631, 1720, 1829, 1833, 1839, 1869, 1870, 1999]. C# [1049, 1871, 21, 22, 23, 24, 26, 27, 166, 425, 426, 427, 724, 725, 726, 826, 827, 852, 1037, 1038, 1040, 1043, 1047, 1050, 1053, 1055, 1266, 1837, 1838, 1874, 1908]. Caching [969, 2058]. California [68]. called [2060]. Camera [4, 18]. Camp [1957]. Can [527, 332]. Canada [2064]. Cancer [527, 1178, 232, 884, 895, 931, 1177]. cancers [884]. Canon [1080, 1081, 1082]. Can’t [1298]. Canvas [2145]. Canyon [1445]. Capistrano [1105]. Car [1854]. caractères [769]. Carbon [1290]. caregivers [926, 929]. Carpenter [1146]. cartography [484]. Cascading [1236, 1239, 1241]. Casner [1145]. Cassandra [821]. Catalog [978, 979, 980, 981, 1424, 61, 982, 983]. Cathedral [1617, 1618, 1619]. CD [66, 69, 339, 349, 518, 526, 557, 559, 780, 1303, 1343, 1430, 1431, 1432, 1446, 1450, 1451, 1452, 1453, 1655, 1784, 1786, 1787, 1817, 2031, 2053]. CD-ROM [66, 69, 339, 349, 518, 557, 1303, 1343, 1431, 1432, 1452, 1655]. CDO [734]. center [259]. Century [948, 630, 631]. Certification [2155, 383, 2134]. Certified [1269]. CG [173]. CGI [309, 389, 386, 737, 738]. Change [2157, 2126]. cheap [638]. Checking [369, 374]. Checklist [1368]. checklists [1633]. Chiappa [1145]. Child [927, 930]. Childhood [884, 926, 929, 931]. children [2025, 2026]. China [1357]. Chinese [1110]. Chip [461]. Choosing [909, 1925]. Christian [1145]. CICS [845]. Cisco [20, 89, 185, 186, 407, 408, 1107, 1754]. Citizens [1740]. Civilizations [1197]. CJKV [1110]. Classes [701, 985, 555]. Classic [1656, 1127, 1661]. classification [2112]. Classroom [1671, 1805]. clear [1599]. CLI [1872]. click [704]. ClickOnce [1010]. Client [2082]. Cliff [1143]. Clinical [527, 893]. CLIs [2104]. Clojure [470, 753]. closed [1700]. Closer [1153]. Cloud [2153, 1935]. clouds [233]. cluster [288, 961, 961]. clusters [1800, 1818]. Cocoa [2139, 126, 376, 632, 1291, 1829]. codages [769]. Code [519, 682, 727, 2149, 883, 1261, 2130, 1636, 85, 487, 842, 866, 867, 1000]. coding [691, 1229]. Coding4Fun [507]. COFF [673]. ColdFusion [219, 220]. collecting [1264]. collection [874]. Collections [1317, 1318]. collective [815, 1756]. Colon [895]. color [1892]. colorful [1066]. COM [1097, 1265]. Command [932, 2104, 1397]. Command-line [2104]. commands [615]. Commerce [620, 629, 633, 2081]. Commercial [841]. commodity [961]. commons [1407]. Communicating [1197]. Communication [40, 1030, 525, 891]. Communications [1775, 1998]. Community [577, 578]. Compact [1005, 916]. Companion [339, 1430, 1431, 1432, 325, 543, 560, 567, 1080, 1081, 1650, 1861, 1862, 1082]. company [1066]. compelling [435]. Complete [557, 1012, 1303, 1442, 1443, 1822, 1823, 772, 876, 1229, 1411, 2033]. completely [308]. compliant [58]. Component [1097]. Components [1560, 1561, 1098, 1099, 1925]. composite [2088]. Comprehensive [684, 895, 967]. CompTIA [336]. compulsive [2026]. Compute [1891]. Computer [422, 605, 660, 868, 1013, 1457, 1704, 1847, 1978, 69, 115, 116, 164, 210, 692, 772, 998, 1456, 1853, 2032]. Computers [855, 695]. Computing [67, 419, 421, 490, 1110]. concentré [1167]. concepts [85, 968, 1000, 1714]. concise [1599]. Concurrency [190, 213]. concurrent [77]. Conference [68, 2121, 2122, 2127, 2128, 2129, 1429, 2131, 2132]. Conferences [1485, 1486]. Conferencing [1416, 1449]. Configuration [38, 1001, 1945, 1946, 284, 636, 1835]. configure [1698]. Connected [420, 2079]. Connecting [485, 486, 2153]. Construction [207, 258]. Consuming [1697]. Contact [1197]. contactless [891]. Containerize [1636]. Content [765, 1794, 1882, 2112]. Continent [3]. continuing [397]. Continuity [2157]. Control [86, 179, 19, 109, 302, 306, 307, 573, 574, 695, 945, 946, 1029, 1510, 1916, 1169]. controlling [1758]. Controls [398]. Conversion [1266]. Cookbook [52, 209, 277, 289, 371, 372, 432, 658, 1083, 1152, 1162, 1729, 1796, 1820, 2103, 12, 30, 48, 49, 50, 104, 127, 128, 130, 176, 196, 2118, 237, 262, 290, 296, 297, 300, 2120, 368, 370, 373, 407, 408, 433, 434, 458, 488, 636, 659, 685, 687, 690, 756, 763, 826, 827, 835, 838, 843, 851, 859, 864, 933, 947, 1071, 1084, 1085, 1086, 1150, 1151, 1164, 1247, 1267, 1356, 1407, 1419, 1455, 1477, 1500, 1669, 1730, 1732, 1734, 1735, 1736, 1765, 1766, 1790, 1798, 1839, 1873, 1906, 1908, 1966, 1999]. cool [487, 1468, 1903]. Core [435, 751, 985, 1269, 1271, 1273, 1720, 148, 1188, 1274, 1631, 839, 1827]. Course [388, 389, 304, 1173, 1174]. Covers [2]. Cracking [461]. craft [682]. crash [1173, 1174]. create [1476]. créatifs [2119]. Creating [192, 2119, 435, 801, 1034, 1128, 2151, 1613, 1641, 1678, 1697, 803, 804, 2126, 1882]. Creative [53, 464]. Crime [868]. Crimefighter [868]. cross [876, 688]. cross-reference [876]. Crossing [477]. crunching [2075]. Crusade [2071]. Cryptography [954, 1998, 1999]. CS [1189, 1190]. CS2 [4, 844, 1191]. CS3 [6, 657, 905, 906, 1194, 1195, 1206, 1766, 1805, 1984]. CS3/CS4 [906]. CS4 [906, 1196, 1208, 1862]. CS5 [1209]. csh [430]. CSS [2117, 57, 597, 684, 876, 1073, 1204, 1237, 1238, 1242, 1729, 1730, 1731, 1732, 1762, 1829]. cult [901, 902]. Cultivating [2067]. culture [1458]. CURL [1733]. Currency [1878]. curses [1863, 1864]. custom [2047]. Customer [2067]. Customizing [2100]. Customs [62, 63, 303, 353, 355, 471, 472, 534, 593, 672, 736, 785, 786, 816, 899, 900, 925, 994, 1148, 1258, 1260, 1371, 1696, 2080]. cuts [405]. CVS [1583, 1584, 1916, 1996, 1997]. cyber [263]. Cython [1808].

D [955, 1157, 1580]. DAM [977]. Dancing [2060]. Danger [1444]. DAO [497]. darkroom [464, 1244]. Darkstar [243]. dashboard [525]. Data [898, 923, 969, 1083, 1215, 1507, 1613, 1672, 1738, 1775, 1982, 2075, 176, 313, 327, 525, 619, 852, 953, 972, 1220, 1264, 2065]. Database [375, 630, 631, 970, 1507, 1624, 1626, 1679, 1684, 2070, 716, 922, 1263, 1628, 1676, 2072, 2073, 2101]. databases [1158, 1622]. date [2057]. DBA [974, 1633]. DBAs [241, 441, 1320]. DBI [391]. DCE [849, 1689, 1690, 1772, 1773]. DCOM [1909]. DDD [1171]. Deals [189]. death [630, 631]. Debian [437, 968, 1126, 1180]. Debug [1011]. debugger [583]. debugging [1171]. decisions [1460]. Deep [1507, 1608]. Deering [1145]. defenses [1396]. Definitive [2139, 103, 107, 193, 497, 539, 542, 549, 641, 645, 647, 656, 683, 684, 689, 745, 751, 969, 995, 996, 997, 1236, 1282, 1283, 1292, 1307, 1308, 1309, 1310, 1311, 1324, 1325, 1326, 1406, 1670, 1688, 1824, 1932, 2021, 1993, 217, 218, 284, 287, 347, 387, 562, 570, 635, 642, 651, 686, 821, 832, 840, 1066, 1093, 1239, 1241, 1293, 1312, 1327, 1328, 1475, 2135, 1828, 1835, 2059, 2063]. Degunking [90, 91, 92, 443, 831, 1471, 1517, 1803]. Delphi [1067]. deluxe [1587]. deploy [294, 1010]. deploying [578]. deployment [1093]. depth [375, 1611]. Describing [1613]. Deserve [928]. Design [388, 461, 478, 479, 537, 1140, 1254, 1337, 1340, 1341, 1679, 1684, 1882, 40, 125, 166, 258, 334, 335, 525, 2124, 966, 989, 1225, 1322, 1338, 1343, 1345, 1346, 1411, 1582, 1676, 1713, 2137, 2106]. designer [136, 2119, 1347]. designers [2120]. Designing [132, 267, 268, 406, 575, 730, 764, 803, 804, 845, 907, 960, 1261, 1336, 1695, 1710, 1930, 332, 1333, 1728, 1762, 2090]. desk [876]. Desktop [323, 445, 538, 554, 556, 588, 614, 665, 667, 726, 779, 985, 1172, 1341, 1414, 1420, 1479, 1482, 1594, 1647, 1685, 1814, 2002, 2045, 2046, 2074, 13, 39, 185, 215, 285, 358, 2120, 383, 384, 426, 427, 480, 481, 540, 547, 555, 563, 564, 569, 668, 671, 783, 805, 950, 1067, 1269, 1270, 1274, 1323, 1338, 1413, 1448, 1402, 1498, 1505, 1506, 1574, 1580, 1598, 1683, 1737, 1763, 1783, 1785, 2136, 1827, 1920, 1947, 1970, 1971, 358]. desperate [1107]. desukutoppu [1869]. detection [1609]. Deutsch [1143]. Deutsche [550, 551, 552]. Developer [124, 517, 1319, 1440, 1408, 1634, 10, 82, 436, 467, 643, 796, 990, 1007, 1049, 1130, 1170, 1221, 1635, 1637, 1896, 2035, 2038]. Developers [194, 214, 276, 371, 372, 488, 847, 1127, 195, 275, 379, 393, 438, 752, 2027]. Developing [473, 475, 489, 660, 734, 766, 1265, 1560, 1561, 1677, 1688, 2047, 842, 1473, 1733]. Development [170, 825, 2154, 100, 401, 1315, 1488, 1515, 1776, 1917, 1918, 2113, 2114]. Developmental [2024]. Device [279, 1702, 1703, 318]. Devices [1887, 1232]. DHCP [38]. DHTML [685, 687, 989]. Diagnosis [894, 2024]. Dictionary [477, 1775, 972]. difficult [1468]. Digital [610, 933, 1080, 1081, 1082, 1243, 1245, 1480, 1546, 1855, 1856, 1857, 1858, 1859, 2077, 4, 5, 127, 128, 381, 464, 851, 892, 977, 1574, 1805, 1861, 1892, 1893]. Director [481]. Directory [167, 599, 600, 601, 602, 603, 1095, 392, 47, 48, 51, 52, 864]. disasters [1959]. Discover [109]. discriminating [295]. discussed [882]. Disease [1178]. disorder [2026]. Disorders [2024, 2025]. Disruptive [1253]. Distributed [67, 273, 447, 490, 1507, 730]. Distributing [1690]. ditch [198]. Dive [1507]. diversions [1458]. DNS [31, 32, 33, 34, 35, 36, 37, 992, 993, 1071, 1072]. do [109]. do-it-yourselfer [109]. DocBook [2021]. Doctor [928]. Documentation [124, 1012, 2066, 123, 1905]. Documents [211, 341, 344, 1794]. Does [1891, 1466]. Dog [853]. Doing [1357, 1358, 1359, 1360, 1361, 1362, 1363, 1364, 1365, 1366, 1738, 2064]. Dojo [1705]. DOM [684, 989]. domain [1475, 1476, 35]. domain-specific [1475, 1476]. domination [872]. Domino [1323]. done [1699]. Don’t [704, 2094]. doors [1700]. Down [743]. Download [1011]. downloaders [295]. Dreamweaver [657, 58, 258, 1201, 1202, 1203, 1205, 1206, 1208, 1209, 2074]. Drive [442]. Drive-by [442]. Driven [1519, 2101]. Driver [279]. Drivers [1702, 1703, 318]. driving [216, 1473]. Droid [2146]. drug [83]. Drupal [2140, 245, 1665]. DSLR [1246]. Duntemann [442, 444]. Durchstarten [1557]. DV [79]. DVD [400, 2031]. Dynamic [657, 683, 1805, 173, 435, 803, 804, 846, 38, 224, 684, 686].

e-mail [1493, 1490]. E-world [109]. early [450]. Easy [1745, 109, 2120, 387, 706, 1107, 2031]. easy-to-follow [2120]. easy-to-use [387]. Eating [1842]. eBay [310, 831, 915, 917, 2094]. EC2 [1306]. Eclipse [242, 837, 838, 1171]. economy [1879]. Eddie [1893]. Edge [1967, 418, 1213]. editing [2125]. Edition [539, 720, 771, 981, 1379, 1388, 1420, 1530, 1591, 1592, 1593, 2013, 308, 2123, 667, 668, 722, 1002, 1533, 1544, 1553, 1590, 1538]. editions [398, 1852]. Editor [56, 986, 988, 69, 1649, 1648]. Editors [1663]. Effect [708]. Effective [1034, 1254, 1652, 525, 1599]. Effectively [987]. effects [128, 1574, 1765, 1766]. Efficient [663]. Eight [1709, 1188]. eight-core [1188]. Einführung [1120]. Einsteiger [1642, 95, 1643]. EJB [1779]. Elective [1272, 148]. Electives [1270]. Electronic [349, 599, 600, 601, 602, 603, 484]. Electronics [1641, 109, 2009]. Elektronik [109, 1642]. Elektronik-Projekte [1642]. elektronische [109]. Elements [228, 1192, 1193, 229, 230, 231]. Emacs [252, 257, 253, 254, 255, 256, 678]. Email [987, 1744, 2084, 443]. Embedded [99, 267, 2097, 2099, 2100, 100, 268, 2098]. Embedding [235, 1674, 1675]. Emergent [1967, 418]. Enabled [44]. encodings [770]. Encore [400]. Encounters [744]. Encryption [461]. Encyclopedia [1302, 1303, 1357, 1358, 1359, 1360, 1361, 1362, 1363, 1364, 1365, 1366, 2064, 1574]. Ending [2112]. Engine [1718, 1719]. Engineer [1269]. engineering [196]. engineers [1811]. engines [1473]. Enhanced [1185]. enlargements [1230]. entdecken [109]. Enterprise [281, 456, 491, 563, 825, 961, 1184, 1273, 2155, 2087, 2090, 398, 559, 1160, 1219, 2089, 223, 239, 492, 554, 1276, 1277, 1278, 1280, 1281, 1372, 2061]. entertainment [1711]. Entity [1025, 1026, 1027]. Entries [1865]. environment [718]. environments [171, 401, 1935]. EOS [1081, 1082]. EPUB [637]. ère [769]. erklärt [1557, 882]. ERLANG [77, 937]. Error [489]. essays [1213]. Essential [112, 317, 403, 590, 611, 612, 613, 616, 617, 797, 970, 971, 1175, 1176, 1285, 1286, 1768, 1996, 1997, 2039, 2042, 2057, 57, 460, 1811]. Essentials [21, 22, 273, 336, 462, 712, 713, 747, 1271, 1872, 1910, 1911, 345, 428, 509, 610, 679, 680, 715, 716, 717, 748, 760, 761, 1489, 1503, 1602, 1603, 1830, 1912, 2000]. Estate [936]. Ethernet [1824]. Etiquette [62, 63, 303, 353, 355, 471, 472, 534, 593, 672, 736, 785, 786, 816, 899, 900, 925, 994, 1148, 1258, 1260, 1371, 1696, 2080]. euphonious [334]. European [1142]. evaluation [450]. Even [874]. Event [1304]. every [815, 2034]. everyday [706, 1873, 2075]. Everything [1484, 1842, 1093]. everywhere [1370]. Evil [872, 873]. evolution [397]. exactly [153, 1550, 1555]. Exam [337, 1271, 1272, 1273, 336]. example [85]. Examples [371, 372, 488, 550, 558, 289, 543, 560, 566, 567, 1071, 1714]. Exams [591, 1269, 1275, 148, 149, 1274, 1827]. Excel [2041, 72, 196, 621, 622, 623, 798, 799, 953, 1017, 1129, 1132, 1134, 1135, 1680, 1686, 1727, 1793, 2038]. excellent [682]. Exchange [1970, 1667, 1669]. exemple [1166]. Exim [800]. Expect [1057, 1058]. Expedition [1349]. Experience [575]. Experimental [527]. expert [1243, 1245]. experts [815]. Explained [964, 882, 1557]. exploitation [482]. Explorer [1813]. Explorer‘s [437]. Exploring [13, 1057, 1058, 1348, 1351, 1352]. exposure [381]. express [2030]. expression [1868, 1870]. Expressions [607, 606, 608, 649, 690, 903, 1869, 1870]. Extending [9, 2100]. Extensibility [1677]. Extensible [452, 455]. Extensions [678, 1499]. Extreme [237, 1581, 772, 1708, 292].

F# [1806, 1807]. FaceBook [2110, 28, 29, 1987, 2156]. Facebook-Programmierung [28, 29]. Facelets [1880]. faces [142]. factor [2037]. Fallacy [2158]. Families [1949, 232, 895, 926, 929]. Family [1153, 83]. Fan [160, 875, 1370, 1860, 2007]. Farther [1153]. Fast [168, 170]. Faster [1894]. favorite [622, 792, 1263, 1881]. FBML [1830]. Fearless [189]. Features [516]. Fedora [1184, 1976]. feed [2085]. feeds [766, 1565]. Ferret [93]. fetching [236]. Fi [442, 1698, 444]. field [2109]. Fifth [2158]. Fight [272]. File [1259, 1302, 1303, 1319, 1969, 69, 2029]. FileMaker [298, 1577, 299]. files [657]. Filipino [593]. filmmakers [933]. filmmaking [79, 1157]. filter [1574]. Filtered [86]. filtering [2112]. finances [1517]. Find [1298, 1460]. findability [1288]. Finding [102, 1559, 1706, 2024, 411]. fine [128, 1333]. fine-grained [1333]. finish [79]. Finishing [272]. Firefox [283, 498, 704, 1210]. Firewalls [278, 2116, 1609]. First [1211, 2154, 113, 114, 133, 308, 2124, 597, 710, 723, 1225, 1287, 1515, 1644, 1779, 1780, 1781, 1837, 1838]. Five [3, 64, 2060]. fix [115, 116, 164, 622, 694, 792, 878, 920, 1263, 1646, 1881]. Fixes [205]. Fixing [920, 1263, 1881]. Flash [1283, 1884, 147, 830, 1030, 1086, 1282, 1284, 1293, 1716, 1777, 1983, 1984]. Flashcards [1271, 1272, 1273]. Flex [212, 275, 450, 819, 825, 924, 1087, 1356]. Flickr [120]. Fluent [1599]. fluff [2123]. FO [1481]. follow [2120]. font [19].

Fontes [769]. Fonts [770]. Food [1840, 1842]. Ford [1145]. forensics [897]. form [2037]. Formats [1302, 1303]. Forms [207, 721]. FORTRAN [1091, 934, 935]. Foundation [555, 1759]. Fourth [1429, 874]. FPS [1306]. Frame [211]. FrameMaker [211]. Framework [22, 1910, 1911, 1912, 726, 1005, 1025, 1026, 1027]. frameworks [365, 754]. Francaise [541]. France [899, 1427]. Francisco [1434]. Free [2071, 582, 961, 2043, 1849]. FreeBSD [963, 1012, 1108, 1173]. Freedom [2071]. French [136, 899]. friend [1135]. friendly [1211, 871, 872, 874]. Friends [1949, 926, 929]. frontend [1811]. Frontier [1324]. FrontPage [1154]. FTP [1070]. fulfillment [2061]. full [874]. fun [966]. function [327]. Functional [2027]. Fundamental [701]. Fusion [946]. Future [1145, 1802, 1891, 2158, 852, 931, 1992]. futurist [1213]. fwsnort [1609].

Gadgets [2150]. Galaxy [2147]. gallery [1322]. Game [194, 243, 1229, 1411, 195, 382, 966, 1157]. game-based [1157]. Games [495, 1458, 1473]. Gaming [261, 957, 1711]. GarageBand [1540]. garages [1474]. Gateway [1981]. GDB [1171, 1657]. geek [2061]. Geeks [879, 624, 705, 885, 887, 888, 889, 890, 918, 1369, 1701]. general [1476]. Generation [2158, 5, 308, 365, 395, 1210, 1232, 1316]. Generics [1317, 1318]. geniuses [872]. Geoff [1146]. Geolocation [45]. German [28, 29, 95, 109, 172, 534, 882, 1518, 1557, 1643, 1661, 1722, 1758]. Germany [534]. gestural [1710]. Get [189, 1050, 1153, 1705, 4, 83, 153, 874, 1550, 1555, 1709, 2094]. Getting [135, 206, 420, 468, 819, 869, 891, 928, 1188, 1212, 1336, 2153, 1580, 2009, 2024, 2069, 1714, 1557]. GIF [959]. Gift [708, 742, 744, 528]. GIMP [2125, 1329]. GIS [379]. Glenn [1142]. global [1198, 1146]. Glyphes [769]. GNOME [2035]. GNU [1126, 100, 252, 253, 254, 255, 256, 257, 437, 678, 1092, 1180, 1231, 2015]. GNU/Linux [437, 1180]. Go [424, 938, 1153, 1177, 1600, 1370, 178, 1211, 1636, 1935]. going [694]. Good [627, 1481, 338, 1518]. Google [248, 249, 250, 359, 378, 411, 483, 531, 661, 1002, 1014, 1250, 1251, 1252, 1502, 1718, 1719]. Gopher [1070]. got [589]. gotchas [1874]. Gourmet [1841]. GPL [882, 882]. grained [1333]. Grand [1445]. graphic [2120]. Graphical [1060, 2022, 313]. graphically [40]. Graphics [2143, 277, 495, 1083, 1302, 1303, 1340, 2015, 955, 1346, 1459, 2016]. grassroots [664]. Greasemonkey [1511]. great [11, 866, 867, 2126, 1700, 1709]. GREP [905, 906, 94]. Griff [1758]. Groupware [1979]. grues [874]. GUI [360, 362, 537]. Guide [2139, 62, 63, 103, 107, 124, 187, 193, 276, 303, 309, 337, 353, 355, 471, 472, 474, 486, 516, 534, 539, 542, 549, 593, 609, 632, 640, 641, 645, 647, 648, 656, 672, 675, 689, 736, 745, 751, 773, 785, 786, 789, 816, 899, 900, 925, 930, 932, 943, 944, 969, 978, 979, 980, 981, 994, 995, 996, 997, 1028, 1148, 1178, 1182, 1197, 1236, 1258, 1260, 1261, 1282, 1283, 1292, 1295, 1296, 1307, 1308, 1309, 1310, 1311, 1319, 1324, 1325, 1326, 1340, 1371, 1421, 1423, 1406, 1508, 1520, 1521, 1529, 1589, 1590, 1591, 1592, 1593, 1634, 1657]. Guide [1664, 1670, 1688, 1696, 1772, 1773, 1774, 2155, 1812, 1824, 1847, 1855, 1859, 1932, 1949, 1951, 1964, 2021, 2036, 2080, 1993, 7, 28, 29, 106, 122, 131, 138, 215, 217, 218, 222, 238, 242, 249, 275, 279, 284, 287, 292, 295, 308, 327, 328, 334, 336, 347, 384, 387, 442, 444, 469, 562, 570, 576, 621, 623, 635, 642, 650, 651, 653, 663, 676, 677, 706, 772, 806, 821, 832, 840, 872, 884, 895, 909, 916, 926, 927, 929, 931, 956, 967, 974, 982, 983, 1004, 1005, 1033, 1061, 1093, 1158, 1159, 1211, 1239, 1241, 1293, 1312, 1321]. guide [1327, 1328, 1346, 2135, 1588, 1611, 1622, 1635, 1639, 1699, 1733, 1778, 2137, 1828, 1835, 1856, 1858, 1898, 1913, 1915, 1917, 1925, 1954, 1955, 1958, 2025, 2033, 2035, 2040, 2059, 2063, 2109]. guidebook [705]. GUIs [1156]. guitar [787]. gut [256, 1557, 1871, 105, 739, 1518, 1648, 1653, 2006]. Gutsy [181, 182, 1841].

Hackers [692, 315]. Hacking [395, 482, 624, 960, 1369]. Hacks [272, 999, 1156, 1469, 1846, 15, 57, 80, 81, 119, 120, 121, 147, 150, 175, 248, 250, 261, 293, 348, 409, 410, 411, 460, 483, 484, 530, 579, 580, 581, 604, 661, 693, 696, 699, 750, 798, 799, 813, 817, 818, 915, 917, 957, 975, 1014, 1074, 1075, 1125, 1136, 1210, 1222, 1240, 1314, 1480, 1501, 1504, 1505, 1511, 1604, 1606, 1607, 1711, 1723, 1757, 1767, 1810, 1826, 1845, 1854, 1857, 1924, 1973, 2001, 2018, 2105]. Hadoop [1254, 2063]. Halo [272]. Handbook [110, 111, 172, 211, 414, 415, 416, 788, 789, 868, 1813, 1982, 173]. Handbuch [172]. handheld [1711]. Hands [728, 705, 1574, 2137]. Hands-on [728, 705, 1574, 2137]. Hard [1745, 198]. Hardcore [1791, 1792]. Hardening [20]. Hardware [267, 624, 1775, 1921, 2103, 164, 268, 961, 1920, 1922, 1925]. Harnessing [469, 1253]. Harvard [2131, 2132]. Haskell [1461]. Hawaii [264]. Haystack [1706]. HBase [656]. HD [1541]. HDR [846]. HDRI [173]. Head [113, 114, 133, 2124, 597, 710, 723, 1211, 1223, 1225, 1287, 1515, 1644, 1779, 1780, 1781, 2154, 1837, 1838]. Healthcare [928]. Healthy [1847]. Heart [279]. Heaven [3]. Help [527, 609, 751, 767, 1003, 2024, 83, 2026]. helping [2025]. Hibernate [467, 468, 469]. High [419, 421, 1036, 1750, 1800, 1811, 1847, 2108, 2111, 173, 791, 846, 867, 1230]. high-level [867]. high-performance [791]. high-rolling [1230]. highly [961]. Hinden [1145]. Hints [720]. history [1066]. Hive [260]. Hodgkin [894]. Holding [1177]. holistic [232]. Home [878, 1096, 1222, 1369, 1530, 1538, 1004, 1240, 1321]. honest [2061]. Hong [736, 1358, 1433]. Hooking [1907]. Hospital [927, 930]. Host [38]. hound [697, 722]. HP [1650]. HP-UX [1650]. HTML [110, 111, 236, 597, 683, 684, 686, 801, 876, 1073, 1307, 1308, 1309, 1310, 1311, 1312, 1339, 1340, 1342, 1344, 1346, 1674, 1675, 1762, 1829, 2095]. HTML5 [2145, 1512]. HTTP [689, 2083]. hucksters [1230]. Huitema [1145]. Humor [991]. Huston [1146]. Hydrocephalus [1949]. hyogen [1869]. Hypertext [2083].

I/O [775, 776, 784]. iBook [1860]. IDE [242, 1677]. ideas [487, 692]. ideation [1093]. Identity [1967, 418, 2077]. idiomatic [178]. IDS [329]. iDVD [1532, 1537, 1541, 1547, 1551]. IFC [1506]. IFS [1259]. Ihre [1661]. iLife [1542]. illuminating [334]. illustrated [967, 1262, 1853]. Illustrations [136, 136]. im [1758]. image [2125]. ImageReady [2]. images [6, 1606, 1965]. imaging [173]. imake [429, 431]. IMAP [1300]. IMovie [1532, 1537, 1541, 1547, 1522, 1525, 1551]. implementation [1476, 2106]. Implementations [1812]. Implementing [577, 958]. import [2065]. improve [1874]. improvement [1503]. in-depth [1611]. Incident [1995]. including [1722]. independence [909]. independent [933]. InDesign [1189, 1195, 2120, 844, 903, 904, 905, 906]. India [589, 900, 1428]. Indian [900]. indirect [2109]. Indonesia [303]. Indonesian [303]. Industrial [65, 999, 119, 147, 150, 248, 579, 580, 693, 798, 813, 915, 975, 1314, 1604, 1810, 1846, 1973]. industrial-strength [119, 147, 150, 248, 579, 580, 693, 798, 813, 915, 975, 1314, 1604, 1810, 1846, 1973]. Industry [1415]. Information [98, 101, 102, 525, 1070, 1110, 1143, 1694, 1695, 411, 1109]. informed [1460]. Infrastructure [1299]. initialization [973]. injection [1396]. innovation [145, 2090]. input [1999]. Ins [1677, 514, 1349]. Insider [187]. Insight [1127]. install [1698]. Installable [1259]. Integrated [276]. Integrating [755, 1727, 1577]. integration [762]. Intel [1409, 1631]. Intellectual [1064]. Intelligence [2158, 1756]. Intelligent [1608]. intelligente [1509]. Interactive [1057, 334, 1058, 1913]. interactivity [1777]. Interface [1582, 435]. Interface-oriented [1582]. Interfaces [1060, 2022, 1710, 1930]. Intermediate [1749]. Internal [8]. Internals [1319, 1507, 137, 1470]. International [1429]. Internationalization [385]. Internet [35, 485, 967, 1145, 2131, 2132, 1733, 1813, 2032, 61, 78, 167, 278, 308, 420, 486, 598, 628, 634, 638, 694, 695, 751, 881, 897, 936, 978, 979, 980, 981, 982, 983, 1070, 1303, 1367, 1368, 2153, 1549, 1594, 1688, 1740, 1760, 1764, 1822, 1823, 1979, 2028, 2084, 2116]. interviews [1142, 1143, 1144, 1145, 1146, 1147]. Intrinsics [1378, 1379, 1380, 1388, 1422, 536, 1392]. Introducing [59, 424, 1397]. Introduction [212, 1408, 1658, 1681, 1975, 226, 842, 846, 960, 1853]. invaders [695]. invasions [1978]. Inventor [2138]. investing [150]. IOS [185, 186, 408, 1754, 43, 44, 45]. IP [1147, 89, 185, 854, 855, 856, 857, 858, 967, 1149, 1272]. iPad [44, 162, 2141]. iPhone [41, 42, 44, 295, 435, 1552, 1829, 2113, 2114]. iPhoto [1526, 1531, 1536, 1543, 1548]. IPod [159, 157, 158, 160, 161, 163, 901, 1006, 1370, 1845]. iptables [1585, 1609]. IPv6 [747, 748, 1301]. IRC [1314]. Island [264]. Israel [1696]. Israeli [1696]. Issue [64, 65, 501, 502, 2121, 503, 504, 505, 2122, 1381, 2127, 1384, 1385, 1386, 2128, 1389, 1390, 1391, 2129, 1393, 1417]. Italian [672]. Italy [247, 672]. iTunes [157, 159, 1845]. iWork [466].

J2EE [335]. J2ME [1947]. jōhō [1109]. Jabber [9]. Jaguar [720]. Jakarta [269, 270, 271, 1407, 1790]. Japan [2158, 472, 655, 1359]. Japanese [472, 1109, 1110, 2157]. Java [1354, 1792, 1871, 10, 67, 87, 234, 235, 237, 243, 280, 294, 309, 361, 364, 370, 371, 372, 373, 374, 382, 385, 390, 451, 456, 457, 469, 473, 476, 487, 490, 491, 492, 538, 540, 541, 543, 547, 544, 545, 546, 548, 550, 551, 552, 554, 555, 556, 557, 558, 559, 560, 561, 563, 564, 566, 567, 568, 569, 594, 595, 701, 702, 703, 729, 730, 752, 773, 774, 775, 776, 777, 778, 782, 784, 791, 829, 860, 862, 863, 879, 898, 954, 955, 1061, 1103, 1104, 1141, 1155, 1216, 1217, 1218, 1219, 1220, 1221, 1224]. Java [1226, 1232, 1234, 1235, 1279, 1317, 1318, 1348, 1349, 1350, 1351, 1352, 1353, 1355, 1398, 1399, 1400, 1401, 1403, 1405, 1406, 1410, 1499, 1500, 1506, 1573, 1624, 1625, 1626, 1628, 1638, 1769, 1770, 1780, 1781, 1791, 1792, 1801, 1869, 1870, 1894, 1895, 1948, 2027, 2075, 2115]. JavaBeans [239, 223, 474, 475, 1276, 1277, 1278, 1280, 1281, 1372]. JavaHelp [1034]. JavaScript [1642, 60, 78, 209, 2143, 338, 460, 539, 542, 549, 553, 562, 565, 570, 684, 685, 687, 803, 804, 2149, 876, 904, 989, 1207, 1287, 1567, 1611, 1829, 1831, 1913, 2108, 28, 276]. JavaServer [139, 140, 141, 142, 143]. JBoss [1159, 1637]. JDBC [87, 88, 1624, 1625, 1626]. JDS [13]. Jeff [1144, 442, 444]. Jiao [1664]. Jini [1402]. Job [1559, 589]. jobs [1661]. John [1144]. Johnson [892]. Journal [499, 64, 65, 1417, 1429, 1457, 1458, 1459]. journalism [664]. jQuery [488]. JRuby [458]. JSON2 [224]. JSP [113, 114, 1500, 2047]. Julia [423, 998]. July [68]. jumpstart [1008]. Jun [1147]. Juniper [1812]. JUnit [131, 860]. JUNOS [636, 1160, 2155]. Just [198, 2061, 2123]. JXTA [1404]. Jython [1489].

Karn [1147]. keep [709]. Kerberos [635]. Kernel [201, 202, 199, 200, 203, 960, 976]. Key [1417]. keys [1882]. Kickstart [1179]. Killer [382, 1156]. kinds [1474]. Kingdom [786]. kings [1230]. Kit [207, 771, 2013]. Kite [688]. Klassische [1661]. Knoppix [1604, 1605, 1607]. know [815, 1093, 2034]. knowledge [1811]. Kochbuch [1509]. Kommandozeilenvirtuosen [399]. kommentiert [882]. Kong [736, 1358, 1433]. Konzept [364]. Korea [654, 925]. Korean [925, 1110]. Korn [399, 69, 1691, 1692, 1693]. Korn-Shell [399]. Kowack [1142]. Kurz [1653, 105, 256, 739, 1518, 1648, 1871, 2006]. Kylix [1899].

Land [477, 1601]. Language [165, 168, 425, 702, 703, 1078, 1476, 1608, 1685, 1720, 515, 521, 522, 524, 571, 839, 865, 1397, 1557, 1683, 1687, 2112, 452, 455]. Languages [935, 1832, 1475, 1476, 1898]. LANs [406]. Large [406, 1695, 1488]. Large-Scale [406, 1695]. larger [232]. Latches [8]. Law [881]. Lawyers [881]. LDAP [265]. LDraw [328]. LDraw.Org [328]. Lead [1316]. Learn [1516, 1188]. learned [316]. Learning [40, 41, 69, 84, 129, 178, 210, 244, 252, 253, 257, 327, 389, 376, 380, 529, 532, 618, 794, 795, 986, 988, 1038, 1041, 1050, 1051, 1052, 1053, 1077, 1114, 1117, 1121, 1122, 1123, 1124, 1126, 1141, 1180, 1181, 1182, 1183, 1184, 1248, 1290, 1291, 1330, 1331, 1332, 1340, 1346, 1353, 1354, 1355, 1495, 1496, 1567, 1572, 1608, 1613, 1616, 1623, 1663, 1678, 1691, 1692, 1693, 1716, 1739, 1741, 1742, 1745, 1746, 1748, 1751, 1778, 1797, 1851, 1852, 1886, 1900, 1901, 1902, 1904, 1909, 1942, 1943, 1944, 2022, 315, 657, 1777, 1805, 1898]. Legal [167]. Lego [1721, 328, 956]. leicht [109]. Leopard [19, 302, 1553, 1958]. Less [991]. lesson [657]. Lessons [1671]. Letting [1177]. leukemia [926, 929]. Level [190, 867]. lex [1031, 1168]. Libraries [1143, 487, 2047]. Library [537, 548, 561, 1111, 1440, 210]. Licensing [1849]. Life [160, 743, 911, 1445, 247, 528, 1974, 2060, 2061, 1601]. lifecycle [1105]. Lighter [1894]. Lighting [381]. Lightroom [5, 1244]. Like [422, 998]. line [2104]. Lingo [480]. Links [1303]. LINQ [25, 852]. lint [369]. Linux [363, 1432, 1642, 1661, 13, 14, 104, 105, 106, 117, 118, 122, 137, 199, 200, 201, 202, 203, 215, 216, 225, 246, 318, 383, 437, 526, 579, 662, 705, 706, 805, 806, 807, 885, 886, 943, 944, 961, 976, 1094, 1126, 1180, 1181, 1182, 1183, 1184, 1185, 1431, 1505, 2134, 1585, 1606, 1609, 1617, 1618, 1619, 1641, 1702, 1703, 1734, 1735, 1736, 1783, 1785, 1786, 1788, 1789, 2136, 1800, 1803, 1804, 2137, 1818, 1873, 1963, 1964, 1965, 1969, 1976, 1980, 2001, 2023, 2034, 2044, 2049, 2050, 2051, 2052, 2053, 2054, 2097, 2098]. Linux-basierte [1642]. Linux-Powered [1641]. Linux-Tasks [1661]. Listproc [1743]. Lists [1743, 1754]. LISTSERV [1743]. Lite [339]. little [1960]. Live [232, 1852]. Living [1178, 1177]. Load [197, 2111]. Local [86, 578]. Locating [1794]. Location [44]. Locks [8, 190]. Logging [1304]. Logons [877]. longer [232]. Look [1481, 1460]. Lösungen [399]. Lotus [1323]. loud [817]. lousy [589]. Love [1559, 2096, 316, 2061]. Loving [1559]. Low [190, 867]. Low-Level [190, 867]. LPI [383, 2134]. Luxury [2067]. LWP [236]. Lymphomas [894]. Lynch [1143].

Mac [90, 91, 112, 377, 409, 410, 435, 573, 644, 720, 722, 879, 887, 888, 889, 890, 902, 1009, 1035, 1214, 1321, 1523, 1527, 1528, 1533, 1534, 1544, 1545, 1553, 1646, 1701, 1708, 1834, 1900, 1901, 1902, 1904, 1950, 1951, 1953, 1954, 1955, 1956, 1958, 1969]. Machine [315, 361, 364, 1234, 1235, 866, 1853]. Machines [1891, 1887]. machinima [1157]. Macintosh [96, 97, 415, 477, 1028, 2008]. Macintosh/Windows [477]. Macromedia [1884]. Macros [1678, 1680, 1681, 1686]. MacRuby [2139]. Macs [1409]. made [706, 2031]. MAGIX [1587]. Mail [599, 600, 601, 602, 603, 734, 800, 1493, 1490]. Mailing [1743]. Maintenance [1001]. Majordomo [1743]. Make [1231, 1418, 1888, 1889]. Makeovers [1724, 1708]. Makers [1474]. Making [644, 870, 894, 1460, 1481, 1601, 1745, 1829, 2020, 528, 1474]. Malaysia [62]. Malaysian [62]. Malicious [727, 695]. Mamas [182]. Man [1447]. Management [251, 588, 883, 1001, 1062, 1305, 1499, 144, 304, 977, 1063, 1107, 1700, 1758, 1836, 1892]. Manager [342, 1259, 590]. Managing [38, 89, 171, 329, 877, 1070, 1179, 1231, 1300, 1418, 1454, 1612, 1627, 1666, 1667, 1668, 1743, 1819, 1843, 1844, 1876, 1888, 1889, 1941, 1980, 2089]. manifesto [1213]. Manual [96, 97, 2141, 163, 340, 342, 343, 333, 465, 500, 506, 536, 639, 2146, 2147, 2148, 808, 809, 810, 811, 812, 965, 2151, 1201, 1202, 2152, 1373, 1374, 1375, 1376, 1377, 1378, 1379, 1380, 1382, 1383, 1387, 1388, 1392, 1395, 1422, 1522, 1523, 1524, 1525, 1526, 1528, 1530, 1534, 1556, 1725, 1726, 1890, 1994, 2156, 2066, 151, 152, 157, 158, 159, 161, 162, 221, 222, 229, 230, 231, 282, 298, 299, 310, 311, 466, 681, 731, 733, 1096, 1128, 1129, 1131, 1132, 1133, 1134, 1135, 1154, 1204, 1205, 1207, 1246, 1252, 1472, 1532, 1533, 1535, 1537, 1538]. manual [1539, 1540, 1541, 1542, 1543, 1544, 1545, 1547, 1548, 1549, 1551, 1552, 1553, 1555, 1610, 1983, 1984, 1985, 1986, 1987, 2008]. MAPI [734]. Mapping [484, 1262]. MapPoint [1928]. MapReduce [1254]. maps [483, 661]. mark [1601]. marketing [2110]. Markup [452, 455]. Mason [1674, 1675]. master [1721]. Mastering [76, 592, 606, 607, 608, 711, 719, 963, 1060, 1088, 1255, 1256, 1282, 1456, 1521, 1934, 5, 130, 1931]. Masterminds [155]. masters [1792]. Mathematica [1152]. Matter [156]. Maven [1093, 1170]. MCSD [591]. MCSE [148, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1827]. Me [912, 1524]. measure [109]. measurement [1504]. Media [1716, 259, 664]. medical [1460]. Medium [1336]. Memory [8]. Message [1279, 1638, 824]. Messages [489]. Messaging [734, 969, 9]. messen [109]. Metastatic [1178, 1177]. Methods [390]. Metro [726]. Mexican [1148]. Mexico [1148, 1360, 1426]. MH [1490, 1493]. microcontroller [1188]. Microcontrollers [2153]. microformats [1875]. microprocessors [1853]. microservices [1333]. Microsoft [1758, 198, 337, 414, 590, 1269, 1471, 1478, 1667, 1758, 1774, 1813, 1970, 2043]. Midst [1891]. Migrating [934, 935]. Mike [1146]. Mind [750, 1826]. Mindstorms [1721, 956]. Mining [1706, 1707]. Misadventure [991]. mishaps [1959]. Missing [96, 97, 2141, 163, 333, 465, 2146, 2147, 2148, 2151, 1201, 1202, 2152, 1522, 1523, 1524, 1525, 1526, 1528, 1530, 1534, 1556, 2156, 151, 152, 157, 158, 159, 161, 162, 221, 222, 229, 230, 231, 282, 298, 299, 310, 311, 466, 681, 731, 733, 1096, 1128, 1129, 1131, 1132, 1133, 1134, 1135, 1154, 1204, 1205, 1207, 1252, 1472, 1532, 1533, 1535, 1537, 1538, 1539, 1540, 1541, 1542, 1543, 1544, 1545, 1547, 1548, 1549, 1551, 1552, 1553, 1555, 1610, 1983, 1984, 1985, 1986, 1987, 2008]. Mitchell [1261]. mixing [1845]. Mobile [727, 1322, 869, 1232, 1711, 1147]. MobileMe [945]. mod perl [134, 585, 1833]. Modding [1708, 772]. Model [497, 1677, 1065, 2065]. Modeling [969, 797]. modern [989, 1264, 1622]. moderne [1120]. modes [127]. Modules [1833, 1746]. Monad [1397]. money [1601, 1712]. MongoDB [287, 288, 2076]. Mongrel [402]. monitor [294]. monitoring [108, 405]. Mono [436]. Monterey [68]. Moodle [304, 305]. Mosaic [414, 415, 416]. Most [720, 4, 115, 116, 164, 528, 622, 694, 792, 878, 920, 1066, 1263, 1646, 1709, 1881, 1978]. Mother [183]. Motif [1593, 500, 506, 537, 810, 812, 1379, 1388, 1590, 1592]. movement [1622]. Mozilla [192, 463]. MP3 [745]. MPI [1800]. MSH [1397]. mSQL [2102]. MTA [387]. Multi [883, 1631, 171]. Multi-core [1631]. Multi-Platform [883, 171]. Multimedia [495, 1963, 1964, 1606, 1728]. multiple [842].

Multithreaded [301]. Murai [1147]. music [644]. Musings [1617, 1618, 1619]. MX [1203, 220, 1202, 1283, 1284]. My [345, 589, 573]. MySQL [380, 432, 433, 434, 448, 449, 791, 1470, 1627, 1629, 1630, 1750, 1886, 2066, 2070, 2072, 2073, 2101, 2102, 2111]. mysteries [279]. myths [145].

Nagios [108, 405]. Name [35]. NaN [1664]. NAS [1570]. nation [630, 631]. Native [390, 1935]. Natural [165, 1608]. Navigation [575]. Need [1484, 153, 1093, 1550, 1555]. Needles [1706]. needs [450]. Neogeography [1975]. Nepal [939]. Nero [2031]. Net [881, 1671]. Net8 [1945, 1946]. NetActivism [1740]. NetBeans [193]. Netcraft [620]. Netduino [2009]. NetLearning [1760]. Netnews [789]. NetResearch [102]. Netscape [1506]. NetSuccess [936]. NetTravel [1764]. NetWare [1919]. Network [394, 404, 2144, 751, 773, 777, 778, 854, 856, 857, 858, 943, 944, 1074, 1075, 1146, 1227, 1228, 1305, 1595, 1799, 1998, 2079, 108, 122, 137, 313, 405, 509, 574, 782, 861, 878, 1301, 1715, 149]. Networked [1143]. Networking [246, 855, 1142, 1147, 1271, 1446, 1451, 1967, 284, 418, 878, 1096, 1698, 1735, 1736, 1835, 1999]. Networks [89, 577, 578, 598, 599, 600, 601, 602, 603, 641, 855, 1144, 1752, 1753, 1981, 185, 642, 869, 2106, 1812]. neue [1557]. never [1699]. News [1744]. Next [365, 5, 308, 395, 1210]. Next-generation [365, 5, 1210]. NFC [891]. NFS [1843, 1844]. NG [1989]. Nicholas [708]. Nihongo [1109]. NIO [829]. NIS [1843, 1844]. No [2123, 991, 1315, 198]. Node [2149, 2076]. Node.js [847]. Noel [1145]. Nokia [2105]. Non [894, 705]. non-geeks [705]. Non-Hodgkin [894]. nonsense [1315]. NoSQL [1622]. Notable [98]. Notebook [437, 10, 136, 2119, 436, 467, 643, 796, 990, 1007, 1049, 1130, 1170, 1221, 1347, 1637, 1896, 2038]. Novell [588]. NSA [1185]. NT [857, 1272, 1273, 1305, 1919, 36, 613, 614, 877, 1001, 1003, 1233, 1304, 1319, 1367, 1368, 1442, 1443, 1482, 1666, 1690]. NT/2000 [1367, 1368]. numérique [769]. Numpy [214]. NUnit [759]. Nutshell [17, 39, 185, 285, 357, 384, 398, 445, 480, 491, 538, 540, 541, 547, 545, 546, 550, 551, 552, 554, 556, 558, 591, 665, 666, 667, 669, 670, 674, 714, 721, 779, 805, 807, 914, 932, 985, 1067, 1078, 1172, 1269, 1270, 1275, 1337, 1341, 1412, 1414, 1420, 1441, 1404, 1482, 1594, 1647, 1659, 1662, 1685, 1763, 1782, 1793, 1814, 1815, 1921, 2045, 2046, 2074, 16, 23, 26, 31, 32, 33, 69, 126, 148, 149, 186, 191, 358, 383, 426, 427, 446, 448, 449, 481, 492, 543, 555, 560, 563, 564, 566, 567, 568, 569, 668, 671]. nutshell [698, 757, 780, 781, 783, 802, 806, 850, 872, 919, 950, 951, 952, 976, 1035, 1068, 1079, 1138, 1158, 1163, 1165, 1214, 1274, 1323, 1338, 1345, 1413, 1448, 2133, 1479, 1498, 1506, 1514, 1576, 2134, 1598, 1650, 1683, 1737, 1783, 1785, 1788, 1789, 2136, 1816, 1827, 1885, 1920, 1922, 1947, 1948, 1953, 1956, 1970, 1971, 1972, 544].

O [775, 776, 784]. OAuth [206]. Object [497, 758, 1000, 1115, 1225, 1713]. Object-Oriented [758, 1115, 1000, 1225, 1713]. Objective [376, 440, 1829]. Objective-C [376, 440, 1829]. Objects [711, 898, 1672, 1746]. Observability [246]. obsession [18]. Obsessive [2026].

Obsessive-compulsive [2026]. Obstacles [1771]. O’Dell [1146]. Odyssey [3, 688]. office [1004, 1321, 2040, 96, 97, 731, 1015, 1018, 1471, 2008]. Official [123, 124, 1905, 328, 706, 2035]. one [1189, 1190, 1191, 1192, 1193, 1194, 1195, 1196]. one-on-one [1189, 1190, 1191, 1192, 1193, 1194, 1195, 1196]. ones [1466]. Online [46, 102, 150, 1143, 694, 1198, 1978, 187, 384]. OO [1120]. OO-Programmierung [1120]. Open [396, 1064, 2089, 304, 487, 582, 2125, 869, 1185, 1619, 2009, 2113, 2114, 358, 396, 397, 439, 1617, 1618, 1849, 1991]. open-source [487]. OpenBSD [963, 1106, 1174]. OpenBTS [869]. OpenCV [210]. openMosix [1800]. OpenOffice.org [2043]. OpenSSL [1998]. Operating [1883, 1942, 1943, 69, 1495, 1496, 1944]. Operations [590, 1146]. OPML [135]. Opportunities [1771]. optimal [909]. Optimization [2062, 942, 2111]. Optimizing [495, 519, 1249]. Options [894]. ORACLE [319, 87, 240, 241, 439, 441, 447, 478, 510, 511, 512, 513, 514, 515, 516, 517, 518, 519, 520, 521, 522, 523, 524, 645, 646, 647, 648, 649, 651, 652, 711, 712, 713, 714, 715, 716, 717, 740, 741, 893, 970, 971, 972, 973, 974, 984, 1076, 1139, 1199, 1200, 1249, 1255, 1256, 1257, 1294, 1320, 1450, 1408, 1572, 1573, 1633, 1914, 1945, 1946]. Oracle8 [479, 712, 713]. Oracle8i [8, 516, 712]. Oracle9 [713]. O’Reilly [46, 208, 620, 886, 1771]. O’Reilly/Netcraft [620]. Organ [528]. Organizers [1521]. Oriented [758, 1115, 1000, 1225, 1582, 1713]. OSCAR [1800]. OSF [810, 1379, 1423, 1590, 1592, 1593]. OSF/1 [1423]. OSF/Motif [810, 1592, 1593, 1379, 1590]. Other [1254, 606]. Our [1891, 1887]. Outfitting [1631]. Outlook [1758, 356, 676, 1020, 1758, 1885]. Overnight [309]. Overview [214, 1415, 1967, 418, 1714]. Own [1484, 1485, 58, 1073, 1476, 1486, 2101].

P2P [418, 1415, 1967]. P3P [331]. Pack [1442, 1443]. Packaged [2088]. Packages [511, 513]. packet [1715]. Pages [141, 139, 143, 236, 803, 804, 1913, 140, 1261]. pain [60]. painters [692]. Painting [1568]. Palm [1529, 833, 1136, 1521, 1634, 1635, 2079]. PalmPilot [1520, 1521]. panoramic [177]. Panther [112, 377, 410, 722, 888, 1533, 1834, 1902, 1953, 1954]. Paper [991]. Parallel [1139]. Parallelism [1631]. parameters [973]. paranoid [1106]. parent [884]. Parents [930, 927]. Paris [1438]. Parrot [1603]. parsing [236]. parts [338]. Passion [184]. passive [2109]. Passport [62, 63, 303, 353, 355, 471, 472, 534, 593, 672, 736, 785, 786, 816, 899, 900, 925, 994, 1148, 1258, 1260, 1371, 1696, 2067, 2080]. patches [823]. Patients [1949, 232, 895]. pattern [1322]. Patterns [1140, 1254, 166, 335, 2124, 1322, 1476, 1713, 1831]. PayPal [1810]. PC [80, 92, 115, 116, 164, 695, 1775, 1854, 1920, 1921, 1922, 1923, 1925, 1926, 1927, 1959]. pcket [275]. PCs [772, 1610, 2037]. PDF [1846]. PDFs [1697]. Peer [1253]. Peer-to-Peer [1253]. Penny [2067]. people [332, 664, 1474]. perfect [1923, 1925, 1926]. Perforce [2078]. Performance [246, 319, 419, 421, 609, 740, 940, 941, 1090, 1313, 1406, 1769, 750, 791, 1036, 1249, 1750, 1770, 1800, 1811, 2108, 2111]. Perl [1457, 1674, 1871, 2011, 68, 171, 233, 236, 251, 274, 289, 290, 293, 314, 330, 389, 391, 441, 583, 592, 606, 737, 771, 791, 839, 886, 990, 1060, 1452, 1456, 1457, 1458, 1459, 1468, 1479, 1602, 1603, 1614, 1615, 1675, 1739, 1741, 1742, 1745, 1746, 1748, 1749, 1751, 1782, 1784, 1787, 1825, 1833, 1869, 1870, 1933, 1934, 2002, 2003, 2004, 2005, 2006, 2012, 2013, 2014, 2015, 2016, 2022, 2082, 2010]. Perl/TK [1459, 1060, 2022]. persistence [469]. Personal [855, 70, 71, 72, 73, 74, 75, 115, 116, 356, 1517, 1978]. Pervasive [2024]. pest [695]. Peter [1143, 1145]. PEXlib [1890]. PGP [627]. PHIGS [639, 965]. Phil [1147]. Philippines [593, 1364]. phone [18]. Photo [1347, 1006, 1765, 1766]. photographers [5, 127, 128, 173, 381, 2125, 851, 977, 1574, 1862]. Photography [1855, 1859, 846, 892, 1243, 1245, 1546, 1857, 1858, 1861, 1892, 1893]. Photomontages [2119]. photos [177]. Photoshop [1190, 1191, 1192, 1194, 1196, 1805, 1, 2, 4, 5, 6, 127, 128, 136, 228, 229, 230, 231, 2119, 841, 851, 1193, 1347, 1412, 1414, 1574, 1765, 1766, 1862, 1893]. PHP [1871, 28, 233, 350, 380, 393, 791, 818, 847, 850, 1000, 1021, 1022, 1023, 1024, 1315, 1488, 1733, 1768, 1796, 1797, 1798, 1869, 1870, 1962, 2070, 2072, 2073, 2101]. PHP/CURL [1733]. PHPUnit [138]. Phrasebook [477]. Physics [194]. Pi [1643, 1643]. Pick [1295]. pills [1230]. Pilot [1521]. Pincher [2067]. Pipes [1579]. pitfalls [922]. PL [518, 1450, 510, 511, 512, 514, 515, 516, 517, 519, 520, 521, 522, 523, 524, 1320, 1408, 1572]. PL/SQL [518, 1450, 510, 511, 512, 514, 515, 516, 517, 519, 520, 521, 522, 523, 524, 1320, 1408, 1572]. Planet [189, 1447]. Planning [578, 1771, 1065]. Platform [883, 1967, 171, 418]. Platforms [477, 842, 1882]. play [109]. playing [2030]. Playlist [160]. PMP [710]. PNG [1670]. Pocket [62, 63, 94, 254, 255, 276, 286, 303, 353, 355, 425, 452, 454, 455, 471, 472, 514, 533, 534, 553, 565, 584, 585, 593, 617, 648, 672, 675, 736, 741, 785, 786, 816, 899, 900, 925, 984, 994, 1023, 1028, 1059, 1113, 1116, 1148, 1199, 1237, 1258, 1260, 1266, 1339, 1342, 1344, 1371, 1529, 1575, 1583, 1597, 1651, 1654, 1657, 1696, 1855, 1859, 1950, 1951, 2003, 2004, 2005, 2080, 2083, 24, 25, 27, 88, 106, 131, 138, 140, 146, 215, 238, 241, 242, 249, 270, 292, 440, 515, 521, 522, 524, 583, 586, 615, 621, 623, 646]. pocket [649, 650, 652, 653, 676, 677, 700, 759, 836, 916, 972, 973, 974, 1005, 1016, 1021, 1061, 1069, 1089, 1118, 1200, 1238, 1242, 1257, 1284, 1329, 1513, 1584, 1585, 1605, 1629, 1630, 1633, 1649, 1687, 1850, 1856, 1858, 1868, 1870, 1954, 1955, 1958, 1968, 2040]. Podcasting [817]. Poland [63]. Policy [56]. Polish [63]. Politics [461]. PolyForm [207]. popular [304, 1478]. porn [1230]. Portability [429, 431]. Portable [360, 362, 1357, 1358, 1359, 1360, 1361, 1362, 1363, 1364, 1365, 1366, 2064, 842, 1033]. Portal [55, 958, 54]. Porting [847, 1011, 2100]. POSIX [1033]. POSIX.1 [1033]. POSIX.4 [625]. Possible [1745]. postfix [823, 387, 824]. PostgreSQL [848, 2091, 2092, 2093]. Postings [1744]. Power [174, 557, 711, 720, 1253, 1299, 1491, 1494, 1562, 1564, 69, 82, 674, 697, 722, 1035, 1492, 1655, 1793, 1801, 1885, 1899, 1977]. Powered [1641]. powerful [606, 1188]. PowerPoint [1881, 1985, 1986, 70]. PowerShell [834, 835, 836]. PPP [1876]. Practical [134, 366, 401, 499, 626, 628, 634, 930, 1456, 1462, 1463, 1464, 1465, 1467, 1563, 1586, 1715, 2155, 1892, 1979, 2078, 2091, 663, 927, 931, 1106, 1159, 1639, 1913, 28, 29]. Practice [190, 682]. Practices [519, 637, 314, 456, 691, 1628]. practitioners [375]. Pragmatic [294, 860, 1916, 1917, 1898, 1915, 1169]. Pratique [1350, 1673, 2092, 2093]. Praxishandbuch [28, 29]. Praxiskurs [399]. Prefactoring [1581]. Preferred [2067]. Presence [1967, 418]. Presentation [1759, 1881]. Pretty [627]. Primer [2104]. Principles [312, 125, 691]. Print [1481, 1969]. Printing [1595, 1580]. Privacy [331, 627, 633, 630, 631, 1978]. Private [793, 1752, 1753]. Pro [298, 299, 333, 1535, 1539]. Problems [1218, 2103, 19, 1468, 1715, 2075]. procedure [791]. Proceedings [68, 2121, 2122, 2127, 2128, 2129, 1429]. Process [1503, 797]. Processing [165, 1110, 1139, 1608, 235, 891, 1109]. Processor [1631, 792]. Producing [582]. production [2120, 933]. productive [587]. products [2126]. Professional [561, 1436, 2086, 398, 933]. program [1466, 1516, 1881]. Programmation [774]. Programmer [340, 341, 535, 1394, 587, 815, 876, 1033, 1915]. Programmers [538, 767, 1091, 1490, 1508, 289, 540, 1493, 1506, 1611]. Programmeur [474]. Programmierreich [1349]. Programmiersprache [1557]. Programmierung [1120, 1661, 28, 29]. Programming [9, 22, 42, 44, 77, 78, 87, 99, 100, 168, 169, 170, 174, 219, 220, 246, 260, 269, 271, 301, 354, 360, 362, 386, 391, 498, 508, 510, 511, 512, 516, 520, 523, 537, 594, 595, 596, 605, 618, 625, 639, 724, 725, 726, 734, 735, 737, 738, 753, 755, 758, 767, 773, 777, 778, 808, 809, 810, 811, 812, 833, 845, 849, 862, 863, 924, 935, 949, 1022, 1024, 1025, 1026, 1027, 1030, 1037, 1039, 1040, 1042, 1043, 1044, 1045, 1046, 1047, 1048, 1050, 1054, 1055, 1056, 1087, 1092, 1098, 1099, 1100, 1101, 1102, 1112, 1115]. Programming [1119, 1232, 1265, 1282, 1306, 1334, 1373, 1375, 1378, 1382, 1388, 1392, 1457, 1462, 1463, 1464, 1465, 1478, 1573, 1586, 1614, 1624, 1626, 1634, 1652, 1678, 1679, 1681, 1684, 1718, 1719, 1756, 1759, 1774, 1806, 1807, 1809, 1825, 1848, 1863, 1864, 1883, 1915, 1928, 2004, 2010, 2011, 2012, 2014, 2015, 2041, 2055, 2056, 2068, 2079, 2082, 2084, 28, 29, 41, 155, 178, 330, 382, 470, 509, 571, 663, 728, 782, 791, 922, 1094, 1335, 1456, 1467, 1476, 1557, 1599, 1635, 1676, 1682, 1713, 1714, 1817, 1898, 1999, 2009, 2016, 2027, 2038, 237, 292]. Programs [369, 1057, 1883, 374, 424, 1033, 1058, 1466, 1874]. Project [179, 2104, 144, 294, 450, 487, 582, 705, 709, 1836, 73]. Project-Based [2104, 705]. Projects [109, 1231, 1418, 1641, 1671, 1888, 1889, 624, 1369, 1639, 1777]. Projekte [109, 1642]. Propeller [1188]. Property [1064]. protect [695]. Protecting [598]. Protection [727]. Protocol [1145, 1376, 1383, 1395, 1725, 1726, 38, 1305, 2083, 1981]. Protocols [751, 967]. Prototype [60, 1620]. Proven [1083]. psad [1609]. PSP [1711].

Pthreads [1335, 1334]. Public [793]. publishers [2120]. Publishing [46, 349, 388, 768, 844, 1213]. pushing [1230]. pyparsing [1212]. Python [1111, 1120, 1600, 1642, 1871, 130, 165, 365, 422, 662, 767, 791, 896, 923, 1036, 1111, 1112, 1113, 1114, 1115, 1116, 1117, 1118, 1119, 1121, 1122, 1123, 1124, 1162, 1163, 1164, 1165, 1166, 1167, 1215, 1264, 1599, 1719, 1869, 1870, 1982, 2075]. PyTorch [1608].

Qmail [823, 1032]. Qt [360, 362, 1268]. QuarkXPress [1413]. Quick [205, 208, 445, 538, 554, 556, 665, 667, 779, 985, 1172, 1341, 1414, 1420, 1479, 1482, 1594, 1647, 1685, 1814, 2045, 2046, 2074, 39, 185, 285, 358, 383, 426, 427, 480, 481, 540, 547, 555, 563, 564, 569, 576, 668, 671, 674, 783, 805, 834, 950, 1035, 1067, 1269, 1270, 1274, 1323, 1338, 1413, 1448, 1402, 1498, 1506, 1598, 1620, 1660, 1683, 1737, 1763, 1783, 1785, 2136, 1793, 1827, 1885, 1920, 1947, 1970, 1971]. Quickbase [311]. QuickBooks [151, 152]. Quicken [153, 154]. QuickTime [10]. Quiz [707].

R [16, 17, 277, 327, 663, 728, 1083, 1755, 1906, 2065]. R6 [1430]. racing [1473]. RAD [1478]. RADIUS [793]. RAID [1980]. rails [1161, 1897, 1917, 146, 393, 459, 496, 752, 822, 1105, 1455, 1519, 1621, 1851, 1852, 1918]. range [173, 846]. Raspberry [1643]. RAW [4, 6]. RCS [179]. RDF [1563]. Reach [1484]. Reaching [1198]. Readability [1581]. reader [222]. Reading [1865]. Real [625, 880, 936, 1218, 1461, 178, 1230, 1715]. Real-World [1218, 178, 1715]. REALbasic [1325, 1326]. Reality [44, 43]. reaping [2089]. Rebel [1080, 1081, 1082]. Recipes [130, 1083, 1999, 2120]. reconnaissance [2109]. Recording [787]. Recovery [1569, 1571]. rectal [895]. Red [1181, 1182, 1183, 1184]. RedHat [1432]. Redis [969]. Refactoring [494]. Reference [340, 343, 254, 255, 286, 323, 425, 445, 452, 454, 455, 497, 500, 506, 514, 533, 536, 538, 548, 553, 554, 556, 557, 565, 584, 585, 614, 617, 648, 665, 667, 683, 684, 701, 702, 703, 741, 779, 932, 965, 970, 971, 984, 985, 1023, 1059, 1113, 1116, 1172, 1199, 1237, 1266, 1295, 1303, 1339, 1341, 1342, 1344, 1374, 1376, 1377, 1379, 1380, 1383, 1387, 1395, 1414, 1420, 1422, 1479, 1482, 1575, 1583, 1594, 1597, 1647, 1651, 1654, 1685, 1725, 1726, 1814, 1890, 1950, 1994, 2002, 2003, 2004, 2005, 2045, 2046, 2066, 2074, 2083, 2115, 24, 25]. reference [27, 39, 88, 94, 140, 146, 185, 241, 270, 285, 358, 383, 426, 427, 440, 480, 481, 515, 521, 522, 524, 540, 547, 555, 563, 564, 569, 583, 586, 615, 646, 649, 652, 668, 671, 674, 686, 700, 759, 783, 805, 834, 836, 876, 916, 950, 967, 972, 973, 1016, 1021, 1035, 1067, 1069, 1089, 1111, 1118, 1200, 1238, 1242, 1257, 1269, 1270, 1274, 1284, 1323, 1329, 1338, 1413, 1448, 1402, 1475, 1498, 1506, 1513, 1574, 1584, 1585, 1598, 1605, 1620, 1629, 1630, 1633, 1649, 1660, 1683, 1687, 1737, 1763, 1783, 1785, 2136, 1793, 1850, 1827, 1868, 1870]. reference [1885, 1920, 1947, 1968, 1970, 1971, 806]. references [1746]. Registry [1508, 1666, 1668]. Regular [607, 690, 1868, 1870, 606, 608, 649, 903, 1869]. Reguläre [1871]. relational [375]. RELAX [1989]. Release [535, 665, 671, 1296, 1379, 1380, 1394, 1432]. Releasing [1731]. Reliable [1981, 424, 1935]. remoting [1293]. Repairing [1927]. replication [2111]. Report [46, 620, 1415, 1771]. reporting [760, 761]. RepRap [1580]. required [1827]. Research [461, 1771]. Researcher [167]. Réseau [774]. Resource [499, 501, 502, 2121, 503, 504, 505, 2122, 557, 684, 771, 1381, 2127, 1384, 1385, 1386, 2128, 1389, 1390, 1391, 2129, 1393, 2013]. Resources [793, 1303, 1967, 418]. Response [1995, 1609]. RESTful [1640]. RESTful.NET [572]. Restore [1003]. results [11]. retouching [841, 851, 1347]. Retro [957]. Reusable [1261]. Revolution [396, 820]. Revolutionary [1617, 1618, 1619]. rewards [2089]. RFID [679, 680, 891]. Rich [1697]. Richard [2071]. rifarensu [1869]. right [1073]. Rim [1445]. ripping [1845]. Rise [1812]. risks [2089]. RJS [496]. RMAN [984]. RMI [729, 730]. Road [181, 991, 1439, 1840, 247, 1004, 1321]. Rob [1142]. robots [956]. rock [2007]. Rocks [1800]. rolling [1230]. ROM [66, 69, 339, 349, 518, 557, 1303, 1343, 1431, 1432, 1452, 1655]. Romance [2096]. Romkey [1144]. Root [873]. rootkits [960]. router [1107]. Routers [20, 89, 1107]. Routing [1149, 2155, 1160]. RPC [174, 273, 1774, 1848]. RPM [1179]. RPM-based [1179]. RSS [765, 766, 2085]. RT [2000]. RTF [238]. Ruby [1871, 2139, 85, 262, 532, 533, 571, 605, 707, 1172, 1870, 1897, 1915]. run [582]. Running [291, 363, 377, 2149, 1009, 1431, 1432, 1957, 2023, 2049, 2050, 2052, 2054, 848, 1512, 1705, 1897]. Runtime [276]. rush [1223]. Russia [1260]. Russian [1260]. Rust [168, 169, 170, 190, 2104].

S3 [1306]. Safe [168, 170]. Safety [1877]. Samba [266, 453, 454, 1968, 1969]. San [1434]. SANs [1570]. SAP [240]. SATAN [598]. Save [3]. Saving [1712, 636]. SAX2 [227]. say [198]. Scala [828]. Scalable [814, 2149, 424]. Scale [406, 1695, 1488]. Scaling [288]. SCCS [179]. Schema [1988]. schemas [1065]. Schematron [1990]. Schiller [1144]. School [948]. Schritt [1557]. Science [605, 1457, 1738, 1982, 2117, 1456, 2065]. scientific [196]. Scientist [422, 998]. Scipy [214]. SCO [358]. Scott [1261]. scrapers [1733]. scraping [1264]. screen [1733]. ScreenOS [2118]. Script [1664]. scriptaculous [60]. Scripting [904, 1115, 1299, 1656, 1661, 1832, 1960]. Scripts [1076, 1658, 1468, 1903]. search [531, 2061]. Searchable [557]. Second [1601]. seconds [216]. Secrets [461, 1792, 283, 1700]. Secure [117, 620, 691, 1998, 1999, 332, 387, 103, 107]. Securing [367, 718, 793, 1367, 1368, 2048]. Security [156, 313, 332, 626, 628, 629, 633, 634, 849, 1013, 1399, 1403, 1497, 1558, 1704, 1877, 1914, 1998, 104, 118, 329, 368, 394, 596, 598, 843, 963, 1074, 1075, 1227, 1228, 1645, 1768, 2116, 149, 1144, 1185]. sed [413, 1654, 1653, 69, 412, 417, 1651]. Seiki [1869]. Selbstbau [109]. Selecting [403]. Selenium [226]. Self [1613]. Self-Describing [1613]. SELINUX [1185]. selves [1887]. sendmail [321, 322, 323, 324, 325, 326, 859, 320]. Sense [894]. Sensor [44]. Sensors [2153, 42]. Separation [1581]. Sequence [1158]. Server [84, 197, 204, 367, 368, 715, 760, 761, 794, 795, 2149, 861, 993, 1261, 1273, 1372, 1667, 1716, 1786, 1919, 1969, 1972, 2068, 2081, 49, 112, 118, 526, 579, 762, 1973, 2001, 2053, 243, 1030, 1669, 1712, 1970]. Server-Side [2149]. Servers [117, 1265, 1368, 1367]. Service [1279, 281, 1632, 1638]. Services [8, 273, 508, 1070, 1097, 1306, 1614, 1771, 1809, 1848, 280, 762, 1100, 1101, 1102, 1161, 2126, 1578, 1640, 1682, 1935, 1948, 2087, 345, 880]. Servlet [862, 863, 1500]. servlets [113, 114]. SETI [1197]. Setting [1002]. setup [288]. setups [1893]. Seven [1622, 1898]. sharding [288]. Shared [1872]. SharePoint [7, 2039, 2040, 2042, 2135]. sharing [2029]. sheets [1346, 1236, 1239, 1241]. Shell [399, 755, 1330, 1331, 1656, 1658, 1661, 1691, 1692, 1693, 69, 1332, 1397, 1903, 103, 107, 1664]. Shell-Programmierung [1661]. Ship [1639]. Shockwave [1728]. Shopper [189]. shori [1109]. Short [405, 1518, 2060]. should [815, 2034]. Shuffle [160]. Side [2149]. Silence [2109]. silicon [875]. Silverlight [59, 2057, 2069]. Simple [1305]. SimpleDB [1306]. Singapore [994, 1361]. Singaporean [994]. sinking [709]. Site [251, 959, 12, 1073, 1504]. Sites [814, 1695, 1128, 1811, 1882]. six [1343, 1882]. Skills [660]. Skype [1767]. Slash [291]. slow [705]. Small [11, 2037]. Smart [1240, 1860, 1756]. smarter [663]. SmartList [1743]. smartphone [2105]. Smarty [1488]. Smileys [1717, 1867]. SMS [1632]. SNMP [1175, 1176, 1305]. Snort [329, 1419]. so-called [2060]. SOA [2090]. SOAP [273, 476, 1809]. Social [1706, 1707, 1706]. Society [2131, 2132]. Software [429, 431, 1011, 1092, 1416, 1449, 1596, 2015, 2071, 2103, 40, 77, 334, 582, 636, 842, 961, 1515, 1639, 1836, 1849, 2125]. Solaris [671, 886, 1647, 2036]. solid [884]. Solution [1822, 1823]. Solutions [208, 289, 371, 372, 488, 1071, 1218, 2103, 1899]. solve [1468, 1715, 2075]. Son [1965]. Sony [1246]. soul [1887]. Source [179, 1012, 1064, 1872, 2066, 304, 487, 582, 2125, 869, 1185, 1619, 2009, 2089, 396, 439, 1617, 1618, 1849, 1991]. Sources [396, 397]. sous [1965]. South [1258]. Southern [688]. space [2060]. Spain [1186, 1371]. Spam [1230, 1744, 443, 2112]. SpamAssassin [1747]. Spanish [1371]. Speaking [1611]. specific [1475, 1476]. Specifications [1417]. Speeches [98]. Spidering [813]. spiders [236, 1733]. spielen [109]. Spirit [264]. Spirits [744]. spot [83]. Sprache [361, 364]. spreadsheet [622]. Spring [643, 1896]. SQL [518, 1450, 84, 129, 133, 348, 493, 494, 510, 511, 512, 514, 515, 516, 517, 519, 520, 521, 522, 523, 524, 650, 653, 741, 760, 761, 762, 922, 949, 950, 951, 952, 971, 1255, 1256, 1267, 1320, 1396, 1408, 1572, 1820, 1961, 2068]. SQL*Loader [647]. SQL*Plus [645, 646, 648, 652, 651]. SQLAlchemy [317]. SQLite [749]. SQLJ [1573]. SQS [1306]. Squid [2059]. SSH [103, 107]. Stallman [2071]. Stamping [1744]. Standard [1111, 1591, 1002, 1033]. standards [58]. Starship [2061]. Start [1473, 79]. Started [1050, 1336, 2153, 153, 206, 468, 819, 869, 891, 1188, 1212, 1550, 1555, 1557, 1580, 1714, 2009, 2069]. starters [153, 282, 733, 1132, 1133, 1135, 1472, 1550, 1555, 1986]. Startups [172, 172]. State [620, 824]. state-of-the-art [824]. States [1365]. Statisphere [1437]. statistical [2112]. Statistics [191, 604, 1083, 723]. Status [2067]. Staying [1847]. Steal [2029, 2032]. Step [632, 327, 469, 1557]. Step-by-Step [632, 327, 469]. Stephen [892]. steuern [109]. Steve [1145]. STL [1069]. Stockman [1145, 1146]. stop [83, 316]. Stopping [1744]. Store [1829]. stored [791]. Stories [184, 264, 743, 1444, 1445, 1447, 1842, 247, 2060]. story [1230]. Strange [477]. Strangers [477]. Strategien [1758]. Strategies [1758]. Streamlined [537]. Strength [65, 999, 119, 147, 150, 248, 579, 580, 693, 798, 813, 915, 975, 1314, 1604, 1810, 1846, 1973]. Strings [688]. Struts [269, 270, 271, 1790]. Studio [719, 959, 400, 841, 1728, 81, 1884]. Studio.Net [438]. Study [1664]. stuff [2123]. Style [1236, 1239, 1241, 1346, 2095]. Subclassing [1907]. Subject [2126]. substance [2095]. Subversion [306, 307, 1169, 1510]. Successful [1596, 334, 582, 1639]. Suite [53]. super [2034]. super-user [2034]. Supercharged [2143]. Superhighway [101]. Supplement [535, 1394]. Supplementary [341, 344]. Survival [640, 1444, 1847]. Surviving [1959]. survivors [931]. SuSE [225]. SVG [462]. SVR4 [354, 1647]. Swift [1557, 1557]. Swing [451, 1103, 1104, 457, 1156]. Switching [1534, 1545, 2017, 638, 704]. Sword [3]. SWT [796]. syncing [302]. syndication [765, 1565]. Syntax [648]. System [56, 171, 342, 279, 351, 352, 499, 535, 611, 612, 613, 616, 617, 1062, 1090, 1298, 1313, 1319, 1368, 1394, 1416, 1421, 1442, 1443, 1482, 1508, 1589, 1593, 1883, 1942, 1943, 14, 69, 108, 265, 304, 598, 662, 968, 1003, 1063, 1071, 1094, 1495, 1496, 1682, 1944, 35, 357, 416, 665, 668, 669, 671, 809, 1259, 1296, 1375, 1376, 1377, 1420, 1588, 1590, 1591, 1592]. Système [2051]. Systems [99, 170, 354, 447, 1254, 1269, 1507, 1742, 2097, 100, 332, 1179, 1333, 2098].

T1 [640]. Tab [2147]. tables [1762]. tag [233, 2047]. Taiwan [353, 1362]. Taiwanese [353]. Take [19, 302, 573, 574, 945, 946, 705]. take-it-slow [705]. Taking [1682, 60]. Tales [180, 181, 182, 247, 264, 655, 746, 853, 939, 1186, 1187, 1425, 1426, 1427, 1428, 1433, 1434, 1438, 1439, 1761, 1840, 1841, 2096, 2061, 183, 742]. Talk [638, 870]. taming [1606]. Tapp [1893]. Tasks [1661]. Taste [1840]. TC [399]. TC-Shell [399]. Tcl [790, 1057, 1058, 1597, 1598]. Tcl-based [1057, 1058]. Tcl/TK [1598, 790, 1597]. TCP [854, 855, 856, 857, 858, 967, 1272]. TCP/IP [854, 855, 856, 857, 858, 967, 1272]. tcsh [430]. Teachers [1760]. teaching [304]. Tech [1847]. Technical [2121, 2122, 2127, 2128, 2129]. Techniken [1758]. technique [1721]. Techniques [312, 313, 606, 636, 933, 968, 1000, 1243, 1245, 1713, 1758, 1873]. Technologies [1253]. Technology [1137]. telephones [638]. téléphonie [1991]. Telephony [1155, 1802, 1992]. tell [2028, 2029, 2032]. template [274]. templates [496]. TERMCAP [1865, 1866]. terminfo [1866]. Terms [1775]. Test [216, 1519, 754]. Testing [828, 938, 226, 843, 860, 990]. Testosterone [1447]. Thai [2080]. Thailand [1425, 2080]. theater [1222]. their [232]. theory [375, 966]. There [991]. Things [1745, 115, 116, 164, 622, 694, 792, 815, 870, 878, 920, 1263, 1474, 1646, 1881, 2153]. Think [422, 423, 998, 198]. thinking [867]. thirty [1882]. thirty-six [1882]. Threading [1631]. Threads [1398, 1400, 1401, 1405]. tidy [2065]. Tiger [889, 1221, 1544, 1904, 1955, 1956]. Tight [1177]. Time [1062, 1063, 636, 1712, 1758, 1758]. time-saving [636]. Tipps [1758]. Tips [181, 272, 479, 720, 948, 999, 1156, 1845, 57, 119, 147, 150, 248, 411, 460, 484, 579, 580, 693, 798, 813, 817, 915, 975, 1210, 1314, 1604, 1606, 1711, 1758, 1810, 1846, 1873, 1973]. Title [2148, 2152]. TiVo [975]. TK [1459, 1598, 790, 1059, 1060, 1597, 2022]. TM [1416, 1435, 1437, 1823, 2086, 207, 1436]. TOAD [1199, 1200]. TOC [2142]. Today [2157]. Toilet [991]. Tomcat [217, 218]. Tool [1478, 5]. Toolkit [536, 886, 1057, 1378, 1379, 1380, 1388, 1392, 1422, 274, 1058, 1502]. Tools [272, 312, 403, 537, 790, 999, 1156, 1297, 1299, 1491, 1494, 1562, 1564, 1799, 2004, 69, 82, 100, 119, 147, 150, 248, 328, 329, 378, 394, 411, 484, 579, 580, 606, 693, 798, 813, 817, 915, 975, 1158, 1210, 1314, 1492, 1604, 1606, 1655, 1711, 1801, 1810, 1845, 1846, 1973, 1977]. Topics [123, 1905]. trainer [70, 71, 72, 73, 74, 75, 356]. Transact [949, 1820]. Transact-SQL [949, 1820]. Transactions [620]. Transcending [1891]. Transfer [800, 2083]. transform [6, 2065]. transformations [1931]. transplants [528]. transport [824]. Travel [181, 184, 688, 742, 991, 1153, 1877, 2067]. Traveled [991]. Traveler [180, 181, 182, 247, 264, 655, 746, 853, 939, 1186, 1187, 1425, 1426, 1427, 1428, 1433, 1434, 1438, 1439, 1761, 1840, 1841, 2096, 183, 742]. travelers [1764]. Treatment [894]. Treatments [527]. Treo [875, 1136]. Trials [527]. tricks [57, 460]. Trouble [1444]. Troubleshooting [1028, 1799, 1945, 1946, 1699]. True [184, 264, 743, 744, 1444, 1445, 1447, 1842, 247, 2060]. Trust [156]. tumor [884]. Tuning [319, 740, 741, 940, 941, 1090, 1313, 1769, 1770, 1961]. tutorial [543, 560, 567]. tutorials [657, 1805]. twill [226]. Twisted [509].

Ubuntu [1469]. UDDI [273]. UI [1322]. Ultimate [1520, 1521]. UML [39, 40, 1248, 1513, 1514]. uncertain [2126]. understand [109]. understandable [1557]. Understanding [137, 199, 200, 201, 202, 203, 673, 893, 1109, 1137, 1155, 1470, 1689, 1849, 866]. unflinchingly [2061]. Unicode [964]. Unit [754, 860]. United [786, 1365]. Unix [399, 1661, 66, 69, 351, 352, 354, 358, 360, 626, 628, 666, 667, 668, 670, 671, 1011, 1033, 1091, 1298, 1420, 1442, 1443, 1491, 1492, 1494, 1495, 1496, 1569, 1647, 1655, 1902, 1904, 1942, 1943, 1944, 2013, 241, 316, 362, 387, 634, 662, 887, 888, 889, 890, 1106, 1562, 1564, 1650, 1659, 1662, 1701, 1900, 1901, 1969]. Unix-Shell [399]. Unix/Linux [1661]. Unix/Linux-Tasks [1661]. Unobtrusive [1795]. unofficial [956]. unreliable [1935]. Unwanted [1744]. unwired [1004, 1321, 2044]. up-to-date [2057]. UpdatePanel [1029]. Upgrading [1962, 1927]. USA [471, 1365]. usability [332]. Use [881, 936, 1010, 1740, 1760, 332, 387, 1188, 1460, 1698, 1764, 1873, 2106]. useful [1468]. Usenet [788, 1938, 789, 1936, 1939, 1454, 1819, 1937, 1940, 1941]. User [343, 344, 575, 789, 911, 978, 979, 980, 981, 1060, 1233, 1297, 1421, 1589, 1590, 1591, 1592, 1593, 1847, 2022, 7, 308, 435, 674, 982, 983, 1035, 1588, 1793, 1885, 2034, 871, 872, 874]. Users [1182, 1490, 1508, 706, 1493]. Using [28, 2140, 245, 266, 304, 305, 351, 352, 366, 403, 430, 453, 463, 673, 923, 987, 1570, 1608, 1612, 1627, 1665, 1875, 1876, 1936, 1937, 1938, 1939, 1940, 1969, 4, 58, 411, 989, 1073, 1169, 1681, 1715, 1762, 1916, 2075, 2101]. Utilities [208, 1257]. utopia [989, 1762]. UUCP [1454, 1612, 1936, 1937, 1938, 1939, 1940, 1941]. UX [1650].

V [671, 668, 1420]. V2.0 [2086]. v3 [576]. validation [1999]. valley [820]. VB [1078, 1127, 1677]. VB.NET [985, 1266, 1683, 1685, 1687, 1874]. VBA [1078, 1681, 1686, 2041]. VBScript [285, 286, 1077, 1079]. Version [306, 307, 535, 541, 665, 809, 810, 1169, 1373, 1374, 1375, 1376, 1377, 1421, 1451, 1452, 1510, 1823, 1394, 1916, 1588]. verständlich [1557]. verstehen [109]. Vi [1648, 69, 986, 988, 1649, 1663]. Vi-Editor [1648]. Video [1587, 657, 933, 1480, 1606, 1805, 1856]. Vietnam [355]. Vietnamese [355, 1110]. VII [1521]. Vim [1663]. Virtual [279, 328, 361, 364, 1234, 1235, 1299, 1752, 1753, 1712]. Virtualisierungslösungen [1509]. Virus [727]. viruses [443]. vision [210]. Vista [337, 346, 347, 698, 700, 718, 921, 1554, 1555, 1828]. visual [525, 81, 398, 438, 591, 719, 734, 735, 755, 758, 833, 1008, 1041, 1044, 1048, 1049, 1127, 1130, 1265, 1477, 2133, 1478, 1677, 1682, 1907, 2030]. visualization [313]. visualize [2065]. Visualizing [619]. VMware [946, 1299, 1966, 2033]. VOCAL [366]. Voices [396]. VoIP [366, 2017, 2018]. Vol [2001, 1219]. Volume [64, 65, 866, 867, 1395, 1417]. vs [1600]. VxDs [279].

Waits [8]. warfare [263]. Warrior [2144, 1812, 404, 1497]. was [589]. way [1073, 1107]. ways [1709, 1874]. WCF [244, 1100, 1101, 1102]. Web [54, 64, 65, 204, 1264, 1417, 1429, 1484, 1756, 226, 365, 379, 791, 937, 1073, 1210, 1707, 1882, 1, 2, 12, 55, 65, 125, 132, 156, 207, 236, 251, 273, 280, 331, 388, 389, 379, 508, 526, 548, 575, 577, 618, 620, 629, 633, 726, 738, 803, 804, 814, 843, 876, 880, 907, 940, 941, 959, 989, 1070, 1128, 1137, 1161, 1262, 1289, 1306, 1315, 1316, 1336, 1337, 1338, 1340, 1341, 1343, 1345, 1346, 1416, 1417, 1440, 1449, 1408, 1459, 1485, 1486, 1488, 1502, 1504, 1568, 1577, 1578, 1614, 1640, 1671, 1694, 1695, 1706, 1724, 1728, 1771]. Web [1786, 1809, 1811, 1817, 1848, 1832, 1913, 1917, 1918, 1948, 1974, 2015, 2053, 2058, 2070, 2072, 2073, 2082, 2116]. Web-Based [1671]. WebBoard [1416, 1449, 1487]. Webbots [1733]. Weblog [403]. WebLogic [1292, 1372]. Weblogs [291]. WebMaster [1441, 1814, 1815, 1816]. Webmin [316]. website [58, 2101, 1435, 1436, 1483, 1484, 2086, 942, 2151, 2062]. websites [11]. WebSphere [223]. weeks [1622, 1898]. well [1557]. well-explained [1557]. Welt [109]. went [589]. werden [399]. Wheelchair [911, 909].

Wheels [911]. where [379, 1460]. Who [1877]. Whole [61, 978, 979, 980, 981, 308, 982, 983]. Wi [442, 1698, 444]. Wi-Fi [442, 1698, 444]. Wicked [487, 1468, 1903]. Wide [64, 65, 738, 1070, 1417, 1429, 1694, 1695]. Wikipedia [221, 222]. Wild [1187]. Win32 [301, 360, 362, 767, 771, 1682, 1742]. Window [357, 416, 499, 535, 665, 669, 809, 1296, 1375, 1376, 1377, 1394, 1421, 1588, 1589, 1590, 1591, 1592, 1593]. Windows [337, 1272, 36, 38, 49, 50, 56, 75, 82, 116, 205, 208, 216, 259, 279, 328, 333, 346, 347, 367, 368, 386, 414, 477, 489, 495, 609, 613, 614, 615, 693, 696, 697, 698, 699, 700, 721, 726, 727, 755, 767, 794, 795, 801, 822, 834, 835, 836, 857, 861, 877, 908, 910, 912, 913, 914, 916, 918, 919, 920, 921, 982, 992, 993, 1001, 1003, 1004, 1009, 1010, 1045, 1095, 1193, 1233, 1259, 1273, 1275, 1304, 1305, 1319, 1367, 1368, 1442, 1443, 1448, 1409, 1482, 1508, 1524, 1530, 1535, 1538, 1539, 1550, 1554, 1555, 1556, 1666, 1668, 1682, 1690]. Windows [1699, 1737, 1759, 1804, 1828, 1919, 1960, 1971, 1972, 1973, 2106, 2107]. wire [2109]. Wireless [577, 578, 580, 581, 618, 641, 1141, 642, 1698]. Wireshark [1715]. Wiretap [461]. Wisdom [181, 815]. Within [1439]. without [1762, 1829]. wizardry [259]. WML [618]. WMLScript [618]. Woman [180, 184, 688]. Women [181, 1187, 1877]. won’t [2028, 2029, 2032]. word [792, 74, 674, 675, 677, 732, 733, 792, 1019, 1678, 1681, 1723, 2043]. WordPress [1722, 1722]. WordPress-Buch [1722]. Work [211, 1507, 1680, 2020, 1159]. Workbook [223, 517, 1372]. workflow [53, 1893]. Working [928, 1466]. works [6, 2034]. Workstation [1001, 2033]. World [3, 180, 183, 184, 625, 708, 853, 1218, 1447, 1847, 1878, 77, 109, 178, 411, 872, 880, 1066, 2126, 1461, 1715, 1804, 64, 65, 738, 1070, 1289, 1417, 1429, 1694, 1695]. worrying [316]. Wrangling [923]. Write [842, 866, 867]. Writer [2043]. Writing [360, 362, 678, 1261, 1680, 1681, 1686, 1772, 1773, 1833, 1865, 1883, 2079, 2104, 236, 682, 867, 1033]. WSDL [273]. WWW [86, 1429].

X [112, 377, 410, 888, 889, 1533, 1553, 1834, 1904, 1953, 1954, 1955, 1956, 1958, 2139, 97, 357, 409, 416, 435, 499, 501, 502, 2121, 503, 504, 505, 2122, 535, 536, 665, 669, 2146, 720, 722, 809, 879, 887, 890, 1009, 1214, 1296, 1297, 1321, 1346, 1375, 1376, 1377, 1378, 1379, 1380, 1381, 1383, 2127, 1384, 1385, 1386, 1388, 2128, 1389, 1390, 1391, 1392, 2129, 1393, 1394, 1395, 1421, 1422, 1430, 1527, 1528, 1544, 1588, 1589, 1590, 1591, 1592, 1593, 1701, 1708, 1725, 1726, 1900, 1901, 1902, 1950, 1951, 1969, 1977]. X11 [1296, 1379, 1380]. XAML [1138]. Xandros [706]. Xbox [2007]. XEN [1509, 1509]. XForms [428, 463]. XHTML [597, 1310, 1311, 1312, 1344]. Xi [1664]. Xlib [1373, 1374, 1375, 1377, 1382, 1387]. xmh [1490, 1493]. XML [9, 188, 235, 273, 312, 452, 455, 530, 768, 779, 780, 781, 783, 844, 896, 1015, 1065, 1216, 1217, 1218, 1220, 1224, 1226, 1294, 1315, 1453, 1481, 1613, 1615, 1616, 1794, 1848, 1850, 1931, 1988]. XML-RPC [273, 1848]. Xmt [537]. XNA [1623]. XP [1535, 1538, 1539, 50, 75, 693, 696, 697, 913, 914, 916, 918, 919, 920, 1004, 1530, 1550, 1699]. XPath [1794]. XPointer [1794]. XQuery [2019]. XS [1081]. XS/1000D [1081]. XSi [1082, 1080]. XSi/450D [1082]. XSL [1481]. XSL-FO [1481]. XSLT [234, 235, 529, 1016, 1150, 1151, 1929, 1931]. Xue [1664]. XView [808, 809, 811, 1994].

yacc [1031, 1168]. Yahoo [121, 1579]. Year [1763]. Years [64]. yourselfer [109]. Yum [1179].

Zeit [1758]. Zeitmanagement [1758]. ZENworks [588]. Zero [284, 1835, 2107]. Zhi [1664]. zum [399]. Zune [1709].

References

Aaland:1998:PW

[1] Mikkel Aaland. Photoshop for the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-350-2. xiii + 197 pp. LCCN T385 .A22 1998; T385 .A325 1998. US$29.95. URL http://www.ora.

com/catalog/photostudio/;   http://www.oreilly.com/catalog/
9781565923508;   http://www.oreilly.com/catalog/webphoto.

Aaland:1999:PWC

[2] Mikkel Aaland. Photoshop for the Web: Covers Photoshop 5.5 and ImageReady 2.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-641-2. xiv + 230 + 64 pp. LCCN T385 .A2325 1999;

T385 .A325 1999. US$29.95. URL http://www.oreilly.com/catalog/
9781565926417;   http://www.oreilly.com/catalog/photoweb2.

Aaland:1999:SHF

[3] Mikkel Aaland. The Sword of Heaven: a Five Continent Odyssey to Save the World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-44-9. 300 (est.) pp. LCCN BL2220.3b.A35 1999. US$24.00.

URL http://www.oreilly.com/catalog/9781885211446;   http://
www.oreilly.com/catalog/sword.

Aaland:2006:PCR

[4] Mikkel Aaland. Photoshop CS2 RAW: using Adobe Camera Raw, Bridge, and Photoshop to get the most out of your digital camera. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00851-1. xvii + 206 pp. LCCN

TR267; TR267 .A35 2006eb. URL http://www.oreilly.com/catalog/
9780596008512.

Aaland:2007:PLA

[5] Mikkel Aaland. Photoshop Lightroom adventure: mastering Adobe’s nextgeneration tool for digital photographers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10099-X. xvii + 332 pp. LCCN TR267.5.A355.

URL http://www.oreilly.com/catalog/9780596100995.

Aaland:2008:PCR

[6] Mikkel Aaland. Photoshop CS3 RAW: transform your RAW images into works of art. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51052-7. xvii + 252 pp. LCCN TR267.5.A3. URL http://www.oreilly.com/

catalog/9780596510527.

Acker:2005:SUG

[7] Bryan Acker, Tyler Davey, and Robert McGovern, editors. SharePoint user’s guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00908-9 (paperback). x + 136 pp. LCCN TK5105.875.I6 S48 2005. URL http://www.oreilly.

com/catalog/9780596009083.

Adams:1999:OIS

[8] Steve Adams. Oracle8i Internal Services for Waits, Latches, Locks, and Memory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-598-X. x + 120 pp. LCCN QA76.9.D3 A33914 1999. US$19.95.

URL http://www.oreilly.com/catalog/9781565925984;   http://
www.oreilly.com/catalog/orinternals.

Adams:2002:PJ

[9] D. J. Adams. Programming Jabber: Extending XML messaging. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00202-5. xxi + 455 pp. LCCN QA76.76.H94 A268 2002. US$39.95. URL

http://safari.oreilly.com/0596002025;   http://www.oreilly.
com/catalog/9780596002022;   http://www.oreilly.com/catalog/
jabber.

Adamson:2005:QJD

[10] Chris Adamson. QuickTime for Java: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00822-8. xix + 233 pp. LCCN TR899 .A33 2005; TR899; TR899 .A33 2005eb. URL http://

www.oreilly.com/catalog/9780596008222.

Addison:2005:SWG

[11] Doug (Doug R.) Addison. Small websites, great results. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-932111-90-5. xx + 279 pp. LCCN TK5105.888 .A33 2005; TK5105.8 .A33 2005. US$34.99. URL

http://www.oreilly.com/catalog/9781932111903.

Addison:2006:WSC

[12] Doug (Doug R.) Addison. Web site cookbook. Web programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10109-0. xv + 261 pp. LCCN TK5105.888 .A33 2006eb; TK5105.888; TK5105.888 .A33

2006. URL http://www.oreilly.com/catalog/9780596101091.

Adelstein:2004:EJL

[13] Tom Adelstein and Sam Hiser. Exploring the JDS Linux desktop. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00752-3 (paperback). xiv + 390 pp. LCCN QA76.76.O63 A353 2004. URL

http://www.oreilly.com/catalog/9780596007522;   http://www.
oreilly.com/catalog/jds/index.html.

Adelstein:2007:LSA

[14] Tom Adelstein and Falko Timme. Linux system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00952-6 (paperback). xiii + 279 pp. LCCN QA76.76.O63; QA76.76.O63 A34 2007eb. URL http://www.

oreilly.com/catalog/9780596009526.

Adler:2005:BH

[15] Joseph Adler. Baseball hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00942-9. 256 pp. LCCN GV877eb; GV877. URL

http://www.oreilly.com/catalog/9780596009427.

Adler:2010:RN

[16] Joseph Adler. R in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-80170-X. xx + 611 pp. LCCN QA276.45.R3eb; QA276.45.R3; Internet.

Adler:2012:RN

[17] Joseph Adler. R in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2012. ISBN 1-4493-1208-X (paperback). xix + 699 pp. LCCN ???? EUR 41.00.

Aitken:2004:CPO

[18] Peter G. Aitken. Camera phone obsession. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2004. ISBN 1-932111-96-4. xviii + 252 pp. LCCN TK6570.M6 A38 2004. URL http://www.oreilly.com/

catalog/9781932111965.

Aker:2008:TCF

[19] Sharon Zardetto Aker. Take control of font problems in Leopard. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-52-1. ???? pp. LCCN QA76.76.O63 A39 2008.

Akin:2002:HCR

[20] Thomas Akin. Hardening Cisco Routers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00166-5. xiii + 173 pp. LCCN TK5105.543

.A47 2002. US$24.95. URL http://www.oreilly.com/catalog/
9780596001667;   http://www.oreilly.com/catalog/hardcisco.

Albahari:2001:CE

[21] Ben Albahari, Peter Drayton, and Brad Merrill. C# Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2001. ISBN 0-596-00079-0. xiv + 199 pp. LCCN QA76.73.C154 A418 2001. US$19.95.

URL http://www.oreilly.com/catalog/9780596000790;   http://
www.oreilly.com/catalog/csharpess.

Albahari:2002:CEP

[22] Ben Albahari, Peter Drayton, and Brad Merrill. C# Essentials: Programming the .NET Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00315-3. xii + 202 pp. LCCN QA76.73.C154 A418 2002. US$24.95. URL

http://safari.oreilly.com/0596003153;   http://www.oreilly.
com/catalog/9780596003159;   http://www.oreilly.com/catalog/
csharpess2.

Albahari:2007:CN

[23] Joseph Albahari, Ben Albahari, and Peter Drayton. C# 3.0 in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-52757-8 (paperback). xviii + 838 pp. LCCN QA76.73.C154 D73 2007; QA76.73.C154; QA76.73.C154 D73 2007eb. URL http://www.

oreilly.com/catalog/9780596527570.

Albahari:2008:CPR

[24] Joseph Albahari and Ben Albahari. C# 3.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51922-2 (paperback). xi + 230 pp. LCCN QA76.73.C154 A4185 2008.

Albahari:2008:LPR

[25] Joseph Albahari and Ben Albahari. LINQ pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51924-9 (paperback). ix + 161 pp. LCCN

QA76.73.L228 A43 2008. URL http://www.loc.gov/catdir/toc/
fy0805/2008299898.html.

Albahari:2010:CN

[26] Joseph Albahari and Ben Albahari. C# 4.0 in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2010. ISBN 0-596-80095-9. xviii + 1025 pp. LCCN ????

Albahari:2010:CPR

[27] Joseph Albahari and Ben Albahari. C# 4.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 1-4493-9401-9 (paperback). ???? pp. LCCN ????

Alber:2014:PFP

[28] Stephan Alber, Klaus Breyer, and Kornelius Nägele. Praxishandbuch Facebook-Programmierung. (German) [Practical guide to Facebook programming: Using Facebook with Javascript and PHP]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 3-95561-794-7. vi + 320 pp. LCCN ????

URL http://d-nb.info/1048697606/04.

Alber:2015:PFP

[29] Stephan Alber, Klaus Breyer, and Kornelius Nägele. Praxishandbuch Facebook-Programmierung. (German) [Practical guide to Facebook programming]. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 3-95561-796-3. ???? pp.

LCCN ????  URL http://d-nb.info/1065918607/34;   http:/
/nbn-resolving.de/urn:nbn:de:101:1-2015013026877;   http://
www.oreilly.de/.

Albing:2007:BC

[30] Carl Albing, J. P. Vossen, and Cameron Newham. Bash cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52678-4 (paperback). xxi + 598 pp. LCCN QA76.76.O63 A39885 2007. URL http://

www.oreilly.com/catalog/9780596526788.

Albitz:1992:DBN

[31] Paul Albitz and Cricket Liu. DNS and BIND in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1992. ISBN 1-56592-010-4 (cover), 0-565-92010-3. xxviii + 381 pp. LCCN TK5105.875.I57 A43 1992. US$29.95. URL http://www.oreilly.com/

catalog/9781565920101.

Albitz:1993:DBN

[32] Paul Albitz. DNS and BIND in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, corrected edition, March 1993. ISBN ???? xxviii + 381 pp. LCCN ????

Albitz:1994:DBN

[33] Paul Albitz and Cricket Liu. DNS and BIND in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1994. ISBN 1-56592-010-4. xxviii + 381 pp. LCCN TK5105.875.I57A43 1992.

Albitz:1997:DB

[34] Paul Albitz and Cricket Liu. DNS and BIND. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-236-0. xviii + 418 pp. LCCN TK5105.875.I57 A43

1997. US$32.95. URL http://www.ora.com/catalog/dns2/;
http://www.oreilly.com/catalog/9781565922365;   http://www.
oreilly.com/catalog/dns2.

Albitz:1998:DB

[35] Paul Albitz and Cricket Liu. DNS and BIND: Domain Name System and Berkeley Internet Name Domain. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1998. ISBN 1-56592-512-2. xvii + 482 pp. LCCN TK5105.875.I57 A42 1998. US$32.95. URL

http://www.ora.com/catalog/dns3/;   http://www.oreilly.com/
catalog/9781565925120;   http://www.oreilly.com/catalog/dns3.

Albitz:1998:DWN

[36] Paul Albitz, Matt Larson, and Cricket Liu. DNS on Windows NT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-511-4. xvi + 329 pp. LCCN TK5105.875.I57 A4325 1998; TK5105.875.I57 A44 1998.

US$34.95. URL http://www.oreilly.com/catalog/9781565925113;
 http://www.oreilly.com/catalog/dnswinnt/.

Albitz:2001:DB

[37] Paul Albitz and Cricket Liu. DNS and BIND. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2001. ISBN 0-596-00158-4. xvii + 601 pp. LCCN TK5105.875.I57 A43 2001. US$44.95.

URL http://www.oreilly.com/catalog/9780596001582;   http://
www.oreilly.com/catalog/dns4.

Alcott:2001:DWM

[38] Neall Alcott. DHCP for Windows 2000: Managing the Dynamic Host Configuration Protocol. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-838-5. x + 277 pp. LCCN TK5105.585 .A44 2001. US$34.95.

URL http://www.oreilly.com/catalog/9781565928381;   http://
www.oreilly.com/catalog/dhcpwin2000.

Alhir:1998:UND

[39] Sinan Si Alhir. UML in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-448-7. xiii + 273 pp. LCCN QA76.76.D47 A424 1998; QA76.9.O35 A44 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565924482;   http://
www.oreilly.com/catalog/umlnut.

Alhir:2003:LU

[40] Sinan Si Alhir. Learning UML: Communication software design graphically. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00344-7. xvi + 234 pp. LCCN QA76.76.D47 A4235 2003. URL http://www.oreilly.

com/catalog/9780596003449.

Allan:2010:LIP

[41] Alasdair Allan. Learning iPhone programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 1-4493-8005-0, 0-321-65870-1. xv + 356 pp. LCCN ????

Allan:2010:PIS

[42] Alasdair Allan. Programming iPhone sensors. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-8265-7 (paperback). ???? pp. LCCN ????

Allan:2011:ARI

[43] Alasdair Allan. Augmented reality in iOS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-0850-3 (paperback). 1 v. pp. LCCN ????

Allan:2011:ISP

[44] Alasdair Allan. iOS 4 Sensor Programming: Augmented Reality and Location Enabled iPhone and iPad Apps. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8265-7. 320 (est.) pp. LCCN ????

Allan:2012:GI

[45] Alasdair Allan. Geolocation in iOS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-0844-9 (paperback). viii + 102 pp. LCCN ????

Allen:1996:OOP

[46] O’Reilly Online Publishing Report, 1996. ISSN 1087-4216. US$295.00.

URL http://www.oreilly.com/catalog/. O’Reilly & Associates, Inc.,

103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800.

Allen:2003:AD

[47] Robbie Allen and Alistair G. Lowe-Norris. Active Directory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00466-4. xvii + 665 pp. LCCN QA76.76.O63 A454 2003. US$44.95 US, CDN$69.95,

UK£31.95. URL http://www.oreilly.com/catalog/9780596004668;
 http://www.oreilly.com/catalog/actdir2/.

Allen:2003:ADC

[48] Robbie Allen. Active Directory cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00464-8. xxvi + 593 pp. LCCN

TK5105.595 .A4497 2003. URL http://www.oreilly.com/catalog/
9780596004644.

Allen:2005:WSC

[49] Robbie Allen. Windows server cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00633-0 (paperback). xxiv + 670 pp. LCCN QA76.76.O63 A444

2005. URL http://www.oreilly.com/catalog/9780596006334.

Allen:2005:WXC

[50] Robbie Allen and Preston Gralla. Windows XP cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00725-6. xxii + 653 pp. LCCN

QA76.76.O63 A5545 2005. URL http://www.oreilly.com/catalog/
9780596007256.

Allen:2006:AD

[51] Robbie Allen, Alistair G. Lowe-Norris, and Joe Richards. Active Directory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-10173-2. 824 pp. LCCN TK5105.595 A45 2006eb; TK5105.595. URL

http://www.oreilly.com/catalog/9780596101732.

Allen:2006:ADC

[52] Robbie Allen and Laura E. Hunter. Active Directory Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10202-X. xxviii + 962 pp. LCCN TK5105.595 .A4497 2006. EUR 48.00. URL http://

www.oreilly.com/catalog/9780596102029.

Alspach:2006:ACS

[53] Jennifer Alspach, Shari Nakano, Steve Samson, et al. Adobe Creative Suite 2 workflow. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10236-4. 16 + 614 pp. LCCN T385 .A462 2006; T385 .A462 2006eb; T385. URL http://www.

oreilly.com/catalog/9780596102364.

AlZabir:2007:BAN

[54] Omar Al Zabir. Building an ASP.NET Web 2.0 Portal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51050-0. LCCN TK5105.8885.A26; TK5105.8885.A26eb. URL http://www.oreilly.

com/catalog/9780596510503.

AlZabir:2008:BWP

[55] Omar Al Zabir. Building a Web 2.0 Portal with ASP.NET 3.5. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51050-0. xvi + 290 pp. LCCN TK5105.8885.A26 A5 2008.

Anderson-Redick:2000:WSP

[56] Stacey Anderson-Redick. Windows System Policy Editor. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2000. ISBN 1-56592-649-8. xii + 533 pp. LCCN QA76.76.O63 A532 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565926493;   http://
www.oreilly.com/catalog/winsyspe.

Andrew:2004:CAE

[57] Rachel Andrew. The CSS anthology: 101 essential tips, tricks and hacks. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2004. ISBN 0-9579218-8-8 (paperback). xiv + 391 pp. LCCN TK5105.888 A54 2004. URL

http://www.oreilly.com/catalog/9780957921887.

Andrew:2005:BYO

[58] Rachel Andrew. Build your own standards compliant website using Dreamweaver 8. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9752402-3-4. xiv + 335 pp. LCCN TK5105.8885.D74 A53 2005eb;

TK5105.8.D74 A53 2005. URL http://www.oreilly.com/catalog/
9780975240236.

Anglin:2007:IS

[59] Todd Anglin. Introducing Silverlight 1.1. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51583-9. LCCN QA76.575; QA76.575 .A54 2007eb. URL http://www.oreilly.com/

catalog/9780596515836.

Angus:2006:PST

[60] Chris Angus. Prototype and scriptaculous: taking the pain out of JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52919-8. LCCN QA76.73.J38 A54 2006eb; QA76.73.J38. URL http://www.

oreilly.com/catalog/9780596529192.

Anonymous:1993:WIC

[61] The Whole Internet catalog, 1993. ISSN 1072-043X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800.

Anonymous:1996:PMY

[62] Anonymous, editor. Passport Malaysia: Your Pocket Guide to Malaysian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-36-4. 96 pp. LCCN ???? US$6.95. URL http://

www.oreilly.com/catalog/9781885073365.

Anonymous:1996:PPY

[63] Anonymous, editor. Passport Poland: Your Pocket Guide to Polish Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-33-X. 96 pp. LCCN HD58.7 .K528 1999. US$6.95. URL

http://www.oreilly.com/catalog/9781885073334.

Anonymous:1996:WWWa

[64] Anonymous, editor. World Wide Web Journal: Volume 1, Issue 3: The Web After Five Years. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1996. ISBN 1-56592-210-7. vii + 218 pp. LCCN ???? US$24.95. URL

http://www.oreilly.com/catalog/9781565922105.

Anonymous:1996:WWWb

[65] Anonymous, editor. World Wide Web Journal: Volume 1, Issue 4: Building an Industrial Strength Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-211-5. 252 pp. LCCN ???? US$24.95. URL http:

//www.oreilly.com/catalog/9781565922112.

Anonymous:1998:UCB

[66] Anonymous. The UNIX CD bookshelf: 6 bestselling books on CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-001-5. ???? pp. LCCN QA76.76.O63 U5453 1992. Includes: UNIX in a nutshell: a desktop quick reference for System V Release 4 and Solaris 2.0, UNIX power tools, Sed & awk, Learning the Korn shell, Learning Vi, and Learning the UNIX operating system.

Anonymous:19xx:DCJ

[67] Anonymous. Distributed Computing for Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN ???? ???? pp. LCCN ???? ????

Anonymous:2000:PPC

[68] Anonymous. Proceedings of the Perl Conference 4.0: July 17–20, 2000, Monterey, California. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 0-596-00013-8. viii + 135 pp. LCCN QA76.73.P22 P475 2000. URL

http://www.oreilly.com/catalog/9780596000134.

Anonymous:2000:UCB

[69] Anonymous. The UNIX CD bookshelf [computer file]: 6 bestselling books on CD-ROM: UNIX power tools; UNIX in a nutshell; Learning the UNIX operating system; sed and awk; Learning the vi editor; Learning the Korn shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 2.1. edition, February 2000. ISBN 0-596-00000-6. 624 pp. LCCN QA76.76.O63.

US$69.95. URL http://www.oreilly.com/catalog/9780596000004;
 http://www.oreilly.com/catalog/unixcdbs2. Includes CD-ROM.

Anonymous:2004:PPT

[70] Anonymous. Powerpoint 2003 personal trainer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00855-4. xiii + 315 pp. LCCN T385 .P6767 2004. URL

http://www.oreilly.com/catalog/9780596008550.

Anonymous:2005:APT

[71] Anonymous. Access 2003 personal trainer. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00937-2. xiii + 348 pp. LCCN QA76.9.D3 A135 2005eb; QA76.9.D3; QA76.9.D3 A135 2005. URL http://www.oreilly.

com/catalog/9780596009373.

Anonymous:2005:EPT

[72] Anonymous. Excel 2003 personal trainer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00853-8. xiii + 464 pp. LCCN HF5548.4.M532; HF5548.4.M523 E929855 2005; HF5548.4.M532 E93 2004; HF5548.4.M532 E93 2004eb.

URL http://www.oreilly.com/catalog/9780596008536.

Anonymous:2005:PPT

[73] Anonymous. Project 2003 personal trainer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00854-6 (paperback). xv + 216 pp. LCCN HD69.P75 P72 2005.

URL http://www.oreilly.com/catalog/9780596008543.

Anonymous:2005:WPT

[74] Anonymous. Word 2003 personal trainer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00936-4 (paperback). xvii + 433 pp. LCCN Z52.5.M52 C98 2005.

URL http://www.oreilly.com/catalog/9780596009366.

Anonymous:2005:WXP

[75] Anonymous. Windows XP personal trainer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00862-7. xv + 452 pp. LCCN QA76.76.W56 W56 2004. URL

http://www.oreilly.com/catalog/9780596008628.

Antonopoulos:2015:MB

[76] Andreas M. Antonopoulos. Mastering Bitcoin. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4493-7404-2, 1-4919-0260-4. xxi + 272 pp. LCCN HG1710 .A58 2014.

Armstrong:2007:PES

[77] Joe Armstrong. Programming ERLANG: software for a concurrent world. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 1-934356-00-X (paperback). xiii + 515 pp. LCCN

QA76.73.E75 A765 2007. URL http://www.oreilly.com/catalog/.

Arnett:1996:IPJ

[78] Nick Arnett and David Land. Internet Programming with JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-56592-205-0. 250 pp. LCCN ???? US$24.95. URL http://www.oreilly.com/

catalog/9781565922051.

Aronson:2006:DFS

[79] Ian David Aronson. DV filmmaking: from start to finish. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00848-1. xviii + 290 pp. LCCN TR860; TR860 .A76 2006eb; TR860 .A76 2006. URL

http://www.oreilly.com/catalog/9780596008482.

Aspinwall:2005:PH

[80] Jim Aspinwall. PC hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00748-5. xvii + 285 pp. LCCN QA76.5 .A77 2005; QA76.5 .A875 2005; QA76.5; QA76.5 .A875 2005eb. URL http://www.oreilly.com/

catalog/9780596007485.

Avery:2005:VSH

[81] James Avery. Visual Studio hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00847-3. xx + 478 pp. LCCN TK5105.8885.M57 A94 2005. URL http:

//www.oreilly.com/catalog/9780596008475.

Avery:2007:WDP

[82] James Avery and Jim Holmes. Windows developer power tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52754-3. xlii + 1263 pp. LCCN QA76.76.D47; QA76.76.D47 A74 2007eb. URL http://www.oreilly.

com/catalog/9780596527549.

Babbit:2000:ADA

[83] Nikki Babbit. Adolescent drug and alcohol abuse: how to spot it, stop it, and get help for your family. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-755-9. xx + 276 pp. LCCN HV5824.Y68 B27 2000. URL http://www.oreilly.com/

catalog/9781565927551.

Bagui:2006:LSS

[84] Sikha Bagui and Richard Earp. Learning SQL on SQL Server 2005. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10215-1. xvi + 324 pp. LCCN QA76.73.S67 B34 2006eb; QA76.73.S67. URL http://

www.oreilly.com/catalog/9780596102159.

Baird:2007:REC

[85] Kevin C. Baird. Ruby by example: concepts and code. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-148-4. xxiii + 281 pp.

LCCN QA76.64 .B27 2007. URL http://www.loc.gov/catdir/toc/
ecip0717/2007018653.html;   http://www.oreilly.com/catalog/
9781593271480.

Baker:1995:LCF

[86] Brenda S. Baker and Eric Grosse. Local control over filtered WWW access. In Fourth International World Wide Web Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1995. URL http://www.w3.org/pub/

Conferences/WWW4/Papers/117/. Boston, MA.

Bales:2002:JPO

[87] Donald Bales. Java Programming with Oracle JDBC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00088-X. xv + 477 pp. LCCN QA76.73.J38 B354 2002. US$39.95. URL http://safari.oreilly.com/

059600088X;   http://www.oreilly.com/catalog/9780596000882;
 http://www.oreilly.com/catalog/jorajdbc.

Bales:2003:JPR

[88] Donald Bales. JDBC pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00457-5. vi + 145 pp. LCCN QA76.73.J38 B355 2003.

URL http://www.oreilly.com/catalog/9780596004576.

Ballew:1997:MIN

[89] Scott M. Ballew. Managing IP Networks with Cisco Routers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-320-0. xiv + 334 pp. LCCN TK5105.585 .B35 1997.

US$29.95. URL http://www.ora.com/catalog/cisco/;   http://
www.oreilly.com/catalog/9781565923201;   http://www.oreilly.
com/catalog/cisco.

Ballew:2004:DYM

[90] Joli Ballew. Degunking your Mac. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2004. ISBN 1-932111-94-8. xxvi + 268 pp. LCCN QA76.76.S64

B35 2004. URL http://www.oreilly.com/catalog/9781932111941.

Ballew:2005:DYM

[91] Joli Ballew. Degunking your Mac. Paraglyph Press, Inc., Scottsdale, AZ, USA, tiger edition, 2005. ISBN 1-933097-05-1. xxvii + 396 pp. LCCN QA76.76.S64 B35 2005. URL http://www.oreilly.com/

catalog/9781933097053.

Ballew:2005:DYP

[92] Joli Ballew and Jeff Duntemann. Degunking your PC. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-03-5. xxix + 370 pp. LCCN QA76.76.S64 B35 2005. URL http://www.oreilly.com/

catalog/9781933097039.

Balmain:2007:F

[93] David Balmain. Ferret. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52785-3. LCCN QA76.9.D3; QA76.9.D3 B35 2007; QA76.9.D3 B35 2007eb; Internet.

Bambenek:2009:GPR

[94] John Bambenek and Agnieszka Klus. grep: Pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15360-0 (paperback). v + 75 pp. LCCN QA76.76.U84 B36 2009.

Banzi:2015:AE

[95] Massimo Banzi and Michael Shiloh. Arduino für Einsteiger. (German) [Arduino for beginners]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2015. ISBN 3-95875-055-9. 232 pp. LCCN ????

Barber:2001:OMM

[96] Nan Barber and David Reynolds. Office 2001 for Macintosh: The Missing Manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, May 2001. ISBN 0-596-00081-2. x + 637 pp. LCCN HF5548.4.M525 B366 2001; HF5548.4.M525 B37 2001. US$29.95.

URL http://www.oreilly.com/catalog/9780596000813;   http://
www.oreilly.com/catalog/msoffmacmm.

Barber:2002:OXM

[97] Nan Barber, Tonya Engst, and David Reynolds. Office X for Macintosh: The Missing Manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2002. ISBN 0-596-00332-3. xii + 714 pp. LCCN HF5548.4.M525 B368 2002. US$29.95.

URL http://www.oreilly.com/catalog/9780596003326;   http://
www.oreilly.com/catalog/offxtmm.

Barlow:1994:NSI

[98] John Perry Barlow. Notable Speeches of the Information Age. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1994. ISBN 1-56592-992-6.

LCCN ???? US$9.95. URL http://www.oreilly.com/catalog/

9781565929920. USENIX Conference Keynote Address, January 17, 1994; San Francisco, CA. Audio tape: 90 minutes.

Barr:1999:PES

[99] Michael Barr. Programming Embedded Systems in C and C++. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-354-5. xvii + 174 pp. LCCN QA76.73.C153 B375 1999; TK7895.E42 B37 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565923546;   http://
www.oreilly.com/catalog/embsys/.

Barr:2006:PES

[100] Michael Barr and Anthony J. Massa. Programming embedded systems: with C and GNU development tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00983-6 (paperback). xxi + 301 pp. LCCN TK7895.E42 B37 2006eb; TK7895.E42. URL http://www.loc.gov/

catdir/enhancements/fy0715/2006287300-d.html;   http://www.
loc.gov/catdir/toc/fy0709/2006287300.html;   http://www.oreilly.
com/catalog/9780596009830.

Barrett:1996:BIS

[101] Daniel J. Barrett. Bandits on the Information Superhighway. What you need to know. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-156-9. xiii + 229 pp. LCCN HV6773 .B37 1996. US$17.95. URL

http://www.ora.com/gnn/bus/ora/item/bandits.html;   http://
www.oreilly.com/catalog/9781565921566;   http://www.oreilly.
com/catalog/bandits.

Barrett:1997:NFI

[102] Daniel J. Barrett. NetResearch: Finding Information Online. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-245-X. xiii + 186 pp. LCCN TK5105.875 .I57 B37 1997. US$24.95. URL http://www.ora.

com/catalog/netresearch/;   http://www.oreilly.com/catalog/
9781565922457;   http://www.oreilly.com/catalog/netresearch.

Barrett:2001:SSS

[103] Daniel J. Barrett and Richard E. Silverman. SSH: The Secure Shell: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00011-1. xv + 540 pp. LCCN QA76.76.O63 B369 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596000110;   http://
www.oreilly.com/catalog/sshtdg/; http://www.snailbook.com/.

Barrett:2003:LSC

[104] Daniel J. Barrett, Richard E. Silverman, and Robert G. Byrnes. Linux security cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00391-9. xvii + 311 pp. LCCN TK5105.59 .B36 2003. URL http://

www.oreilly.com/catalog/9780596003913.

Barrett:2004:LKG

[105] Daniel J. Barrett and Torsten Wilhelm. Linux — kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 3-89721-501-2 (paperback). 204 pp. LCCN ???? EUR 9.90. URL http://www.

gbv.de/dms/ilmenau/toc/388679999.PDF.

Barrett:2004:LPG

[106] Daniel J. Barrett. Linux pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00628-4. viii + 191 pp. LCCN QA76.76.O63 B3685 2004. URL http:/

/www.oreilly.com/catalog/9780596006280.

Barrett:2005:SSS

[107] Daniel J. Barrett, Richard E. Silverman, and Robert G. Byrnes. SSH: The Secure Shell: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00895-3. 672 (est.) pp. LCCN QA76.76.O63 B369 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596008956;   http://
www.oreilly.com/catalog/sshtdg2/.

Barth:2006:NSN

[108] Wolfgang Barth. Nagios: system and network monitoring. Open Source Press, Munich, Germany, 2006. ISBN 1-59327-070-4. 462 pp. LCCN TK5105.5 .B374 2006; TK5105.5 .B374 2006eb. URL http://www.

oreilly.com/catalog/9781593270704.

Bartmann:2014:EWA

[109] Erik Bartmann. Die elektronische Welt mit Arduino entdecken: [mit dem Arduino messen, steuern und spielen; Elektronik leicht verstehen; 44 Arduino-Projekte für den Selbstbau]. (German) [Discover the E-world with Arduino: measure, control and play with the Arduino; easy to understand electronics; 44 Arduino Projects for the do-it-yourselfer]. O’Reillys basics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2014. ISBN 3-95561-115-9. xxii + 1051 pp. LCCN ???? URL http://d-nb.info/

1048697568/04.

Bartok:1995:HH

[110] ?. Bartok. HTML Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-119-4. ???? pp. LCCN ???? URL http://www.oreilly.

com/catalog/9781565921191.

Bartok:19xx:HH

[111] ?. Bartok. HTML Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 1-56592-119-4. ???? pp. LCCN ????

Bartosh:2005:EMX

[112] Michael Bartosh and Ryan Faas. Essential Mac OS X Panther server administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00635-7 (paperback). xxii + 822 pp. LCCN QA76.76.O63 B38 2005. URL http:

//www.oreilly.com/catalog/9780596006358.

Basham:2004:HFS

[113] Bryan Basham, Kathy Sierra, and Bert Bates. Head first servlets and JSP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00540-7, 0-596-00540-7. xxxii + 854 pp. LCCN QA76.3 .B384 2004.

Basham:2008:HFS

[114] Bryan Basham, Kathy Sierra, and Bert Bates. Head first servlets and JSP. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2008. ISBN 0-596-51668-1. xxxii + 879 pp. LCCN QA76.3 .B377 2008eb; QA76.3; QA76.3 .B377 2008; Internet.

Bass:2004:PAH

[115] Steve Bass. PC annoyances: how to fix the most annoying things about your personal computer. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00593-8. xxii + 176 pp. LCCN QA76.5 .B319 2003. URL http:/

/www.oreilly.com/catalog/9780596005931.

Bass:2005:PAH

[116] Steve Bass. PC annoyances: how to fix the most annoying things about your personal computer, Windows, and more. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00882-1. xix + 230 pp. LCCN QA76.5 .B38 2005. URL

http://www.oreilly.com/catalog/9780596008826.

Bauer:2002:BSS

[117] Michael D. Bauer. Building Secure Servers with Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00217-3. xv + 430 pp. LCCN TK5105.59

.B38 2002. US$44.95. URL http://www.oreilly.com/catalog/
9780596002176;   http://www.oreilly.com/catalog/bssrvrlnx.

Bauer:2005:LSS

[118] Michael D. Bauer. Linux server security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00670-5. xvii + 522 pp. LCCN TK5105.59 .B38 2005; TK5105.59 .B38 2005eb; TK5105.59. URL http:

//www.oreilly.com/catalog/9780596006709.

Bausch:2003:AHI

[119] Paul Bausch. Amazon hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00542-3. xix + 280 pp. LCCN HF5548.323.A4 B38 2003. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/;   http://www.loc.gov/catdir/
toc/fy045/2004272811.html;   http://www.oreilly.com/catalog/
9780596005429.

Bausch:2006:FH

[120] Paul Bausch and Jim Bumgardner. Flickr hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10245-3. xxv + 335 pp. LCCN

TR267; TR267 .B38 2006eb. URL http://www.oreilly.com/catalog/
9780596102456.

Bausch:2006:YH

[121] Paul Bausch. Yahoo! hacks. Hacks series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00945-3 (paperback). xxiii + 464 pp. LCCN ZA4234.Y33 B38 2006. URL http://www.oreilly.com/

catalog/9780596009458.

Bautts:2005:LNA

[122] Tony Bautts, Terry Dawson, and Gregor N. Purdy. Linux network administrator’s guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00548-2. xxii + 338 pp. LCCN QA76.76.O63; QA76.76.O63 K566 2005; QA76.76.O63 K566 2005eb. URL http://

www.oreilly.com/catalog/9780596005481.

BDT:1998:ATO

[123] The Be Development Team. Be Advanced Topics: Official documentation for the BeOS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1998. ISBN 1-56592-396-0. xi + 365 pp. LCCN QA76.76.O63 B42 1998. US$39.95.

URL http://www.oreilly.com/catalog/9781565923966;   http://
www.oreilly.com/catalog/beadv. Includes BeSpecific 5 CD-ROM.

Be:1997:DGO

[124] The Be Development Team. The Be Developer’s Guide: The Official Documentation for the BeOS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-287-5. xi + 934 pp. LCCN QA76.76.O63

B43 1997. US$49.95. URL http://www.ora.com/catalog/be/;
http://www.oreilly.com/catalog/9781565922877;   http://www.
oreilly.com/catalog/bedev.

Beaird:2007:PBW

[125] Jason Beaird. The principles of beautiful Web design. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2007. ISBN 0-9758419-6-3. xii + 168 pp. LCCN TK5105.888 .B42 2007. URL http://www.oreilly.com/

catalog/9780975841969.

Beam:2003:CN

[126] Michael Beam and James Duncan Davidson. Cocoa in a nutshell. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00462-1. xvii + 545 pp. LCCN QA76.64 .B398 2003. URL http://www.

oreilly.com/catalog/9780596004620.

Beardsworth:2005:PBM

[127] John Beardsworth. Photoshop blending modes cookbook for digital photographers. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10020-5. 176 pp. LCCN T385 .B4237 2005 .B439 2005; TR267; T385 .B4237 2005eb; T385. URL http://www.oreilly.

com/catalog/9780596100209.

Beardsworth:2006:PFA

[128] John Beardsworth. Photoshop fine art effects cookbook for digital photographers. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10062-0. 176 pp. LCCN T385 .B4238 2006eb; T385 .B4238

2006; T385. URL http://www.oreilly.com/catalog/9780596100629.

Beaulieu:2005:LS

[129] Alan Beaulieu. Learning SQL. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00727-2. xiii + 289 pp. LCCN QA76.73.S67 B42 2005; QA76.73.S67 B43 2005; QA76.73.S67; QA76.73.S67 B43 2005eb. URL http://www.oreilly.

com/catalog/9780596007270.

Beazley:2014:PCR

[130] David M. Beazley and Brian K. (Brian Kenneth) Jones. Python cookbook: Recipes for mastering Python 3. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2014. ISBN 1-4493-4037-7 (paperback), 1-4493-5736-9 (e-book). xvi + 687 pp. LCCN QA76.73.P98 B43 2013eb.

Beck:2004:JPG

[131] Kent Beck. JUnit pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00743-4 (paperback). vi + 84 pp. LCCN QA76.76.T48 B43

2004. URL http://www.oreilly.com/catalog/9780596007430.

Beggs:2001:DWA

[132] Josh Beggs and Dylan Thede. Designing Web Audio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2001. ISBN 1-56592-353-7. xiii + 382 pp. LCCN TK5105.8863 .B44 2001. US$34.95.

URL http://www.oreilly.com/catalog/9781565923539;   http://
www.oreilly.com/catalog/sound.

Beighley:2007:HFS

[133] Lynn Beighley. Head first SQL. Head first series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52684-9, 0-596-52684-9. xxxv + 571 pp. LCCN QA76.73.S67 B45 2007; QA76.73.S67; QA76.73.S67 B45 2007eb. URL http://www.

oreilly.com/catalog/9780596526849.

Bekman:2003:PM

[134] Stas Bekman and Eric Cholet. Practical mod perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00227-0. xxvii + 893 pp. LCCN

TK5105.8885.A63 B45 2003. URL http://www.oreilly.com/catalog/
9780596002275.

Bellinger:2006:GAO

[135] Amy Bellinger. Getting acquainted with OPML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52822-1. LCCN QA76.76.H94; QA76.76.H94

B45 2006eb. URL http://www.oreilly.com/catalog/9780596528225.

Bengal:2005:IAP

[136] Bengal, editor. Illustrations avec Photoshop. (French) [Illustrations with Photoshop: a designer’s notebook]. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00859-7. 92 pp. LCCN T385 .B454 2005. URL http://www.oreilly.

com/catalog/9780596008598.

Benvenuti:2006:ULN

[137] Christian Benvenuti. Understanding Linux network internals. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00255-6. xxiv + 1035 pp. LCCN QA76.76.O63; QA76.76.O63 B46 2006eb; QA76.76.O63 B46 2006. URL

http://www.oreilly.com/catalog/9780596002558.

Bergmann:2006:PPG

[138] Sebastian Bergmann. PHPUnit pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10103-1. vii + 79 pp. LCCN QA76.73.P224

B47 2006eb; QA76.73.P224. URL http://www.oreilly.com/catalog/
9780596101039.

Bergsten:2001:JP

[139] Hans Bergsten. JavaServer pages. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-746-X. xviii + 552 pp. LCCN QA76.73.J38 B49 2001; TK5105.8885.J38 B37 2001. URL http://www.oreilly.com/ catalog/9781565927469. Editors: Robert Eckstein and Paula Ferguson.

Bergsten:2001:JPP

[140] Hans Bergsten. JavaServer Pages: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00231-9. v + 81 pp. LCCN TK5105.8885.J38 B473 2001. US$11.95.

URL http://www.oreilly.com/catalog/9780596002312;   http://
www.oreilly.com/catalog/javaserverpr.

Bergsten:2002:JP

[141] Hans Bergsten. JavaServer Pages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00317-X. xxii + 660 pp. LCCN TK5105.8885.J38 B47 2002. US$44.95. URL

http://safari.oreilly.com/059600317X;   http://www.oreilly.
com/catalog/9780596003173;   http://www.oreilly.com/catalog/
jserverpages2.

Bergsten:2004:JF

[142] Hans Bergsten. JavaServer faces. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00539-3. xiv + 589 pp. LCCN TK5105.8885.J38 B46

2004. URL http://www.oreilly.com/catalog/9780596005399.

Bergsten:2004:JP

[143] Hans Bergsten. JavaServer pages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00563-6. xxii + 740 pp. LCCN

TK5105.8885.J38 B47 2004. URL http://www.oreilly.com/catalog/
9780596005634.

Berkun:2005:APM

[144] Scott Berkun. The art of project management. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00786-8. xii + 488 pp. LCCN HD69.P75

B47 2005. URL http://www.oreilly.com/catalog/9780596007867.

Berkun:2007:MI

[145] Scott Berkun. The myths of innovation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN (hardcover), 0-596-52705-5 (hardcover), 978-059-652-7. xiii + 176 pp. LCCN T173.8 .B457 2007; T173.8; T173.8 .B4688 2007.

URL http://www.oreilly.com/catalog/9780596527051.

Berry:2008:RPR

[146] Eric Berry. Rails pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52070-0. x + 186 pp. LCCN QA76.73.R83

B47 2008. URL http://www.loc.gov/catdir/enhancements/fy0915/
2009278937-d.html.

Bhangal:2004:FH

[147] Sham Bhangal. Flash hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00645-4. xxvii + 466 pp. LCCN

TR897.7 .B48797 2004. URL http://www.oreilly.com/catalog/
9780596006457.

Bhardwaj:2007:MCE

[148] Pawan K. Bhardwaj and Roger A. Grimes. MCSE core elective exams in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10229-1. xii + 571 pp. LCCN QA76.76.O63; QA76.76.O63 B45 2007eb. URL

http://www.oreilly.com/catalog/9780596102296.

Bhardwaj:2007:NSE

[149] Pawan K. Bhardwaj. A+, Network+, Security+ exams in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52824-8. xii + 798 pp. LCCN TK7885.54 B48 2007eb; TK7885.54. URL http://www.

oreilly.com/catalog/9780596528249.

Biafore:2004:OIH

[150] Bonnie Biafore. Online investing hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00677-2. xviii + 483 pp.

LCCN HG4515.95 .B53 2004. URL http://www.oreilly.com/catalog/
9780596006778.

Biafore:2005:QMM

[151] Bonnie Biafore. QuickBooks 2005: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00901-1. xiv + 592 pp. LCCN HF5679 .B53 2005. URL http://

www.oreilly.com/catalog/9780596009014.

Biafore:2006:QMM

[152] Bonnie Biafore. QuickBooks 2006: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10184-8. xiiiv + 608 pp. LCCN HF5679 .B53 2006; HF5679 .B53 2006eb; HF5679. URL http://www.oreilly.com/

catalog/9780596101848.

Biafore:2006:QSE

[153] Bonnie Biafore. Quicken 2006 for starters: exactly what you need to get started. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10127-9. xvii + 407 pp. LCCN

HG179 .B53 2006eb; HG179. URL http://www.oreilly.com/catalog/
9780596101275.

Biafore:2008:Q

[154] Bonnie Biafore. Quicken 2009. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52248-7. xvi + 557 pp. LCCN HG179 .B53 2008.

Biancuzzi:2009:MP

[155] Federico Biancuzzi and Shane Warden, editors. Masterminds of programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-51517-0 (paperback). xii + 480 pp. LCCN QA76.6 .M37 2009.

Biddle:1997:WSM

[156] C. Bradford Biddle, Simson Garfinkel, John Gilmore, Rohit Khare, Cricket Liu, Lincoln Stein, et al. Web Security: a Matter of Trust. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-329-4. 282

pp. LCCN ???? US$29.95. URL http://www.ora.com/catalog/
wjsum97/;   http://www.oreilly.com/catalog/9781565923294;
http://www.oreilly.com/catalog/wjsum97.

Biersdorfer:2003:IIM

[157] J. D. Biersdorfer and David Pogue. iPod and iTunes: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2003. ISBN 0-596-00658-6. x + 349 pp.

LCCN ML74.4.I48 B54 2004. URL http://www.oreilly.com/catalog/
9780596006587.

Biersdorfer:2003:IMM

[158] J. D. Biersdorfer. iPod: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00477-X. x + 331 pp. LCCN ML74.4.I48 B54 2003. URL http://

www.oreilly.com/catalog/9780596004774.

Biersdorfer:2005:IIM

[159] J. D. Biersdorfer, David Pogue, and Peter Meyers. IPod and iTunes: the missing manual. Missing manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, third edition, 2005. ISBN 0-596-00877-5 (paperback). 402 pp. LCCN ML74.4.I48 B54 2005. URL http://www.oreilly.com/

catalog/9780596008772.

Biersdorfer:2005:ISF

[160] J. D. Biersdorfer. iPod Shuffle Fan Book: Life Is a Playlist. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10070-1. 96 (est.) pp. LCCN http://www.oreilly.com/catalog/ipodshufflefb/. US$9.95, CDN$13.95,

UK£6.95. URL http://www.oreilly.com/catalog/9780596100704.

Biersdorfer:2007:IMM

[161] J. D. Biersdorfer and David Pogue. iPod: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, sixth edition, 2007. ISBN 0-596-51491-3. xv + 278 pp. LCCN ML74.4.I48. URL http://www.oreilly.

com/catalog/9780596514914.

Biersdorfer:2010:IMM

[162] J. D. Biersdorfer and David Pogue. iPad: the missing manual. The missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-8784-5. xvii + 299 pp. LCCN ????

Biersdorfer:2012:IMM

[163] J. D. Biersdorfer and David Pogue. iPod: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1619-0. 352 pp. LCCN ???? EUR 18.00.

Bigelow:2005:PHA

[164] Stephen J. Bigelow. PC hardware annoyances: how to fix the most annoying things about your computer hardware. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00715-9. xiii + 246 pp. LCCN TK7887.5 .B52 2004; TK7887.5; TK7887.5 .B52 2004eb. URL http://www.oreilly.com/

catalog/9780596007157.

Bird:2009:NLP

[165] Steven Bird, Ewan Klein, and Edward Loper. Natural Language Processing with Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-51649-5. xx + 479 pp. LCCN QA76.73.P98 B57 2009.

Bishop:2007:CDP

[166] J. M. (Judith Mary) Bishop. C# design patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52773-X (paperback). xxi + 290 pp. LCCN

Internet. URL http://www.oreilly.com/catalog/9780596527730.

Blackman:1996:LRI

[167] Josh Blackman. The Legal Researcher’s Internet Directory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-160-7. ???? pp.

LCCN JK1764.S38 1996. URL http://www.oreilly.com/catalog/

9781565921603. Same ISBN as [1740]. Did this book appear?

Blandy:2015:RPL

[168] Jim Blandy. The Rust Programming Language: Fast, Safe, and Beautiful. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-2544-2. ???? pp. LCCN ???? Video file (1h10m).

Blandy:2017:PR

[169] Jim Blandy and Jason Orendorff. Programming Rust. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-2728-3 (paperback), 1-4919-2727-5, 1-4919-2723-2 (e-book), 1-4919-2725-9 (e-book). xx + 598 pp. LCCN

QA76.73.R88 B53 2017. URL http://proquest.safaribooksonline.
com/9781491927274.

Blandy:2021:PRF

[170] Jim Blandy, Jason Orendorff, and Leonora F. S. Tindall. Programming Rust: Fast, Safe Systems Development. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2021. ISBN 1-4920-5259-0. xix + 713 pp. LCCN QA76.73.R87 B58 2021.

Blank-Edelman:2000:PSA

[171] David N. Blank-Edelman. Perl for System Administration: Managing multi-platform environments with Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2000. ISBN 1-56592-609-9. xiv + 430 pp. LCCN QA76.73.P22 B43 2000.

US$34.95. URL http://www.oreilly.com/catalog/9781565926097;
 http://www.oreilly.com/catalog/perlsysadm.

Blank:2014:HS

[172] Steve Blank and Bob Dorf. Das Handbuch für Startups. (German) [The Startups Handbook]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 3-95561-812-9. xxxv + 467 pp. LCCN

???? URL http://d-nb.info/1051456061/04;   http://swbplus.
bsz-bw.de/bsz412834766cov.htm.

Bloch:2007:HHH

[173] Christian Bloch. The HDRI handbook: high dynamic range imaging for photographers and CG artists. Rocky Nook, Santa Barbara, CA, USA, 2007. ISBN 1-933952-05-9. 344 pp. LCCN TR267 .B594 2007. URL

http://www.oreilly.com/catalog/9781933952055.

Bloomer:1992:PPR

[174] John Bloomer. Power Programming with RPC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1992. ISBN 0-937175-77-3. xxxii + 459 pp. LCCN QA76.9.D5 B55 1991. US$29.95. URL http://www.oreilly.com/

catalog/9780937175774;   http://www.oreilly.com/catalog/rpc.

Bluttman:2005:AH

[175] Ken Bluttman. Access hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00924-0. xvi + 334 pp. LCCN QA76.9.D3 B58 2005. URL

http://www.oreilly.com/catalog/9780596009243.

Bluttman:2007:ADA

[176] Ken Bluttman and Wayne S. Freeze. Access data analysis cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10122-8 (paperback). xiii + 351 pp. LCCN QA76.9.D3; QA76.9.D3 B5948 2007eb;

QA76.9.D3 B5948 2007. URL http://www.oreilly.com/catalog/
9780596101220.

Bodin:2005:PPP

[177] Bertrand Bodin, editor. Assembling panoramic photos. A designer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00975-5. 96 pp.

LCCN TR661 .P46 2005. URL http://www.oreilly.com/catalog/
9780596009755.

Bodner:2021:LGI

[178] Jon Bodner. Learning Go: an idiomatic approach to real-world Go programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2021. ISBN 1-4920-7721-6. xvii + 352 pp. LCCN QA76.73.G63 B64 2021.

Bolinger:1995:ARS

[179] Don Bolinger and Tan Bronson. Applying RCS and SCCS: From Source Control to Project Control. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-117-8. xxiii + 501 pp. LCCN QA76.76.D49 B64 1995; QA76.9.F5 B65 1995. US$29.95.

URL http://www.oreilly.com/catalog/9781565921177;   http://
www.oreilly.com/catalog/rcs.

Bond:1995:TTW

[180] Marybeth Bond, editor. Travelers’ Tales: a Woman’s World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1995. ISBN 1-885211-06-6. xxi + 391 pp. LCCN PN6071.T7 .B58 1995. US$17.95.

URL http://www.oreilly.com/catalog/9781885211064;   http://
www.oreilly.com/catalog/woman.

Bond:1996:TTG

[181] Marybeth Bond, editor. Travelers’ Tales: Gutsy Women: Travel Tips and Wisdom for the Road. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885211-15-5. 100 pp. LCCN PN6071.T7 .B58 1995. US$7.95.

URL http://www.oreilly.com/catalog/9781885211156;   http://
www.oreilly.com/catalog/gutsy.

Bond:1997:TTG

[182] Marybeth Bond. Travelers’ Tales: Gutsy Mamas. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-20-1. 100 (est.) pp.

LCCN ???? US$7.95. URL http://www.ora.com/catalog/mamas/
;   http://www.oreilly.com/catalog/9781885211200;   http://
www.oreilly.com/catalog/mamas.

Bond:1998:TTM

[183] Marybeth Bond and Pamela Michael. Travelers’ Tales: a Mother’s World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1998. ISBN 1-885211-26-0. 234 (est.) pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9781885211262;   http://
www.oreilly.com/catalog/mothers.

Bond:1999:WPT

[184] Marybeth Bond and Pamela Michael. A Woman’s Passion for Travel: More True Stories from a Woman’s World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-36-8. 328 pp. LCCN G465.W645 1999.

US$17.95. URL http://www.oreilly.com/catalog/9781885211361;
 http://www.oreilly.com/catalog/passion.

Boney:2002:CIN

[185] James Boney. Cisco IOS in a Nutshell: a desktop quick reference for IOS on IP networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-942-X. xii + 592 pp. LCCN TK5105.8.C57 B65 2002; TK5105.5 .B66

2002. US$34.95. URL http://safari.oreilly.com/156592942X;
http://www.oreilly.com/catalog/9781565929425;   http://www.
oreilly.com/catalog/cisiosnut.

Boney:2005:CIN

[186] James Boney. Cisco IOS in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00869-4 (paperback). xiii + 779 pp. LCCN

TK5105.8.C57 B65 2005. URL http://www.oreilly.com/catalog/
9780596008697;   http://www.oreilly.com/catalog/cisconut2/.

Booker:1997:IGA

[187] Meg Booker and Songline Studios. The Insider’s Guide to America Online. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1997. ISBN 1-56592-373-1. 416 (est.) pp. LCCN ???? US$24.95. URL http://www.oreilly.com/

catalog/9781565923737;   http://www.oreilly.com/catalog/aol.

Bornstein:2004:NX

[188] Niel M. Bornstein. .NET and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00397-8. xiv + 455 pp. LCCN QA76.76.H94 B67 2004.

URL http://www.oreilly.com/catalog/9780596003975.

Borrus:1999:FSH

[189] Kathy Borrus. The Fearless Shopper: How to Get the Best Deals on the Planet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-39-2. 296 pp. LCCN TX335.B635 1999. US$14.95.

URL http://www.oreilly.com/catalog/9781885211392;   http://
www.oreilly.com/catalog/shopper.

Bos:2023:RAL

[190] Mara Bos. Rust Atomics and Locks: Low-Level Concurrency in Practice. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2023. ISBN 1-0981-1944-4. xvii + 230 pp. LCCN QA76.73.R87 B67 2023.

Boslaugh:2008:SN

[191] Sarah Boslaugh and Paul A. Watters. Statistics in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51049-7. xxi + 452 pp. LCCN QA276.12 .B67 2008; QA276.12 .B67 2008eb; QA276.12.

Boswell:2002:CAM

[192] David Boswell, Brian King, Ian Oeschger, Pete Collins, and Eric Murphy, editors. Creating applications with Mozilla. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00052-9. xvii + 454 pp. LCCN QA76.625

.C73 2002. US$39.95. URL http://www.oreilly.com/catalog/
9780596000523;   http://www.oreilly.com/catalog/mozilla.

Boudreau:2003:NDG

[193] Tim Boudreau, Jesse Glick, Simeon Greene, Vaughn Spurlin, and Jack J. Woehr. NetBeans: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00280-7. xxiv + 646 pp. LCCN QA76.73.J38

N47 2003. US$44.95. URL http://www.oreilly.com/catalog/
9780596002800;   http://www.oreilly.com/catalog/netbeans.

Bourg:2002:PGD

[194] David M. Bourg. Physics for Game Developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00006-5. xvi + 326 pp. LCCN QC23.2

.B68 2002. US$39.95. URL http://www.oreilly.com/catalog/
9780596000066;   http://www.oreilly.com/catalog/physicsgame.

Bourg:2004:AGD

[195] David M. Bourg and Glenn Seemann. AI for game developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00555-5. xv + 373 pp. LCCN QC23.2 .B67 2004; QA76.76.C672 B68 2004; QA76.76.C672

B68 2004eb; QA76.76.C672. URL http://www.oreilly.com/catalog/
9780596005559.

Bourg:2006:ESE

[196] David M. Bourg. Excel scientific and engineering cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00879-1 (paperback). xv + 424 pp. LCCN HF5548.4.M523 B68 2006; HF5548.4.M523; HF5548.4.M523

B68 2006eb. URL http://www.oreilly.com/catalog/9780596008796.

Bourke:2001:SLB

[197] Tony Bourke. Server Load Balancing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00050-2. xiv + 175 pp. LCCN QA76.9.C55 B675 2001.

US$34.95. URL http://www.oreilly.com/catalog/9780596000509;
 http://www.oreilly.com/catalog/serverload.

Bove:2005:JSN

[198] Tony Bove. Just say no to Microsoft: how to ditch Microsoft and why it’s not as hard as you think. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-064-X. xvi + 243 pp. LCCN QA76.76.O63 B682

2005. URL http://www.oreilly.com/catalog/9781593270643.

Bovet:2000:ULK

[199] D. (Daniele) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B665 2001.

Bovet:2001:ULK

[200] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B665 2001; QA76.76.O63 B683

2001. URL http://www.oreilly.com/catalog/9780596000028.

Bovet:2002:ULK

[201] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002.

ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003.

US$49.95. URL http://www.oreilly.com/catalog/linuxkernel2.

Bovet:2003:ULK

[202] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003.

US$49.95. URL http://www.oreilly.com/catalog/9780596002138;
 http://www.oreilly.com/catalog/linuxkernel2.

Bovet:2006:ULK

[203] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00565-2 (paperback). xvi + 923 pp. LCCN QA76.76.O63

B683 2006. URL http://www.oreilly.com/catalog/9780596005658.

Bowen:2007:WNA

[204] Richard Cooper Bowen. What’s new in Apache Web Server 2.2? O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52927-9. LCCN QA76.6.A62;

QA76.6.A62 B69 2007eb. URL http://www.oreilly.com/catalog/
9780596529277.

Boyce:2001:WQF

[205] Jim Boyce. Windows 2000 Quick Fixes. Quick fixes. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00017-0. xviii + 285 pp. LCCN QA76.76.O63 B6867 2001; QA76.76.O63 B69 2001. US$29.95.

URL http://www.oreilly.com/catalog/9780596000172;   http://
www.oreilly.com/catalog/qfwin2000.

Boyd:2012:GSO

[206] Ryan Boyd. Getting started with OAuth 2.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1784-7, 1-4493-1160-1 (paperback). x + 66 pp. LCCN ????

Boynton:1996:PTW

[207] John Robert Boynton. PolyForm (TM): Web Forms Construction Kit. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1996. ISBN 1-56592-182-8. 146

pp. LCCN ???? US$149.00. URL http://www.oreilly.com/catalog/
9781565921825;   http://www.oreilly.com/catalog/polyform. In-

cludes two diskettes.

Bracewell:1998:OUQ

[208] Mark Bracewell and David A. Karp. O’Reilly Utilities: Quick Solutions for Windows 98 Annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-549-1. ???? pp. LCCN ???? US$39.95.

URL http://www.oreilly.com/catalog/9781565925496;   http://
www.oreilly.com/catalog/utilwin98. Includes CD-ROM.

Bradenbaugh:1999:JAC

[209] Jerry Bradenbaugh. JavaScript Application Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-577-7. xiv + 462 pp. LCCN QA76.73.J39 B73 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565925779;   http://
www.oreilly.com/catalog/jscook.

Bradski:2012:LOC

[210] Gary R. Bradski and Adrian Kaehler. Learning OpenCV: [computer vision with the OpenCV library]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2012. ISBN 1-4493-1465-1 (paperback). 300 pp. LCCN ????

Branagan:1994:FHB

[211] Linda Branagan and Mike Sierra. The Frame Handbook: Building FrameMaker Documents That Work. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1994. ISBN 1-56592-009-0. xxxiii + 506 pp. LCCN ????

US$24.95. URL http://www.oreilly.com/catalog/9781565920095.

Braunstein:2007:IF

[212] Roger Braunstein. Introduction to Flex 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-55003-0. LCCN QA76.625; QA76.625 .B73

2007eb. URL http://www.oreilly.com/catalog/9780596550035.

Breshears:2009:AC

[213] Clay Breshears. The art of concurrency. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-80242-0. xiii + 285 pp. LCCN QA76.642 .B74 2009; Internet.

Bressert:2012:SNO

[214] Eli Bressert. Scipy and Numpy: An Overview for Developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-0546-6. 80 pp. LCCN ????

Brickner:2005:LDP

[215] David Brickner. Linux desktop pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10104-X. xvi + 181 pp. LCCN QA76.76.O63 B7423 2005.

Brickner:2005:TDL

[216] David Brickner. Test driving Linux: from Windows to Linux in 60 seconds. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00754-X. xvii + 341 pp. LCCN QA76.76.O63 B744 2005. URL http://www.oreilly.

com/catalog/9780596007546.

Brittain:2003:TDG

[217] Jason Brittain and Ian F. Darwin. Tomcat: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00318-8. xiii + 303 pp. LCCN

TK5105.8885.A63 B75 2003. URL http://www.oreilly.com/catalog/
9780596003180.

Brittain:2007:TDG

[218] Jason Brittain and Ian F. Darwin. Tomcat: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-10106-6 (paperback). xvi + 476 pp. LCCN TK5105.8885.A63 B75 2008. URL

http://www.oreilly.com/catalog/9780596101060.

Brooks-Bilson:2001:PC

[219] Rob Brooks-Bilson. Programming ColdFusion. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-698-6. xvii + 953 pp. LCCN QA76.9.W43

B76 2001. US$49.95. URL http://www.oreilly.com/catalog/
9781565926981;   http://www.oreilly.com/catalog/coldfusion.

Brooks-Bilson:2003:PCM

[220] Rob Brooks-Bilson. Programming ColdFusion MX. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00380-3. xviii + 1115 pp. LCCN QA76.9.W43 B76 2003. URL http://www.oreilly.

com/catalog/9780596003807.

Broughton:2008:WMM

[221] John Broughton. Wikipedia: the missing manual. The missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51516-2. xxi + 477 pp. LCCN TK5105.8882 .B76 2008; TK5105.8882. URL http://

www.oreilly.com/catalog/9780596515164.

Broughton:2008:WRG

[222] John Broughton. Wikipedia reader’s guide: the missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52174-X. 62 pp. LCCN AE5 .B86.

Brown:2002:WAW

[223] Kyle Brown. WebSphere 4.0 AEs Workbook for Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00418-4. 186 (est.) pp. LCCN ???? US$24.95.

URL http://www.oreilly.com/catalog/9780596004187;   http://
www.oreilly.com/catalog/websphereworkbook.

Brown:2006:DAA

[224] Tracy Steven Brown. Dynamic Apache with Ajax and JSON2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52839-6. LCCN TK5105.8885.A63 B76 2006eb; TK5105.8885.A63. URL http://www.

oreilly.com/catalog/9780596528393.

Brown:2006:SL

[225] Chris Brown. SuSE Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10183-X. xiii + 430 pp. LCCN QA76.76.O63 B76 2006.

US$39.99. URL http://www.oreilly.com/catalog/9780596101831.

Brown:2007:ITW

[226] C. Titus Brown, Gheorghe Gheorghiu, and Jason Huggins. An introduction to testing web applications with twill and Selenium. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52780-2. LCCN QA76.76.T48

B76 2007eb; QA76.76.T48. URL http://www.oreilly.com/catalog/
9780596527808.

Brownell:2002:S

[227] David Brownell. SAX2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00237-8. ix + 228 pp. LCCN QA76.76.H94 B789

2002. US$29.95. URL http://safari.oreilly.com/0596002378;
http://www.oreilly.com/catalog/9780596002374;   http://www.
oreilly.com/catalog/sax2.

Brundage:2005:PE

[228] Barbara (Barbara K.) Brundage. Photoshop Elements 3. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00453-2. xxiv + 499 pp. LCCN TR267 .B786 2005. URL http://www.oreilly.com/

catalog/9780596004538.

Brundage:2006:PEMa

[229] Barbara (Barbara K.) Brundage. Photoshop elements 4: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10158-9. xiii + 493 pp. LCCN TR267.5.A33 B78 2006eb; TR267.5.A33

B78 2006; TR267.5.A33. URL http://www.oreilly.com/catalog/
9780596101589.

Brundage:2006:PEMb

[230] Barbara (Barbara K.) Brundage. Photoshop elements 5: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52728-4. xvii + 554 pp. LCCN TR267.5.A33; TR267.5.A33 B79 2006eb. URL http:

//www.oreilly.com/catalog/9780596527280.

Brundage:2008:PEM

[231] Barbara (Barbara K.) Brundage. Photoshop elements 6: the missing manual. Missing manual series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51444-1 (paperback). xvi + 567 pp. LCCN TR267.5.A33

B783 2008. URL http://www.oreilly.com/catalog/9780596514440.

Buchholz:2001:LLL

[232] William M. Buchholz and Susan W. Buchholz. Live longer, live larger: a holistic approach for cancer patients and their families. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-845-8. xiv + 389 pp. LCCN RC263 .B755 2001. URL http://www.oreilly.com/

catalog/9781565928459.

Bumgardner:2006:BTC

[233] Jim Bumgardner. Building tag clouds in Perl and PHP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52794-2. 48 pp. LCCN QA76.73.P22

B86 2006eb; QA76.73.P22. URL http://www.oreilly.com/catalog/
9780596527945.

Burke:2001:JX

[234] Eric M. Burke. Java and XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00143-6. xvi + 510 pp. LCCN QA76.73.J38 B88 2001.

Burke:2001:JXE

[235] Eric M. Burke. Java and XSLT: Embedding XML processing into Java applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2001. ISBN 0-596-00143-6. xvi + 510 pp. LCCN QA76.73.J38 B88 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596001438;   http://
www.oreilly.com/catalog/javaxslt.

Burke:2002:PLF

[236] Sean M. Burke. Perl and LWP: fetching Web pages, parsing HTML, writing spiders and more. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00178-9. xv + 242 pp. LCCN QA76.73.P22 B87 2002. URL http:

//www.oreilly.com/catalog/9780596001780.

Burke:2003:JEP

[237] Eric M. Burke and Brian M. Coyner. Java Extreme Programming cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00387-0. xii + 275 pp. LCCN QA76.73.J38 B873 2003. URL http://www.oreilly.

com/catalog/9780596003876.

Burke:2003:RPG

[238] Sean M. Burke. RTF pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00475-3. vi + 149 pp. LCCN QA76.9.T48 B87 2003.

URL http://www.oreilly.com/catalog/9780596004750.

Burke:2006:EJ

[239] Bill Burke and Richard Monson-Haefel. Enterprise JavaBeans 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-00978-X. xxv + 732 pp. LCCN QA76.73.J38 M65 2006. URL http://

www.loc.gov/catdir/enhancements/fy0715/2006285201-d.html.

Burleson:1999:OSA

[240] Donald K. Burleson. Oracle SAP Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-696-X. xiii + 197 pp. LCCN QA76.9.D3

B8753 1999. US$32.95. URL http://www.oreilly.com/catalog/
9781565926967;   http://www.oreilly.com/catalog/sapadm.

Burleson:2001:UOD

[241] Donald K. Burleson. Unix for Oracle DBAs: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00066-9. v + 104 pp. LCCN QA76.9.D3 B8754 2001. URL http://www.oreilly.com/

catalog/9780596000660.

Burnette:2005:EIP

[242] Ed Burnette. Eclipse IDE: pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10065-5. ix + 117 pp. LCCN QA76.73.J38 B875

2005. URL http://www.oreilly.com/catalog/9780596100650.

Burns:2007:DJG

[243] Brendan Burns. Darkstar: The Java Game Server. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51484-0. LCCN QA76.73.J38 B87 2007eb; QA76.73.J38. URL http://www.oreilly.

com/catalog/9780596514846.

Bustamante:2007:LW

[244] Michele Leroux Bustamante. Learning WCF. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-10162-7, , 0-596-10162-7. xxiv + 582 pp. LCCN QA76.9.D5; QA76.9.D5 B87 2007eb; QA76.9.D5 B87 2007. URL http://www.

oreilly.com/catalog/9780596101626.

Byron:2009:UD

[245] Angela Byron, Addison Berry, Nathan Haug, and Jeff Eaton, editors. Using Drupal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-51580-4 (paperback). xxi + 464 pp. LCCN TK5105.8885.D78 U84 2009.

Calavera:2020:LOB

[246] David Calavera and Lorenzo Fontana. Linux Observability with BPF: Advanced Programming for Performance Analysis and Networking. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2020. ISBN 1-4920-5020-2. xiii + 162 pp.

LCCN QA76.774.L46. URL http://proquest.safaribooksonline.

com/?fpi=9781492050193. Foreword by Jessie Frazelle.

Calcagno:1997:TTI

[247] Ann Calcagno, editor. Travelers’ Tales Italy: true stories of life on the road. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-16-3. xxii + 437 pp. LCCN DG430.2 .I83

1998. US$17.95. URL http://www.ora.com/catalog/italy/;
http://www.oreilly.com/catalog/9781885211163;   http://www.
oreilly.com/catalog/italy.

Calishain:2003:GHI

[248] Tara Calishain and Rael Dornfest. Google hacks: [100 industrial-strength tips and tools]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00447-8. xxi + 329 pp. LCCN TK5105.885.G66 C35 2003. URL http:

//www.oreilly.com/catalog/9780596004477.

Calishain:2003:GPG

[249] Tara Calishain, Rael Dornfest, and D. J. Adams. Google pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00550-4. ix + 129 pp. LCCN ZA4251.G66 C35 2003. URL http://www.oreilly.com/

catalog/9780596005504.

Calishain:2005:GH

[250] Tara Calishain and Rael Dornfest. Google hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00857-0. xxxiv + 443 pp. LCCN

TK5105.885.G66 C35 2005. URL http://www.oreilly.com/catalog/
9780596008574.

Callender:2002:PWS

[251] John (John Bryan) Callender. Perl for Web Site Management. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-647-1. xviii + 508 pp. LCCN QA76.73.P76 C39 2002. US$34.95. URL http://safari.oreilly.com/

1565926471;   http://www.oreilly.com/catalog/9781565926479;
 http://www.oreilly.com/catalog/perlwsmng.

Cameron:1991:LGE

[252] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-84-6. xxvii + 411 pp. LCCN QA76.76.T49 C35 1991. URL http://www.oreilly.com/

catalog/9780937175842.

Cameron:1996:LGE

[253] Debra Cameron, Bill Rosenblatt, and Eric S. Raymond. Learning GNU Emacs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, September 1996. ISBN 1-56592-152-6. xxiv + 533 pp. LCCN QA76.76.T49 C35 1996.

US$29.95. URL http://www.oreilly.com/catalog/9781565921528;
 http://www.oreilly.com/catalog/gnu2.

Cameron:1998:GEP

[254] Debra Cameron. GNU Emacs Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-496-7. iii + 58 pp. LCCN QA76.76.T49

C348 1998. US$6.95. URL http://www.oreilly.com/catalog/gnupr.

Cameron:1999:GEP

[255] Debra Cameron. GNU Emacs Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-496-7. iii + 58 pp. LCCN QA76.76.T49 C34 1999; QA76.76.T49 C348 1998. US$6.95.

URL http://www.oreilly.com/catalog/9781565924963;   http://
www.oreilly.com/catalog/gnupr.

Cameron:2000:GEK

[256] Debra Cameron. GNU Emacs — kurz & gut. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 3-89721-211-0. 60 pp. LCCN ???? German translation of [255].

Cameron:2005:LGE

[257] Debra Cameron, James Elliot, Marc Loy, Eric Raymond, and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00648-9. xxii + 509 pp. LCCN QA76.76.T49 C35 2005. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/;   http://www.oreilly.com/catalog/
9780596006488.

Campbell:2006:DDC

[258] Marc Campbell. Dreamweaver 8: design and construction. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10163-5. xiv + 295 pp. LCCN TK5105.8885.D74eb; TK5105.8885.D74 C36 2006; TK5105.8885.D74.

URL http://www.oreilly.com/catalog/9780596101633.

Campbell:2006:WMC

[259] Tony (George Anthony) Campbell. Windows media center wizardry. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52784-

5. LCCN QA76.76.O63. URL http://www.oreilly.com/catalog/
9780596527846.

Capriolo:2012:PH

[260] Edward Capriolo, Dean Wampler, and Jason Rutherglen. Programming Hive. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1933-5 (paperback). xvii + 328 pp. LCCN ????

Carless:2004:GH

[261] Simon Carless. Gaming hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00714-0. xxiii + 436 pp. LCCN QA76.76.C672 C37 2004; QA76.76.C672 C37 2004eb; QA76.76.C672. URL http://www.oreilly.

com/catalog/9780596007140.

Carlson:2006:RC

[262] Lucas Carlson and Leonard Richardson. Ruby cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52369-6. xxx + 873 pp. LCCN QA76.64

.C37 2006. URL http://www.oreilly.com/catalog/9780596523695.

Carr:2009:ICW

[263] Jeffrey Carr. Inside cyber warfare. Mapping the cyber underworld. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-80215-3. xviii + 212 pp. LCCN ????

Carroll:1999:TTH

[264] Rick Carroll and Marcie Carroll. Travelers’ Tales: Hawaii: True Stories of the Island Spirit. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-35-X. 416 pp. LCCN DU623.Z5H36 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211354;   http://
www.oreilly.com/catalog/hawaii.

Carter:2003:LSA

[265] Gerald Carter. LDAP system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 1-56592-491-6. xii + 294 pp. LCCN TK5105.5725

.C37 2003. URL http://www.oreilly.com/catalog/9781565924918.

Carter:2007:US

[266] Gerald Carter, Jay Ts, and Robert Eckstein. Using Samba. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00769-8. 432 (est.) pp. LCCN

QA76.9.C55 E38 2007. URL http://www.oreilly.com/catalog/
9780596007690;   http://www.oreillynet.com/catalog/samba3/.

Catsoulis:2002:DEH

[267] John Catsoulis. Designing Embedded Hardware. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00362-5. xviii + 298 pp. LCCN TK7895.E42

C38 2003. US$39.95. URL http://www.oreilly.com/catalog/
9780596003623;   http://www.oreilly.com/catalog/dbhardware.

Catsoulis:2005:DEH

[268] John Catsoulis. Designing embedded hardware. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00755-8. xvi + 377 pp. LCCN TK7895.E42 C37 2005. URL http://www.oreilly.com/

catalog/9780596007553.

Cavaness:2002:PJS

[269] Chuck Cavaness. Programming Jakarta Struts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00328-5. xvii + 441 pp. LCCN QA76.76.A65

C39 2003. US$39.95. URL http://www.oreilly.com/catalog/
9780596003289;   http://www.oreilly.com/catalog/jakarta.

Cavaness:2003:JSP

[270] Chuck Cavaness and Brian Keeton. Jakarta Struts pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00519-9. viii + 134 pp. LCCN QA76.76.A65 C385 2003. URL http://www.oreilly.

com/catalog/9780596005191.

Cavaness:2004:PJS

[271] Chuck Cavaness. Programming Jakarta Struts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00651-9. xvii + 450 pp. LCCN QA76.76.A65 C39 2004. US$44.95, CDN$65.95,

UK£31.95. URL http://www.oreilly.com/catalog/0596006519/;
http://www.oreilly.com/catalog/9780596006518.

Cawood:2005:HHT

[272] Stephen Cawood. Halo 2 Hacks: Tips & Tools for Finishing the Fight. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10059-0. xxiii + 243 pp. LCCN GV1469.35.H35eb. US$24.95, CDN$34.95, UK£17.50.

URL http://www.oreilly.com/catalog/9780596100599;   http://
www.oreilly.com/catalog/halo2hks/index.html.

Cerami:2002:WSE

[273] Ethan Cerami. Web Services Essentials: Distributed Applications with XML-RPC, SOAP, UDDI & WSDL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00224-6. xiii + 288 pp. LCCN TK5105.88813 .C47

2002. US$29.95. URL http://safari.oreilly.com/0596002246;
http://www.oreilly.com/catalog/9780596002244;   http://www.
oreilly.com/catalog/webservess.

Chamberlain:2003:PTT

[274] Darren Chamberlain, David Cross, and Andy Wardley. Perl template toolkit. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00476-1. xiii + 574 pp. LCCN QA76.73.P22 C35 2004. URL http://www.oreilly.

com/catalog/9780596004767.

Chambers:2007:AAF

[275] Mike Chambers, Rob Dixon, and Jeff Swartz. Apollo for Adobe Flex developers pcket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51391-7 (paperback). xvi + 121 pp. LCCN QA76.625; QA76.625eb. URL http://www.

oreilly.com/catalog/9780596513917.

Chambers:2007:AIR

[276] Mike Chambers, Daniel Dura, and Kevin Hoyt. Adobe Integrated Runtime (Air): for Javascript Developers Pocket Guide. Adobe developer library. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51519-7. xix + 154 pp. LCCN QA76.625; QA76.625 .C42 2007eb. URL http://www.

oreilly.com/catalog/9780596515195.

Chang:2018:RGC

[277] Winston Chang. R Graphics Cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2018. ISBN 1-4919-7860-0 (paperback), 1-4919-7855-4 (e-book), 1-4919-7857-0 (e-book), 1-4919-7859-7. xiii + 425 pp. LCCN QA276.45.R3 C46 2018.

Chapman:1995:BIF

[278] D. Brent Chapman and Elizabeth D. Zwicky. Building Internet Firewalls. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-124-0. xxvi + 517 pp. LCCN TK5105.59 .C49 1995; TK5105.875

.I57 C43 1995. US$32.95. URL http://www.oreilly.com/catalog/
9781565921245;   http://www.oreilly.com/catalog/fire. Minor

corrections, Nov. 1995.

Chappell:1996:WVG

[279] Geoff Chappell and Ralph Shnelvar. Windows 95 VxDs: a guide through the mysteries of the Virtual Device Driver System at: the Heart of Windows 95 and Windows 3.x. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1996. ISBN 1-56592-195-X. 350 pp. LCCN ???? US$29.95. URL

http://www.oreilly.com/catalog/9781565921955.

Chappell:2002:JWS

[280] David A. Chappell and Tyler Jewell. Java Web services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00269-6. xii + 262 pp. LCCN

QA76.73.J38 C4772 2002. URL http://www.oreilly.com/catalog/

9780596002695. Using Java in service-oriented architectures.

Chappell:2004:ESB

[281] David A. Chappell. Enterprise service bus. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00675-6 (Standard printing), 0-596-00814-7 (Sonic Software printing). xxiii + 247 pp. LCCN QA76.9.A73 C384 2004. URL http://www.

oreilly.com/catalog/9780596006754.

Chase:2006:ASM

[282] Kate Chase and Scott D. Palmer. Access 2003 for starters: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00665-9. xiv + 379 + 1 pp. LCCN HF5548.4.M525 C43 2006eb; HF5548.4.M525. URL

http://www.oreilly.com/catalog/9780596006655.

Cheah:2005:FS

[283] Chu Yeow Cheah. Firefox secrets. A need-to-know guide. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9752402-4-2. xvi + 297 pp. LCCN TK5105.883.F55 C44 2005. URL http://www.oreilly.com/

catalog/9780975240243.

Cheshire:2006:ZCN

[284] Stuart Cheshire and Daniel H. (Daniel Howard) Steinberg. Zero configuration networking: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10100-7 (paperback). xxiii + 226 pp. LCCN QA76.76.C69 C48 2006. URL http://www.oreilly.com/

catalog/9780596101008.

Childs:2000:VND

[285] Matt Childs, Paul Lomax, and Ronald Petruska. VBScript in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-720-6. xiii + 495 pp. LCCN QA76.73.V27 C48 2000.

US$29.95. URL http://www.oreilly.com/catalog/9781565927209;
 http://www.oreilly.com/catalog/vbscriptian.

Childs:2001:VPR

[286] Matt Childs, Paul Lomax, and Ronald Petrusha. VBScript Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2001. ISBN 0-596-00126-6. xii + 106 pp. LCCN QA76.73.V27 C44 2001; QA76.73.V27 C483 2001.

US$12.95. URL http://www.oreilly.com/catalog/9780596001261;
 http://www.oreilly.com/catalog/vbscriptpr.

Chodorow:2010:MDG

[287] Kristina Chodorow and Michael Dirolf. MongoDB: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-8156-1 (paperback), 1-4493-8157-X. xvii + 193 pp. LCCN HF5548.38.O64 .C462 2010.

Chodorow:2011:SMS

[288] Kristina Chodorow. Scaling MongoDB: [sharding, cluster setup, and administration]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-0321-8 (paperback), 1-4493-0412-5 (e-book), 1-4493-0387-0 (e-book). ix + 49 pp. LCCN QA76.9.D3 C463 2011.

Christiansen:1998:PC

[289] Tom Christiansen and Nathan Torkington. Perl Cookbook: Solutions and examples for Perl programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-243-3. xxxiv + 757 pp. LCCN QA76.73.P22 C38 1998.

US$39.95. URL http://www.oreilly.com/catalog/9781565922433;
 http://www.oreilly.com/catalog/cookbook/.

Christiansen:2003:PC

[290] Tom Christiansen and Nathan Torkington. Perl cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00313-7. xxxiv + 927 pp. LCCN QA76.73.P22 C38 2003. URL http://www.oreilly.

com/catalog/9780596003135.

Chromatic:2002:RWS

[291] Chromatic, Brian Aker, and Dave Krieger. Running Weblogs with Slash. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00100-2. xviii + 262 pp. LCCN TK5105.8883 .C57

2002. US$34.95. URL http://safari.oreilly.com/0596001002;

http://www.oreilly.com/catalog/9780596001001;   http://www.
oreilly.com/catalog/runblogslash.

Chromatic:2003:EPP

[292] Chromatic. Extreme Programming pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00485-0. xv + 90 pp. LCCN QA76.76.D47

C56 2003. URL http://www.oreilly.com/catalog/9780596004859.

Chromatic:2006:PH

[293] Chromatic, Damian Conway, and Curtis Poe. Perl hacks. Hacks series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52674-1. xx + 273 pp. LCCN QA76.73.P22 C382 2006; QA76.73.P22

C47 2006eb; QA76.73.P22. URL http://www.oreilly.com/catalog/
9780596526740.

Clark:2004:PPA

[294] Mike Clark. Pragmatic project automation: how to build, deploy, and monitor Java applications, volume 3 of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-3-9. xiv + 161 pp. LCCN QA76.76.D47 C53 2004. URL http://www.oreilly.

com/catalog/9780974514031.

Clark:2009:BIA

[295] Josh Clark. Best iPhone Apps: the guide for discriminating downloaders. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-80617-5. x + 228 pp. LCCN TK6570.M6 C53 2009.

Coar:2004:AC

[296] Ken A. L. Coar and Richard Cooper Bowen. Apache cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00191-6. xviii + 234 pp. LCCN TK5105.8885.A63 C6297 2004. URL http://www.oreilly.com/

catalog/9780596001919.

Coar:2007:AC

[297] Ken A. L. Coar and Richard Cooper Bowen. Apache cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52994-5 (paperback). xviii + 285 pp. LCCN Internet.

Coffey:2005:FPM

[298] Geoff Coffey and Susan Prosser. FileMaker Pro 8: the missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00579-2. xxiv + 736 pp. LCCN QA76.9.D3; QA76.9.D3 C64 2005; QA76.9.D3 C64 2005eb. URL http://www.oreilly.com/

catalog/9780596005795.

Coffey:2007:FPM

[299] Geoff Coffey and Susan Prosser. Filemaker Pro 9: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51413-1. xx + 778 pp. LCCN QA76.9.D3 C6774 2007eb; QA76.9.D3; QA76.9.D3 C6774 2007. URL http://www.oreilly.com/

catalog/9780596514136.

Cogswell:2005:CC

[300] Jeffrey M. Cogswell et al., editors. C++ cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00761-2. 504 pp. LCCN QA76.73.C153 C12

2005. URL http://www.oreilly.com/catalog/9780596007614.

Cohen:1998:WMP

[301] Aaron Cohen and Mike Woodring. Win32 Multithreaded Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-296-4. xv + 705 pp. LCCN QA76.76.O63 C633 1998. US$39.95. URL http://www.ora.

com/catalog/multithread/;   http://www.oreilly.com/catalog/
9781565922969;   http://www.oreilly.com/catalog/multithread.

Cohen:2008:TCS

[302] Michael E. Cohen. Take control of syncing in Leopard. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-51-3. ???? pp. LCCN QA76.76.O63 C64 2008.

Cole:1996:PIY

[303] Gregory Cole, editor. Passport Indonesia: Your Pocket Guide to Indonesian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-37-2. 96 pp. LCCN HD58.7.C622 1997. US$6.95.

URL http://www.oreilly.com/catalog/9781885073372.

Cole:2005:UMT

[304] Jason R. Cole. Using Moodle: teaching with the popular open source course management system. O’Reilly Community Press, Sebastopol, CA, USA, 2005. ISBN 0-596-00863-5. xv + 219 pp. LCCN LB1028.68 .C65

2005. URL http://www.oreilly.com/catalog/9780596008635.

Cole:2008:UM

[305] Jason R. Cole and Helen Foster. Using Moodle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-52918-X (paperback). xiii + 266 pp. LCCN LB1028.68 .C65 2007. URL http://www.oreilly.com/

catalog/9780596529185.

Collins-Sussman:2004:VCS

[306] Ben Collins-Sussman, Brian W. Fitzpatrick, and C. Michael Pilato. Version control with Subversion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00448-6. xvii + 299 pp. LCCN QA76.6

.C6274 2004. URL http://svnbook.red-bean.com/;   http://
www.onlamp.com/pub/a/onlamp/2004/08/19/subversiontips.html;
 http://www.oreilly.com/catalog/9780596004484.

Collins-Sussman:2008:VCS

[307] Ben Collins-Sussman, Brian W. Fitzpatrick, and C. Michael Pilato. Version control with Subversion. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51033-0. xxiv + 404 pp. LCCN QA76.6 .C6274 2008.

Conner-Sax:1999:WIN

[308] Kiersten Conner-Sax and Ed Krol. The whole Internet: the next generation: a completely new edition of the first and best user’s guide to the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-428-2. xiv + 542 pp. LCCN TK5105.875.I57 C658 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565924284;   http://
www.oreilly.com/catalog/twi3.

Conner:1997:OGJ

[309] Nauman Conner. Overnight Guide to Java, CGI, Animation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-283-2. ???? pp. LCCN ????

Conner:2005:EMM

[310] Nancy Conner. eBay: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00644-6 (paperback). xxii + 439

pp. LCCN HF5476 MUE. URL http://www.oreilly.com/catalog/
9780596006440;   http://www.oreilly.com/catalog/ebaytmm/.

Conner:2006:QMM

[311] Nancy Conner. Quickbase: the missing manual. Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52960-0. xvii + 398 pp. LCCN QA76.625 .Q85 2007. URL http://www.oreilly.com/

catalog/9780596529604.

Connolly:1997:XPT

[312] Dan Connolly, editor. XML: Principles, Tools, and Techniques, volume 2(4) of The World Wide Web Journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1997. ISBN 1-56592-349-9. ISSN 1085-2301. ix + 248 pp. LCCN QA76.76.H94 C66 1997. US$29.95.

URL http://www.amazon.com/exec/obidos/ISBN%3D1-56592-349-
9/8788-6902305-556196; http://www.ora.com/catalog/wjfall97/
;   http://www.oreilly.com/catalog/9781565923492;   http://
www.oreilly.com/catalog/wjfall97.

Conti:2007:SDV

[313] Greg Conti. Security data visualization: graphical techniques for network analysis. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-143-3. xx + 244 pp. LCCN TK5105.59 .C5792 2007. URL http://

www.loc.gov/catdir/enhancements/fy0728/2007019524-d.html;
http://www.loc.gov/catdir/toc/ecip0717/2007019524.html;   http:
//www.oreilly.com/catalog/9781593271435.

Conway:2005:PBP

[314] Damian Conway. Perl best practices. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00173-8. xxii + 517 pp. LCCN QA76.73.P22 C65

2005. URL http://www.oreilly.com/catalog/9780596001735.

Conway:2012:MLH

[315] Drew Conway and John Myles White. Machine learning for hackers. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2012. ISBN 1-4493-0371-4. xiii + 303 pp. LCCN QA76.9.A43 C674 2012. URL http://www.loc.

gov/catdir/enhancements/fy1307/2012277057-b.html;   http://
www.loc.gov/catdir/enhancements/fy1307/2012277057-d.html;
http://www.loc.gov/catdir/enhancements/fy1307/2012277057-t.
html.

Cooper:2003:BWH

[316] Joe (R. Joe) Cooper. The book of Webmin, or, How I learned to stop worrying and love Unix. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-92-1. xxviii + 281 pp. LCCN QA76.9.U83 C6597 2003eb.

URL http://www.oreilly.com/catalog/9781886411920.

Copeland:2008:ES

[317] Rick Copeland. Essential SQLAlchemy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51614-2. xi + 215 pp. LCCN QA76.9.W43 C67 2008.

Corbet:2005:LDD

[318] Jonathan Corbet, Alessandro Rubini, and Greg Kroah-Hartman. Linux device drivers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00590-3. xviii + 615 pp. LCCN QA76.76.D49 R92 2005; QA76.76.D49 R92 2005eb; QA76.76.D49. URL http://www.oreilly.

com/catalog/9780596005900.

Corrigan:1993:OPT

[319] Peter Corrigan and Mark Gurry. ORACLE Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-048-1. xxxv + 603 pp. LCCN QA76.9.D3 C674 1993. US$34.95. URL

http://www.oreilly.com/catalog/9781565920484.

Costales:1993:S

[320] Bryan Costales, Eric Allman, and Neil Rickert. sendmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1993. ISBN 1-56592-056-2. xxxvi + 792 pp. LCCN HE6239.E54 C67 1993. US$32.95. URL http://

www.oreilly.com/catalog/9781565920569. From Samuel Ko: Highly

recommended. This is really an encyclopedia on sendmail. The authors did a great job in de-mysterifying this complex email routing and delivering program . . .. The book is divided into four parts. Part One is a tutorial on understanding sendmail from the ground up . . .. Part Two deals with practical topics in sendmail administration . . .. Part Three is a comprehensive reference section . . .. Part Four consists of appendices with details . . ..

Costales:1994:S

[321] Bryan Costales. sendmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1994. ISBN 1-56592-056-2. 830 pp. US$32.95. While not strictly an Internet book this tome focuses on one thing: the UNIX program sendmail, which is a huge part of how electronic mail moves around on the Internet. Mainly for system administrators, the book shows how to use every function, mode and mood of sendmail to get your e-mail where it’s going. A great, if single-minded, book.

Costales:1997:S

[322] Bryan Costales and Eric Allman. sendmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-222-0. xxv + 1021 pp. LCCN HE6239.E54 C67 1997. US$39.95. URL http://www.ora.com/catalog/

sendmail2/;   http://www.oreilly.com/catalog/9781565922228;
 http://www.oreilly.com/catalog/sendmail2.

Costales:1997:SDR

[323] Bryan Costales. sendmail Desktop Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-278-6. viii + 68 pp. LCCN

???? US$6.95. URL http://www.ora.com/catalog/sendmailqr/;
http://www.oreilly.com/catalog/9781565922785;   http://www.
oreilly.com/catalog/sendmailqr.

Costales:2003:S

[324] Bryan Costales and Eric Allman. Sendmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 1-56592-839-3. xxiv + 1205 pp. LCCN

TK5105.74.S44+. US$59.95. URL http://www.oreilly.com/catalog/
9781565928398;   http://www.oreilly.com/catalog/sendmail3.

Costales:2004:SC

[325] Bryan Costales, Gregory Neil Shapiro, Claus Assmann, and George Jansen. Sendmail 8.13 companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00845-7 (paperback). x + 179 pp. LCCN TK5105.74.S44 C67 2004. URL

http://www.oreilly.com/catalog/9780596008451.

Costales:2008:S

[326] Bryan Costales, George Jansen, Claus Assmann, and Gregory Neil Shapiro. Sendmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2008. ISBN 0-596-51029-2 (paperback). xxiv + 1,282 pp. LCCN

TK5105.74.S44 C57 2008. URL http://www.oreilly.com/catalog/
9780596510299.

Cotton:2013:LRS

[327] Richard Cotton. Learning R: [a step-by-step function guide to data analysis]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2013. ISBN 1-4493-5710-5 (paperback), 1-4493-5718-0, 1-4493-5719-9. xix + 377 pp. LCCN QA276.45.R3 C68 2013.

Courtney:2003:VLO

[328] Tim Courtney, Steve Bliss, and Ahui Herrera. Virtual LEGO: the official LDraw.Org guide to LDraw tools for Windows. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-94-8. xx + 449 pp. LCCN TS2301.T7 C69 2003eb. US$39.95. URL http://www.oreilly.com/

catalog/9781886411944.

Cox:2004:MSS

[329] Kerry Cox and Christopher Gerg. Managing security with Snort and IDS tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00661-6 (paperback). xvi + 269 pp.

LCCN TK5105.59 .C65 2004. URL http://www.oreilly.com/catalog/
9780596006617.

Cozens:2005:APP

[330] Simon Cozens. Advanced Perl programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00456-7 (paperback). xiii + 281 pp. LCCN

QA76.73.P22 C69 2005. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://proquest.safaribooksonline.com/0596004567;
 http://www.oreilly.com/catalog/9780596004569.

Cranor:2002:WPP

[331] Lorrie Faith Cranor. Web Privacy with P3P. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00371-4. xx + 321 pp. LCCN TK5105.59

.C72 2002. US$39.95. URL http://www.oreilly.com/catalog/
9780596003715;   http://www.oreilly.com/catalog/webprivp3p.

Cranor:2005:SUD

[332] Lorrie Faith Cranor and Simson Garfinkel. Security and usability: designing secure systems that people can use. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00827-9 (paperback). xviii + 714 pp. LCCN QA76.9.A25 S3533 2005. US$44.95, CDN$62.95, UK£31.95.

URL http://www.oreilly.com/catalog/9780596008277;   http://
www.oreilly.com/catalog/securityusability/.

Crawford:2000:WPM

[333] Sharon Crawford. Windows 2000 Pro: The Missing Manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, November 2000.

ISBN 0-596-00010-3. xii + 443 pp. LCCN QA76.76.O63 C72743 2000.

US$19.95. URL http://www.oreilly.com/catalog/9780596000103;
 http://www.oreilly.com/catalog/win2promm.

Crawford:2003:AID

[334] Chris Crawford. The Art of interactive design: a euphonious and illuminating guide to building successful software. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-84-0. xix + 385 pp. LCCN QA76.76.I59 C73 2003; QA76.76.I59 C73 2003eb. URL http://www.

oreilly.com/catalog/9781886411845.

Crawford:2003:JDP

[335] William Crawford and Jonathan Kaplan. J2EE design patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00427-3. xvi + 350 pp. LCCN QA76.73.J38 C74 2003. URL http://www.oreilly.com/

catalog/9780596004279.

Crayton:2007:CEE

[336] Christopher A. Crayton. CompTIA A+ Essentials 220-601 exam guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51372-0. LCCN QA76.3; QA76.3 .C73 2007eb. URL http://www.oreilly.com/

catalog/9780596513726.

Crayton:2007:MWV

[337] Christopher A. Crayton. Microsoft Windows Vista. Exam 70-620 Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51069-1. LCCN QA76.28 .C73 2007eb; QA76.28. URL http://www.oreilly.

com/catalog/9780596510695.

Crockford:2008:JGP

[338] Douglas Crockford. JavaScript: the good parts. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51774-2. xiii + 153 pp. LCCN QA76.73.J39 C763 2008.

CSRG:1994:CRC

[339] Computer Systems Research Group, UC Berkeley. 4.4BSD-Lite CD-ROM Companion. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1994. ISBN 1-56592-081-3 (domestic edition), 1-56592-092-9 (international edition). 112 pp. LCCN QA76.8.U65F67 .F684 1994. US$40.00. URL http://

www.oreilly.com/catalog/9781565920811.

CSRG:1994:PRM

[340] Computer Systems Research Group, UC Berkeley. 4.4BSD Programmer’s Reference Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1994. ISBN 1-56592-078-3. 886 pp. LCCN QA76.8.U65 F682 1994. US$30.00. URL

http://www.oreilly.com/catalog/9781565920781.

CSRG:1994:PSD

[341] Computer Systems Research Group, UC Berkeley. 4.4BSD Programmer’s Supplementary Documents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1994. ISBN 1-56592-079-1. 596 pp. LCCN QA76.8.U65 F683 1994.

US$30.00. URL http://www.oreilly.com/catalog/9781565920798.

CSRG:1994:SMM

[342] Computer Systems Research Group, UC Berkeley. 4.4BSD System Manager’s Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1994. ISBN 1-56592-080-5. 804 pp. LCCN QA76.8.U65 F68 1994. US$30.00. URL

http://www.oreilly.com/catalog/9781565920804.

CSRG:1994:URM

[343] Computer Systems Research Group, UC Berkeley. 4.4BSD User’s Reference Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1994. ISBN 1-56592-075-9. 905 pp. LCCN QA76.8.U65 F67 1994. US$30.00. URL

http://www.oreilly.com/catalog/9781565920750.

CSRG:1994:USD

[344] Computer Systems Research Group, UC Berkeley. 4.4BSD User’s Supplementary Documents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1994.

ISBN 1-56592-076-7. 712 pp. LCCN QA76.8.U65 F672 1994. US$30.00.

URL http://www.oreilly.com/catalog/9781565920767.

Culbert:2002:NMS

[345] James Culbert. .NET My Services essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00279-3. ???? pp. LCCN QA76.76.M52 C85

2002. URL http://www.oreilly.com/catalog/9780596002794.

Culp:2006:WNW

[346] Brian Culp. What’s New in Windows Vista? O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-51031-4. LCCN QA76.76.O63; QA76.76.O63

C85 2006eb. URL http://www.oreilly.com/catalog/9780596510312.

Culp:2007:WVA

[347] Brian Culp. Windows Vista administration: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52959-7 (paperback). xviii + 780 pp. LCCN QA76.76.O63 C85 2007. URL http://www.oreilly.

com/catalog/9780596529598.

Cumming:2007:SH

[348] Andrew Cumming and Gordon Russell. SQL hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52799-3. xxi + 386 pp. LCCN QA76.73.S67 C85 2007eb; QA76.73.S67. URL http://www.oreilly.

com/catalog/9780596527990.

Cunningham:1996:EPC

[349] Steve Cunningham and Judson Rosebush. Electronic Publishing on CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1996. ISBN 1-56592-209-3. xviii + 382 pp. LCCN Z286.E43 C86 1996; Z480.O68 C86 1996.

US$36.95. URL http://www.oreilly.com/catalog/9781565922099.

Includes CD-ROM.

Curioso:2007:AP

[350] Andrew G. Curioso. Ajax with PHP 5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51403-4. LCCN TK5105.8885.A52 C87

2007eb; TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596514037.

Curry:1989:UCU

[351] Dave Curry. Using C on the UNIX System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1989. ISBN 0-937175-23-4. 250 pp. LCCN QA76.73.C15 C87 1989. US$24.95. URL http://www.oreilly.com/

catalog/9780937175231.

Curry:1990:UCU

[352] David Curry. Using C on the UNIX System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-23-4. xii + 379 pp. LCCN QA76.76.O63 H68 1992. US$24.95. This book is directed to (would-be) system programmers.

Curry:1996:PTY

[353] Jeffrey E. Curry, editor. Passport Taiwan: Your Pocket Guide to Taiwanese Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-27-5. 96 pp. LCCN HD58.7.C847 1997.

US$6.95. URL http://www.oreilly.com/catalog/9781885073273.

Curry:1996:USP

[354] David A. (David Allan) Curry. UNIX Systems Programming for SVR4. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1996. ISBN 1-56592-163-1. xxi + 596 pp. LCCN QA76.76.O63 C865 1996. US$34.95.

URL http://www.oreilly.com/catalog/9781565921634;   http://
www.oreilly.com/catalog/usp.

Curry:1997:PVY

[355] Jeff Curry and ‘Jim’ Chinh T. Nguyen. Passport Vietnam: Your Pocket Guide to Vietnamese Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885073-25-9. 96 pp. LCCN HD58.7.C87 1997.

US$6.95. URL http://www.oreilly.com/catalog/9781885073259.

CustomGuide:2005:OPT

[356] CustomGuide, Inc. Outlook 2003 personal trainer. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00935-6. xiv + 320 pp. LCCN

HF5548.4.M5255 C98 2005. URL http://www.oreilly.com/catalog/
9780596009359.

Cutler:1992:XWS

[357] Ellie Cutler, Daniel Gilly, and Tim O’Reilly, editors. The X Window System in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1992. ISBN 1-56592-017-1. 424 pp. LCCN QA76.76.W56 X2 1992. US$9.95.

Cutler:1994:SUN

[358] Ellie Cutler and The staff of O’Reilly & Associates. SCO UNIX in a nutshell: a desktop quick reference for SCO UNIX and Open Desktop. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1994. ISBN 1-56592-037-6. xix + 568 pp. LCCN QA76.76.O63 C93 1994.

US$24.95. URL http://www.oreilly.com/catalog/9781565920378;
 http://www.oreilly.com/catalog/sco.

Cutroni:2007:GA

[359] Justin Cutroni. Google Analytics. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51496-4. LCCN TK5105.885.G66; TK5105.885.G66 C88 2007eb. URL http://www.

oreilly.com/catalog/9780596514969.

Dalheimer:1999:PQW

[360] Matthias Kalle Dalheimer. Programming with Qt: Writing Portable GUI applications on UNIX and Win32. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-588-2. xix + 361 pp. LCCN QA76.9.U83 D355 1999.

US$32.95. URL http://www.oreilly.com/catalog/9781565925885;
 http://www.oreilly.com/catalog/prowqt/.

Dalheimer:19xx:JVM

[361] Matthias K. Dalheimer. Java Virtual Machine, Sprache, Konzept. . .. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 3-930673-73-8 (??invalid

ISBN??). LCCN ???? 29 DM. URL http://www.oreilly.de/german/
essential/javavm/index.html.

Dalheimer:2002:PQW

[362] Matthias Kalle Dalheimer. Programming with Qt: Writing Portable GUI applications on Unix and Win32. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00064-2. xviii + 499 pp. LCCN QA76.9.U83

D355 2002. US$39.95. URL http://www.oreilly.com/catalog/
9780596000646;   http://www.oreilly.com/catalog/prowqt2.

Dalheimer:2006:RL

[363] Matthias Kalle Dalheimer and Matt Welsh. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-00760-4 (paperback). xviii + 951 pp. LCCN QA76.76.O63 R855 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0715/2007270110-d.html;
 http://www.oreilly.com/catalog/9780596007607.

Dallmeier:1997:JVM

[364] Matthias K. Dallmeier. Java Virtual Machine, Sprache, Konzept, Architektur. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 3-930673-73-8. 209 pp. LCCN ???? 25 DM. URL http://www.ora.de/. Includes CD-ROM.

Daly:2007:NGW

[365] Liza Daly. Next-generation web frameworks in Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51371-2. LCCN QA76.73.P98

L59 2007eb; QA76.73.P98. URL http://www.oreilly.com/catalog/
9780596513719.

Dang:2002:PVU

[366] Luan Dang, Cullen Jennings, and David Kelly. Practical VoIP Using VOCAL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00078-2. xvii + 502 pp. LCCN TK5105.8865 .D35 2002. US$44.95.

URL http://www.oreilly.com/catalog/9780596000783;   http://
www.oreilly.com/catalog/voip.

Danseglio:2005:SWS

[367] Mike Danseglio. Securing Windows Server 2003. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00685-3. xvi + 426 pp. LCCN TK5105.59 .D37 2005 D36 2003; QA76.9.A25; QA76.9.A25 D36 2003; TK5105.59

.D37 2005eb; TK5105.59. URL http://www.oreilly.com/catalog/
9780596006853.

Danseglio:2005:WSS

[368] Mike Danseglio and Robbie Allen. Windows Server 2003 security cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00753-1. xxiii + 494 pp. LCCN QA76.76.O63; QA76.76.O63eb. URL http://www.

oreilly.com/catalog/9780596007539.

Darwin:1988:CCP

[369] Ian F. Darwin. Checking C Programs with lint. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1988. ISBN 0-937175-30-7. x + 72 pp. LCCN QA76.73.C15 D37 1990. US$12.95.

URL http://www.oreilly.com/catalog/9780937175309;   http://
www.oreilly.com/catalog/lint.

Darwin:2001:JC

[370] Ian F. Darwin. Java cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00170-3. 850 pp. LCCN QA76.73.J38 D348 2001.

Darwin:2001:JCS

[371] Ian Darwin. Java Cookbook: Solutions and Examples for Java Developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 0-596-00170-3 (paperback). xxix + 850 pp. LCCN QA76.73.J38 D348 2001. US$44.95.

URL http://www.oreilly.com/catalog/9780596001704;   http://
www.oreilly.com/catalog/javacook.

Darwin:2003:JCS

[372] Ian Darwin. Java Cookbook: Solutions and Examples for Java Developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00701-9. xxix + 829 pp. LCCN QA76.73.J38 D348 2004. US$49.95, CDN$72.95, UK£35.50. URL http:/

/examples.oreilly.com/javacook2/;   http://www.oreilly.com/
catalog/0596007019/;   http://www.oreilly.com/catalog/9780596007010.

Darwin:2004:JC

[373] Ian F. Darwin. Java cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00701-9. xxiv + 829 pp. LCCN QA76.73.J38 D348 2004. URL

http://www.oreilly.com/catalog/9780596007010.

Darwin:2007:CJP

[374] Ian F. Darwin. Checking Java programs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51023-3. 54 pp. LCCN QA76.73.J38eb; QA76.73.J38. URL http:

//www.oreilly.com/catalog/9780596510237.

Date:2005:DDR

[375] C. J. Date. Database in depth: relational theory for practitioners. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10012-4. xviii + 208 pp. LCCN QA76.9.D3 D384 2005. URL http://www.oreilly.com/

catalog/9780596100124.

Davidson:2002:LCO

[376] James Duncan Davidson. Learning Cocoa with Objective-C. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00301-3. xxii + 358 pp. LCCN QA76.64 .D39

2002. US$34.95. URL http://safari.oreilly.com/0596003013;
http://www.oreilly.com/catalog/9780596003012;   http://www.
oreilly.com/catalog/learncocoa2.

Davidson:2006:RMX

[377] James Duncan Davidson. Running Mac OS X Panther. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00913-5 (paperback). 372 pp. LCCN QA76.76.O63 D3498 2006.

URL http://www.oreilly.com/catalog/9780596009137;   http://
www.oreillynet.com/catalog/runmacx2/.

Davis:2006:GAT

[378] Harold Davis. Google advertising tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10108-2. 366 pp. LCCN HF6146.I58; HF6146.I58 D38 2006eb; HF6146.I58 D38 2006. URL http://www.oreilly.com/

catalog/9780596101084.

Davis:2007:GWD

[379] Scott Davis. GIS for web developers: adding where to your Web applications. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 0-9745140-9-8 (paperback), (pbk.,, ). xiii + 254 pp.

LCCN G70.212 .D38 2007. URL http://www.oreilly.com/catalog/
9780974514093.

Davis:2007:LPM

[380] Michele E. Davis and Jon A. (John Andrew) Phillips. Learning PHP and MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51401-8 (paperback). xiii + 411 pp. LCCN QA76.73.P224 D38 2007.

URL http://www.loc.gov/catdir/toc/fy0803/2008271806.html;
http://www.oreilly.com/catalog/9780596514013.

Davis:2008:LED

[381] Harold Davis. Lighting and exposure for digital photographers. Practical artistry. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52988-0. 176 pp. LCCN TR590 .D38 2008.

Davison:2005:KGP

[382] Andrew Davison. Killer game programming in Java. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00730-2. xxiv + 969 pp. LCCN QA76.76.C672 D38

22005. URL http://www.oreilly.com/catalog/9780596007300.

Dean:2001:LLC

[383] Jeffrey Dean. LPI Linux certification in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 1-56592-748-6. xvi + 551 pp. LCCN QA76.76.O63

D435 2001. US$39.95. URL http://www.oreilly.com/catalog/
9781565927483;   http://www.oreilly.com/catalog/lpicertnut.

Degenhart:1998:AND

[384] Curt Degenhart and Jen Muehlbauer. AOL in a Nutshell: a desktop guide to America Online. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1998. ISBN 1-56592-424-X. xxiii + 510 pp. LCCN QA76.57.A43 D44 1998.

US$24.95. URL http://www.oreilly.com/catalog/9781565924246;
 http://www.oreilly.com/catalog/aolnut.

Deitsch:2001:JI

[385] Andrew Deitsch and David Czarnecki. Java Internationalization. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00019-7. xvi + 444 pp. LCCN QA76.73.J38 D454 2001.

URL http://www.oreilly.com/catalog/9780596000196;   http://
www.oreilly.com/catalog/javaint/.

Denny:1996:CPW

[386] Bob Denny and Linda Mui. CGI Programming for Windows. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-215-8. 250 pp. LCCN ????

US$29.95. URL http://www.oreilly.com/catalog/9781565922150.

Includes diskette.

Dent:2004:PDG

[387] Kyle D. Dent. Postfix: the definitive guide: a secure and easy-to-use MTA for Unix. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00212-2. xv + 260 pp. LCCN TK5105.74.P66 D46 2004.

URL http://www.oreilly.com/catalog/9780596002121;   http://
www.oreilly.com/catalog/postfix/.

DES:1998:WPD

[388] Digital Education Systems and O’Reilly and Associates. Web Publishing and Design (Web Course). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 5, 1998. ISBN ???? ???? pp. LCCN ???? US$249. URL http://www.

oreilly.com/catalog/webpubwc.

DES:1999:LPC

[389] Digital Education Systems. Learning Perl and CGI (Web Course). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN ???? ???? pp. LCCN

???? US$249. URL http://www.oreilly.com/catalog/lperlwc.

Descartes:1998:JNM

[390] Descartes. Java Native Methods. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-56592-345-6. LCCN ???? ????

Descartes:2000:PPD

[391] Alligator Descartes and Tim Bunce. Programming the Perl DBI. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-699-4. xvi + 346 pp. LCCN QA76.73.P22 D47 2000. US$34.95. URL http://www.oreilly.com/

catalog/9781565926998.

Desmond:2009:AD

[392] Brian Desmond et al., editors. Active directory. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2009. ISBN 0-596-15517-4. 864 (est.) pp. LCCN TK5105.7 A28 2008.

DeVries:2008:RPD

[393] Derek DeVries and Mike Naberezny. Rails for PHP developers. The pragmatic programmers; The facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, 2008. ISBN 1-934356-04-2 (paperback). xx + 406 pp. LCCN QA76.73.P224 D483 2008. URL http://www.oreilly.com/

catalog/.

Dhanjani:2005:NST

[394] Nitesh Dhanjani and Justin Clarke. Network security tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00794-9 (paperback). xiii + 324 pp. LCCN

QA76.9.A25 D48 2005. URL http://www.oreilly.com/catalog/
9780596007942.

Dhanjani:2009:HNG

[395] Nitesh Dhanjani, Billy Rios, and Brett Hardin. Hacking: the next generation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-80630-2. xiii + 279 pp. LCCN QA76.9.A25 D436 2009.

DiBona:1999:OSV

[396] Chris DiBona, Sam Ockman, and Mark Stone, editors. Open Sources: Voices from the Open Source Revolution. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-582-3. viii + 272 pp. LCCN QA76.754 .O63 1999; QA76.76.S46 O64 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565925823;   http://
www.oreilly.com/catalog/opensources/. Includes a chapter The

GNU Operating System and the Free Software Movement, by Richard Stallman, and a chapter Future of Cygnus Solutions: An Entrepreneur’s Account, by Michael Tiemann.

DiBona:2005:OSC

[397] Chris DiBona, Danese Cooper, and Mark Stone, editors. Open Sources 2.0: the continuing evolution. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00802-3. xl + 445 pp. LCCN QA76.76.S46 O643 2005. URL

http://www.oreilly.com/catalog/9780596008024.

Dictor:1999:VBC

[398] Evan S. Dictor. Visual Basic Controls in a Nutshell: the controls of the professional and enterprise editions. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-294-8. xii + 748 pp. LCCN QA76.76.D47 D52 1999.

US$24.95. URL http://www.ora.com/catalog/visualnut/;   http:
//www.oreilly.com/catalog/9781565922945.

Dietze:2011:PUS

[399] Martin Dietze. Praxiskurs Unix-Shell: [werden Sie zum Kommandozeilenvirtuosen, behandelt Bourne-Shell, Korn-Shell, bash and TC-Shell, mit Aufgaben und Lösungen]. O’Reilly basics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 3-89721-565-9. xxiii + 282 pp. LCCN ???? EUR 19.90.

Dixon:2004:AED

[400] Douglas Dixon. Adobe Encore DVD: in the studio. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00600-4. xi + 321 pp. LCCN TK7882.D93 D578 2004. URL http://www.oreilly.com/

catalog/9780596006006.

Doar:2005:PDE

[401] Matthew Doar. Practical development environments. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00796-5 (paperback). 352 (est.) pp. LCCN QA76.76.D47 D63 2005.

URL http://uclibs.org/PID/83501;   http://www.oreilly.com/

catalog/9780596007966;   http://www.oreilly.com/catalog/practicalde/

.

Dobbs-Sciortino:2006:M

[402] Angel Dobbs-Sciortino. Mongrel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52854-X. LCCN TK5105.8883. URL http://www.

oreilly.com/catalog/9780596528546.

Doctorow:2002:EBS

[403] Cory Doctorow, Rael Dornfest, J. Scott Johnson, Shelley Powers, Benjamin Trott, and Mena G. Trott, editors. Essential Blogging: Selecting and Using Weblog Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00388-9. xiii + 244 pp. LCCN TK5105.8884 .E88

2002. US$29.95. URL http://safari.oreilly.com/0596003889;
http://www.oreilly.com/catalog/9780596003883;   http://www.
oreilly.com/catalog/essblogging.

Donahue:2007:NW

[404] Gary A. Donahue. Network warrior. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10151-1 (paperback). xx + 576 pp. LCCN TK5105.5eb;

TK5105.5. URL http://www.oreilly.com/catalog/9780596101510.

Dondich:2006:SCN

[405] Taylor Dondich. Short cuts network monitoring with Nagios. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52819-1. LCCN QA76.9.E94.

URL http://www.oreilly.com/catalog/9780596528195.

Dooley:2002:DLS

[406] Kevin Dooley. Designing Large-Scale LANs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00150-9. xi + 385 pp. LCCN TK5105.7 .D66

2002. US$39.95. URL http://safari.oreilly.com/0596001509;
http://www.oreilly.com/catalog/9780596001506;   http://www.
oreilly.com/catalog/lgscalelans.

Dooley:2003:CC

[407] Kevin Dooley and Ian J. Brown. Cisco cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00367-6. xix + 885 pp. LCCN TK5105.543

.D66 2003. URL http://www.oreilly.com/catalog/9780596003678.

Dooley:2007:CIC

[408] Kevin Dooley and Ian J. Brown. Cisco IOS cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second (revised and updated) edition, 2007. ISBN 0-596-52722-5. 1207 pp. LCCN TK5105.543; TK5105.543 .D66 2007;

TK5105.543 .D66 2007eb. URL http://www.oreilly.com/catalog/
9780596527228.

Dornfest:2003:MXH

[409] Rael Dornfest and Kevin Hemenway. Mac OS X hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00460-5. xxii + 406 pp. LCCN

QA76.76.O63 D675 2003. URL http://www.oreilly.com/catalog/
9780596004606.

Dornfest:2004:MXP

[410] Rael Dornfest and James Duncan Davidson. Mac OS X Panther hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00718-3. xxii + 566 pp. LCCN

QA76.76.O63 D6753 2004. URL http://www.oreilly.com/catalog/
9780596007188.

Dornfest:2006:GHT

[411] Rael Dornfest, Paul Bausch, and Tara Calishain. Google hacks: tips and tools for finding and using the world’s information. Hacks series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2006. ISBN 0-596-52706-3. xxxii + 510 pp. LCCN TK5105.885.G66; TK5105.885.G66 D67 2006eb. URL http://

www.oreilly.com/catalog/9780596527068.

Dougherty:1990:SA

[412] Dale Dougherty. sed and awk. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1990. ISBN 0-937175-59-5. 414 pp. LCCN QA76.76.U84 D69 1991.

US$27.95. URL http://www.oreilly.com/catalog/9780937175590.

Dougherty:1991:SA

[413] Dale Dougherty. sed & awk. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-59-5. xxii + 394 pp. LCCN QA76.76.U84 D69 1991. URL

http://www.oreilly.com/catalog/9780937175590.

Dougherty:1994:MHMa

[414] Dale Dougherty and Richard Koman. The Mosaic Handbook for Microsoft Windows. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1994. ISBN 1-56592-094-5. xxiii + 204 pp. LCCN TK5105.875.I57 D686 1994; TK5105.883.M67 D68 1994. US$29.95. URL http://www.

oreilly.com/catalog/9781565920941.

Dougherty:1994:MHMb

[415] Dale Dougherty and Richard Koman. The Mosaic Handbook for the Macintosh. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1994. ISBN 1-56592-096-1. xxiii + 171 pp. LCCN TK5105.875.I57 D68 1994; TK5105.883.M67 D682 1994. US$29.95. URL http://www.

oreilly.com/catalog/9781565920965.

Dougherty:1994:MHX

[416] Dale Dougherty, Richard Koman, and Paula M. Ferguson. The Mosaic Handbook for the X Window System. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1994. ISBN 1-56592-095-3. xxiii + 262 pp. LCCN TK5105.875.I57 D687 1994; TK5105.883.M67 D684 1994.

US$30.00. URL http://www.oreilly.com/catalog/9781565920958.

Dougherty:1997:SA

[417] Dale Dougherty and Arnold Robbins. sed and awk. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, February 1997. ISBN 1-56592-225-5. xxii + 407 pp. LCCN QA76.76.U84 D69 1997. US$29.95.

URL http://www.oreilly.com/catalog/9781565922259;   http://
www.oreilly.com/catalog/sed2.

Dougherty:2001:PNO

[418] Dale Dougherty and Richard Koman, editors. 2001 P2P networking overview: the emergent P2P platform of presence, identity, and edge resources. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00185-1. iii + 301 pp. LCCN TK5105.525 .A162 2001. URL http://www.oreilly.

com/catalog/9780596001858.

Dowd:1993:HPC

[419] Kevin Dowd. High Performance Computing. RISC architectures, optimization and benchmarks; A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1993. ISBN 1-56592-032-5. xxv + 371 pp. LCCN QA76.88 .D6 1993; QA76.9.A73 D68 1993. US$25.95. URL http://www.oreilly.

com/catalog/9781565920323.

Dowd:1996:GCI

[420] Kevin Dowd. Getting Connected: The Internet at 56K and Up. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-56592-154-2 (pig cover), 1-56592-203-4 (lion cover). xii + 410 pp. LCCN TK5105.875.I57 D69 1996. US$29.95, CDN$42.95.

URL http://www.oreilly.com/catalog/9781565921542;   http://
www.oreilly.com/catalog/getconnect.

Dowd:1998:HPC

[421] Kevin Dowd and Charles R. Severance. High Performance Computing. RISC architectures, optimization and benchmarks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 1998. ISBN 1-56592-312-X. xviii + 446 pp. LCCN QA76.9.A73 D69 1998. US$29.95.

URL http://www.oreilly.com/catalog/9781565923126;   http://
www.oreilly.com/catalog/hpc2.

Downey:2015:TPH

[422] Allen B. Downey. Think Python: How to Think Like a Computer Scientist. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-3936-2. ???? pp. LCCN ????

Downey:2019:TJ

[423] Allen B. Downey and Ben Lauwens. Think Julia. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4920-4500-4 (e-book), 1-4920-4503-9. LCCN QA76.73.J85 L38 2019.

Doxsey:2016:IGB

[424] Caleb Doxsey. Introducing Go : build reliable, scalable programs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2016. ISBN 1-4783-5582-4, 1-4919-4195-2 (paperback). x + 111 pp. LCCN QA76.73.G63 D69 2016. URL http://proquest.tech.

safaribooksonline.de/9781491941997.

Drayton:2002:CLP

[425] Peter Drayton, Ben Albahari, and Ted Neward. C# Language Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00429-X. ix + 118 pp. LCCN QA76.73.C155 D72 2002. US$12.95.

URL http://www.oreilly.com/catalog/9780596004293;   http://
www.oreilly.com/catalog/csharplangpr.

Drayton:2002:CND

[426] Peter Drayton, Ben Albahari, and Ted Neward. C# in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00181-9. xxii + 832 pp. LCCN QA76.73.C154 D73 2002. US$39.95. URL

http://safari.oreilly.com/0596001819;   http://www.oreilly.
com/catalog/9780596001810;   http://www.oreilly.com/catalog/
csharpnut.

Drayton:2003:CND

[427] Peter Drayton, Ben Albahari, and Ted Neward. C# in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00526-1. xxiii + 897 pp. LCCN QA76.73.C154

D73 2003. URL http://www.oreilly.com/catalog/9780596005269.

Dubinko:2003:XE

[428] Micah Dubinko. XForms essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00369-2. xiv + 215 pp. LCCN TK5105.888 .D84 2003.

URL http://www.oreilly.com/catalog/9780596003692.

DuBois:1993:SPI

[429] Paul DuBois. Software Portability with imake. Practical software engineering; A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1993. ISBN 1-56592-055-4. xxiii + 365 pp. LCCN QA76.76.D47 D83 1993. US$27.95. URL http://www.oreilly.com/

catalog/9781565920552.

DuBois:1995:UCT

[430] Paul DuBois. Using csh and tcsh. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1995. ISBN 1-56592-132-1. xviii + 221 pp. LCCN QA76.76.O63 D82 1995. US$29.95.

URL http://www.oreilly.com/catalog/9781565921320;   http://
www.oreilly.com/catalog/tcsh.

DuBois:1996:SPI

[431] Paul DuBois. Software Portability with imake. Practical software engineering; A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 1996. ISBN 1-56592-226-3. xix + 387 pp. LCCN QA76.76.D47 D83 1996. US$32.95.

URL http://www.oreilly.com/catalog/9781565922266;   http://
www.oreilly.com/catalog/imake2.

DuBois:2003:MC

[432] Paul DuBois. MySQL Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00145-2. xxvii + 992 pp. LCCN QA76.73.S67 D587 2003.

US$49.95. URL http://www.oreilly.com/catalog/9780596001452;
 http://www.oreilly.com/catalog/mysqlckbk.

DuBois:2006:MC

[433] Paul DuBois. MySQL cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52708-X. LCCN ????

DuBois:2007:MC

[434] Paul DuBois. MySQL cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52708-X (paperback). xxvi + 948 pp. LCCN QA76.73.S67 D82 2007. URL http://www.oreilly.com/

catalog/9780596527082.

Dudney:2008:CAM

[435] Bill Dudney. Core animation for Mac OS X and the iPhone: Creating compelling dynamic user interface. Pragmatic Bookshelf, Raleigh, NC, USA, 2008. ISBN 1-934356-10-7. ix + 182 pp. LCCN ????

Dumbill:2004:MDN

[436] Edd Dumbill and Niel M. Bornstein. Mono: a developer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00792-2. xxii + 278 pp. LCCN

QA76.76.M52 D85 2004. URL http://www.oreilly.com/catalog/
9780596007928.

Dumbill:2005:DGL

[437] Edd Dumbill. Debian GNU/Linux: An Explorer‘s Notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00883-X. 256 (est.) pp.

LCCN ???? EUR 24.00. URL http://www.oreilly.com/catalog/
9780596008833.

Dunaway:2002:BVS

[438] Robert B. Dunaway. The book of Visual Studio.Net: aguide for developers. No Starch Press, San Francisco, CA, USA, 2002. ISBN 1-886411-69-7 (paperback). xix + 369 pp. LCCN TK5105.8885.M57 D86 2002eb. URL

http://www.oreilly.com/catalog/9781886411692.

Duncan:2001:OOS

[439] Andy Duncan and Sean Hull. Oracle and Open Source. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00018-9. xvi + 406 pp. LCCN QA76.9.D3 D845 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596000189;   http://
www.oreilly.com/catalog/oracleopen.

Duncan:2002:OCP

[440] Andrew Duncan. Objective-C: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00423-0. v + 122 pp. LCCN QA76.64

.D85 2002. US$12.95. URL http://www.oreilly.com/catalog/
9780596004231;   http://www.oreilly.com/catalog/objectcpr.

Duncan:2002:POD

[441] Andrew Duncan and Jared Still. Perl for Oracle DBAs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00210-6. xvi + 602 pp. LCCN QA76.73.P22 D86 2002. US$44.95. URL http://safari.oreilly.com/

0596002106;   http://www.oreilly.com/catalog/9780596002107;
 http://www.oreilly.com/catalog/oracleperl.

Duntemann:2003:JDD

[442] Jeff Duntemann. Jeff Duntemann’s Drive-by Wi-Fi guide. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2003. ISBN 1-932111-74-3 (paperback). xviii + 446 pp. LCCN TK5105.78 .D86 2003. URL http://www.

oreilly.com/catalog/.

Duntemann:2004:DYE

[443] Jeff Duntemann. Degunking your email, spam, and viruses. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2004. ISBN 1-932111-93-X. xxv + 334 pp. LCCN TK5105.73 .D86 2004. URL http://www.oreilly.com/

catalog/9781932111934.

Duntemann:2004:JDW

[444] Jeff Duntemann. Jeff Duntemann’s wi-fi guide. Paraglyph Press, Inc., Scottsdale, AZ, USA, second edition, 2004. ISBN 1-932111-88-3. xx + 498 pp. LCCN TK5105.78 D86 2004; TK5105.78. URL http://www.

oreilly.com/catalog/9781932111880.

Duthie:2002:ANN

[445] G. Andrew Duthie and Raja Mani. ASP.NET in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00116-9. xv + 785 pp. LCCN TK5105.8885.A26 D88 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596001162;   http://
www.oreilly.com/catalog/aspdotnetnut.

Duthie:2003:ANN

[446] G. Andrew Duthie and Matthew MacDonald. ASP .NET in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00520-2. xvi + 979 pp. LCCN TK5105.8885.A26 D88 2003. URL http:

//www.oreilly.com/catalog/9780596005207.

Dye:1999:ODS

[447] Charles Dye. Oracle Distributed Systems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-432-0. xv + 530 pp. LCCN QA76.73.S674

D94 1999. US$39.95. URL http://www.oreilly.com/catalog/
9781565924321;   http://www.oreilly.com/catalog/ordistsys/.

Dyer:2005:MN

[448] Russel J. T. Dyer. MySQL in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00789-2. xxiv + 321 pp. LCCN QA76.73.S67 D93 2005. URL

http://www.oreilly.com/catalog/9780596007898.

Dyer:2008:MN

[449] Russell J. T. Dyer. MySQL in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51433-6. xvi + 545 pp. LCCN QA76.73.S67 D93 2008.

E:2007:FEE

[450] EffectiveUI Inc . Flex early evaluation: assessing Flex and your project needs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51441-7. LCCN QA76.625; QA76.625 E35 2007eb. URL http://www.oreilly.

com/catalog/9780596514419.

Eckstein:1998:JS

[451] Robert Eckstein, Marc Loy, and Dave Wood. Java Swing. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-455-X. xxviii + 1227 pp. LCCN QA76.73.J38 E27 1998; QA76.73.J38 E35 1998.

US$44.95. URL http://www.oreilly.com/catalog/9781565924550;
 http://www.oreilly.com/catalog/jswing/.

Eckstein:1999:XPR

[452] Robert Eckstein. XML Pocket Reference: Extensible Markup Language. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-709-5. iii + 107 pp. LCCN QA76.76.H94 E25 1999. US$8.95.

URL http://www.oreilly.com/catalog/9781565927094;   http://
www.oreilly.com/catalog/xmlpr.

Eckstein:2000:US

[453] Robert Eckstein, David Collier-Brown, and Peter Kelly. Using Samba. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-449-5. xi + 398 pp. LCCN QA76.9.C55 E267 2000; QA76.9.C55 E267 2000; TK5105.5 .E35 2000. US$34.95. URL http://www.oreilly.com/ catalog/9781565924499. Includes CD-ROM with complete mirror of Samba site with Samba 2.0.5a, including sources, documentation, binaries, and supplementary tools.

Eckstein:2001:SPR

[454] Robert Eckstein, David Collier-Brown, and Peter Kelly. Samba Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00099-5. 78 pp. LCCN QA76.76.O63 E266

2001. US$9.95. URL http://safari.oreilly.com/0596000995;
http://www.oreilly.com/catalog/9780596000998;   http://www.
oreilly.com/catalog/sambapr.

Eckstein:2001:XPR

[455] Robert Eckstein and Michel Casabianca. XML Pocket Reference: Extensible Markup Language. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, April 2001. ISBN 0-596-00133-9. vi + 96 pp. LCCN QA76.76.H94

E25 2001. US$12.95. URL http://www.oreilly.com/catalog/
9780596001339;   http://www.oreilly.com/catalog/xmlpr2. Cov-

ers XSTL.

Eckstein:2002:JEB

[456] Robert Eckstein, editor. Java Enterprise best practices. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00384-6. x + 277 pp. LCCN QA76.73.J38

J3662 2002. URL http://www.oreilly.com/catalog/9780596003845.

Eckstein:2002:JS

[457] Robert Eckstein and Marc Loy. Java Swing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 E27 2002.

Edelson:2009:JC

[458] Justin Edelson and Henry Liu. JRuby cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15506-9. 250 (est.) pp. LCCN QA76.73.R83 E33 2008.

Ediger:2007:AR

[459] Brad Ediger. Advanced Rails. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51032-2 (paperback). xiii + 342 pp. LCCN Internet. URL

http://www.oreilly.com/catalog/9780596510329.

Edwards:2006:JAE

[460] James Edwards and Cameron Adams. The JavaScript anthology: 101 essential tips, tricks and hacks. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2006. ISBN 0-9752402-6-9. xvi + 588 pp. LCCN QA76.73.J39

E392 2006. URL http://www.oreilly.com/catalog/9780975240267.

EFF:1998:CSE

[461] Electronic Frontier Foundation. Cracking DES: Secrets of Encryption Research, Wiretap Politics & Chip Design. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1998. ISBN 1-56592-520-3. 272 pp. LCCN

QA76.9.A25 C783. US$29.95. URL http://www.eff.org/descracker/
;   http://www.oreilly.com/catalog/9781565925205;   http://
www.sunworld.com/swol-07-1998/swol-07-if.html?072098a#2.

Eisenberg:2002:SE

[462] J. David Eisenberg. SVG Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00223-8. xvii + 335 pp. LCCN T385 .E38555 2002; T385 .E47 2002. US$34.95. URL http://safari.oreilly.com/

0596002238;   http://www.oreilly.com/catalog/9780596002237;
 http://www.oreilly.com/catalog/svgess.

Eisenberg:2006:UXM

[463] J. Eisenberg. Using XForms with Mozilla. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-55004-9. LCCN QA76.625; QA76.625 .E34

2006eb. URL http://www.oreilly.com/catalog/9780596550042.

Eismann:2007:CDD

[464] Katrin Eismann and Seán Duggan. The creative digital darkroom. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10047-7. ix + 409 pp. LCCN TR267; TR267 .E377 2007eb. URL http://www.oreilly.

com/catalog/9780596100476.

Elferdink:2000:AMM

[465] Jim Elferdink and David Reynolds. AppleWorks 6: The Missing Manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, May 2000. ISBN 1-56592-858-X. xii + 440 pp. LCCN QA76.76.I57 E45

2000. US$19.95. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.missingmanual.com/cds/;   http://www.
oreilly.com/catalog/9781565928589;   http://www.oreilly.com/
catalog/aplwrksmm.

Elferdink:2005:IMM

[466] Jim Elferdink. iWork ’05: the missing manual. Missing manuals. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-10037-X. xxiii + 379 pp. LCCN QA76.8.M3; QA76.8.M3 E44 2005eb; QA76.8.M3 E44 2005. URL http://www.oreilly.com/

catalog/9780596100377.

Elliott:2004:HDN

[467] James Elliott. Hibernate: a developer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00696-9. xi + 176 pp. LCCN QA76.9.D26 E43 2004. URL

http://www.oreilly.com/catalog/9780596006969.

Elliott:2006:GSH

[468] James Elliott. Getting started with Hibernate 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52818-3. LCCN QA76.73.J38

E45 2006eb; QA76.73.J38. URL http://www.oreilly.com/catalog/
9780596528188.

Elliott:2008:HHS

[469] James Elliott, Ryan Fowler, and Tim OB́rien. Harnessing Hibernate: [a step-by-step guide to Java persistence]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51772-6. xiv + 363 pp. LCCN QA76.76.M54 E46 2008. US$39.99.

Emerick:2012:CP

[470] Chas Emerick, Brian Carper, and Christophe Grand. Clojure programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-9470-1. xviii + 607 pp. LCCN QA76.73.C565 E538 2012. URL http://www.loc.

gov/catdir/enhancements/fy1211/2012405367-b.html;   http://
www.loc.gov/catdir/enhancements/fy1211/2012405367-d.html;
http://www.loc.gov/catdir/enhancements/fy1211/2012405367-t.
html.

Engel:1995:PUY

[471] Dean Engel and Larry K. Peterson. Passport USA: Your Pocket Guide to American Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1995. ISBN 1-885073-15-1. 96 pp. LCCN HD58.7.E4455 1996. US$6.95. URL http://www.oreilly.com/

catalog/9781885073150.

Engel:1996:PJY

[472] Dean Engel, Ken Murakami, and Patrick Bray. Passport Japan: Your Pocket Guide to Japanese Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-885073-17-8. 96 pp. LCCN

HD58.7.E445 1996. US$6.95. URL http://www.oreilly.com/catalog/
9781885073174.

Englander:1997:DJB

[473] Robert Englander. Developing Java Beans. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-289-1. xvi + 298 pp. LCCN QA76.73.J38 E545 1997; QA76.73.J38 E64

1997. US$29.95. URL http://www.ora.com/catalog/javabeans/;
http://www.oreilly.com/catalog/9781565922891;   http://www.
oreilly.com/catalog/javabeans.

Englander:1997:JGD

[474] Robert Englander. JavaBeans Guide du Programmeur. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1997. ISBN 2-84177-038-

9. LCCN ???? 220 FF. URL http://www.editions-oreilly.fr/
international/france/prog/java-beans.html.

Englander:19xx:DJ

[475] Robert Englander. Developing JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-47-8. LCCN ????

3,700 yen. URL http://www.ohmsha.co.jp/data/books/contents/4-

900900-47-8.htm. Japanese translation.

Englander:2002:JS

[476] Robert Englander. Java and SOAP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00175-4. xv + 258 pp. LCCN QA76.76.H94 E54

2002. US$39.95. URL http://safari.oreilly.com/0596001754;
http://www.oreilly.com/catalog/9780596001759;   http://www.
oreilly.com/catalog/javasoap.

Engst:1999:CPM

[477] Adam C. Engst and David Pogue. Crossing Platforms: A Macintosh/ Windows Phrasebook: a Dictionary for Strangers in a Strange Land. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-539-4. v + 321 pp. LCCN QA76.76.O63 E54 1999; QA76.76.O63 E65 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565925397;   http://
www.oreilly.com/catalog/platforms.

Ensor:1997:OD

[478] Dave Ensor and Ian Stevenson. Oracle Design. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-268-9. xix + 526 pp. LCCN QA76.9.D26 E57

1997. US$39.95. URL http://www.ora.com/catalog/oracledes/;
http://www.oreilly.com/catalog/9781565922686;   http://www.
oreilly.com/catalog/oracledes.

Ensor:1997:ODT

[479] Dave Ensor and Ian Stevenson. Oracle8 Design Tips. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-361-8. xiii + 115 pp. LCCN QA76.9.D26 E575 1997. US$14.95. URL

http://www.ora.com/catalog/oracledt/;   http://www.oreilly.
com/catalog/9781565923614;   http://www.oreilly.com/catalog/

oracledt. Text edited by Deborah Russell.

Epstein:1998:LND

[480] Bruce A. Epstein. Lingo in a Nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-493-2. xxii + 609 pp. LCCN QA76.73.L22 E67 1998; QA76.73.L22

E78 1998. US$24.95. URL http://www.oreilly.com/catalog/
9781565924932;   http://www.oreilly.com/catalog/lingonut/.

Epstein:1999:DND

[481] Bruce A. Epstein. Director in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-382-0. xxvii + 628 pp. LCCN QA76.575 .E67 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565923829;   http://
www.oreilly.com/catalog/directnut/.

Erickson:2008:HAE

[482] Jon Erickson. Hacking: the art of exploitation. No Starch Press, San Francisco, CA, USA, second edition, 2008. ISBN 1-59327-144-1. x + 472 pp. LCCN QA76.9.A25 E75 2008. URL http://www.oreilly.com/

catalog/9781593271442.

Erle:2005:GMH

[483] Schuyler Erle and Rich Gibson. Google maps hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10161-9. 256 pp. LCCN TK5105.885.G66; TK5105.885.G66eb. URL http://www.oreilly.com/

catalog/9780596101619.

Erle:2005:MHT

[484] Schuyler Erle, Rich Gibson, and Jo Walsh. Mapping hacks: tips and tools for electronic cartography. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00703-5. xxxvi + 525 pp. LCCN GA139 .E75 2005. URL http://www.oreilly.

com/catalog/9780596007034.

Estrada:1993:CI

[485] Susan Estrada. Connecting to the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK 5105.875 I57 E82c 1993. US$15.95.

Estrada:1993:CIB

[486] Susan Estrada. Connecting to the Internet: a Buyer’s Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK 5105.875 I57 E82c 1993. US$15.95. URL http://

www.oreilly.com/catalog/9781565920613. This small book focuses on

choosing the best type of network connection for your personal, school or business needs, and how to get the best price for the type of access you require. Explains the differences between SLIP, PPP, ISDN, X.25 and other options. Includes an extensive list of Internet service providers. This is a single- purpose book, telling how to choose a connection and get online. It doesn’t try to teach you how to use the ’net once you’re there. That is graceful in its simplicity.

Eubanks:2005:WCJ

[487] Brian D. Eubanks. Wicked cool Java: code bits, open-source libraries, and project ideas. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-061-5. xvii + 224 pp. LCCN QA76.73.J38 E92 2005; QA76.73.J38

E92 2005eb. URL http://www.oreilly.com/catalog/9781593270612.

Experts:2009:JCS

[488] jQuery Community Experts, editor. jQuery cookbook: Solutions & Examples for jQuery Developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15977-3. xxii + 451 pp. LCCN QA76.73.J39 J68 2010.

Ezzell:1998:DWE

[489] Ben Ezzell. Developing Windows Error Messages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-356-1. xi + 240 pp. LCCN QA76.76.D47 E99 1998; QA76.9.E77 E99

1998. US$39.95. URL http://www.ora.com/catalog/error/;
http://www.oreilly.com/catalog/9781565923560;   http://www.
oreilly.com/catalog/error.

Farley:1998:JDC

[490] Jim Farley. Java Distributed Computing. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-206-9. xiv + 368 pp. LCCN QA76.9.D5 F37 1998. US$32.95. URL http://www.ora.com/catalog/

javadc/;   http://www.oreilly.com/catalog/9781565922068.

Farley:2002:JEN

[491] Jim Farley, William Crawford, and David Flanagan. Java Enterprise in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00152-5. xiii + 971 pp. LCCN QA76.73.J38 J366

2002. US$39.95. URL http://safari.oreilly.com/0596001525;
http://www.oreilly.com/catalog/9780596001520;   http://www.
oreilly.com/catalog/jentnut2.

Farley:2006:JEN

[492] Jim Farley and William Crawford. Java Enterprise in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-10142-2 (paperback). xviii + 871 pp. LCCN QA76.73.J38 F375 2006.

URL http://www.oreilly.com/catalog/9780596101428.

Faroult:2006:AS

[493] Stéphane Faroult and Peter Robson. The art of SQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00894-5. xvi + 349 pp. LCCN QA76.73.S67 F376 2006eb; QA76.73.S67. URL http://www.oreilly.

com/catalog/9780596008949.

Faroult:2008:RSA

[494] Stéphane Faroult and Pascal L’Hermite. Refactoring SQL applications. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51497-2. xv + 277 pp. LCCN QA76.73.S67 F37 2008eb; QA76.73.S67 F37 2008; QA76.73.S67; Internet.

Farquhar:2000:OWG

[495] David L. Farquhar. Optimizing Windows for Games, Graphics and Multimedia. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 2000. ISBN 1-56592-677-3. xiii + 278 pp. LCCN QA76.76.O63 F375 2000. US$24.95.

URL http://www.oreilly.com/catalog/9781565926776;   http://
www.oreilly.com/catalog/win9x.

Fauser:2006:RTR

[496] Cody Fauser. RJS templates for Rails. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52809-4. LCCN QA76.73.R72. URL http://www.

oreilly.com/catalog/9780596528096.

Feddema:2000:DOM

[497] Helen Bell Feddema. DAO Object Model: The Definitive Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-435-5. xiv + 389 pp. LCCN QA76.9.D3 F434 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565924352;   http://
www.oreilly.com/catalog/progacdao. Includes CD-ROM.

Feldt:2007:PF

[498] Kenneth C. Feldt. Programming Firefox. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10243-7. xvi + 494 pp. LCCN TK5105.883.F55 F35 2007eb; TK5105.883.F55. URL http://www.

oreilly.com/catalog/9780596102432.

Ferguson:1992:XRP

[499] Paula Ferguson, editor. The X Resource: a Practical Journal of the X Window System, 1992. ISSN 1058-5591. URL http://www.oreilly. com/catalog/. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800. Publication of this journal ceased at volume 17. A complete bibliography can be found at

https://www.math.utah.edu/pub/tex/bib/index.html#xres.

Ferguson:1993:MRM

[500] Paula Ferguson and David Brennan. Motif Reference Manual, volume 6B of The definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1993. ISBN 1-56592-038-4. xii + 908 pp. LCCN QA76.76.W56 F47 1993. US$34.95. URL http://www.

loc.gov/catdir/enhancements/fy0915/94185493-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0915/94185493-d.html;

http://www.oreilly.com/catalog/9781565920385;   http://www.
oreilly.com/catalog/v6b.

Ferguson:1994:XRIa

[501] Paula M. Ferguson, editor. The X Resource: Issue 11. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1994. ISBN 1-56592-068-6. 264 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565920682.

Ferguson:1994:XRIb

[502] Paula M. Ferguson, editor. The X Resource: Issue 12. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1994. ISBN 1-56592-069-4. 222 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565920699.

Ferguson:1995:XRIb

[503] Paula Ferguson, editor. The X Resource: Issue 14. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1995. ISBN 1-56592-122-4. 208 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565921221.

Ferguson:1995:XRIc

[504] Paula Ferguson, editor. The X Resource: Issue 15. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1995. ISBN 1-56592-140-2. 174 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565921405.

Ferguson:1996:XRIa

[505] Paula Ferguson, editor. The X Resource: Issue 16. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-166-6. 184 pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9781565921665.

Ferguson:2000:MRM

[506] Paula M. Ferguson and Dan Heller. Motif Reference Manual: For Motif 2.1, volume 6B of The definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, February 2000. ISBN 1-56592-654-4. xiii + 1154 pp. LCCN QA76.76.W56 F47 2000. US$49.95.

URL http://www.oreilly.com/catalog/9781565926547;   http://
www.oreilly.com/catalog/motifref2.

Fernandez:2008:C

[507] Dan Fernandez and Brian Peek. Coding4Fun. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15512-3. ???? pp. LCCN QA76.6 .F47 2008.

Ferrara:2002:PNW

[508] Alex Ferrara and Matthew MacDonald. Programming .NET Web Services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00250-5. xvi + 396 pp. LCCN QA76.625 .F49 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596002503;   http://
www.oreilly.com/catalog/prognetws.

Fettig:2006:TNP

[509] Abe Fettig and Glyph Lefkowitz. Twisted network programming essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10032-9. xix + 213 pp. LCCN QA76.73.P98 F48 2005eb; QA76.73.P98. URL http:

//www.oreilly.com/catalog/9780596100322.

Feuerstein:1995:OPS

[510] Steven Feuerstein. Oracle PL/SQL Programming. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-142-9. xxvii + 885 pp. LCCN QA76.73.S67 F49 1995. US$39.95. URL

http://www.ora.com/gnn/bus/ora/item/oraclep.html;   http://
www.oreilly.com/catalog/9781565921429. “Database management

systems”–Cover.. Includes 3 1/2 in. disk covering PL/SQL v1.1-2.3..

Feuerstein:1996:AOP

[511] Steven Feuerstein. Advanced Oracle PL/SQL Programming with Packages. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel:

+1 617 354 5800, 1996. ISBN 1-56592-238-7. xxvi + 661 pp. LCCN QA76.73.P258 F48 1996. US$42.95. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://www.
ora.com/catalog/advoracle/;   http://www.oreilly.com/catalog/
9781565922389;   http://www.oreilly.com/catalog/advoracle.

Feuerstein:1997:OPS

[512] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-335-9. xxxiii + 987 pp. LCCN QA76.73.S67 F49

1997. US$46.95. URL http://www.ora.com/catalog/oraclep2/;
http://www.oreilly.com/catalog/9781565923355;   http://www.
oreilly.com/catalog/oraclep2.

Feuerstein:1998:OBP

[513] Steven Feuerstein, Charles Dye, and John Beresniewicz. Oracle Built-in Packages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-56592-375-8. xxi + 931 pp. LCCN QA76.73.S67 F48 1998; QA76.9.D3 F49 1998. US$46.95. URL

http://www.oreilly.com/catalog/9781565923751;   http://www.
oreilly.com/catalog/oraclebip. Covers Oracle8.

Feuerstein:1998:OPS

[514] Steven Feuerstein, John Beresniewicz, and Chip Dawes. Oracle PL/ SQL Built-ins Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-456-8. vi + 72 pp. LCCN QA76.73.S67 F487 1998.

US$6.95. URL http://www.oreilly.com/catalog/9781565924567;
http://www.oreilly.com/catalog/orbipr.

Feuerstein:1999:OPSa

[515] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-457-6. v + 94 pp. LCCN QA76.73.P258 F488 1999; QA76.9.D3

F493 1999. US$7.95. URL http://www.oreilly.com/catalog/
9781565924574;   http://www.oreilly.com/catalog/orbilangpr/.

Feuerstein:1999:OPSb

[516] Steven Feuerstein. Oracle PL/SQL Programming: Guide to Oracle8i Features. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-675-7. xvi + 248 pp. LCCN QA76.73.P258 F489 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565926752;   http://
www.oreilly.com/catalog/ornewfeatures.

Feuerstein:2000:OPSa

[517] Steven Feuerstein and Andrew Odewahn. Oracle PL/SQL Developer’s Workbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-674-9. xix + 569 pp. LCCN QA76.73.P258 F4875 2000; QA76.9.D3 F492 2000. US$36.95. URL

http://www.oreilly.com/catalog/9781565926745;   http://www.
oreilly.com/catalog/ordevworkbook.

Feuerstein:2000:OPSb

[518] Steven Feuerstein. The Oracle PL/SQL CD bookshelf: 7 bestselling books on CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 1.0 edition, 2000. ISBN 1-56592-675-7 (paperback). LCCN Disk 1608 Protect. URL

http://www.oreilly.com/catalog/9781565926752.

Feuerstein:2001:OPS

[519] Steven Feuerstein. Oracle PL/SQL Best Practices: Optimizing Oracle Code. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00121-5. xix + 182 pp. LCCN QA76.9.D3 F49 2001; QA76.9.D3 F4916 2001.

US$19.95. URL http://www.oreilly.com/catalog/9780596001216;
 http://www.oreilly.com/catalog/orbestprac.

Feuerstein:2002:OPS

[520] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00381-1. xxiii + 989 pp. LCCN QA76.73.S67 F49

2002. US$54.95. URL http://safari.oreilly.com/0596003811;
http://www.oreilly.com/catalog/9780596003814;   http://www.
oreilly.com/catalog/oraclep3.

Feuerstein:2003:OPS

[521] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00472-9. vi + 121 pp. LCCN QA76.9.D3 F493

2003. URL http://www.oreilly.com/catalog/9780596004729.

Feuerstein:2004:OPS

[522] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00680-2. viii + 147 pp. LCCN ????

Feuerstein:2005:OP

[523] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2005. ISBN 0-596-00977-1 (paperback). xxiv + 1171 pp. LCCN QA76.73.S67 F49 2005.

URL http://www.oreilly.com/catalog/9780596009779;   http://
www.oreilly.com/catalog/oraclep4/.

Feuerstein:2008:OPS

[524] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL language: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2008. ISBN 0-596-51404-2 (paperback). vii + 170 pp. LCCN QA76.9.D3. URL

http://www.oreilly.com/catalog/9780596514044.

Few:2006:IDD

[525] Stephen Few. Information dashboard design: the effective visual communication of data. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10016-7. viii + 211 pp. LCCN HD30.213; HD30.213 F49 2006eb.

URL http://www.oreilly.com/catalog/9780596100162.

Figgin:2003:LWS

[526] Stephen Figgin, Aaron Weber, Ellen Siever, Matt Welsh, George Reese, Ben Laurie, Peter Laurie, Rasmus Lerdorf, Stas Bekman, and Eric Cholet, editors. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 2.0. edition, 2003. ISBN 0-596-00529-

6. LCCN QA76.76.O63. URL http://www.oreilly.com/catalog/
9780596005290.

Finn:1999:CCT

[527] Robert Finn. Cancer Clinical Trials: Experimental Treatments & How They Can Help You. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-566-1. xvi + 206 pp. LCCN RC267 .F56 1999.

US$14.95. URL http://www.oreilly.com/catalog/9781565925663;
 http://www.oreilly.com/catalog/cancerct.

Finn:2000:OTM

[528] Robert Finn. Organ transplants: making the most of your gift of life. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-634-X. xv + 311 pp. LCCN RD120.75 .F55 2000. URL http:/

/www.oreilly.com/catalog/9781565926349.

Fitzgerald:2003:LX

[529] Michael Fitzgerald. Learning XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00327-7. xiii + 352 pp. LCCN QA76.73.X58 F58

2003. US$34.95, CDN$54.95. URL http://press.oreilly.com/pub/
pr/1126;   http://www.oreilly.com/catalog/9780596003272.

Fitzgerald:2004:XH

[530] Michael Fitzgerald. XML hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00711-6. xviii + 460 pp. LCCN QA76.76.H94; QA76.76.H94 F473 2004; QA76.76.H94 F4725 2004; QA76.76.H94 F58 2004; QA76.76.H94 F473 2004eb. URL http://www.oreilly.com/

catalog/9780596007119.

Fitzgerald:2007:GAS

[531] Michael Fitzgerald. Google Ajax search API. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52953-8. LCCN TK5105.8885.A52 F58

2007eb; TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596529536.

Fitzgerald:2007:LR

[532] Michael Fitzgerald. Learning Ruby. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52986-4 (paperback). xv + 238 pp. LCCN QA76.64 .F58 2007; QA76.64 .F58 2007eb; QA76.64. URL http://www.oreilly.

com/catalog/9780596529864.

Fitzgerald:2007:RPR

[533] Michael Fitzgerald. Ruby Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51481-6. vi + 170 pp. LCCN QA76.73.R82; QA76.73.R82 F58 2007. URL http://www.oreilly.com/

catalog/9780596514815.

Flamini:1996:PGY

[534] Roland Flamini. Passport Germany: Your Pocket Guide to German Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-885073-20-8. 96 pp. LCCN HD58.7.F58 1996. US$6.95.

URL http://www.oreilly.com/catalog/9781885073204.

Flanagan:1991:PSR

[535] David Flanagan. Programmer’s Supplement for Release 5 of the X Window System, Version 11. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-86-2. xx + 367 pp. LCCN QA76.76.W56 F5 1991. URL http:

//www.oreilly.com/catalog/9780937175866.

Flanagan:1992:XTI

[536] David Flanagan. X Toolkit Intrinsics Reference Manual, volume 5 of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1992. ISBN 1-56592-007-4. xiii + 899 pp. LCCN QA76.76.W56 N94 1992. URL http://www.oreilly.com/

catalog/9781565920071;   http://www.oreilly.com/catalog/v5.

Flanagan:1994:MTS

[537] David Flanagan. Motif Tools: Streamlined GUI Design and Programming with the Xmt Library. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1994. ISBN 1-56592-044-9. xxxvii + 984 pp. LCCN QA76.76.W56 F53 1994. US$54.95 (CD-ROM included). URL http://www.oreilly.com/

catalog/9781565920446.

Flanagan:1995:JND

[538] David Flanagan. Java in a Nutshell: a Desktop Quick Reference for Java Programmers. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-183-6. xix + 438 pp. LCCN QA76.73.J38F553 1996.

US$19.95. URL http://www.ora.com/info/java.

Flanagan:1996:JDG

[539] David Flanagan. JavaScript: The Definitive Guide: Beta Edition. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, beta edition, August 1996. ISBN 1-56592-193-3. xv + 454 pp. LCCN QA76.73.J38 F555 1996; QA76.73.J39 F53 1996. US$29.95. URL http://www.oreilly.

com/catalog/9781565921931.

Flanagan:1996:JN

[540] David Flanagan. Java in a Nutshell: a desktop quick reference for Java programmers. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-183-6. xix + 438 pp. LCCN QA76.73.J38F553

1996. US$19.95. URL http://www.ora.com/info/java;   http://
www.oreilly.com/catalog/9781565921832.

Flanagan:1996:JNV

[541] David Flanagan. Java in a Nutshell, Version Française. Editions O’Reilly International Thomson, 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1996. ISBN 2-84177-009-5. xxii + 492 pp. LCCN ???? 220 FF. URL http://www.

editions-oreilly.fr/international/france/prog/javanut.html;

http://www.ora.de/. French translation by Serge Chaumette, Georges Eyrolles and Alain Miniussi. Includes CD-ROM.

Flanagan:1997:JDG

[542] David Flanagan. JavaScript: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-234-4. xiv + 647 pp. LCCN QA76.73.J39 F53 1997; QA76.73.V38 F53 1997.

URL http://www.oreilly.com/catalog/9781565922341;   http://
www.oreilly.com/catalog/jscript2.

Flanagan:1997:JEN

[543] David Flanagan. Java examples in a nutshell: a tutorial companion to Java in a nutshell. A Nutshell handbook; The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-371-5. xiv + 397 pp. LCCN QA76.73.J38 F552 1997.

US$24.95. URL http://www.ora.com/catalog/jenut/;   http://
www.oreilly.com/catalog/9781565923713;   http://www.oreilly.
com/catalog/jenut.

Flanagan:1997:JNb

[544] David Flanagan. Java in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, October 1997. ISBN 3-89721-100-9 (??invalid ISBN??).

LCCN ???? 49 DM. URL http://www.oreilly.de/german/java/

javanut2/. Deutsche Ausgabe.

Flanagan:1997:JNc

[545] David Flanagan. Java in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 4-900900-08-7. ???? pp. LCCN ???? 3900 yen. URL http:/

/www.ora.com/catalog/jscriptbeta/author.html;   http://www.
ora.com/international/catalog/japanese/inet/java.html;   http:
//www.ora.com/international/japan/;   http://www.oreilly.co.

jp/BOOK/java.htm. Japanese translation.

Flanagan:1997:JNd

[546] David Flanagan et al. Java in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, deluxe edition, 1997. ISBN 1-56592-304-9. xvi + 610 pp. LCCN QA76.73.J38 J38 1997. US$69.95. URL http://www.

ora.com/catalog/javadeluxe/;   http://www.ora.com/catalog/{
javadeluxe}/;   http://www.oreilly.com/catalog/9781565923041;
 http://www.oreilly.com/catalog/javadeluxe.

Flanagan:1997:JNDa

[547] David Flanagan. Java in a Nutshell: a desktop quick reference. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-262-X, 1-56592-273-5. xvi + 610 pp. LCCN QA76.73.J38 F553 1997. US$24.95. URL http:

//www.ora.com/catalog/javanut2/noframes.html;   http://www.
oreilly.com/catalog/9781565922624;   http://www.oreilly.com/
catalog/doctor;   http://www.oreilly.com/catalog/javanut2.

Flanagan:1997:JRL

[548] David Flanagan, Mark Grand, Pat Niemeyer, Josh Peck, and John Zukowski. Java Reference Library on the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-359-6. 4000 (est.) pp. LCCN

???? US$59.95. URL http://www.ora.com/catalog/javarlw/;
http://www.oreilly.com/catalog/9781565923591;   http://www.
oreilly.com/catalog/javarlw.

Flanagan:1998:JDG

[549] David Flanagan. JavaScript: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, June 1998. ISBN 1-56592-392-8. xiv + 776 pp. LCCN QA76.73.J39 F53 1998. US$39.95.

URL http://www.oreilly.com/catalog/9781565923928;   http://
www.oreilly.com/catalog/jscript3.

Flanagan:1998:JEN

[550] David Flanagan. Java Examples in a Nutshell — Deutsche Ausgabe. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1998. ISBN 3-89721-112-2 (??invalid ISBN??). LCCN ???? 49 DM. URL http://www.oreilly.

de/german/java/jenut/.

Flanagan:1998:JNDa

[551] David Flanagan. Java in a Nutshell, Deutsche Ausgabe. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 3-930673-46-0 (??invalid ISBN??).

???? pp. LCCN ???? 39 DM. URL http://www.ora.com/gnn/bus/
ora/item/javanut.author.html;   http://www.ora.de/.

Flanagan:1998:JNDb

[552] David Flanagan. Java in a Nutshell (Deutsche Ausgabe), 2. Auflage. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 3-89721-100-9. xvi +

638 pp. LCCN ???? 49 DM. URL http://www.oreilly.de/german/
java/javanut2/.

Flanagan:1998:JPR

[553] David Flanagan. JavaScript Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-521-1. v + 89 pp. LCCN QA76.73.J39

F52 1998. US$7.95. URL http://www.oreilly.com/catalog/
9781565925212;   http://www.oreilly.com/catalog/jscriptpr.

Flanagan:1999:JEN

[554] David Flanagan, Jim Farley, William Crawford, and Kris Magnusson, editors. Java Enterprise in a Nutshell: a Desktop Quick Reference. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-483-5. xvi + 604 pp. LCCN QA76.73.J38 J366 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565924833;   http://
www.oreilly.com/catalog/jentnut.

Flanagan:1999:JFC

[555] David Flanagan. Java Foundation Classes in a nutshell: a desktop quick reference. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-488-6. xiii + 731 pp. LCCN QA76.73.J38 F5524 1999.

US$29.95. URL http://www.oreilly.com/catalog/9781565924888;
 http://www.oreilly.com/catalog/jfcnut.

Flanagan:1999:JNE

[556] David Flanagan. Java in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, November 1999. ISBN 1-56592-487-8. xviii + 648 pp. LCCN QA76.73.J38 F553 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565924871;   http://
www.oreilly.com/catalog/javanut3.

Flanagan:1999:JPR

[557] David Flanagan. Java Power Reference: a Complete Searchable Resource on CD-ROM. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-589-0. xi + 51 pp. LCCN QA76.76.J38. US$19.95.

URL http://www.oreilly.com/catalog/9781565925892;   http://
www.oreilly.com/catalog/jpower/. Includes CD-ROM.

Flanagan:19xx:JEN

[558] David Flanagan. Java Examples in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-09-5. LCCN ???? 4,300 yen. Japanese translation.

Flanagan:2000:JEC

[559] David Flanagan, David Flanagan, Richard Monson-Haefel, Jason Hunter, Scott Oaks, and Jim Farley. The Java enterprise CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-850-4 (set). LCCN QA76.73.J38 J3656 2000. US$89.95.

URL http://www.oreilly.com/catalog/9781565928503;   http://
www.oreilly.com/catalog/javacdbs. Title from disc label Bonus book

has title: Java Enterprise in a nutshell: a desktop quick reference / David Flanagan … [et al.]. 1999 Contents: Java enterprise in a nutshell – Java in a nutshell, 3rd ed. – Java foundation classes in a nutshell – Enterprise JavaBeans, 2nd ed. – Java servlet programming – Java security – Java distributed computing.

Flanagan:2000:JEN

[560] David Flanagan. Java examples in a nutshell: a tutorial companion to Java in a nutshell. The Java series; In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, October 2000. ISBN 0-596-00039-1. xvi + 566 pp. LCCN QA76.73.J38 F5518 2000. US$29.95.

URL http://www.oreilly.com/catalog/9780596000394;   http://
www.oreilly.com/catalog/jenut2.

Flanagan:2000:JPL

[561] David Flanagan, Jim Farley, William Crawford, and Kris Magnusson. Java Professional Library. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 2000. ISBN 0-596-00107-X. LCCN ????

URL http://www.oreilly.com/catalog/9780596001070;   http://
www.oreilly.com/catalog/javaprolib.

Flanagan:2002:JDG

[562] David Flanagan. JavaScript: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2002. ISBN 0-596-00048-0. xvii + 916 pp. LCCN QA76.73.J39 F53 1998. URL http://www.oreilly.com/

catalog/9780596000486.

Flanagan:2002:JEN

[563] David Flanagan, Jim Farley, and William Crawford. Java Enterprise in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00152-5. xiii + 971 pp. LCCN QA76.73.J38 J366 2002. US$39.95. URL http://safari.oreilly.com/

0596001525;   http://www.oreilly.com/catalog/9780596001520;
 http://www.oreilly.com/catalog/jentnut2.

Flanagan:2002:JND

[564] David Flanagan. Java in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2002. ISBN 0-596-00283-1. xxi + 969 pp. LCCN QA76.73.J38 F553 2002. URL http://www.oreilly.

com/catalog/9780596002831. Covers Java 1.4.

Flanagan:2002:JPR

[565] David Flanagan. JavaScript Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00411-7. vii + 127 pp. LCCN QA76.73.J39 F525 2002eb. US$14.95.

URL http://www.oreilly.com/catalog/9780596004118;   http://
www.oreilly.com/catalog/jscriptpr2.

Flanagan:2004:JENa

[566] David Flanagan. Java examples in a nutshell. A nutshell handbook; The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00620-9. xiv + 701 pp. LCCN QA76.73.J38 F485 2004; QA76.73.J38 F552 2004eb; QA76.73.J38 F552 2004; QA76.73.J38 F53

2004. URL http://www.oreilly.com/catalog/9780596006204.

Flanagan:2004:JENb

[567] David Flanagan. Java examples in a nutshell: a tutorial companion to Java in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00585-7. xiv + 701 pp. LCCN QA76.73.J38

F5518 2004. URL http://www.oreilly.com/catalog/9780596005856.

Flanagan:2005:JN

[568] David Flanagan. Java in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fifth edition, 2005. ISBN 0-596-00773-6. xxiv + 1224 pp. LCCN QA76.73.J38 F553 2005.

URL http://www.oreilly.com/catalog/9780596007737.

Flanagan:2005:JND

[569] David Flanagan. Java in a nutshell: [a desktop quick reference]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2005. ISBN 0-596-00773-6. xxiv + 1224 pp. LCCN QA76.73.J38 F553 2005. URL http://www.oreilly.

com/catalog/9780596007737.

Flanagan:2006:JDG

[570] David Flanagan. JavaScript: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-10199-6. xxii + 994 pp. LCCN QA76.73.J39 F53 2006eb; QA76.73.J39 F53 2006; QA76.73.J39. URL

http://www.oreilly.com/catalog/9780596101992.

Flanagan:2008:RPL

[571] David Flanagan and Yukihiro Matsumoto. The Ruby programming language. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51617-7. xi

+ 429 pp. LCCN QA76.64. URL http://www.oreilly.com/catalog/
9780596516178.

Flanders:2009:RN

[572] Jon Flanders. RESTful.NET. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15502-6. 304 (est.) pp. LCCN TK5105.88813 F53 2008.

Fleishman:2008:TCB

[573] Glenn Fleishman. Take control of back to my Mac. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-46-7. ???? pp. LCCN QA76.76.O63 F54 2008.

Fleishman:2008:TCY

[574] Glenn Fleishman. Take control of your 802.11n AirPort network. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-50-5. ???? pp. LCCN TK5105.78 .F56 2008.

Fleming:1998:WND

[575] Jennifer Fleming. Web Navigation: Designing the User Experience. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-351-0. xv + 253 pp. LCCN TK5105.888 .F64 1998 CRB. US$34.95. URL http:/

/www.ora.com/catalog/navigation/;   http://www.oreilly.com/
catalog/9781565923515.

Flesner:2007:AVY

[576] Andy Flesner. AutoIt v3: your quick guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51512-X. LCCN HF5548; HF5548 .F54 2007eb.

URL http://www.oreilly.com/catalog/9780596515126.

Flickenger:2002:BWC

[577] Rob Flickenger. Building Wireless Community Networks: Implementing the Wireless Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00204-1. x + 125 pp. LCCN TK5105.85 .F58

2002. US$24.95. URL http://safari.oreilly.com/0596002041;
http://www.oreilly.com/catalog/9780596002046;   http://www.
oreilly.com/catalog/wirelesscommnet.

Flickenger:2003:BWC

[578] Rob Flickenger. Building Wireless Community Networks: Planning and deploying local wireless networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00502-4. xii + 168 pp. LCCN TK5105.85 .F58 2003. US$24.95. URL http://safari.oreilly.com/

0596005024;   http://www.oreilly.com/catalog/9780596005023;
 http://www.oreilly.com/catalog/wirelesscommnet2/.

Flickenger:2003:LSH

[579] Rob Flickenger. Linux server hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00461-3. xvii + 221 pp. LCCN QA76.9.C55 F58 2003. URL http://www.oreilly.

com/catalog/9780596004613.

Flickenger:2003:WHI

[580] Rob Flickenger. Wireless hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00559-8. xv + 286 pp. LCCN TK5103.2 .F55 2003. URL http://www.oreilly.com/

catalog/9780596005597.

Flickenger:2006:WH

[581] Rob Flickenger and Roger Weeks. Wireless hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10144-9. xxii + 440 pp. LCCN TK5103.2 .F55 2006.

URL http://www.loc.gov/catdir/toc/fy0606/2006272661.html;
http://www.oreilly.com/catalog/9780596101442.

Fogel:2006:POS

[582] Karl Fogel. Producing open source software: how to run a successful free software project. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00759-0. xx + 279 pp. LCCN QA76.76.D47 F64 2006. URL http:

//www.oreilly.com/catalog/9780596007591.

Foley:2004:PDP

[583] Richard Foley. Perl debugger: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00503-2. vi + 135 pp. LCCN QA76.73.P22

P476 2004. URL http://www.oreilly.com/catalog/9780596005030.

Ford:2000:APR

[584] Andrew Ford. Apache Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2000. ISBN 1-56592-706-0. iii + 107 pp. LCCN TK5105.8885.A63 F67 2000. US$9.95.

URL http://www.oreilly.com/catalog/9781565927063;   http://
www.oreilly.com/catalog/apachepr.

Ford:2001:MPR

[585] Andrew Ford. mod perl Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00047-2. vii + 83 pp. LCCN TK5105.8885.A63

F673 2001. US$9.95. URL http://www.oreilly.com/catalog/
9780596000479;   http://www.oreilly.com/catalog/modperlpr.

Ford:2008:APR

[586] Andrew Ford. Apache 2 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-15489-5. ???? pp. LCCN TK5105.8885.A63 F67 2008.

Ford:2008:PP

[587] Neal Ford. The productive programmer. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51978-8. xiii + 206 pp. LCCN QA76.755 .F67 2008.

Foster:2000:DMN

[588] Gerald Foster. Desktop Management with Novell ZENworks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2000. ISBN 1-56592-711-7. xii + 271 pp. LCCN QA76.9.U83 F68 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565927117;   http://
www.oreilly.com/catalog/novellzen.

Fowler:2005:MJW

[589] Chad Fowler. My job went to India: and all I got was this lousy book. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-1-8. viii + 185 pp. LCCN QA76.25 .F69 2005.

Fox:2006:EMO

[590] Chris Fox. Essential Microsoft Operations Manager. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00953-4. xxi + 355 pp. LCCN

QA76.9.C55eb; QA76.9.C55. URL http://www.oreilly.com/catalog/
9780596009533.

Foxall:2000:MNV

[591] James D. Foxall. MCSD in a Nutshell: The Visual Basic Exams. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-752-4. xviii + 612 pp. LCCN QA76.3 .F69 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565927520;   http://
www.oreilly.com/catalog/mcsdnut.

Foy:2007:MP

[592] Brian D. Foy. Mastering Perl. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52724-1 (paperback). xvii + 322 pp. LCCN QA76.73.P22 F69 2007. URL

http://www.oreilly.com/catalog/9780596527242.

Francia:1996:PPY

[593] Luis Francia, editor. Passport Philippines: Your Pocket Guide to Filipino Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-40-2. 96 pp. LCCN HD58.7.F696 1997. US$6.95.

URL http://www.oreilly.com/catalog/9781885073402.

Freedman:1996:JP

[594] Edie (?) Freedman and Rinaldo DiGiorgio. Java Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-185-2. ???? pp. LCCN ????

???? URL http://www.ora.com/.

Freedman:1996:JPD

[595] Edie (?) Freedman and ?. DiGiorgio. Java Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-185-2. ???? pp. LCCN ????

URL http://www.oreilly.com/catalog/9781565921856.

Freeman:2003:PNS

[596] Adam Freeman and Allen Jones. Programming .NET security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00442-7. xviii + 693 pp.

LCCN TK5105.59 .F68 2003. URL http://www.oreilly.com/catalog/
9780596004422.

Freeman:2006:HFH

[597] Elisabeth Freeman and Eric Freeman. Head first HTML with CSS and XHTML. Head first series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10197-X. xxxv + 658 pp. LCCN QA76.76.H94 F74 2006; QA76.76.H94; QA76.76.H94 F74 2006eb. URL http://www.oreilly.

com/catalog/9780596101978.

Freiss:1998:PNS

[598] Martin Freiss. Protecting Networks with SATAN: Internet security for system administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-425-8. xiv + 112 pp. LCCN TK5105.59 .F7413 1998.

US$19.95. URL http://www.oreilly.com/catalog/9781565924253;
 http://www.oreilly.com/catalog/satan.

Frey:1989:DEM

[599] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1989. ISBN 0-937175-39-0. xv + 284 pp. LCCN HE6239.E54 F73 1989.

US$26.95. URL http://www.oreilly.com/catalog/9780937175392. A

quick desk reference to many networks, with two pages on each one.

Frey:1989:VDE

[600] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1989. ISBN 0-937175-39-0. xv + 284 pp. LCCN HE6239.E54 F73 1989.

Frey:1993:DEM

[601] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1993. ISBN 1-56592-031-7. xvi + 443 pp.

LCCN HE6239.E54 F78 1993. US$24.95. URL http://www.oreilly. com/catalog/9781565920316. This book is a reference work. It details the various computer networks with gateways to the Internet. It is common to cite such works as “indispensable”: in fact, most users, and even site managers, muddle along quite happily without it. Quick reference “electronic” versions exist of very similar documents, which provide the addressing schemes for the more common network and commercial service gateways. Frey and Adams have, however, put together a very complete and interesting reference, and I do suggest it to anyone managing, or using, extensive email correspondence. As a user of electronic mail, or the manager of a small Internet node or UUCP site, it would be hard to say that you “need” this book. If, however, you are at all interested in the topic of email, you will find this fascinating and useful.

Frey:1994:DEM

[602] Donnalyn Frey and Rick Adams. !%@:: A Directory of Electronic Mail Addressing and Networks. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, June 1994. ISBN 1-56592-046-5. xix + 640 pp. LCCN HE6239.E54 F78 1994. US$9.95. URL http://www.oreilly.

com/catalog/9781565920460.

Frey:1994:VDE

[603] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 1994. ISBN 1-56592-046-5. 662 pp. LCCN HE6239.E54 F73 1989. US$9.95.

Frey:2006:SH

[604] Bruce Frey. Statistics hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10164-3. xvii + 336 pp. LCCN QA276 .F74 2006eb; QA276.

URL http://www.oreilly.com/catalog/9780596101640.

Frieder:2012:CSP

[605] Ophir Frieder, Gideon Frieder, and David Grossman. Computer Science Programming Basics in Ruby. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-5597-8. 180 pp. LCCN ???? EUR 20.00.

Friedl:1997:MRE

[606] Jeffrey E. F. Friedl. Mastering regular expressions: powerful techniques for Perl and other tools. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-257-3. xxiv + 342 pp. LCCN QA76.73.P22 F75 1997; QA76.9.D3G728

1997. US$29.95. URL http://www.ora.com/catalog/regex/;
http://www.oreilly.com/catalog/9781565922570;   http://www.
oreilly.com/catalog/regex.

Friedl:2002:MRE

[607] Jeffrey E. F. Friedl. Mastering Regular Expressions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00289-0. xxii + 460 pp. LCCN QA76.73.P22 F75 2002; QA76.9.T48 F75 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596002893;   http://
www.oreilly.com/catalog/regex2.

Friedl:2006:MRE

[608] Jeffrey E. F. Friedl. Mastering regular expressions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-52812-4 (paperback). xxiv + 515 pp. LCCN QA76.9.T48 F75 2006. URL http://

www.loc.gov/catdir/enhancements/fy0715/2007272426-d.html;
http://www.oreilly.com/catalog/9780596528126.

Friedman:2002:WPG

[609] Mark Friedman and Odysseas Pentakalos. Windows 2000 Performance Guide: Help for Windows 2000 Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-466-5. xv + 700 pp. LCCN QA76.76.O63 F774 2002. US$44.95. URL

http://safari.oreilly.com/1565924665;   http://www.oreilly.
com/catalog/9781565924666;   http://www.oreilly.com/catalog/
w2kperf.

Fries:2005:DAE

[610] Bruce Fries. Digital audio essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00856-2 (paperback). xiv + 357 pp. LCCN

TK5105.8863 .F75 2005. URL http://www.oreilly.com/catalog/
9780596008567.

Frisch:1991:ESA

[611] Æleen Frisch. Essential System Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-74-9 (paperback), 0-937175-80-3 (paperback). xxiii + 440 pp. LCCN QA76.76.O63 F78 1992. URL http:

//www.oreilly.com/catalog/9780937175804.

Frisch:1995:ESA

[612] Æleen Frisch. Essential System Administration. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, September 1995. ISBN 1-56592-127-5. xxvii + 758 pp. LCCN QA76.76.O63 F78 1995.

US$34.95. URL http://www.oreilly.com/catalog/9781565921276;
 http://www.oreilly.com/catalog/esa2.

Frisch:1998:EWN

[613] Æleen Frisch. Essential Windows NT System Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-274-3. xvi + 467 pp. LCCN QA76.76.O63 F65 1998. US$34.95. URL

http://www.ora.com/catalog/esawinnt/;   http://www.oreilly.
com/catalog/9781565922747;   http://www.oreilly.com/catalog/
esawinnt.

Frisch:1998:WND

[614] Æleen Frisch. Windows NT Desktop Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1998. ISBN 1-56592-437-1. iii + 60 pp. LCCN QA76.76.O63 F748 1998. US$6.95.

URL http://www.oreilly.com/catalog/9781565924376;   http://
www.oreilly.com/catalog/winntdr.

Frisch:2001:WCP

[615] Æleen Frisch. Windows 2000 commands: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2001. ISBN 0-596-00148-7. vi + 110 pp. LCCN QA76.76.O63 F7837 2001. US$9.95.

URL http://www.oreilly.com/catalog/9780596001483;   http://
www.oreilly.com/catalog/win2commpr.

Frisch:2002:ESAa

[616] Æleen Frisch. Essential System Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00343-9. xxiv + 1149 pp. LCCN QA76.76.O63 F75 2002; QA76.76.O63 F78 2002. US$54.95,

CDN$85.95, UK£38.95. URL http://www.oreilly.com/catalog/
9780596003432;   http://www.oreilly.com/catalog/esa3.

Frisch:2003:ESAb

[617] Æleen Frisch. Essential System Administration Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00449-4. vi + 137 pp. LCCN QA76.76.O63 F782 2003. US$14.95.

URL http://www.oreilly.com/catalog/9780596004491;   http://
www.oreilly.com/catalog/esapr.

Frost:2000:LWW

[618] Martin Frost. Learning WML and WMLScript: Programming the Wireless Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-947-0. xvii + 179 pp. LCCN TK5105.5865 .F76 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565929470;   http://
www.oreilly.com/catalog/learnwml.

Fry:2008:VD

[619] Ben Fry. Visualizing data. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51455-7. xiii + 366 pp. LCCN T385 .F79 2008; T385. URL http://www.oreilly.

com/catalog/9780596514556.

Frye:1997:ONR

[620] Curtis Frye and Richard Peck. The O’Reilly/Netcraft Report on Secure Transactions: The State of Web Commerce. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISSN 1092-0560. 120 pp. LCCN

???? URL http://www.ora.com/catalog/marketres/;   http://
www.oreilly.com/catalog/. Includes CD-ROM.

Frye:2003:EPG

[621] Curtis Frye. Excel: pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00534-2. viii + 111 pp. LCCN HF5548.4.M523 F785

2003. URL http://www.oreilly.com/catalog/9780596005344.

Frye:2005:EAH

[622] Curtis Frye. Excel annoyances: how to fix the most annoying things about your favorite spreadsheet. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00728-0. xvii + 235 pp. LCCN HF5548.4.M523 F79 2004eb; HF5548.4.M523

F79 2004; HF5548.4.M523. URL http://www.oreilly.com/catalog/
9780596007287.

Frye:2008:EPG

[623] Curtis Frye. Excel 2007 pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51452-2. vi + 165 pp. LCCN HF5548.4.M523 F79 2007eb; HF5548.4.M523. URL http://www.

oreilly.com/catalog/9780596514525.

Fullam:2004:HHP

[624] Scott Fullam. Hardware hacking projects for geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00314-5. xiv + 331 pp. LCCN TK9965 .F85

2004. URL http://www.oreilly.com/catalog/9780596003142.

Gallmeister:1995:PPR

[625] Bill O. Gallmeister. POSIX.4: Programming for the Real World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1995. ISBN 1-56592-074-0. xviii + 548 pp. LCCN QA76.54 .G34 1995; QA76.76.O63 G34 1995. US$29.95.

URL http://www.oreilly.com/catalog/9781565920743;   http://
www.oreilly.com/catalog/posix4.

Garfinkel:1991:PUS

[626] Simson Garfinkel and Gene Spafford. Practical UNIX Security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1991. ISBN 0-937175-72-2. 512 pp. LCCN QA76.76.O63 G38 1992. US$29.95. URL http://www.oreilly.com/

catalog/9780937175729.

Garfinkel:1995:PPG

[627] Simson Garfinkel. PGP: Pretty Good Privacy. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-098-8. xxxiii + 393 pp. LCCN QA76.9.A25G36 1995. US$24.95.

Garfinkel:1996:PUI

[628] Simson Garfinkel and Gene Spafford. Practical UNIX & Internet Security. Computer security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1996. ISBN 1-56592-148-8. xxix + 971 pp. LCCN QA76.76.O63

G38 1996. US$39.95. URL http://www.ora.com/gnn/bus/ora/item/
pus2.html;   http://www.oreilly.com/catalog/puis.

Garfinkel:1997:WSC

[629] Simson Garfinkel and Gene Spafford. Web Security & Commerce. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-269-7. xx + 483 pp. LCCN TK5105.59 .G37 1997. US$32.95.

URL http://www.oreilly.com/catalog/9781565922693;   http://
www.oreilly.com/catalog/websec. Covers Java, JavaScript, and Ac-

tiveX.

Garfinkel:2000:DND

[630] Simson Garfinkel. Database nation: the death of privacy in the 21st Century. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-653-6. vii + 312 pp. LCCN JC596.2.U5 G37 2000. URL http://www.oreilly.

com/catalog/9781565926530.

Garfinkel:2001:DND

[631] Simson Garfinkel. Database nation: the death of privacy in the 21st Century. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2001. ISBN 0-596-00105-3. vii + 326 pp. LCCN JC596.2.U5 G37 2001. US$16.95.

URL http://www.oreilly.com/catalog/9780596001056;   http://
www.oreilly.com/catalog/dbnationtp.

Garfinkel:2002:BCA

[632] Simson Garfinkel and Michael K. Mahoney. Building Cocoa Applications: a Step-by-Step Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00235-1. xxv + 620 pp. LCCN QA76.64 .G36 2002; QA76.76.O63 G37

2002. US$44.95. URL http://safari.oreilly.com/0596002351;
http://www.oreilly.com/catalog/9780596002350;   http://www.
oreilly.com/catalog/buildcocoa.

Garfinkel:2002:WSP

[633] Simson Garfinkel and Gene Spafford. Web Security, Privacy & Commerce. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00045-6. xxviii + 756 pp. LCCN TK5105.59 .G37

2002. US$44.95. URL http://safari.oreilly.com/0596000456;
http://www.oreilly.com/catalog/9780596000455;   http://www.
oreilly.com/catalog/websec2.

Garfinkel:2003:PUI

[634] Simson Garfinkel, Gene Spafford, and Alan Schwartz. Practical Unix & Internet Security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00323-4. xxix + 954 pp. LCCN QA76.76.O63 G38 2003.

US$54.95. URL http://www.oreilly.com/catalog/9780596003234;
 http://www.oreilly.com/catalog/puis3.

Garman:2003:KDG

[635] Jason Garman. Kerberos: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00403-6. xiv + 253 pp. LCCN TK5105.59

.G375 2003. URL http://www.oreilly.com/catalog/9780596004033.

Garrett:2006:JCT

[636] Aviva Garrett. JUNOS cookbook: time-saving techniques for JUNOS software configuration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10014-0. xxii + 657 pp. LCCN TK5105.543; TK5105.543 .G37

2006eb. URL http://www.oreilly.com/catalog/9780596100148.

Garrish:2012:EBP

[637] Matt Garrish and Markus Gylling. EPUB 3 Best Practices. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2914-4. 350 pp. LCCN ???? EUR 24.50.

Gaskin:2005:TCS

[638] James E. Gaskin. Talk is cheap: switching to Internet telephones. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00960-7 (paperback). xvi + 248 pp. LCCN TK5105.8865 .G37 2005. URL http://

www.oreilly.com/catalog/9780596009601.

Gaskins:1992:PPMa

[639] Tom Gaskins. PHIGS Programming Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 0-937175-85-4 (paperback), 0-937175-92-7 (hardcover). xxxix + 908 pp. LCCN QA76.76.W56 G37 1992. US$42.95 (softcover), US$52.95 (hardcover). URL http://www.oreilly.com/

catalog/9780937175859.

Gast:2001:TSG

[640] Matthew Gast. T1: a Survival Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00127-4. xiv + 288 pp. LCCN TK5103.15 .G37 2001.

US$29.95. URL http://www.oreilly.com/catalog/9780596001278;
 http://www.oreilly.com/catalog/t1survival.

Gast:2002:WND

[641] Matthew Gast. 802.11 Wireless Networks: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00183-

5. xvii + 443 pp. LCCN TK5103.2 .G37 2002. US$44.95. URL

http://safari.oreilly.com/0596001835;   http://www.oreilly.
com/catalog/802dot11;   http://www.oreilly.com/catalog/9780596001834.

Gast:2005:WND

[642] Matthew Gast. 802.11 wireless networks: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-10052-3 (paperback). xxi + 630 pp. LCCN TK5103.2 .G385 2005. URL http:/

/www.oreilly.com/catalog/9780596100520.

Gehtland:2005:SDN

[643] Justin Gehtland and Bruce A. Tate. Spring: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00910-0. xv + 184 pp. LCCN QA76.73.S67 T38 2005. URL http://www.oreilly.com/

catalog/9780596009106.

Gemmell:2005:MMA

[644] Keith Gemmell. Making music on the Apple Mac. PC, Merton, UK, 2005. ISBN 1-870775-95-3 (paperback). 103 pp. LCCN MT723 .G46

2004; MT723 .G46 2005. URL http://www.oreilly.com/catalog/
9781870775953.

Gennick:1999:OSD

[645] Jonathan Gennick. Oracle SQL*Plus: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-578-

5. xxi + 502 pp. LCCN QA76.73.S67 G46 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565925786;   http://
www.oreilly.com/catalog/orsqplus/.

Gennick:2000:OSP

[646] Jonathan Gennick. Oracle SQL*Plus: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2000. ISBN 1-56592-941-1. iv + 90 pp. LCCN QA76.73.S67 G46 2000; QA76.9.D3 G455 2000. US$9.95.

URL http://www.oreilly.com/catalog/9781565929418;   http://
www.oreilly.com/catalog/orsqlpluspr.

Gennick:2001:OSD

[647] Jonathan Gennick and Sanjay Mishra. Oracle SQL*Loader: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 1-56592-948-9. xvii + 251 pp. LCCN QA76.73.S67 G44 2001. US$34.95.

URL http://www.oreilly.com/catalog/9781565929487;   http://
www.oreilly.com/catalog/orsqlloader.

Gennick:2002:OSP

[648] Jonathan Gennick. Oracle SQL*Plus Pocket Reference: a Guide to SQL*Plus Syntax. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00441-9. vi + 113 pp. LCCN QA76.73.S675 G462

2002. US$12.95. URL http://safari.oreilly.com/0596004419;
http://www.oreilly.com/catalog/9780596004415;   http://www.
oreilly.com/catalog/orsqlpluspr2.

Gennick:2003:ORE

[649] Jonathan Gennick and Peter Linsley. Oracle regular expressions: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2003. ISBN 0-596-00601-2. iv + 60 pp. LCCN QA76.9.T48 G46 2003. URL http://www.oreilly.com/

catalog/9780596006013.

Gennick:2004:SPG

[650] Jonathan Gennick. SQL pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00512-1. vi + 154 pp. LCCN QA76.73.S67 G47 2004.

URL http://www.oreilly.com/catalog/9780596005122.

Gennick:2005:OSD

[651] Jonathan Gennick. Oracle SQL*Plus: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00746-9. xxvi + 554 pp. LCCN QA76.73.S67 G46 2005eb; QA76.73.S67; QA76.73.S67

G46 2005. URL http://www.oreilly.com/catalog/9780596007461.

Gennick:2005:OSP

[652] Jonathan Gennick. Oracle SQL*Plus: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00885-6. 146 pp. LCCN QA76.9.D3

G455 2000. URL http://www.oreilly.com/catalog/9780596008857.

Gennick:2006:SPG

[653] Jonathan Gennick. SQL pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52688-1. viii + 184

pp. LCCN QA76.73.S67. URL http://www.oreilly.com/catalog/
9780596526887.

Genzberger:1994:KB

[654] Christine Genzberger et al. Korea Business. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1994. ISBN 0-9631864-4-2. 332 pp. LCCN HC467.K6198583 1994.

US$24.95. URL http://www.ora.com/catalog/koreabiz/;   http:/
/www.oreilly.com/catalog/9780963186447.

George:1998:TTJ

[655] Donald W. George and Amy G. Carlson. Travelers’ Tales Japan. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-885211-04-X. xxi + 411 pp. LCCN

DS812 .J36 1999. US$17.95. URL http://www.oreilly.com/catalog/
9781885211040;   http://www.oreilly.com/catalog/japan.

George:2011:HDG

[656] Lars George. HBase: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9610-0 (paper), 1-4493-1577-1 (e-book). xxviii + 522 pp. LCCN QA76.9.D5 G46 2011. URL http://proquest.

safaribooksonline.com/9781449314682.

Gerantabee:2007:DLD

[657] Fred Gerantabee. Dynamic learning: Dreamweaver CS3: with video tutorials and lesson files. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51057-8. xv + 400 pp. LCCN TK5105.8885.D74. URL http://www.oreilly.com/

catalog/9780596510572.

Getz:2002:AC

[658] Ken Getz, Paul Litwin, and Andy Baron. Access Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00084-7. xxvii + 687 pp. LCCN QA76.9.D3 G4828 2002. US$49.95. URL http://safari.oreilly.com/

0596000847;   http://www.oreilly.com/catalog/9780596000844;
 http://www.oreilly.com/catalog/accesscook.

Getz:2004:AC

[659] Ken Getz, Paul Litwin, and Andy Baron. Access cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00678-0. xxviii + 810 pp. LCCN

QA76.9.D3 G4828 2002. URL http://www.oreilly.com/catalog/
9780596006785.

Gibas:2001:DBC

[660] Cynthia Gibas and Per Jambeck. Developing Bioinformatics Computer Skills. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 1-56592-664-1. xv + 427 pp. LCCN QH324.2 .G53 2001. US$34.95.

URL http://www.oreilly.com/catalog/9781565926646;   http://
www.oreilly.com/catalog/bioskills.

Gibson:2006:GMH

[661] Rich Gibson and Schuyler Erle. Google maps hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10161-9 (paperback). xxv + 337 pp. LCCN ZA4234.G64 G52 2006; G70.212 .G53 2006; ; GA139 .G53 2006.

URL http://www.oreilly.com/catalog/9780596101619.

Gift:2008:PUL

[662] Noah Gift and Jeremy M. Jones. Python for Unix and Linux system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51582-0. xix + 433 pp. LCCN QA76.73.P98 G54 2008.

Gillespie:2016:ERP

[663] Colin (Colin Stevenson) Gillespie and Robin Lovelace. Efficient R programming: a practical guide to smarter programming. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2016. ISBN 1-4919-5078-1 (paperback), 1-4919-5072-2 (e-book), 1-4919-5075-7 (e-book), 1-4919-5073-0 (e-book). xiii + 203 pp. LCCN

QA276.45.R3 G55 2016. URL http://proquest.safaribooksonline.
com/9781491950777.

Gillmor:2004:WMG

[664] Dan Gillmor. We the media: grassroots journalism by the people, for the people. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00733-7. xviii + 299 pp. LCCN PN4784.O62 G55 2004. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/;   http://www.oreilly.com/catalog/
9780596007331;   http://www.oreilly.com/catalog/wemedia/book/

.

Gilly:1990:XWS

[665] Daniel Gilly and Tim O’Reilly. The X Window System in a Nutshell: Desktop Quick Reference for Version 11 Release 4 of the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-24-2. xi + 367 pp. LCCN QA76.76.W56 X84x. URL http://www.oreilly.com/

catalog/9780937175248.

Gilly:1992:UN

[666] Daniel Gilly and the staff of O’Reilly and Inc. Associates. UNIX in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1992. ISBN 1-56592-001-5. xv + 400 pp. LCCN QA76.76.O63 G55 1992. US$9.95.

Gilly:1992:UNB

[667] Daniel Gilly and The staff of O’Reilly & Associates, Inc. UNIX in a Nutshell. Berkeley edition—a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 0-937175-20-X. 272 pp. LCCN QA76.76.O63 U544 1989. US$19.50. URL http://www.oreilly.com/

catalog/9780937175200.

Gilly:1992:UNS

[668] Daniel Gilly and The staff of O’Reilly & Associates. UNIX in a nutshell: System V edition: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 1992. ISBN 1-56592-001-5. 444 pp. LCCN QA76.76.O63 G55 1992; QA76.76.O63

U5453 1992. US$9.95. URL http://www.oreilly.com/catalog/
9781565920019;   http://www.oreilly.com/catalog/unutv.

Gilly:1992:XWS

[669] Daniel Gilly and Tim O’Reilly. The X Window System in a Nutshell. The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1992. ISBN 1-56592-017-1. xi + 409 pp. LCCN QA76.76.W56 X24 1992. US$19.95. URL http://www.oreilly.com/ catalog/9781565920170. Updated for X11 R5 by Ellie Cutler.

Gilly:1998:UN

[670] Daniel Gilly et al. UNIX in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, deluxe edition, July 1998. ISBN 1-56592-406-1. 444 (est.) pp. LCCN ????

US$69.95. URL http://www.oreilly.com/catalog/9781565924062;
 http://www.oreilly.com/catalog/unixcd.

Gilly:1998:UND

[671] Daniel Gilly. UNIX in a nutshell: a desktop quick reference for System V Release 4 and Solaris 2.0. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-001-5. LCCN QA76.76.O63 G565 1998.

URL http://www.oreilly.com/catalog/9781565920019.

Gioseffi:1996:PIY

[672] Claudia Gioseffi, editor. Passport Italy: Your Pocket Guide to Italian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1996. ISBN 1-885073-34-8. 96 pp. LCCN HD58.7.G57 1996.

US$6.95. URL http://www.oreilly.com/catalog/9781885073341.

Gircys:1988:UUC

[673] Gintaras R. Gircys. Understanding and Using COFF. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1988. ISBN 0-937175-31-5. 196 pp. LCCN QA76.8.U65 G57 1988. US$21.95. URL http://www.oreilly.

com/catalog/9780937175316.

Glenn:2000:WNP

[674] Walter J. Glenn. Word 2000 in a Nutshell: a power user’s quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-489-4. xvi + 491 pp. LCCN Z52.5.M52 G57 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565924895;   http://
www.oreilly.com/catalog/word2000ian.

Glenn:2002:WPG

[675] Walter Glenn. Word Pocket Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00445-1. viii + 143 pp. LCCN Z52.5.M52 G573 2002.

US$12.95. URL http://www.oreilly.com/catalog/9780596004453;
 http://www.oreilly.com/catalog/wordpg.

Glenn:2003:OPG

[676] Walter J. Glenn. Outlook pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00444-3. vi + 186 pp. LCCN HD69.T54 G584 2003.

URL http://www.oreilly.com/catalog/9780596004446.

Glenn:2004:WPG

[677] Walter J. Glenn. Word pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004.

ISBN 0-596-00684-5 (paperback). vi + 170 pp. LCCN Z52.5.M52 G58

2004. URL http://www.oreilly.com/catalog/9780596006846.

Glickstein:1997:WGE

[678] Bob Glickstein. Writing GNU Emacs Extensions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-261-1. xviii + 215 pp. LCCN QA76.76.T49G56 1997.

Glover:2005:RE

[679] William R. Glover and Himanshu Bhatt. RFID essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00944-5. 272 pp. LCCN TK6553.

URL http://www.oreilly.com/catalog/9780596009441.

Glover:2006:RE

[680] Bill Glover and Himåaòmśu Bhaòtòta. RFID essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00944-5 (paperback). xiii + 260 pp. LCCN TK6553 .G56 2006. URL http://www.oreilly.com/

catalog/9780596009441.

Goldstein:2005:AMM

[681] Adam Goldstein. AppleScript: the missing manual. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00850-3 (paperback). xxii + 325 pp. LCCN QA76.73.A67 G65 2005. URL http://www.oreilly.com/

catalog/9780596008505.

Goodliffe:2006:CCP

[682] Pete Goodliffe. Code craft: the practice of writing excellent code. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-119-0. xxxvi + 580 pp. LCCN QA76.6 .G656 2007; QA76.6 .G656 2006eb. URL

http://www.oreilly.com/catalog/9781593271190.

Goodman:1998:DHD

[683] Danny Goodman. Dynamic HTML: The Definitive Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1998. ISBN 1-56592-494-0. xii + 1073 pp. LCCN QA76.76.H94 G657 1998. US$39.95.

URL http://www.oreilly.com/catalog/9781565924949;   http://
www.oreilly.com/catalog/dhtmlref.

Goodman:2002:DHD

[684] Danny Goodman. Dynamic HTML: The Definitive Reference: a Comprehensive Resource for HTML, CSS, DOM & JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00316-

1. xiii + 1401 pp. LCCN QA76.76.H94 G657 2002. US$59.95.

URL http://www.oreilly.com/catalog/9780596003166;   http://
www.oreilly.com/catalog/dhtmlref2.

Goodman:2003:JDC

[685] Danny Goodman. JavaScript and DHTML cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00467-2. xvii + 520 pp. LCCN QA76.73.J39 G63 2003. URL http://www.oreilly.com/

catalog/9780596004675.

Goodman:2007:DHD

[686] Danny Goodman. Dynamic HTML: the definitive reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-52740-3 (paperback). xi + 1307 pp. LCCN QA76.76.H94; QA76.76.H94 G657 2007eb. URL http://www.loc.gov/

catdir/enhancements/fy0715/2007273943-d.html;   http://www.
loc.gov/catdir/toc/fy0710/2007273943.html;   http://www.oreilly.
com/catalog/9780596527402.

Goodman:2007:JDC

[687] Danny Goodman. JavaScript and DHTML cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51408-5 (paperback). xx + 582 pp. LCCN QA76.73.J39 G63 2007eb; QA76.73.J39

G63 2007; QA76.73.J39. URL http://www.oreilly.com/catalog/
9780596514082.

Gough:1999:KSS

[688] Laurie Gough. Kite Strings of the Southern Cross: a Woman’s Travel Odyssey. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-30-9. 400 (est.) pp. LCCN G155.C3G68 1999. US$24.00.

URL http://www.oreilly.com/catalog/9781885211309;   http://
www.oreilly.com/catalog/kite/.

Gourley:2002:HDG

[689] David Gourley, Brian Totty, Marjorie Sayer, Sailu Reddy, and Anshu Aggarwal. HTTP: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-509-2. xviii + 635 pp. LCCN QA76.76.H94

G695 2002. US$44.95. URL http://www.oreilly.com/catalog/
9781565925090;   http://www.oreilly.com/catalog/httptdg.

Goyvaerts:2012:REC

[690] Jan Goyvaerts and Steven Levithan. Regular expressions cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2012. ISBN 1-4493-1943-2 (paperback). xiv + 594 pp. LCCN NLS PB8.212.757/6.

Graff:2003:SCP

[691] Mark Graff and Kenneth R. Van Wyk. Secure coding: principles and practices. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00242-4. xx + 202 pp. LCCN QA76.9.A25 G69 2003. URL http://www.

oreilly.com/catalog/9780596002428.

Graham:2004:HPB

[692] Paul Graham. Hackers and painters: big ideas from the computer age. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00662-4. xi + 258 pp. LCCN HD8039.D37; HD8039.D37 G73 2004eb; HD8039.D37 G73

2004. URL http://www.oreilly.com/catalog/9780596006624.

Gralla:2003:WXH

[693] Preston Gralla. Windows XP hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00511-3. xviii + 392 pp. LCCN QA76.76.O63 G7225 2003. URL http://www.

oreilly.com/catalog/9780596005115.

Gralla:2005:IAH

[694] Preston Gralla. Internet annoyances: how to fix the most annoying things about going online. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00735-3 (paperback). xiii + 239 pp. LCCN TK5105.875.I57 G723 2005. URL http://

www.oreilly.com/catalog/9780596007355.

Gralla:2005:PPC

[695] Preston Gralla. PC pest control: protect your computers from malicious Internet invaders. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00926-7. vii + 275 pp. LCCN QA76.9.A25 G73 2005. URL http:

//www.oreilly.com/catalog/9780596009267.

Gralla:2005:WXH

[696] Preston Gralla. Windows XP hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00918-6. xxi + 548 pp. LCCN QA76.76.O63 G7225 2005.

URL http://www.oreilly.com/catalog/9780596009182.

Gralla:2005:WXP

[697] Preston Gralla. Windows XP power hound. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00619-5 (paperback). xxi + 359 pp. LCCN QA76.76.O63; QA76.76.O63 G73 2005; QA76.76.O63 G73 2005eb; QA76.76.O63 G716 2004;

QA76.76.O63 G72285 2005. URL http://www.oreilly.com/catalog/
9780596006198.

Gralla:2006:WVN

[698] Preston Gralla. Windows Vista in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52707-1. xvi + 732 pp. LCCN QA76.76.O63; QA76.76.O63 G7 2006eb. URL http://www.oreilly.

com/catalog/9780596527075.

Gralla:2007:BBW

[699] Preston Gralla. Big book of Windows hacks. Hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52835-3. 647 pp. LCCN QA76.76.O63; QA76.76.O63 G72237 2007eb. URL http://www.

oreilly.com/catalog/9780596528355.

Gralla:2007:WVP

[700] Preston Gralla. Windows Vista pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52808-6. x + 182 pp. LCCN QA76.76.O63; QA76.76.O63 G73 2007; QA76.76.O63 G73 2007eb. URL http://www.

oreilly.com/catalog/9780596528089.

Grand:1997:JFC

[701] Mark Grand and Jonathan Knudsen. Java Fundamental Classes Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1997. ISBN 1-56592-241-7. xxi + 1090 pp. LCCN QA76.73.J38 G728 1997. US$44.95.

URL http://www.oreilly.com/catalog/9781565922419;   http://
www.oreilly.com/catalog/javafund.

Grand:1997:JLRa

[702] Mark Grand. Java Language Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1997. ISBN 1-56592-204-2. xiv + 448 pp. LCCN QA76.73.J38 G73 1997. US$24.95. URL http://www.oreilly.com/

catalog/9781565922044.

Grand:1997:JLRb

[703] Mark Grand. Java Language Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-326-X. xiv + 475 pp. LCCN QA76.73.J38 G73 1997b. US$32.95. URL http:

//www.ora.com/catalog/javalang2/;   http://www.oreilly.com/
catalog/9781565923263;   http://www.oreilly.com/catalog/javalang2.

Granneman:2005:DCB

[704] Scott Granneman. Don’t click on the blue e!: switching to Firefox. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00939-9 (paperback). xiv + 267 pp. LCCN TK5105.883.F55 G73 2005. URL http://

www.oreilly.com/catalog/9780596009397.

Grant:2004:LNG

[705] Rickford Grant. Linux for non-geeks: a hands-on, project-based, takeit-slow guidebook. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-034-8. xx + 335 pp. LCCN QA76.76.O63 G723 2004eb.

URL http://www.loc.gov/catdir/toc/ecip0411/2003024732.html;
 http://www.oreilly.com/catalog/9781593270346.

Grant:2005:LME

[706] Rickford Grant. Linux made easy: the official guide to Xandros 3 for everyday users. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-057-7. xxvi + 463 pp. LCCN QA76.76.O63 G7245 2005; QA76.76.O63 .G7245 2005eb. URL http://www.oreilly.com/

catalog/9781593270575.

Gray:2006:BRQ

[707] James Edward Gray. Best of Ruby Quiz. Pragmatic programmers LLC, Raleigh, NC, USA, 2006. ISBN 0-9766940-7-7. viii + 298 pp.

LCCN QA76.64 .G73 2006. URL http://www.oreilly.com/catalog/
9780976694076.

Green:1999:NEB

[708] Reg Green. The Nicholas Effect: a Boy’s Gift to the World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-597-1. vii + 246 pp. LCCN RD129.5 .G74 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565925977;   http://
www.oreilly.com/catalog/nicholas/.

Greene:2006:HKY

[709] Jennifer Greene and Andrew Stellman. How to keep your boss from sinking your project. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52802-7. LCCN QA76.758; QA76.758 G74 2006eb. URL http://

www.oreilly.com/catalog/9780596528027.

Greene:2007:HFP

[710] Jennifer Greene and Andrew Stellman. Head first PMP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10234-8. xxxiii + 658 pp. LCCN HD69.P75 G735 2007eb; HD69.P75 G735 2007; HD69.P75. URL

http://www.oreilly.com/catalog/9780596102340.

Greenwald:1997:MOP

[711] Rick Greenwald and Robert E. Hoskin. Mastering Oracle Power Objects. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-239-5. xx + 485 pp. LCCN QA76.9.D3 G728

1997. US$39.95. URL http://www.ora.com/catalog/oraclepo/;
http://www.oreilly.com/catalog/9781565922396;   http://www.
oreilly.com/catalog/oraclepo.

Greenwald:1999:OEO

[712] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle Essentials: Oracle8 and Oracle8i. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-708-7. xvi + 355 pp. LCCN QA76.9.D3 G7283 1999; QA76.9.D3 G74 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565927087;   http://
www.oreilly.com/catalog/oressentials.

Greenwald:2001:OEO

[713] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle Essentials: Oracle9 i, Oracle8 i and Oracle8. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 2001. ISBN 0-596-00179-7. xvii + 364 pp. LCCN QA76.9.D3 G72833 2001; QA76.9.D3 G74 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596001797;   http://
www.oreilly.com/catalog/oressentials2.

Greenwald:2002:ON

[714] Rick Greenwald and David C. Kreines. Oracle in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00336-6. xvii + 906 pp. LCCN QA76.9.D3 G74 2003. US$49.95.

URL http://www.oreilly.com/catalog/9780596003364;   http://
www.oreilly.com/catalog/oracleian.

Greenwald:2004:OAS

[715] Rick Greenwald, Robert Stackowiak, and Donald Bales. Oracle Application Server 10g essentials. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00621-7 (paperback). xvii + 262 pp. LCCN QA76.9.D3 G7282 2004. URL

http://www.oreilly.com/catalog/9780596006211.

Greenwald:2004:OEO

[716] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle essentials: Oracle database 10g. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00585-7. xvii + 374 pp. LCCN

QA76.9.D3 G72834 2004. URL http://www.oreilly.com/catalog/
9780596005856.

Greenwald:2007:OE

[717] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2007. ISBN 0-596-51454-9. xviii + 386 pp. LCCN QA76.9.D3. URL http://www.

oreilly.com/catalog/9780596514549.

Gregory:2007:SVE

[718] Peter H. Gregory. Securing the Vista environment. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51430-1. LCCN QA76.9.A25;

QA76.9.A25 .G74 2007eb. URL http://www.oreilly.com/catalog/
9780596514303.

Griffiths:2003:MVS

[719] Ian Griffiths, Jon Flanders, and Chris Sells. Mastering Visual Studio .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00360-9. xii + 397 pp. LCCN QA76.76.M52 G73 2003. US$39.95.

URL http://www.oreilly.com/catalog/9780596003609;   http://
www.oreilly.com/catalog/mastvsnet.

Griffiths:2003:MXH

[720] Rob Griffiths. Mac OS X Hints: Jaguar Edition: The 500 Most Amazing Power Tips. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00451-6. xvi + 461 pp. LCCN QA76.76.O63 G755 2003. US$24.95, CDN$38.95, UK£17.50. URL

http://www.oreilly.com/catalog/9780596004514;   http://www.
oreilly.com/catalog/macxhints/.

Griffiths:2003:NWF

[721] Ian Griffiths and Matthew Adams. .NET Windows Forms in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00338-2. xi + 879 pp. LCCN QA76.76.D47 G775 2003. US$49.95, CDN$77.95, UK£35.50.

URL http://www.oreilly.com/catalog/9780596003388;   http://
www.oreilly.com/catalog/netwinformian. Contains one CD-ROM.

Griffiths:2004:MXP

[722] Rob Griffiths. Mac OS X power hound: Panther edition. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2004. ISBN 0-596-00818-X. xxxv + 536 pp. LCCN QA76.76.O63 G754 2004; QA76.76.O63 G75 2004; QA76.76 O63; QA76.76 G754 2004eb; QA76.76; QA76.76 G754 2004. URL http://www.oreilly.com/

catalog/9780596008185.

Griffiths:2009:HFS

[723] Dawn Griffiths. Head first statistics. Head first series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-52758-6 (paperback). xxxviii + 677 pp. LCCN QA276.12 .G75 2009. URL http://www.loc.gov/

catdir/enhancements/fy0913/2009278026-b.html;   http://www.
loc.gov/catdir/enhancements/fy0915/2009278026-d.html.

Griffiths:2010:PC

[724] Ian Griffiths, Matthew Adams, and Jesse Liberty. Programming C# 4.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, sixth edition, 2010. ISBN 0-596-15983-8. xxi + 830 pp. LCCN ????

Griffiths:2012:PC

[725] Ian Griffiths. Programming C# 5.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2041-4. 886 pp. LCCN ????

Griffiths:2012:PCB

[726] Ian Griffiths. Programming C# 5.0: Building Windows 8 Metro, Web, and Desktop Applications for the .NET 4.5 Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2041-4 (paperback). 800 pp. LCCN ???? EUR 45.00, UK £42.50.

Grimes:2001:MMC

[727] Roger A. Grimes. Malicious Mobile Code: Virus Protection for Windows. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-682-X. xviii + 522 pp. LCCN QA76.76.C68 G75 2001. US$39.95.

URL http://www.oreilly.com/catalog/9781565926820;   http://
www.oreilly.com/catalog/malmobcode.

Grolemund:2014:HPR

[728] Garrett Grolemund. Hands-on programming with R. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 1-4493-5901-9 (paperback). xiii + 230 pp. LCCN QA276.45.R3 G76 2014.

Grosso:2002:JR

[729] William Grosso. Java RMI. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-452-5. xxiii + 545 pp. LCCN QA76.73.J38 G76 2002. Designing and building distributed applications.

Grosso:2002:JRD

[730] William Grosso. Java RMI: Designing and building distributed applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-452-5. xxiii + 545 pp. LCCN QA76.73.J38 G78 2002. URL http://www.oreilly. com/catalog/9781565924529. Designing and building distributed applications.

Grover:2007:OMM

[731] Chris Grover, Matthew MacDonald, and Emily A. Vander Veer. Office 2007: the missing manual. The missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51422-0 (paperback), 0-596-51422-0. xxiv + 853 pp. LCCN HF5548.4.M525 G76 2007; HF5548.4.M525 G76 2007eb; HF5548.4.M525.

URL http://www.oreilly.com/catalog/9780596514228.

Grover:2007:W

[732] Chris Grover. Word 2007. The Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52739-X. xvii + 502 pp. LCCN Z52.5.M52 G752 2007eb; Z52.5.M52 G752 2007; Z52.5.M52. URL http://www.

oreilly.com/catalog/9780596527396.

Grover:2007:WSM

[733] Chris Grover. Word 2007 for starters: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52830-2. xiv + 338 pp. LCCN Z52.5.M52 G76 2007eb; Z52.5.M52. URL http://www.

oreilly.com/catalog/9780596528300.

Grundgeiger:2000:CMP

[734] Dave Grundgeiger. CDO and MAPI Programming with Visual Basic: Developing Mail and Messaging Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-665-X. xvi + 361 pp. LCCN QA76.73.B3 G775 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565926653;   http://
www.oreilly.com/catalog/cdomapi.

Grundgeiger:2002:PVB

[735] Dave Grundgeiger. Programming Visual Basic .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00093-6. xii + 446 pp. LCCN QA76.73.B3 G776 2002. US$39.95. URL http://safari.oreilly.com/

0596000936;   http://www.oreilly.com/catalog/9780596000936;
 http://www.oreilly.com/catalog/progvbdotnet.

Grzeskowiak:1996:PHK

[736] Andrew Grzeskowiak. Passport Hong Kong: Your Pocket Guide to Hong Kong Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-885073-31-3. 96 pp. LCCN HD58.7.G79 1996.

US$6.95. URL http://www.oreilly.com/catalog/9781885073310.

Guelich:2000:CPP

[737] Scott Guelich, Shishir Gundavaram, and Gunther Birznieks. CGI Programming with Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 2000. ISBN 1-56592-419-3. xv + 451 pp. LCCN QA76.73.P22 G84 2000.

US$34.95. URL http://www.oreilly.com/catalog/9781565924192;
 http://www.oreilly.com/catalog/cgi2.

Gundavaram:1996:CPW

[738] Shishir Gundavaram. CGI Programming on the World Wide Web. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-168-2. xiv + 433 pp. LCCN TK5105.565 .G85 1996; TK5105.888 .G86 1996. US$29.95, CDN$42.95. URL http:

//www.ora.com/gnn/bus/ora/item/cgi_prog.html;   http://www.
oreilly.com/catalog/9781565921689;   http://www.oreilly.com/
catalog/cgi.

Gunther:2007:BKG

[739] Karsten Günther. Bash: kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 3-89721-533-0. 144 pp. LCCN ???? EUR 9.90.

Gurry:1996:OPT

[740] Mark Gurry and Peter Corrigan. Oracle Performance Tuning. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, November 1996. ISBN 1-56592-237-9. xxii + 940 pp. LCCN QA76.9.D3 G873 1996.

US$44.95. URL http://www.oreilly.com/catalog/9781565922372;
 http://www.oreilly.com/catalog/oracle2.

Gurry:2002:OST

[741] Mark Gurry. Oracle SQL Tuning: Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00268-8. vi + 102 pp. LCCN QA76.73.S67 G873 2002. US$12.95. URL

http://safari.oreilly.com/0596002688;   http://www.oreilly.
com/catalog/9780596002688;   http://www.oreilly.com/catalog/
orsqltunpr.

Habegger:1998:GTB

[742] Larry Habegger, James O’Reilly, and Sean O’Reilly, editors. The Gift of Travel: The Best of Travelers’ Tales. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-885211-25-2. 250 (est.) pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9781885211255;   http://
www.oreilly.com/catalog/gift.

Habegger:1999:ATS

[743] Larry Habegger and Amy Greimann Carlson, editors. Australia: True Stories of Life Down Under. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-40-6. 375 (est.) pp. LCCN DU105.2.A937 1999.

US$17.95. URL http://www.oreilly.com/catalog/9781885211408;
 http://www.oreilly.com/catalog/australia.

Habegger:1999:GBT

[744] Larry Habegger and Amy Greimann Carlson, editors. The Gift of Birds: True Encounters with Avian Spirits. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-41-4. 352 pp. LCCN QL677.5.G54 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211415;   http://
www.oreilly.com/catalog/birds.

Hacker:2000:MDG

[745] Scot Hacker. MP3: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2000. ISBN 1-56592-661-7. x + 388 pp. LCCN ML74.4.M6 H33 2000. US$29.95. URL http://www.oreilly.com/

catalog/9781565926615;   http://www.oreilly.com/catalog/mp3.

Haddad:1997:TTB

[746] Annette Haddad, Scott Doggett, and Alex Shoumatoff. Travelers’ Tales Brazil. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-11-2. xix + 437 pp. LCCN PS366.T73 B73 1997. US$17.95. URL http://www.ora.

com/catalog/brazil/noframes.html;   http://www.oreilly.com/
catalog/9781885211118;   http://www.oreilly.com/catalog/brazil.

Hagen:2002:IE

[747] Silvia Hagen. IPv6 Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00125-8. xv + 338 pp. LCCN TK5105.585 .H3397

2002. US$39.95. URL http://safari.oreilly.com/0596001258;
http://www.oreilly.com/catalog/9780596001254;   http://www.
oreilly.com/catalog/ipv6ess.

Hagen:2006:IE

[748] Silvia Hagen. IPv6 essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10058-2. 436 pp. LCCN TK5105.585 .H3397 2006eb; TK5105.585 .H3397 2006; TK5105.585. URL http://

www.oreilly.com/catalog/9780596100582.

Haldar:2007:IS

[749] Sibsankar Haldar. Inside SQLite. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-55006-5. LCCN QA76.73.S67

H35 2007eb; QA76.73.S67. URL http://www.oreilly.com/catalog/
9780596550066.

Hale-Evans:2006:MPH

[750] Ron Hale-Evans. Mind performance hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10153-8. xx + 308 pp. LCCN

QP376; QP376 .H26 2006eb. URL http://www.oreilly.com/catalog/
9780596101534.

Hall:2000:ICP

[751] Eric A. Hall. Internet Core Protocols: The Definitive Guide: Help for Network Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2000. ISBN 1-56592-572-6. xx + 449 pp. LCCN TK5105.875.I57

H355 2000. US$39.95. URL http://www.oreilly.com/catalog/
9781565925724;   http://www.oreilly.com/catalog/coreprot. In-

cludes CD-ROM.

Halloway:2007:RJD

[752] Stuart Dabbs Halloway and Justin Gehtland. Rails for Java developers. Pragmatic programmers; Facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 0-9776166-9-X. xix + 311 pp. LCCN

QA76.73.R83 H35 2007. URL http://www.loc.gov/catdir/toc/
fy0709/2007273022.html;   http://www.oreilly.com/catalog/9780977616695.

Halloway:2009:PC

[753] Stuart Dabbs Halloway. Programming Clojure. The pragmatic programmer. Pragmatic Bookshelf, Raleigh, NC, USA, 2009. ISBN 1-934356-33-6 (paperback). xxi + 280 pp. LCCN QA76.62 .H35 2009.

Hamill:2004:UTF

[754] Paul Hamill. Unit test frameworks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00689-6. xii + 198 pp. LCCN QA76.76.T48 H356 2004. URL http://

www.oreilly.com/catalog/9780596006891.

Hamilton:2000:VBS

[755] J. P. Hamilton. Visual Basic Shell Programming: Integrating Applications with the Windows Shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2000. ISBN 1-56592-670-6. xix + 373 pp. LCCN QA76.73.B3 H339 2000.

US$29.95. URL http://www.oreilly.com/catalog/9781565926707;
 http://www.oreilly.com/catalog/vbshell.

Hamilton:2003:ANC

[756] Bill Hamilton. ADO.NET cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00439-7. xvii + 605 pp. LCCN QA76.9.D26 H358

2003. URL http://www.oreilly.com/catalog/9780596004392.

Hamilton:2003:ANN

[757] Bill Hamilton and Matthew MacDonald. ADO.NET in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00361-7. xviii + 599 pp. LCCN QA76.76.O63 H34463 2003. US$44.95,

CDN$69.95, UK£31.95. URL http://www.oreilly.com/catalog/
9780596003616;   http://www.oreilly.com/catalog/adonetian/.

Hamilton:2003:OOP

[758] J. P. Hamilton. Object-Oriented Programming with Visual Basic .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00146-0. xiii + 291 pp. LCCN QA76.64 .H3565

2003. US$34.95. URL http://safari.oreilly.com/0596001460;
http://www.oreilly.com/catalog/9780596001469;   http://www.
oreilly.com/catalog/objectvbnet.

Hamilton:2004:NPR

[759] Bill Hamilton. NUnit pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00739-6. vi + 88 pp. LCCN QA76.76.T48 H36 2004. URL http:

//www.oreilly.com/catalog/9780596007393.

Hamilton:2005:SSR

[760] Bill Hamilton. SQL Server 2005 reporting essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-52941-4. LCCN QA76.9.D3; QA76.9.D3

H36 2006eb. URL http://www.oreilly.com/catalog/9780596529413.

Hamilton:2006:SSR

[761] Bill Hamilton. SQL Server 2005 reporting essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52941-4. LCCN ????

Hamilton:2007:SSI

[762] Bill Hamilton. SQL server integration services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51375-5. LCCN QA76.9.D3; QA76.9.D3 H36

2007eb. URL http://www.oreilly.com/catalog/9780596513757.

Hamilton:2008:ANC

[763] Bill Hamilton. ADO.NET 3.5 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-10140-6. xxii + 955 pp. LCCN QA76.9.D26 H3583 2008.

Hamlin:1998:DA

[764] J. Scott Hamlin. Designing with Animation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-441-X. 250 (est.) pp. LCCN ???? URL

http://www.oreilly.com/catalog/9781565924413.

Hammersley:2003:CSR

[765] Ben Hammersley. Content syndication with RSS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00383-8. xi + 208 pp. LCCN QA76.76.H94

H33 2003. URL http://www.oreilly.com/catalog/9780596003838.

Hammersley:2005:DFR

[766] Ben Hammersley. Developing feeds with RSS and Atom. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00881-3. xiii + 253 pp. LCCN QA76.625

.H355 2005. URL http://www.oreilly.com/catalog/9780596008819.

Hammond:2000:PPW

[767] Mark (Mark J.) Hammond and Andy Robinson. Python Programming on Win32: Help for Windows Programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-621-8. xvii + 652 pp. LCCN QA76.73.P98 H36 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565926219;   http://
www.oreilly.com/catalog/pythonwin32.

Hampton:2004:XPA

[768] Kip Hampton. XML Publishing with AxKit. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00216-5. x + 204 pp. LCCN QA76.76.H94 H334 2004. URL http:

//www.oreilly.com/catalog/9780596002169.

Haralambous:2004:FCG

[769] Yannis Haralambous. Fontes & codages: Glyphes et caractères à l’ère du numérique. O’Reilly, Paris, France, 2004. ISBN 2-84177-273-X.

xx + 990 pp. LCCN ???? URL http://www.oreilly.com/catalog/
9782841772735.

Haralambous:2007:FE

[770] Yannis Haralambous and P. Scott Horne. Fonts and encodings. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-10242-9. xx + 1016 pp. LCCN Z250. URL

http://www.oreilly.com/catalog/9780596102425.

Hardt:1998:PRK

[771] Dick Hardt, Erik Olson, David Futato, and Brian Jepson. Perl Resource Kit – Win32 Edition. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1998. ISBN 1-56592-409-6. 1500 (est.) pp. LCCN QA76.73.P22 P48 1998. US$149.95.

URL http://www.oreilly.com/catalog/9781565924093;   http://
www.oreilly.com/catalog/prkwin32. Includes four books and CD-

ROM.

Hardwidge:2006:BEP

[772] Ben Hardwidge. Building extreme PCs: the complete guide to computer modding. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10136-8, 0-596-10136-8. 191 pp. LCCN TK7887 .H34 2006eb; TK7887 .H34 2006;

TK7887. URL http://www.oreilly.com/catalog/9780596101367.

Harold:1997:JNP

[773] Elliotte Rusty Harold. Java Network Programming Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-227-1. xvii + 422 pp. LCCN QA76.73.J38 H37 1997. US$34.95.

URL http://www.oreilly.com/catalog/9781565922273;   http://
www.oreilly.com/catalog/javanetwk.

Harold:1997:PRA

[774] Elliotte Rusty Harold. Programmation Réseau avec Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1997. ISBN 2-84177-034-6. LCCN ????

220 FF. URL http://www.editions-oreilly.fr/international/
france/prog/javanet.html.

Harold:1999:J

[775] Elliotte Rusty Harold. Java I/O. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-485-1. xxvi + 568 pp. LCCN QA76.73.J38 H372 1999.

US$32.95. URL http://metalab.unc.edu/javafaq/books/javaio/;
 http://www.oreilly.com/catalog/javaio/.

Harold:1999:JIO

[776] Elliotte Rusty Harold. Java I/O. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-485-1. xxvi + 568 pp. LCCN QA76.73.J38

H372 1999. US$32.95. URL http://www.oreilly.com/catalog/
9781565924857;   http://www.oreilly.com/catalog/javaio/.

Harold:19xx:JNP

[777] Elliotte Rusty Harold. Java Network Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-56-7. LCCN ????

4,300 yen. URL http://www.ohmsha.co.jp/data/books/contents/4-

900900-56-7.htm. Japanese translation.

Harold:2000:JNP

[778] Elliotte Rusty Harold. Java Network Programming. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2000. ISBN 1-56592-870-9. xxvi + 731 pp. LCCN QA76.625 .H367 2000; QA76.73.J38 H37 2000.

US$39.95. URL http://www.oreilly.com/catalog/9781565928701;
 http://www.oreilly.com/catalog/javanp2.

Harold:2001:XND

[779] Elliotte Rusty Harold and W. Scott Means. XML in a Nutshell: a Desktop Quick Reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2001. ISBN 0-596-00058-8. xvi + 480 pp. LCCN QA76.76.H94

H36 2001. US$29.95. URL http://www.oreilly.com/catalog/
9780596000585;   http://www.oreilly.com/catalog/xmlnut.

Harold:2002:XCB

[780] Elliotte Rusty Harold, editor. The XML CD bookshelf: XML in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 1.0. edition, 2002. ISBN 0-596-00335-8. 600 (est.) pp. LCCN QA76.76. US$119.95.

URL http://www.oreilly.com/catalog/9780596003357;   http://
www.oreilly.com/catalog/xmlcdbksf. One CD-ROM.

Harold:2002:XN

[781] Elliotte Rusty Harold and W. Scott Means. XML in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00292-0. xvii + 613 pp. LCCN QA76.76.H94 H36 2002. URL

http://www.oreilly.com/catalog/9780596002923.

Harold:2004:JNP

[782] Elliotte Rusty Harold. Java network programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00721-3. xxii + 735 pp. LCCN QA76.625; QA76.625 .H367 2004eb; QA76.625 .H367

2004; QA76.625 .H367 2005. URL http://www.oreilly.com/catalog/
9780596007218.

Harold:2004:XND

[783] Elliotte Rusty Harold and W. Scott Means. XML in a nutshell: [a desktop quick reference]. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00764-7. xix + 689 pp. LCCN QA76.76.H94; Internet; QA76.76.H94 H36 2004eb; QA76.76.H94 H36 2004. URL http:

//www.oreilly.com/catalog/9780596007645.

Harold:2006:J

[784] Elliotte Rusty Harold. Java I/O. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52750-0. xxiv + 700 pp. LCCN QA76.76.J38 H37 2006. US$49.99.

URL http://www.oreilly.com/catalog/9780596527501;   http://
www.oreilly.com/catalog/javaio2/.

Harper:1996:PAY

[785] Tim Harper. Passport Argentina: Your Pocket Guide to Argentine Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-885073-21-6. 96 pp. LCCN HF5389.3.A7 C36 1999.

US$6.95. URL http://www.oreilly.com/catalog/9781885073211.

Harper:1996:PUK

[786] Tim Harper. Passport United Kingdom: Your Pocket Guide to British Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-885073-28-3. 96 pp. LCCN HD58.7.H36925 1996.

US$6.95. URL http://www.oreilly.com/catalog/9781885073280.

Harris:1997:RG

[787] John Harris. Recording the guitar. PC Publishing, Tonbridge, Kent, UK, 1997. ISBN 1-870775-45-7 (paperback). 156 pp. LCCN MT580.H27 R43.

URL http://www.oreilly.com/catalog/9781870775458.

Harrison:1994:UH

[788] Mark Harrison. The Usenet Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1994. ISBN 1-56592-101-1. 250 pp. US$24.95. A user’s guide to netnews.

Harrison:1995:UHU

[789] Mark Harrison. The USENET Handbook: a User’s Guide to Netnews. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1995. ISBN 1-56592-101-1. xiv + 372 pp. LCCN TK5105.875.I57 H534 1995; TK5105.875.U83 H37 1995. US$25.00. URL http://www.oreilly.com/

catalog/9781565921016.

Harrison:1997:TTT

[790] Mark Harrison and John K. Ousterhout. Tcl/Tk Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-218-2. xvi + 653 pp. LCCN QA76.73.T44 H37 1997. US$49.95. URL

http://www.ora.com/catalog/tcltools/;   http://www.oreilly.
com/catalog/9781565922181;   http://www.oreilly.com/catalog/
tcltools.

Harrison:2006:MSP

[791] Guy Harrison and Steven Feuerstein. MySQL stored procedure programming: building high-performance web applications with PHP, Perl, Python, Java & .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10089-2. xxiii + 609 pp. LCCN QA76.73.S67

H377 2006eb; QA76.73.S67. URL http://www.oreilly.com/catalog/
9780596100896.

Hart-Davis:2005:WAH

[792] Guy Hart-Davis. Word annoyances: how to fix the most annoying things about your favorite word processor. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00954-2 (paperback). xiv + 192 pp.

LCCN Z52.5.M52 H37 2005. £912.95. URL http://www.oreilly.com/

catalog/9780596009540.

Hassell:2002:RSP

[793] Jonathan Hassell. RADIUS: Securing Public Access to Private Resources. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00322-6. xiii + 190 pp. LCCN TK5105.55 .H36 2002; TK5105.5825 .H37 2003. US$34.95.

URL http://www.oreilly.com/catalog/9780596003227;   http://
www.oreilly.com/catalog/radius.

Hassell:2004:LWS

[794] Jonathan Hassell. Learning Windows Server 2003. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00624-1 (paperback). xv + 654 pp. LCCN QA76.76.O63 H37 2004. URL http://www.oreilly.com/

catalog/9780596006242.

Hassell:2006:LWS

[795] Jonathan Hassell. Learning Windows Server 2003. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10123-6. xv + 723 pp. LCCN QA76.76.O63; QA76.76.O63 H3745 2006eb; QA76.76.O63 H3745 2006.

URL http://www.oreilly.com/catalog/9780596101237.

Hatton:2004:SDN

[796] Tim Hatton. SWT: a developer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00838-4. xix + 297 pp. LCCN QA76.9.U83 H38 2004. URL http:

//www.oreilly.com/catalog/9780596008383.

Havey:2005:EBP

[797] Michael Havey. Essential business process modeling. Theory in practice. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00843-0. xv + 332

pp. LCCN T58.64 .H37 2005. URL http://www.oreilly.com/catalog/
9780596008437.

Hawley:2004:EH

[798] David E. Hawley and Raina Hawley. Excel hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00625-X (paperback). xviii + 283 pp. LCCN HF5548.4.M523 H39 2004. URL http://www.

oreilly.com/catalog/9780596006259.

Hawley:2007:EH

[799] David E. Hawley and Raina Hawley. Excel hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52834-5 (paperback). xxi + 386 pp. LCCN HF5548.4.M523 H39 2007eb; HF5548.4.M523

H39 2007; HF5548.4.M523. URL http://www.oreilly.com/catalog/
9780596528348.

Hazel:2001:EMT

[800] Philip Hazel. Exim: The Mail Transfer Agent. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2001. ISBN 0-596-00098-7. xviii + 611 pp. LCCN TK5105.73 .H39 2001. US$44.95.

URL http://www.oreilly.com/catalog/9780596000981;   http://
www.oreilly.com/catalog/exim.

Heath:2012:CWA

[801] Robert A. Heath. Creating Windows Applications With HTML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1622-0. ???? pp. LCCN ????

Heineman:2009:AN

[802] George T. Heineman, Stanley Selkow, and Gary Pollice. Algorithms in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-51624-X (paperback). xv + 343 pp. LCCN QA76.9.A43 H45 2009. US$49.99.

Heinle:1997:DJC

[803] Nick Heinle. Designing with JavaScript: creating dynamic Web pages. Web review studio series. Songline Studios, Inc. and O’Reilly & Associates, Inc., Sebastopol, CA, 1997. ISBN 1-56592-300-6. xii + 241 pp. LCCN QA76.73.J39 H45 1997. US$29.95. URL

http://www.ora.com/catalog/designjs/;   http://www.oreilly.
com/catalog/9781565923003;   http://www.oreilly.com/catalog/

designjs. Includes CD-ROM.

Heinle:2002:DJC

[804] Nick Heinle and Bill Pena. Designing with JavaScript: creating dynamic Web pages. O’Reilly Web studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 1-56592-360-X. xi + 216 pp. LCCN QA76.73.J39

H45 2002. URL http://www.oreilly.com/catalog/9781565923607.

Hekman:1997:LND

[805] Jessica Perry Hekman. Linux in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-167-4. xi + 424 pp. LCCN QA76.76.O63 H453 1997. US$19.95.

URL http://www.oreilly.com/catalog/9781565921672;   http://
www.oreilly.com/catalog/linuxnut.

Hekman:1998:LNG

[806] Jessica Perry Hekman, Alain Nadeau, and Jean-Michel Vansteene. Linux in a nutshell: guide de référence. O’Reilly, Paris, France, 1998. ISBN 2-84177-031-1. xi + 424 pp. LCCN ????

Hekmann:1997:LN

[807] Jessica Perry Hekmann and The Staff of O’Reilly and Associates. Linux in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-167-4. xi + 424 pp. LCCN QA76.76.O63 H453 1997. US$19.95.

Heller:1990:XPMa

[808] Dan Heller. XView Programming Manual, volume 7. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-38-2. xxviii + 557 pp. LCCN

QA76.76.W56 D44 v.7 1990. URL http://www.oreilly.com/catalog/
9780937175385.

Heller:1990:XPMb

[809] Dan Heller. XView Programming Manual: for Version 11 of the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-52-8. xxviii + 642 pp. LCCN QA76.76.W56 H355 1990. URL http://

www.oreilly.com/catalog/9780937175521.

Heller:1991:MPM

[810] Dan Heller. Motif Programming Manual for OSF/Motif Version 1.1, volume 6. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-70-6. xl + 990 pp. LCCN QA76.76.W56 H446 1991; QA76.76.W56 H465 1991.

US$39.95. URL http://www.oreilly.com/catalog/9780937175705.

Heller:1991:XPM

[811] Dan Heller and Thomas Van Raalte. XView Programming Manual, volume 7A of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, September 1991. ISBN 0-937175-87-0. xxxvii + 729 pp. LCCN QA76.76.W56 H447 1990. US$34.95.

URL http://www.oreilly.com/catalog/9780937175873. Latest up-

date: August 1993.

Heller:1994:MPM

[812] Dan Heller, Paula M. Ferguson, and David Brennan. Motif Programming Manual, volume 6A. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, February 1994. ISBN 1-56592-016-3. 1016 pp. LCCN QA76.76.W56 H448 1994. US$34.95. URL http://www.oreilly.com/

catalog/9781565920163;   http://www.oreilly.com/catalog/v6a.

Hemenway:2004:SH

[813] Kevin Hemenway and Tara Calishain. Spidering hacks: 100 industrialstrength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00577-6. xix + 402 pp. LCCN QA76.9.D343 H46 2003. URL

http://www.oreilly.com/catalog/9780596005771.

Henderson:2006:BSW

[814] Cal Henderson. Building Scalable Web Sites. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10235-6. xvi + 330 pp. LCCN TK5105.888

.H354 2006. EUR 38.00. URL http://www.oreilly.com/catalog/
9780596102357.

Henney:2010:TEP

[815] Kevlin Henney, editor. 97 things every programmer should know: collective wisdom from the experts. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 0-596-80951-4, 0-321-65870-1. xxiv + 229 pp. LCCN ????

Herrington:1997:PBY

[816] Elizabeth Ann Herrington, editor. Passport Brazil: Your Pocket Guide to Brazilian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885073-18-6. 96 pp. LCCN HD58.7 .H465 1997.

US$6.95. URL http://www.oreilly.com/catalog/9781885073181.

Herrington:2005:PHT

[817] Jack D. Herrington. Podcasting hacks: tips and tools for blogging out loud. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10066-3. 428 pp. LCCN TK5105.887 .H47 2005eb; TK5105.887; TK5105.887 .H47

2005. URL http://www.oreilly.com/catalog/9780596100667.

Herrington:2006:PH

[818] Jack D. Herrington. PHP hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10139-2. xxi + 443 pp. LCCN QA76.73.P224eb; QA76.73.P224; QA76.73.P224 H47 2006. URL http://www.oreilly.

com/catalog/9780596101398.

Herrington:2008:GSF

[819] Jack Herrington and Emily Kim. Getting started with Flex 3. Adobe developer library. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52064-6. xv + 129 pp. LCCN QA76.625 .H47 2008.

Hertzfeld:2005:RV

[820] Andy Hertzfeld, Steve Capps, et al. Revolution in the valley. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00719-1. xxiv + 291 pp. LCCN

HD9696.2.U64 A674 2005. URL http://www.oreilly.com/catalog/
9780596007195.

Hewitt:2010:CDG

[821] Eben Hewitt. Cassandra: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-9041-2 (paperback), 1-4493-9976-2. xxiii + 301 pp. LCCN QA76.9.D3 H4898 2011.

Hibbs:2007:RW

[822] Curt Hibbs and Brian Hogan. Rails on Windows. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52965-1. LCCN QA76.64 .H658 2007eb;

QA76.64. URL http://www.oreilly.com/catalog/9780596529659.

Hildebrandt:2005:APP

[823] Ralf Hildebrandt. Absolute postfix: patches are for Qmail. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-001-1. xxviii + 448 pp. LCCN TK5105.74.P66 H55 2005eb. URL http://www.oreilly.

com/catalog/9781593270018.

Hildebrandt:2005:BPS

[824] Ralf Hildebrandt and Patrick Koetter. The book of Postfix: state-of-the-art message transport. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-001-1. xxviii + 464 pp. LCCN TK5105.74.P66 H55 2005.

URL http://www.oreilly.com/catalog/9781593270018.

Hillerson:2008:EAD

[825] Tony Hillerson, editor. Enterprise Application Development with Flex. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51440-9. 104 (est.) pp. LCCN QA76.625 .E35 2008.

Hilyard:2006:CC

[826] Jay Hilyard and Stephen Teilhet. C# cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10063-9. xxviii + 1154 pp. LCCN QA76.73.C154 T43 2004. URL http://www.oreilly.com/

catalog/9780596003395.

Hilyard:2007:CC

[827] Jay Hilyard and Stephen Teilhet. C# 3.0 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-51610-X (paperback). xxvi + 857 pp. LCCN QA76.73.C154; QA76.73.C154eb. URL http://

www.oreilly.com/catalog/9780596516109.

Hinojosa:2013:TS

[828] Daniel Hinojosa. Testing in Scala. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2013. ISBN 1-4493-1511-9, 1-4493-6034-3 (e-book), 1-4493-6033-5 (e-book). xiv + 148 pp. LCCN QA76.7 .H555 2013. URL http://

proquest.tech.safaribooksonline.de/9781449360313;   http://
www.loc.gov/catdir/enhancements/fy1310/2012554271-b.html;
http://www.loc.gov/catdir/enhancements/fy1310/2012554271-d.
html.

Hitchens:2002:JN

[829] Ron Hitchens. Java NIO. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00288-2. xvii + 282 pp. LCCN QA76.73.J38 H58

2002. US$34.95. URL http://safari.oreilly.com/0596002882;
http://www.oreilly.com/catalog/9780596002886;   http://www.
oreilly.com/catalog/javanio.

Hoekman:2005:FB

[830] Robert Hoekman. Flash out of the box. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00691-8 (paperback). xvi + 236 pp. LCCN TR897.7 .H64 2005.

URL http://www.oreilly.com/catalog/9780596006914.

Holden:2005:DE

[831] Greg Holden. Degunking eBay. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-932111-99-9. xxvi + 280 pp. LCCN HF5478 .H64

2005; HF5478 .H65 2005. URL http://www.oreilly.com/catalog/
9781932111996.

Holdener:2008:ADG

[832] Anthony T. Holdener. Ajax: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52838-8. xx + 957 pp. LCCN TK5105.8885.A52 H65 2008eb; TK5105.8885.A52. URL http://www.

oreilly.com/catalog/9780596528386.

Holmes:2002:PVB

[833] Matt Holmes, Patrick Burton, and Roger Knoell. Programming Visual Basic for the Palm OS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00200-9. 406 pp. LCCN QA76.73.B3 H545

2002. US$39.95. URL http://safari.oreilly.com/0596002009;
http://www.oreilly.com/catalog/9780596002008;   http://www.
oreilly.com/catalog/vbpropalm.

Holmes:2006:WPQ

[834] Lee Holmes. Windows PowerShell quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52813-2. LCCN QA76.66. URL http://

www.oreilly.com/catalog/9780596528133.

Holmes:2008:WPC

[835] Lee Holmes. Windows PowerShell cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52849-3. xxvii + 553 pp. LCCN QA76.76.O63.

URL http://www.oreilly.com/catalog/9780596528492.

Holmes:2008:WPP

[836] Lee Holmes. Windows Powershell pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52178-2. xi + 162 pp. LCCN QA76.76.O63 H6485 2008. URL http://www.loc.gov/catdir/

enhancements/fy0917/2009279013-d.html.

Holzner:2004:E

[837] Steven Holzner. Eclipse. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00641-1. xiv + 317 pp. LCCN QA76.73.J38 H63 2004. URL http://www.oreilly.com/

catalog/9780596006419.

Holzner:2004:EC

[838] Steven Holzner. Eclipse cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00710-8 (paperback). xv + 343 pp. LCCN QA76.73.J38 H635 2004. URL http:

//www.oreilly.com/catalog/9780596007102.

Holzner:2004:PCL

[839] Steven Holzner. Perl Core language. Little black book. Paraglyph Press, Inc., Scottsdale, AZ, USA, second edition, 2004. ISBN 1-932111-92-1. xxx + 495 pp. LCCN QA76.73.P47 H65 2004. URL http://www.

oreilly.com/catalog/9781932111927.

Holzner:2005:ADG

[840] Steven Holzner. Ant: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00609-8. xv + 316 pp. LCCN QA76.73.J38 T57 2005. URL http://www.oreilly.com/

catalog/9780596006099.

Honiball:2005:CPR

[841] Glenn Honiball. Commercial Photoshop retouching: in the studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00849-X. x + 257 pp. LCCN TR267; TR267 .H65 2005eb. URL http://www.oreilly.

com/catalog/9780596008499.

Hook:2005:WPC

[842] Brian Hook. Write portable code: an introduction to developing software for multiple platforms. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-056-9. xx + 248 pp. LCCN QA76.76.C64 H66 2005eb; QA76.76.C64 H66 2005. URL http://www.oreilly.com/

catalog/9781593270568.

Hope:2009:WST

[843] Paco Hope and Ben Walther. Web security testing cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-51483-2. xxiii + 285 pp. LCCN TK5105.88813 H67 2008.

Hopkins:2007:XPI

[844] Dorothy Hopkins. XML publishing with InDesign CS2+. O’Reilly shortcut. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51399-

2. LCCN Z253.532.A34. URL http://www.oreilly.com/catalog/
9780596513993.

Horswill:2000:DPC

[845] John Horswill and Members of the CICS Development Team at IBM Hursley. Designing and Programming CICS Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2000. ISBN 1-56592-676-5. xiv + 398 pp. LCCN QA76.76.D47 H69 2000. US$44.95.

URL http://www.oreilly.com/catalog/9781565926769;   http://
www.oreilly.com/catalog/cics.

Howard:2007:HIH

[846] Jack Howard. HDR: an introduction to high dynamic range photography. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51017-9. LCCN

TR267; TR267 .H69 2007eb. URL http://www.oreilly.com/catalog/
9780596510176.

Howard:2012:NJP

[847] Daniel Howard. Node.js for PHP Developers: Porting PHP to Node.js. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-3360-5. 256 pp. LCCN ???? EUR 24.50.

Hsu:2012:PRH

[848] Regina Obe and Leonard Hsu. PostgreSQL: Up and running. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2632-3, 1-4493-2633-1. xiv + 145 pp. LCCN QA76.9.D3 O24 2012. EUR 16.00. URL http://

proquest.safaribooksonline.com/9781449326326.

Hu:1995:DSP

[849] Wei Hu. DCE Security Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1995. ISBN 1-56592-134-8. xix + 364 pp. LCCN QA76.9.D5 H85 1995.

US$29.95. URL http://www.oreilly.com/catalog/9781565921344.

Hudson:2006:PN

[850] Paul Hudson. PHP in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10067-1 (paperback). xv + 352 pp. LCCN QA76.73.P224

H83 2006. URL http://www.oreilly.com/catalog/9780596100674.

Huggins:2005:PRC

[851] Barry Huggins. Photoshop retouching cookbook for digital photographers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10030-2 (paperback). 176 (est.) pp. LCCN T385 .H84 2005eb. URL http://www.

oreilly.com/catalog/9780596100308.

Hummel:2006:LFD

[852] Joe Hummel. LINQ, the future of data access in C# 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52841-8. 64 pp. LCCN QA76.73.C154; QA76.73.C154 H86 2006eb. URL http://www.oreilly.

com/catalog/9780596528416.

Hunsicker:1997:TTD

[853] Christine Hunsicker, editor. Travelers’ Tales: a Dog’s World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-23-6. 100 (est.) pp.

LCCN ???? US$12.95. URL http://www.ora.com/catalog/dogs/;
http://www.oreilly.com/catalog/9781885211231;   http://www.
oreilly.com/catalog/dogs.

Hunt:1992:TIN

[854] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 0-937175-82-X (paperback). xxii + 471 pp. LCCN TK5105.9 .H86.

Hunt:1995:NPC

[855] Craig Hunt. Networking Personal Computers with TCP/IP: Building TCP/IP Networks. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1995. ISBN 1-56592-123-2. xxi + 383 pp. LCCN TK5105.585 .H86 1995.

US$29.95. URL http://www.oreilly.com/catalog/9781565921238;
 http://www.oreilly.com/catalog/netpc.

Hunt:1998:TIN

[856] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1998. ISBN 1-56592-322-7. xv + 612 pp. LCCN TK5105.55 .H86 1998. US$32.95. URL

http://www.ora.com/catalog/tcp2/;   http://www.oreilly.com/
catalog/9781565923225;   http://www.oreilly.com/catalog/tcp2.

Hunt:1998:WNT

[857] Craig Hunt and Robert Bruce Thompson. Windows NT TCP/IP Network Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-377-4. xvi + 487 pp. LCCN TK5105.585 .H864

1998. US$34.95. URL http://www.ora.com/catalog/wintcp/;
http://www.oreilly.com/catalog/9781565923775;   http://www.
oreilly.com/catalog/wintcp.

Hunt:2002:TIN

[858] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00297-1. xvii + 725 pp. LCCN TK5105.55 .H86 2002. US$44.95. URL http://safari.oreilly.com/

0596002971;   http://www.oreilly.com/catalog/9780596002978;
 http://www.oreilly.com/catalog/tcp3.

Hunt:2003:SC

[859] Craig Hunt. Sendmail cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00471-0. xviii + 388 pp. LCCN TK5105.74.S44 H85 2004.

URL http://www.oreilly.com/catalog/9780596004712.

Hunt:2004:PUT

[860] Andrew Hunt and David Thomas. Pragmatic unit testing: in Java with JUnit, volume 2, [pt. 2] of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-1-2. xv + 159 pp. LCCN QA76.76.T48 H862 2004. URL http://www.oreilly.com/

catalog/9780974514017.

Hunt:2005:WSN

[861] Craig Hunt. Windows Server 2003 network administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00800-7 (paperback). 600 (est.) pp. LCCN TK5105.5 .H86 2005. URL

http://www.oreilly.com/catalog/9780596008000;   http://www.
oreilly.com/catalog/windowsvrnet/.

Hunter:1998:JSP

[862] Jason Hunter and William Crawford. Java Servlet Programming. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-391-X. xvi + 510 pp. LCCN QA76.73.J38 H86 1998. US$32.95.

URL http://www.oreilly.com/catalog/9781565923911;   http://
www.oreilly.com/catalog/jservlet.

Hunter:2001:JSP

[863] Jason Hunter and William Crawford. Java Servlet Programming. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, April 2001. ISBN 0-596-00040-5. xxiii + 753 pp. LCCN QA76.73.J38 H88 2001. US$44.95.

URL http://www.oreilly.com/catalog/9780596000400;   http://
www.oreilly.com/catalog/jservlet2.

Hunter:2009:ADC

[864] Laura E. Hunter and Robbie Allen. Active Directory cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2009. ISBN 0-596-15630-8. ???? pp. LCCN TK5105.595 .H86 2009.

Hyde:2003:AAL

[865] Randall Hyde. The art of assembly language. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-97-2. xix + 903 pp. LCCN

QA76.73.A8 H97 2003eb. URL http://www.oreilly.com/catalog/
9781886411975.

Hyde:2004:WGCa

[866] Randall Hyde. Write great code. Volume 1: understanding the machine. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-003-8. xvi + 410 pp. LCCN QA76.6 .H94. URL http://www.oreilly.com/

catalog/9781593270032.

Hyde:2006:WGCb

[867] Randall Hyde. Write great code. Volume 2: thinking low-level, writing high-level. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-065-8. xix + 620 pp. LCCN QA76.6 .H94 2006eb. URL http://

www.oreilly.com/catalog/9781593270650.

Icove:1995:CCC

[868] David J. Icove, Karl A. Seger, and William R. VonStorch. Computer Crime: a Crimefighter’s Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-086-4. xxi + 437 pp. LCCN HV6773 .I26 1995;

HV6773 .I36 1995. US$24.95. URL http://www.oreilly.com/catalog/
9781565920866;   http://www.oreilly.com/catalog/crime.

Iedema:2015:GSO

[869] Michael Iedema and Harvind Samra. Getting started with OpenBTS: build open source mobile networks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-1065-8, 1-4919-2430-6 (e-book). ???? pp. LCCN

TK5105.597. URL http://alltitles.ebrary.com/Doc?id=11006437;
 http://proquest.safaribooksonline.com/?fpi=9781491924280.

Igoe:2007:MTT

[870] Tom Igoe. Making things talk. Make: projects. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51051-9. 426 pp. LCCN TJ223.P76; TJ223.P76 I46 2007. URL

http://www.oreilly.com/catalog/9780596510510.

Illiad:1999:UF

[871] Illiad. User Friendly. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-673-0. x + 122 pp. LCCN PN6728.U84 I45 1999. US$12.95.

URL http://www.oreilly.com/catalog/9781565926738;   http://
www.oreilly.com/catalog/friendly.

Illiad:2000:EGN

[872] Illiad. Evil geniuses in a nutshell: a User Friendly guide to world domination. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2000. ISBN 1-56592-861-X. ix + 122 pp. LCCN PN6728.U84 I43 2000; PN6728.U74 I55 2000. US$12.95.

URL http://www.oreilly.com/catalog/9781565928619;   http://
www.oreilly.com/catalog/evilgenius.

Illiad:2001:RAE

[873] Illiad. The Root of All Evil. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00193-2. vii + 136 pp. LCCN PN6728.U84 I447 2001.

US$12.95. URL http://www.oreilly.com/catalog/9780596001933;
 http://www.oreilly.com/catalog/rootofevil.

Illiad:2003:EGG

[874] Illiad. Even grues get full: the fourth User Friendly collection. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00566-0. x + 122 pp. LCCN PN6728.U84 I39 2003. URL http://www.oreilly.com/

catalog/9780596005665.

Ishaq:2004:TFB

[875] Jeff Ishaq. Treo fan book: your brain on silicon. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004.

ISBN 0-596-00816-3. xiv + 114 pp. LCCN QA76.8.H36 I84 2004. URL

http://www.oreilly.com/catalog/9780596008161.

IssiCamy:2004:WPD

[876] Lázaro Issi Camy and Joseph Issi Cohen. The Web programmer’s desk reference: a complete cross-reference to HTML, CSS, and JavaScript. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-011-9. xxxiii + 1068 pp. LCCN TK5105.888 .I572 2004eb. URL http://www.

oreilly.com/catalog/9781593270117.

Ivens:2000:MWN

[877] Kathy Ivens. Managing Windows NT Logons. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-637-4. x + 224 pp. LCCN QA76.76.O63 I925 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565926370;   http://
www.oreilly.com/catalog/ntlogon.

Ivens:2005:HNA

[878] Kathy Ivens. Home networking annoyances: how to fix the most annoying things about your home network. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00808-2. xiii + 209 pp. LCCN TK5105.75 .I92 2005eb; TK5105.75 .I92 2005; TK5105.75 .I94 2005; TK5105.75. URL http://

www.oreilly.com/catalog/9780596008086.

Iverson:2003:MXJ

[879] Will Iverson. Mac OS X for Java Geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00400-1. xii + 282 pp. LCCN QA76.76.O63 I936 2003. US$39.95 US, CDN$61.95, UK£28.50.

URL http://www.oreilly.com/catalog/9780596004002;   http://
www.oreilly.com/catalog/macxjvgks/.

Iverson:2005:RWW

[880] Will Iverson. Real world Web services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00642-X (paperback). xi + 207 pp. LCCN TK5105.88813 I96 2004; TK5105.88813; TK5105.88813 I96 2004eb;

TK5105.88813 .I96 2005. URL http://www.oreilly.com/catalog/
9780596006426.

Jacobsen:1997:NLH

[881] Paul S. Jacobsen. Net Law: How Lawyers Use the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-258-1. xx + 233 pp. LCCN KF242.A1 J32 1997. US$29.95. URL http://www.ora.com/catalog/

netlaw/;   http://www.oreilly.com/catalog/9781565922587;   http:
//www.oreilly.com/catalog/netlaw. Includes CD-ROM.

Jaeger:2005:GKE

[882] Till Jaeger et al. Die GPL kommentiert und erklärt. (German) [The GPL discussed and explained]. O’Reilly, Köln, Germany, 2005. ISBN 3-89721-389-3. 182 pp. LCCN ???? URL http://www.oreilly.com/

catalog/9783897213890.

Jameson:1994:MCM

[883] Kevin Jameson. Multi-Platform Code Management. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1994. ISBN 1-56592-059-7. xxxii + 320

pp. LCCN ???? US$39.95. URL http://www.oreilly.com/catalog/
9781565920590.

Janes-Hodder:1999:CCP

[884] Honna Janes-Hodder and Nancy Keene. Childhood cancer: a parent’s guide to solid tumor cancers. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-531-9. xvi + 521 pp. LCCN QZ 275 J33c 1999; RC281.C4 J36 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565925311;   http://
www.oreilly.com/catalog/childcancer/.

Jang:2006:LAG

[885] Michael H. Jang. Linux annoyances for geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00801-5. xv + 484 pp. LCCN QA76.76.O63; QA76.76.O63 J36 2006eb; QA76.76.O63 J36 2006. URL http://www.

oreilly.com/catalog/9780596008017.

Jepson:1999:OPT

[886] Brian Jepson and Larry Wall. O’Reilly Perl Toolkit for Linux and Solaris. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-604-8. 120

pp. LCCN ???? US$49.95. URL http://www.oreilly.com/catalog/
9781565926042;   http://www.oreilly.com/catalog/perltkit/.

Jepson:2002:MXU

[887] Brian Jepson and Ernest E. (Ernest Eric) Rothman. Mac OS X for Unix geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00356-0. xv + 198 pp. LCCN QA76.76.O63 J476

2002. US$24.95. URL http://safari.oreilly.com/0596003560;
http://www.oreilly.com/catalog/9780596003562;   http://www.
oreilly.com/catalog/mosxgeeks.

Jepson:2004:MXP

[888] Brian Jepson and Ernest E. (Ernest Eric) Rothman. Mac OS X Panther for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00607-1. xvii + 363 pp. LCCN QA76.76.O63 J475 2004. URL http://www.

oreilly.com/catalog/9780596006075.

Jepson:2005:MXT

[889] Brian Jepson and Ernest E. (Ernest Eric) Rothman. Mac OS X Tiger for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00912-7. xviii + 395 pp. LCCN QA76.76.O63 J47 2005. URL http://www.

oreilly.com/catalog/9780596009120.

Jepson:2008:MXU

[890] Brian Jepson, Richard Rosen, and Ernest E. Rothman. Mac OS X for Unix geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2008. ISBN 0-596-52062-X. xviii + 406 pp. LCCN QA76.76.O63 R68 2008.

Jepson:2011:GSR

[891] Brian Jepson and Tom Igoe. Getting started with RFID and NFC: contactless communication with Android, Arduino, and processing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-0852-X (paperback). 1 v. pp. LCCN ????

Johnson:2006:SJD

[892] Stephen Johnson. Stephen Johnson on digital photography. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52370-X (paperback). xiii + 305 pp. LCCN TR267; TR267 .J65 2006. URL http://www.oreilly.

com/catalog/9780596523701.

Johnson:2007:UOC

[893] Joan M. Johnson. Understanding Oracle Clinical. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-55005-7. LCCN QA76.9.D3; QA76.9.D3 J64

2007eb. URL http://www.oreilly.com/catalog/9780596550059.

Johnston:1999:NHL

[894] Lorraine Johnston. Non-Hodgkin’s Lymphomas: Making Sense of Diagnosis, Treatment, and Options. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-444-4. xx + 556 pp. LCCN RC280.L9

J64 1999. US$24.95. URL http://www.oreilly.com/catalog/
9781565924444;   http://www.oreilly.com/catalog/nhl/.

Johnston:2000:CRC

[895] Lorraine Johnston. Colon and rectal cancer: a comprehensive guide for patients and families. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-633-1. xvi + 528 pp. LCCN RC280.C6

J64 2000. URL http://www.oreilly.com/catalog/9781565926332.

Jones:2002:PX

[896] Christopher A. Jones and Fred L. Drake. Python and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00128-2. xvi + 360 pp. LCCN QA76.73.P98 J66 2002. URL http://www.oreilly.com/

catalog/9780596001285.

Jones:2005:IF

[897] Robert Jones. Internet forensics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10006-X. xii + 223 pp. LCCN TK5105.875.I57. URL

http://www.oreilly.com/catalog/9780596100063.

Jordan:2003:JDO

[898] David Jordan and Craig L. Russell. Java Data Objects. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00276-9. xxi + 356 pp. LCCN QA76.73.J38 J66 2003. US$39.95 US, CDN$61.95, UK£28.50.

URL http://www.oreilly.com/catalog/9780596002763;   http://
www.oreilly.com/catalog/jvadtaobj/.

Joseph:1996:PFY

[899] Nadine Joseph. Passport France: Your Pocket Guide to French Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-885073-29-1. 96 pp. LCCN HD58.7.J65 1996. US$6.95.

URL http://www.oreilly.com/catalog/9781885073297.

Joshi:1996:PIY

[900] Manoj Joshi. Passport India: Your Pocket Guide to Indian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1996. ISBN 1-885073-23-2. 96 pp. LCCN HD58.7.J655 1996. US$6.95.

URL http://www.oreilly.com/catalog/9781885073235.

Kahney:2005:CI

[901] Leander Kahney. The cult of iPod. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-066-6. ???? pp. LCCN ML74.4.I48 K34

2005. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/toc/ecip0517/2005023404.html;
 http://www.oreilly.com/catalog/9781593270667.

Kahney:2006:CM

[902] Leander Kahney. The cult of Mac. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-122-0 (paperback), 1-886411-83-2 (paperback).

268 pp. LCCN QA76.8.M3 K35 2004; QA76.8.M3 K35 2006. URL http:

//www.oreilly.com/catalog/9781593271220.

Kahrel:2006:AIR

[903] Peter Kahrel. Automating InDesign with regular expressions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52937-6. LCCN Z253.532.A34;

Z253.532.A34 K34 2006eb. URL http://www.oreilly.com/catalog/
9780596529376.

Kahrel:2006:SIJ

[904] Peter Kahrel. Scripting InDesign with JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52817-5. LCCN Z253.532.A34. URL

http://www.oreilly.com/catalog/9780596528171.

Kahrel:2007:GIC

[905] Peter Kahrel. GREP in InDesign CS3. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51706-8. LCCN Z253.532.A34;

Z253.532.A34 K34 2007eb. URL http://www.oreilly.com/catalog/
9780596517069.

Kahrel:2008:GIC

[906] Peter Kahrel. GREP in InDesign CS3/CS4. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15717-7. 53 pp. LCCN Z253.532.A34 K34 2008.

Kalbach:2007:DWN

[907] Jim Kalbach and Jennifer Web navigation Fleming. Designing Web navigation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, updated edition, 2007. ISBN 0-596-52810-8 (paperback). xv + 394 pp. LCCN

TK5105.888eb; TK5105.888. URL http://www.oreilly.com/catalog/
9780596528102.

Karp:1997:WA

[908] David A. Karp. Windows Annoyances. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-266-2. xi + 285 pp. LCCN QA76.76.O63 K366 1997; QA76.76.W56 K37

1997. US$24.95. URL http://www.ora.com/catalog/annoyances/;
http://www.oreilly.com/catalog/9781565922662;   http://www.
oreilly.com/catalog/annoyances.

Karp:1998:CWG

[909] Gary Karp. Choosing a wheelchair: a guide for optimal independence. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-411-8. vi + 180 pp. LCCN RD757.W4 K27 1998; RD757.W4 K37 1998. US$9.95. URL

http://www.oreilly.com/catalog/9781565924116;   http://www.
oreilly.com/catalog/wheelchair.

Karp:1998:WA

[910] David A. Karp. Windows 98 Annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-417-7. xiv + 448 pp. LCCN QA76.76.O63 K3658 1998; QA76.76.O63 K377 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565924178;   http://
www.oreilly.com/catalog/win98annoy.

Karp:1999:LWA

[911] Gary Karp. Life on Wheels: For the Active Wheelchair User. Patientcentered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-253-0. xvi + 549 pp. LCCN RC406.P3 K37 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565922532;   http://
www.oreilly.com/catalog/lifewheels/.

Karp:2001:WMA

[912] David A. Karp. Windows Me Annoyances: Annoyances.org. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2001. ISBN 0-596-00060-X. xvii + 451 pp. LCCN QA76.76.O63 K367 2001; QA 76.76 O63 K37 2001. US$29.95.

URL http://www.oreilly.com/catalog/9780596000608;   http://
www.oreilly.com/catalog/winmeannoy.

Karp:2002:WXA

[913] David A. Karp. Windows XP Annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00416-8. xx + 564 pp. LCCN QA76.76.O63

K3674 2002. US$29.95. URL http://www.oreilly.com/catalog/
9780596004163;   http://www.oreilly.com/catalog/winxpannoy.

Karp:2002:WXN

[914] David Karp and Tim O’Reilly & Troy Mott. Windows XP in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00249-1. 616 pp. LCCN QA76.76.O63 K378

2002. US$29.95. URL http://safari.oreilly.com/0596002491;
http://www.oreilly.com/catalog/9780596002497;   http://www.
oreilly.com/catalog/winxpnut.

Karp:2003:EHI

[915] David A. (David Aaron) Karp. eBay hacks: [100 industrial-strength tips and tools]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00564-4. xxv + 331 pp. LCCN HF5478 .K37 2003. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://www.
loc.gov/catdir/toc/fy045/2004272738.html;   http://www.oreilly.
com/catalog/9780596005641.

Karp:2003:WXP

[916] David A. (David Aaron) Karp. Windows XP pocket reference: compact guide to Windows XP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00425-7. vi + 181 pp. LCCN QA76.76.O63 K3675 2003. US$12.95.

URL http://www.oreilly.com/catalog/9780596004255;   http://
www.oreilly.com/catalog/winxpg. Edited by Tim O’Reilly & Troy

Mott.

Karp:2005:EH

[917] David A. (David Aaron) Karp. eBay hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-10068-X (paperback). xxiv + 438 pp. LCCN HF5478

.K37 2005. URL http://www.oreilly.com/catalog/9780596100681.

Karp:2005:WXA

[918] David A. Karp. Windows XP annoyances for geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00876-7 (paperback). xxiii + 637 pp. LCCN

QA76.76.O63 K3674 2005. URL http://www.oreilly.com/catalog/
9780596008765.

Karp:2005:WXN

[919] David A. (David Aaron) Karp, Tim O’Reilly, Troy Mott, and Richard Cobbett. Windows XP in a nutshell. In a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00900-3. xvii + 663 pp. LCCN QA76.76.O63

K37 2005. URL http://www.oreilly.com/catalog/9780596009007.

Karp:2006:FWX

[920] David A. Karp. Fixing Windows XP annoyances: how to fix the most annoying things about the Windows OS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10053-1. xix + 219 pp. LCCN QA76.76.O63;

QA76.76.O63 K3763 2006eb. URL http://www.oreilly.com/catalog/
9780596100537.

Karp:2007:WVA

[921] David A. (David Aaron) Karp. Windows Vista annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52762-4. xx + 641 pp. LCCN QA76.76.O63 K17 2007.

Karwin:2010:SAA

[922] Bill Karwin. SQL antipatterns: avoiding the pitfalls of database programming. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-55-7 (paperback). xii + 333 pp. LCCN QA76.73.S67 K37 2010.

Kazil:2015:DWU

[923] Jacqueline Kazil and Katharine Jarmul. Data Wrangling Using Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-4881-7. ???? pp. LCCN ????

Kazoun:2007:PF

[924] Chafic Kazoun, Joey Lott, and Peter Hall. Programming Flex 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52689-X. xxvii + 472 pp. LCCN QA76.625 .K39 2007; TR897.7; TR897.7 .K39 2007eb. URL http://

www.oreilly.com/catalog/9780596526894.

Keating:1997:PKY

[925] Kevin Keating, editor. Passport Korea: Your Pocket Guide to Korean Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885073-39-9. 96 pp. LCCN HF5389.K43 1997. US$6.95.

URL http://www.oreilly.com/catalog/9781885073396.

Keene:1997:CLG

[926] Nancy Keene. Childhood leukemia: a guide for families, friends and caregivers. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-191-7. xxiii + 539 pp. LCCN RJ416.L4 K44 1997; RJ416.L4 K446 1998. US$24.95. URL

http://www.ora.com/catalog/leukemia/;   http://www.oreilly.
com/catalog/9781565921917;   http://www.oreilly.com/catalog/
leukemia.

Keene:1997:YCH

[927] Nancy Keene and Rachel Prentice. Your Child in the Hospital: a practical guide for parents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-346-4. vi + 126 pp. LCCN RJ242 .K44

1997. US$8.95. URL http://www.ora.com/catalog/hospital/;
http://www.oreilly.com/catalog/9781565923461;   http://www.
oreilly.com/catalog/hospital.

Keene:1998:WYD

[928] Nancy Keene. Working with Your Doctor: Getting the Healthcare You Deserve. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-273-5. xx + 357 pp. LCCN R727.3 .K26 1998; R727.3 .K44 1998. US$15.95.

URL http://www.oreilly.com/catalog/9781565922730;   http://
www.oreilly.com/catalog/doctor.

Keene:1999:CLG

[929] Nancy Keene. Childhood leukemia: a guide for families, friends and caregivers. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-632-3. xix + 494 pp. LCCN RJ416.L4

K44 1999. US$24.95. URL http://www.oreilly.com/catalog/
9781565926325;   http://www.oreilly.com/catalog/leukemia2.

Keene:1999:YCH

[930] Nancy Keene and Rachel Prentice. Your Child in the Hospital: a Practical Guide for Parents. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-573-4. vi + 160 pp. LCCN

RJ242 .K44 1999. US$11.95. URL http://www.oreilly.com/catalog/
9781565925731;   http://www.oreilly.com/catalog/hospital2/.

Keene:2000:CCS

[931] Nancy Keene, Wendy Hobbie, and Kathy Ruccione. Childhood cancer survivors: a practical guide to your future. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-460-6. xxviii + 482 pp. LCCN RC281.C4 K44 2000. URL http://www.oreilly.com/

catalog/9781565924604.

Kent:2000:ANC

[932] Dorothy Kent. AutoCAD 2000 in a Nutshell: a Command Reference Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2000. ISBN 1-56592-690-0. 637 pp. LCCN T385 .K566 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565926905;   http://
www.oreilly.com/catalog/autocadnut.

Kenworthy:2006:DVP

[933] Christopher Kenworthy. Digital video production cookbook: 100 professional techniques for independent and amateur filmmakers. O’Reilly digital studio. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10031-0. 192 pp. LCCN TR860 .K46 2006; TR860; TR860 .K46 2006eb. URL http://www.

oreilly.com/catalog/9780596100315.

Kerrigan:1993:MF

[934] James F. Kerrigan. Migrating to Fortran 90. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1993. ISBN 1-56592-049-X. xxvi + 361 pp. LCCN QA76.73.F25 K47 1993; QA76.73.F28 K47 1993.

US$27.95. URL http://www.loc.gov/catdir/enhancements/fy0715/
94174237-d.html;   http://www.oreilly.com/catalog/9781565920491.

Kerrigan:1993:MFP

[935] Jim Kerrigan. Migrating to Fortran 90 (Programming Languages). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1993. ISBN 1-56592-049-X.

Kersnar:1996:NHR

[936] Scott Kersnar. NetSuccess: How Real Estate Agents Use the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-56592-213-1. 214 pp. LCCN ???? US$24.95. URL

http://www.oreilly.com/catalog/9781565922136;   http://www.
oreilly.com/catalog/netsuccess.

Kessin:2012:BWA

[937] Zachary Kessin. Building web applications with Erlang. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2066-X (e-book), 1-4493-2065-1 (e-book), 1-4493-2062-7, 1-4493-0996-8. xvi + 133 pp. LCCN QA76.73.E75. URL http://

cdn.totalboox.com/static/covers/ORL/798004101cf496f1-b.jpg;
 http://excerpts.cdn.overdrive.com/FormatType-410/2858-1/433/
F8B/F8/BuildingWebApplicationswithErlangWork9781449320652.epub;
 http://images.contentreserve.com/ImageType-100/2858-1/{433F8BF8-
1081-406E-AC43-B29E5DCB23B9}Img100.jpg;   http://proquest.safaribooksonline.
com/9781449320621;   http://public.eblib.com/choice/publicfullrecord.

aspx?p=939411;   http://site.ebrary.com/id/10758595;   http:
//www.totalboox.com/book/id-8755002142860023537.

Ketelsen:2017:TG

[938] Brian Ketelsen. Testing in Go. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-7276-9, 1-4919-7277-7. 1 video file (25 minutes) pp. LCCN ????

Khadka:1997:TTN

[939] Rajendra S. Khadka, editor. Travelers’ Tales Nepal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-14-7. xxiii + 398 pp. LCCN DS485.N4 T73 1997.

US$17.95. URL http://www.ora.com/catalog/nepal/;   http://
www.oreilly.com/catalog/9781885211149;   http://www.oreilly.
com/catalog/nepal.

Killelea:1998:WPT

[940] Patrick Killelea. Web Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-379-0. xix + 351 pp. LCCN QA76.9.C63 K55 1998; TK5105.888 .K493 1998. US$32.95.

URL http://www.oreilly.com/catalog/9781565923799;   http://
www.oreilly.com/catalog/webpt.

Killelea:2002:WPT

[941] Patrick Killelea. Web Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00172-X. xxi + 456 pp. LCCN TK5105.888 .K493 1998; TK5105.888 .K54 2002. US$44.95.

URL http://www.oreilly.com/catalog/9780596001728;   http://
www.oreilly.com/catalog/webpt2.

King:2008:WO

[942] Andrew B. King. Website optimization. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51508-1. xxiii + 367 pp. LCCN QA76.625 .K56 2008.

Kirch:1995:LNA

[943] Olaf Kirch. Linux Network Administrators’ Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1995. ISBN 1-56592-087-2. xxxi + 335 pp. LCCN QA76.76.O63 K566 1995. US$24.95.

URL http://www.oreilly.com/catalog/9781565920873;   http://
www.oreilly.com/catalog/linag.

Kirch:2000:LNA

[944] Olaf Kirch and Terry Dawson. Linux Network Administrator’s Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 2000. ISBN 1-56592-400-2. xxix + 474 pp. LCCN QA76.76.O63 K566 2000. US$39.95.

URL http://www.oreilly.com/catalog/9781565924000;   http://
www.oreilly.com/catalog/linag2.

Kissell:2008:TCM

[945] Joe Kissell. Take control of MobileMe. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-48-3. 112 (est.) pp. LCCN TK5105.885 K47 2008.

Kissell:2008:TCV

[946] Joe Kissell. Take control of VMware Fusion 2. Take control. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-933671-49-1. ???? pp. LCCN QA76.76.O63 K57 2008.

Kittel:2006:ANC

[947] Michael A. Kittel and Geoffrey T. LeBlond. ASP.NET 2.0 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10064-7. xxii + 989 pp. LCCN TK5105.8885.A26; TK5105.8885.A26 K58 2005; TK5105.8885.A26 K58 2006; TK5105.8885.A26 K58 2005eb.

URL http://www.oreilly.com/catalog/9780596100643.

Klemens:2012:CCC

[948] Ben Klemens. 21st Century C: C Tips from the New School. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2714-1. 256 pp. LCCN ???? EUR 24.50.

Kline:1999:TSP

[949] Kevin E. Kline, Lee Gould, and Andrew Zanevsky. Transact-SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-401-0. xxvi + 807 pp. LCCN QA76.73.S67 K55 1999. US$44.95.

URL http://www.oreilly.com/catalog/9781565924017;   http://
www.oreilly.com/catalog/wintrnssql/. Includes CD-ROM.

Kline:2001:SND

[950] Kevin E. Kline and Daniel Kline. SQL in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-744-3. viii + 214 pp. LCCN QA76.73.S67 K55 2001. US$29.95. URL http://www.

oreilly.com/catalog/9781565927445.

Kline:2004:SN

[951] Kevin E. Kline, Daniel Kline, and Brand Hunt. SQL in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00481-8. xv + 691 pp. LCCN QA76.73.S67 K54 2004; QA76.73.S67 K55 2004; QA76.73.S67; QA76.73.S67 K55 2004eb. URL http://www.

oreilly.com/catalog/9780596004811.

Kline:2008:SN

[952] Kevin E. Kline, Brand Hunt, and Daniel Kline. SQL in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2008. ISBN 0-596-51884-6. xii + 578 pp. LCCN QA76.73.S67 K54 2009.

Knight:2006:ABD

[953] Gerald Knight. Analyzing business data with Excel. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10073-6. xi + 248 pp. LCCN HF5548.4.M523eb; HF5548.4.M523. URL http://www.oreilly.com/

catalog/9780596100735.

Knudsen:1998:JC

[954] Jonathan B. Knudsen. Java Cryptography. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-402-9. xvi + 344 pp. LCCN QA76.73.J38 K59

1999. US$29.95. URL http://www.ora.com/catalog/javacrypt/;
http://www.oreilly.com/catalog/9781565924024;   http://www.
oreilly.com/catalog/javacrypt.

Knudsen:1999:JG

[955] Jonathan Knudsen. Java 2D graphics. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-484-3. xvi + 339 pp. LCCN QA76.73.J38 K685 1999; T385 .K59 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565924840;   http://
www.oreilly.com/catalog/java2d/.

Knudsen:1999:UGL

[956] Jonathan B. Knudsen. The unofficial guide to Lego Mindstorms robots. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-692-7. xv + 247 pp. LCCN TJ211.15 .K558 1999; TJ211.15 .K684 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565926929;   http://
www.oreilly.com/catalog/lmstorms.

Kohler:2006:RGH

[957] Chris Kohler. Retro gaming hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00917-8. xxvi + 473 pp. LCCN

GV1469.3; GV1469.3eb. URL http://www.oreilly.com/catalog/
9780596009175.

Kolb:2007:IBP

[958] Mark A. Kolb. Implementing BEA portal 92. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51402-6. LCCN TK5105.88813;

TK5105.88813 .K65 2007eb. URL http://www.oreilly.com/catalog/
9780596514020.

Koman:1996:GAS

[959] Richard Koman. GIF Animation Studio: Animating Your Web Site. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1996. ISBN 1-56592-230-1. xi + 159 pp. LCCN TK5105.888 .K65 1996. US$39.95.

URL http://www.oreilly.com/catalog/9781565922303;   http://
www.oreilly.com/catalog/gif.

Kong:2007:DBR

[960] Joseph Kong. Designing BSD rootkits: an introduction to kernel hacking. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-142-5. xvii + 136 pp. LCCN QA76.76.O63 K649 2007. URL http://www.

oreilly.com/catalog/9781593271428.

Kopper:2005:LEC

[961] Karl Kopper. The Linux Enterprise Cluster: build a highly available cluster with commodity hardware and free software. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-036-4. xxiv + 430 pp. LCCN QA76.58 .K67 2005; QA76.58 K67 2005eb. URL http://www.

oreilly.com/catalog/9781593270360.

Korf:2003:B

[962] Ian Korf, Mark Yandell, and Joseph Bedell. BLAST. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00299-8. xviii + 339 pp. LCCN QH324.2

.K67 2003. URL http://www.oreilly.com/catalog/9780596002992.

Korff:2005:MFO

[963] Yanek Korff, Paco Hope, and Bruce Potter. Mastering FreeBSD and OpenBSD security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00626-8. xiv + 445 pp. LCCN QA76.76.O63 K67 2005. URL http://

www.oreilly.com/catalog/9780596006266.

Korpela:2006:UE

[964] Jukka Korpela. Unicode Explained. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10121-X. xvii + 658 pp. LCCN QA268 .K67 2006.

US$59.99. URL http://www.oreilly.com/catalog/9780596101213;
 http://www.oreilly.com/catalog/unicode/index.html.

Kosko:1992:PRM

[965] Linda Kosko, editor. PHIGS Reference Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1992. ISBN 0-937175-91-9. xv + 1099 pp. LCCN QA76.76.W56 P45 1992, T385 .P487 199. US$39.95. URL http://www.

oreilly.com/catalog/9780937175910.

Koster:2005:TFG

[966] Raph Koster. A theory of fun for game design. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-932111-97-2. x + 244 pp. LCCN

GV1469.17.S63 K6 2005. URL http://www.oreilly.com/catalog/
9781932111972.

Kozierok:2005:TIG

[967] Charles Kozierok. The TCP/IP guide: a comprehensive, illustrated Internet protocols reference. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-047-X. lxxiv + 1539 pp. LCCN TK5105.585 .K69

2005. URL http://www.loc.gov/catdir/toc/ecip0417/2004008863.
html;   http://www.oreilly.com/catalog/9781593270476.

Krafft:2005:DSC

[968] Martin F. Krafft. The Debian system: concepts and techniques. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-069-0. 605 pp. LCCN QA76.76.O63 K68 2005; QA76.76.O63 K68 2005eb. URL

http://www.oreilly.com/catalog/9781593270698.

Kreibich:2011:RDG

[969] Jay A. Kreibich. Redis: the Definitive Guide Data Modeling, Caching, and Messaging. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9609-7. ???? pp. LCCN ????

Kreines:1999:ODA

[970] David C. Kreines and Brian Laskey. Oracle Database Administration: The Essential Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-516-5. xvii + 560 pp. LCCN QA76.9.D3 K7344 1999. US$36.95.

URL http://www.oreilly.com/catalog/9781565925168;   http://
www.oreilly.com/catalog/oradba/.

Kreines:2000:OSE

[971] David C. Kreines. Oracle SQL: The Essential Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-697-8. xxxiii + 382 pp. LCCN QA76.73.S67 K74 2000. US$39.95.

URL http://www.oreilly.com/catalog/9781565926974;   http://
www.oreilly.com/catalog/orsqlter.

Kreines:2003:ODD

[972] David C. Kreines. Oracle data dictionary pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00517-2. iv + 140 pp. LCCN QA76.9.D3 K7343 2003. URL http://www.oreilly.com/

catalog/9780596005177.

Kreines:2004:OIP

[973] David C. Kreines. Oracle initialization parameters: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00770-1. vi + 112 pp. LCCN

QA76.9.D3 K73445 2004. URL http://www.oreilly.com/catalog/
9780596007706.

Kreines:2005:ODP

[974] David C. Kreines. Oracle DBA pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10049-3. vi + 145 pp. LCCN QA76.9.D3 K744

2004. URL http://www.oreilly.com/catalog/9780596100490.

Krikorian:2003:TH

[975] Raffi Krikorian. TiVo hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00553-9. xxi + 254 pp. LCCN QA76.76.I59 K75 2003. URL http://www.oreilly.com/

catalog/9780596005535.

Kroah-Hartman:2007:LKN

[976] Greg Kroah-Hartman. Linux kernel in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10079-5. xiii + 182 pp. LCCN QA76.76.O63

K786 2007. URL http://www.kroah.com/lkn/;   http://www.loc.
gov/catdir/enhancements/fy0728/2007274361-d.html;   http://
www.loc.gov/catdir/toc/fy0713/2007274361.html;   http://www.
oreilly.com/catalog/9780596100797.

Krogh:2006:DBD

[977] Peter Krogh. The DAM book: digital asset management for photographers. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10018-3. xiv + 280 pp. LCCN T385eb; T385. URL http:/

/www.oreilly.com/catalog/9780596100186.

Krol:1992:WIU

[978] Ed Krol. The Whole Internet User’s Guide & Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K76 1992. US$24.95. URL http:/

/www.oreilly.com/catalog/9781565920255. Perhaps the most ambi-

tious of the new crop of books, this one is both a user’s guide and a catalog of resources in one. It is aimed at graduate students who want to use the Internet for research.

Krol:1993:WIU

[979] Ed Krol. The Whole Internet: User’s Guide and Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1993. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K86 1992. US$24.95.

Krol:1994:WIU

[980] Ed Krol and Michael Kosta Loukides. The Whole Internet User’s Guide and Catalog. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, April 1994. ISBN 1-56592-063-5. xxv + 544 pp. LCCN TK5105.875.I57

K76 1994. US$24.95. URL http://www.oreilly.com/catalog/
9781565920637;   http://www.oreilly.com/catalog/twi2.

Krol:1995:WIU

[981] Ed Krol. The Whole Internet User’s Guide and Catalog, Academic Edition. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1995. ISBN 0-534-50674-7. 609 pp. LCCN TK5105.875.I57K76 1996. US$24.95.

URL http://www.oreilly.com/catalog/9780534506742;   http://
www.oreilly.com/catalog/twi2aca. Adapted by Bruce Klopfenstein.

Krol:1995:WIW

[982] Ed Krol and Paula M. Ferguson. The whole Internet for Windows 95: user’s guide and catalog. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1995. ISBN 1-56592-155-0. xix + 625 pp. LCCN TK5105.875.I57 K7597 1995; TK5105.875.I57 K862 1995. US$24.95. URL http://www.ora.com/gnn/

bus/ora/item/twiwin.html;   http://www.oreilly.com/catalog/
9781565921559;   http://www.oreilly.com/catalog/twiwin.

Krol:1996:WIU

[983] Ed Krol and Bruce Klopfenstein. The whole Internet user’s guide and catalog. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, academic edition, 1996. ISBN 0-534-50674-7. xxix + 609 pp. LCCN TK5105.875.I57K76 1996.

Kuhn:2002:ORP

[984] Darl Kuhn and Scott Schulze. Oracle RMAN: Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00233-5. viii + 111 pp. LCCN QA76.9.B32 K84 2002; QA76.9.D3 K84

2002. US$12.95. URL http://safari.oreilly.com/0596002335;
http://www.oreilly.com/catalog/9780596002336;   http://www.
oreilly.com/catalog/rman.

Kurniawan:2002:VNC

[985] Budi Kurniawan and Ted Neward. VB.NET Core Classes in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00257-2. xiii + 558 pp. LCCN QA76.73.B3 K875 2002.

US$44.95. URL http://www.oreilly.com/catalog/9780596002572;
 http://www.oreilly.com/catalog/vbdotnetcore.

Lamb:1990:LVE

[986] Linda Lamb. Learning the vi Editor. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, October 1990. ISBN 0-937175-67-6. xvi + 173 pp. LCCN QA76.76.O63 L355 1990; QA76.8.U65 L35 1988. US$21.95.

URL http://www.oreilly.com/catalog/9780937175675;   http://
www.oreilly.com/catalog/vi5. A very good guide to vi and ex com-

mands . . .. With a quick reference card . . .. Other good vi stuff (including an online tutorial) is obtained by anonymous ftp from cs.uwp.edu

(in pub/vi).

Lamb:1995:UEE

[987] Linda Lamb and Jerry D. Peek. Using Email Effectively. What you need to know. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1995. ISBN 1-56592-103-8. xii + 146 pp. LCCN TK5105.73 .L36 1995. US$14.95. URL http://

www.oreilly.com/catalog/9781565921030.

Lamb:1998:LVE

[988] Linda Lamb and Arnold Robbins. Learning the vi Editor. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, sixth edition, 1998. ISBN 1-56592-426-6. xvii + 327 pp. LCCN QA76.76.O63 L355 1998.

US$24.95. URL http://www.oreilly.com/catalog/9781565924260;
 http://www.oreilly.com/catalog/vi6.

Langridge:2005:DUM

[989] Stuart Langridge. DHTML utopia: modern Web design using JavaScript and DOM. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9579218-9-6. xiv + 318 pp. LCCN TK5105.888 .L36 2005. URL http:

//www.oreilly.com/catalog/9780957921894.

Langworth:2005:PTD

[990] Ian Langworth and Chromatic. Perl testing: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10092-2. xvii + 180 pp. LCCN QA76.73.P22 L35 2005. URL http://www.oreilly.com/

catalog/9780596100926.

Lansky:1998:TNT

[991] Doug Lansky, editor. There’s No Toilet Paper on the Road Less Traveled: The Best of Travel Humor and Misadventure. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-885211-27-9. 200 (est.) pp. LCCN ????

US$12.95. URL http://www.oreilly.com/catalog/9781885211279;
 http://www.oreilly.com/catalog/humor.

Larson:2001:DW

[992] Matt Larson, Cricket Liu, and Paul Albitz. DNS on Windows 2000. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00230-0. xvi + 334 pp. LCCN TK5105.875.I57 A44 2001; TK5105.875.I57 L346 2001. US$39.95. URL

http://safari.oreilly.com/0596002300;   http://www.oreilly.
com/catalog/9780596002305;   http://www.oreilly.com/catalog/
dnswin2.

Larson:2003:DWS

[993] Matt Larson, Cricket Liu, and Robbie Allen. DNS on Windows Server 2003. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00562-8. xv + 398 pp. LCCN TK 5105.8835 .L37 2003. URL http://

www.oreilly.com/catalog/9780596005627.

Lasky:1997:PSY

[994] Jane E. Lasky. Passport Singapore: Your Pocket Guide to Singaporean Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885073-38-0. 96 pp. LCCN HD58.7.L356 1997. US$6.95.

URL http://www.oreilly.com/catalog/9781885073389.

Laurie:1997:ADG

[995] Ben Laurie and Peter Laurie. Apache: The Definitive Guide. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-250-6. xvi + 255 pp. LCCN QA76.76.A7 L3 1997; TK5105.888

.L39 1997. US$34.95. URL http://www.ora.com/catalog/apache/;
http://www.oreilly.com/catalog/9781565922501;   http://www.
oreilly.com/catalog/apache.

Laurie:1999:ADG

[996] Ben Laurie and Peter Laurie. Apache: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-528-9. xvi + 369 pp. LCCN TK5105.888 .L39 1999; TK5105.8885.A63

L38 1999. US$34.95. URL http://www.ora.com/catalog/apache/;
http://www.oreilly.com/catalog/9781565925281;   http://www.
oreilly.com/catalog/apache.

Laurie:2002:ADG

[997] Ben Laurie and Peter Laurie. Apache: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00203-3. xviii + 568 pp. LCCN TK5105.888 .L39 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596002039;   http://
www.oreilly.com/catalog/apache3.

Lauwens:2019:TJH

[998] Ben Lauwens and Allen Downey. Think Julia: how to think like a computer scientist. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4920-4503-9. xviii + 276

pp. LCCN QA76.73.J85. URL http://proquest.safaribooksonline.
com/?fpi=9781492045021.

Lavigne:2004:BHI

[999] Dru Lavigne. BSD Hacks: 100 Industrial Strength Tips & Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00679-9. xvii + 427 pp. LCCN QA76.76.O63 L38 2004. US$24.95, CDN$36.95. URL http://

www.oreilly.com/catalog/9780596006792.

Lavin:2006:OOP

[1000] Peter Lavin. Object-oriented PHP: concepts, techniques, and code. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-077-1. xx + pp. LCCN QA76.73.P224 L38 2006eb; QA76.73.P224 L38 2006. URL

http://www.oreilly.com/catalog/9781593270773.

Lavy:1999:WNW

[1001] Matthew M. Lavy and Ashley J. Meggitt. Windows NT Workstation Configuration and Maintenance: Automated Workstation Management. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-613-7. xiv + 164 pp. LCCN QA76.76.O63 L373 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565926134;   http://
www.oreilly.com/catalog/ntmaint.

Lawrence:2007:SGA

[1002] William Lawrence. Setting up Google Apps standard edition. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51487-5. LCCN QA76.76.A65

L39 2007eb; QA76.76.A65. URL http://www.oreilly.com/catalog/
9780596514877.

Leber:1998:WNB

[1003] Jody Leber. Windows NT Backup & Restore: Help for system administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-272-7. xiv + 304 pp. LCCN QA76.9.B32 L43

1998. US$29.95. URL http://www.ora.com/catalog/backup/;
http://www.oreilly.com/catalog/9781565922723;   http://www.
oreilly.com/catalog/backup.

Lee:2003:WXU

[1004] Wei Meng Lee. Windows XP unwired: a guide for home, office, and the road. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00536-9. xiv + 267 pp. LCCN TK5105.78 .L44 2003. URL http://www.oreilly.

com/catalog/9780596005368.

Lee:2004:NCF

[1005] Wei Meng Lee. .NET Compact Framework pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00757-4. vi + 105 pp. LCCN QA76.76.M52 L44 2004.

URL http://www.oreilly.com/catalog/9780596007577.

Lee:2005:AAY

[1006] Wei Meng Lee. All about your iPod photo. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. v + 41

pp. LCCN TR267eb; TR267. URL http://www.oreilly.com/catalog/

.

Lee:2005:AND

[1007] Wei Meng Lee. ASP.NET 2.0: a developer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00812-0 (paperback). xx + 326 pp. LCCN

TK5105.8885.A26 L44 2005. URL http://www.oreilly.com/catalog/
9780596008123.

Lee:2005:VBJ

[1008] Wei Meng Lee. Visual Basic 2005 jumpstart. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10071-X. xiv + 197 pp. LCCN QA76.73.B3; QA76.73.B3 L44 2005; QA76.73.B3 L44 2005eb. URL http://www.

oreilly.com/catalog/9780596100711.

Lee:2006:RMX

[1009] Wei-Ming Lee. Running Mac OS X on Windows. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52843-4. LCCN QA76.76.O63. URL

http://www.oreilly.com/catalog/9780596528430.

Lee:2006:UCD

[1010] Wei Meng Lee. Use ClickOnce to deploy Windows applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52755-1. 32 pp. LCCN QA76.76.W56; QA76.76.W56 L44 2006eb. URL http://www.oreilly.

com/catalog/9780596527556.

Lehey:1995:PUS

[1011] Greg Lehey. Porting UNIX Software: From Download to Debug. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1995. ISBN 1-56592-126-7. xix + 515 pp. LCCN QA76.76.O63 L4472 1995.

US$29.95. URL http://www.ora.com/gnn/bus/ora/item/port.html;
 http://www.oreilly.com/catalog/9781565921269;   http://www.
oreilly.com/catalog/port.

Lehey:2003:CFD

[1012] Greg Lehey. The Complete FreeBSD: Documentation from the Source. O’Reilly Community Press, Sebastopol, CA, USA, fourth edition, 2003. ISBN 0-596-00516-4. xxxiii + 679 pp. LCCN QA76.76.O63 L44715 2003. US$44.09, CDN$69.95, UK£31.95. URL http://

press.oreilly.com/commpress.html;   http://www.oreilly.com/
catalog/9780596005160;   http://www.oreilly.com/catalog/cfreebsd/
desc.html.

Lehtinen:2006:CSB

[1013] Rick Lehtinen, Deborah Russell, and G. T. Gangemi. Computer Security Basics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00669-1. xii + 296 pp. LCCN QA76.9.A25 L44 2006. EUR 38.00

(freier Pr.). URL http://www.oreilly.com/catalog/9780596006693.

Lenssen:2008:GAH

[1014] Philipp Lenssen. Google apps hacks. Hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51588-X. xvii + 361 pp. LCCN TK5105.8885.G643 L46 2008.

Lenz:2004:OX

[1015] Evan Lenz, Mary McRae, and Simon St Laurent. Office 2003 XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00538-5. xv + 567 pp. LCCN HF5548.4.M525 L455 2004. URL http://www.oreilly.

com/catalog/9780596005382.

Lenz:2005:XPR

[1016] Evan Lenz. XSLT 1.0 pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10008-6 (paperback). 104 (est.) pp. LCCN QA76.73.X58 L46 2005. URL http://www.oreilly.com/

catalog/9780596100087.

Leonhard:1997:EA

[1017] Woody Leonhard, Lee Hudspeth, and Timothy-James J. Lee. Excel 97 Annoyances. A nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-309-X. xiii + 320 pp. LCCN HF5548.4.M523 L46

1997. US$24.95. URL http://www.ora.com/catalog/excelannoy/;
http://www.oreilly.com/catalog/9781565923096;   http://www.
oreilly.com/catalog/excelannoy.

Leonhard:1997:OA

[1018] Woody Leonhard, Lee Hudspeth, and T. J. Lee. Office 97 Annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-310-3. 200 (est.) pp. LCCN ???? US$24.95. URL http://www.ora.

com/catalog/officeannoy/;   http://www.oreilly.com/catalog/
9781565923102;   http://www.oreilly.com/catalog/officeannoy.

Leonhard:1997:WA

[1019] Woody Leonhard, Lee Hudspeth, and Timothy-James J. Lee. Word 97 Annoyances. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-308-1. xv + 338 pp. LCCN Z52.5.M52 L47

1997. US$24.95. URL http://www.ora.com/catalog/wordannoy/;
http://www.oreilly.com/catalog/9781565923089;   http://www.
oreilly.com/catalog/wordannoy.

Leonhard:1998:OA

[1020] Woody Leonhard, Lee Hudspeth, and Timothy-James Lee. Outlook Annoyances. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1998. ISBN 1-56592-384-7. xiv + 377 pp. LCCN HD69.T54 L46 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565923843;   http://
www.oreilly.com/catalog/outannoy.

Lerdorf:2000:PPR

[1021] Rasmus Lerdorf. PHP pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-769-9. vi + 114 pp. LCCN QA76.73.P224 L47 2000.

URL http://www.oreilly.com/catalog/9781565927698.

Lerdorf:2002:PP

[1022] Rasmus Lerdorf and Kevin Tatroe. Programming PHP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-610-2. xiii + 507 pp. LCCN QA76.73.P224 L48 2002. US$39.95. URL

http://safari.oreilly.com/1565926102;   http://www.oreilly.
com/catalog/9781565926103;   http://www.oreilly.com/catalog/
progphp.

Lerdorf:2002:PPR

[1023] Rasmus Lerdorf. PHP Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00402-8. vi + 132 pp. LCCN QA76.73.P224 L47 2002. US$14.95.

URL http://www.oreilly.com/catalog/9780596004026;   http://
www.oreilly.com/catalog/phppr2.

Lerdorf:2006:PP

[1024] Rasmus Lerdorf, Kevin Tatroe, and Peter MacIntyre. Programming PHP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00681-0. xvi + 521 pp. LCCN QA76.73.P224 L47 2006eb; QA76.73.P224; QA76.73.P224 L47 2006; QA76.73.P224 L48 2006. URL http://www.

oreilly.com/catalog/9780596006815.

Lerman:2009:PEFa

[1025] Julia Lerman. Programming Entity Framework. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15624-3. ???? pp. LCCN ????

Lerman:2009:PEFb

[1026] Julia Lerman. Programming Entity Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-52028-X. xxxiv + 792 pp. LCCN ????

Lerman:2010:PEF

[1027] Julia Lerman. Programming Entity Framework. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2010. ISBN 0-596-80727-9, 0-596-80726-0. ???? pp. LCCN ????

Lerner:2002:MTP

[1028] David Lerner and Aaron Freimark. Macintosh Troubleshooting Pocket Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00443-5. v + 72 pp. LCCN QA76.8.M3 L47 2002. US$12.95.

URL http://www.oreilly.com/catalog/9780596004439;   http://
www.oreilly.com/catalog/mactrblpg.

LeRoy:2007:ANA

[1029] Bertrand Le Roy and Matt Gibbs. ASP.NET AJAX UpdatePanel control. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52747-0. 61 pp. LCCN TK5105.8885.A52 L476 2007eb; TK5105.8885.A52. URL http:/

/www.oreilly.com/catalog/9780596527471.

Lesser:2005:PFC

[1030] Brian Lesser. Programming Flash Communication Server. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00504-0. xxxv + 816 pp. LCCN TK6570.M6 L38 2005.

URL http://www.oreilly.com/catalog/9780596005047.

Levine:1992:LY

[1031] John R. Levine, Tony Mason, and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1992. ISBN 1-56592-000-7. xxii + 366 pp. LCCN QA76.76.U84M37 1992. US$29.95.

URL http://www.oreilly.com/catalog/9781565920002;   http://
www.oreilly.com/catalog/lex.

Levine:2004:Q

[1032] John R. Levine. Qmail. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 1-56592-628-5. xii + 234 pp. LCCN TK5105.74.Q43 L48 2004. URL

http://www.oreilly.com/catalog/9781565926288.

Lewine:1991:PPG

[1033] Donald A. Lewine. POSIX programmer’s guide: writing portable UNIX programs with the POSIX.1 standard. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-73-0. xxvii + 607 pp. LCCN QA76.76.O63 L487 1991b. US$34.95. March 1994 printing with corrections, updates, and December 1991 Appendix G.

Lewis:2000:CEJ

[1034] Kevin Lewis. Creating Effective JavaHelp. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2000. ISBN 1-56592-719-2. xiii + 171 pp. LCCN QA76.73.J38 L495 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565927193;   http://
www.oreilly.com/catalog/creatingjavahelp.

Lewis:2000:MNP

[1035] Rita F. Lewis and Bill Fishman. Mac OS in a nutshell: a power user’s quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2000. ISBN 1-56592-533-5. xvii + 376 pp. LCCN QA76.76.O63 L489 2000. US$24.95.

URL http://www.oreilly.com/catalog/9781565925335;   http://
www.oreilly.com/catalog/macosnut.

Lewis:2009:HPP

[1036] Andrew Lewis. High performance Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15996-X. ???? pp. LCCN ????

Liberty:2001:PC

[1037] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2001. ISBN 0-596-00117-7. xix + 658 pp. LCCN QA76.73.C154

L432 2001. US$39.95. URL http://www.oreilly.com/catalog/
9780596001179;   http://www.oreilly.com/catalog/progcsharp.

Liberty:2002:LC

[1038] Jesse Liberty. Learning C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00376-5. xi + 354 pp. LCCN QA76.73.C154 L43

2002. US$34.95. URL http://safari.oreilly.com/0596003765;
http://www.oreilly.com/catalog/9780596003760;   http://www.
oreilly.com/catalog/learncsharp.

Liberty:2002:PAN

[1039] Jesse Liberty and Dan Hurwitz. Programming ASP.NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00171-1. xiii + 944 pp. LCCN QA76.73.C154 L52 2002; TK5105.8885.A26 L53

2002. US$49.95. URL http://safari.oreilly.com/0596001711;
http://www.oreilly.com/catalog/9780596001711;   http://www.
oreilly.com/catalog/progaspdotnet.

Liberty:2002:PC

[1040] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00309-9. xviii + 629 pp. LCCN QA76.73.C154 L53 2002. US$39.95. URL

http://safari.oreilly.com/0596003099;   http://www.oreilly.
com/catalog/9780596003098;   http://www.oreilly.com/catalog/
progcsharp2.

Liberty:2003:LVB

[1041] Jesse Liberty. Learning Visual Basic .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00386-2. xiii + 303 pp. LCCN QA76.73.B3

L52 2003. US$34.95. URL http://www.oreilly.com/catalog/
9780596003869;   http://www.oreilly.com/catalog/learnvbnet.

Liberty:2003:PAN

[1042] Jesse Liberty and Dan Hurwitz. Programming ASP.NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00487-7. xiii + 988 pp. LCCN TK5105.8885.A26 L53 2003. URL http://www.oreilly.

com/catalog/9780596004873.

Liberty:2003:PC

[1043] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00489-3. xviii + 689 pp. LCCN QA76.76.C154

L53 2003. URL http://www.oreilly.com/catalog/9780596004897.

Liberty:2003:PVB

[1044] Jesse Liberty and Dave Grundgeiger. Programming Visual Basic .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00438-9. xiv + 541 pp. LCCN QA76.73.B3 G776 2003. URL http://

www.oreilly.com/catalog/9780596004385.

Liberty:2004:PNW

[1045] Jesse Liberty and Dan Hurwitz. Programming .NET Windows applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00321-8. xv + 1228 pp. LCCN QA76.76.A65 L52 2004. URL http://www.oreilly.

com/catalog/9780596003210.

Liberty:2005:PAN

[1046] Jesse Liberty and Dan Hurwitz. Programming ASP.NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00916-X. 704 pp. LCCN TK5105.8885.A26; TK5105.8885.A26 L52 2005; TK5105.8885.A26 L52 2005eb. URL http://www.oreilly.com/

catalog/9780596009168.

Liberty:2005:PC

[1047] Jesse Liberty. Programming C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2005. ISBN 0-596-00699-

3. xx + 644 pp. LCCN QA76.73.C154 L53 2005. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/proquest.safaribooksonline.com/0596006993;   http://www.oreilly.
com/catalog/9780596006990.

Liberty:2005:PVB

[1048] Jesse Liberty. Programming Visual Basic 2005. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00949-6. xviii + 548 pp. LCCN QA76.73.B3

L523 2005. URL http://www.oreilly.com/catalog/9780596009496.

Liberty:2005:VCD

[1049] Jesse Liberty. Visual C# 2005: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00799-X. xv + 221 pp. LCCN

QA76.73.C154 L43 2005. URL http://www.oreilly.com/catalog/
9780596007997.

Liberty:2006:LC

[1050] Jesse Liberty and Brian MacDonald. Learning C# 2005: Get Started with C# 2005 and .NET Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 1-59059-589-0, 0-596-10209-7. ???? pp. LCCN ????

URL http://www.oreilly.com/catalog/9781590595893.

Liberty:2007:LAN

[1051] Jesse Liberty, Dan Hurwitz, and Brian MacDonald. Learning ASP.NET 2.0 with Ajax. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51397-6, 0-596-51397-6. xx + 498 pp. LCCN TK5105.8885.A26; QAeb; QA; QA L52 2007. URL

http://www.oreilly.com/catalog/9780596513979.

Liberty:2008:LAN

[1052] Jesse Liberty, Dan Hurwitz, and Brian MacDonald. Learning ASP.NET 3.5. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51845-5. xvii + 588 pp. LCCN QA76.625 .L53 2008.

Liberty:2008:LC

[1053] Jesse Liberty and Brian MacDonald. Learning C#. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-15501-8. 704 (est.) pp. LCCN QA76.73.C153 L53 2008.

Liberty:2008:PAN

[1054] Jesse Liberty, Dan Hurwitz, and Daniel Maharry. Programming ASP.NET 3.5. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2008. ISBN 0-596-15665-0. 904 (est.) pp. LCCN TK5105.8885.A26 L53 2008.

Liberty:2008:PC

[1055] Jesse Liberty and Donald Xie. Programming C# 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2008. ISBN 0-596-52743-8. xviii + 587

pp. LCCN QA76.73.C154. URL http://www.oreilly.com/catalog/
9780596527433.

Liberty:2008:PN

[1056] Jesse Liberty and Alex Horovitz. Programming .NET 3.5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52756-X. xvii + 455 pp. LCCN QA76.76.M52 L53 2008; QA76.76.M52 L53 2008eb; Internet; QA76.76.M52.

Libes:1994:EET

[1057] Don Libes. Exploring Expect: A Tcl-based Toolkit for Automating Interactive Programs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1994. ISBN 1-56592-090-2. xxxiii + 566 pp. LCCN QA76.755 .L52 1995. US$29.95. URL http://www.ora.com/gnn/

bus/ora/item/expect.html;   http://www.oreilly.com/catalog/
9781565920903;   http://www.oreilly.com/catalog/expect.

Libes:1995:EET

[1058] Don Libes. Exploring expect: a Tcl-based toolkit for automating interactive programs. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-090-2. xxxiii + 566 pp. LCCN QA76.755 .L52 1995; QA76.9.I58 L53 1995. US$32.95. URL http://www.ora.com/gnn/

bus/ora/item/expect.html;   http://www.oreilly.com/catalog/
9781565920903;   http://www.oreilly.com/catalog/expect.

Lidie:1998:PTP

[1059] Stephen Lidie. Perl… Tk Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-517-3. iii + 100 pp. LCCN QA76.73.P22 L54 1998. US$9.95.

URL http://www.oreilly.com/catalog/9781565925175;   http://
www.oreilly.com/catalog/perltkpr.

Lidie:2002:MPT

[1060] Stephen Lidie and Nancy Walsh. Mastering Perl/Tk: Graphical User Interfaces in Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-716-8. xx + 746 pp. LCCN QA76.73.L53 W363

2002. US$44.95. URL http://safari.oreilly.com/1565927168;
http://www.oreilly.com/catalog/9781565927162;   http://www.
oreilly.com/catalog/mastperltk.

Liguori:2008:JPG

[1061] Robert Liguori and Patricia Liguori. Java pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51419-0. xiii + 177 pp. LCCN QA76.73.J38 L545 2008.

Limoncelli:2005:ATM

[1062] Tom Limoncelli. The Art of Time Management for System Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00783-3 (paperback), 0-596-51950-8 (e-book), 0-596-55303-X (e-book). xxiii + 200 pp. LCCN HD69.T54 L546 2006.

Limoncelli:2006:TMS

[1063] Tom Limoncelli. Time management for system administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00783-3. xxiii + 200 pp. LCCN HD69.T54 L546 2006.

URL http://www.loc.gov/catdir/toc/fy0608/2006275166.html;
http://www.oreilly.com/catalog/9780596007836.

Lindberg:2008:IPO

[1064] Van Lindberg. Intellectual Property and Open Source. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51796-3 (paperback). xv + 371 pp. LCCN K1519.C6 L56 2008.

Linton:2007:BSP

[1065] Jennifer Linton. Beyond schemas: planning your XML model. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52770-5. LCCN QA76.76.H94;

QA76.76.H94 L56 2007eb. URL http://www.oreilly.com/catalog/
9780596527709.

Linzmayer:2004:ACD

[1066] Owen W. Linzmayer and Owen W. Apple confidential Linzmayer. Apple confidential 2.0: the definitive history of the world’s most colorful company. No Starch Press, San Francisco, CA, USA, second edition, 2004. ISBN 1-59327-010-0 (paperback). x + 323 pp. LCCN HD9696.2.U64 A674 2004. URL http://www.loc.gov/

catdir/toc/ecip047/2003017567;   http://www.loc.gov/catdir/
toc/ecip047/2003017567.html;   http://www.oreilly.com/catalog/
9781593270100.

Lischner:2000:DND

[1067] Ray Lischner. Delphi in a Nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2000. ISBN 1-56592-659-5. xi + 561 pp. LCCN QA76.76.D47 L56 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565926592;   http://
www.oreilly.com/catalog/delphi.

Lischner:2003:CN

[1068] Ray Lischner. C++ in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00298-X. xiv + 791 pp. LCCN

QA76.73.C153 L582 2003. URL http://www.oreilly.com/catalog/
9780596002985.

Lischner:2003:SPR

[1069] Ray Lischner. STL: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00556-3. vii + 120 pp. LCCN QA76.73.C153 L585

2004. URL http://www.oreilly.com/catalog/9780596005566.

Liu:1994:MII

[1070] Cricket Liu, Jerry Peek, Russ Jones, Bryan Buus, and Adrian Nye. Managing Internet Information Services: World Wide Web, Gopher, FTP, and more. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1994. ISBN 1-56592-062-7. xxxvi + 630 pp. LCCN TK5105.875.I57 M36 1994.

US$29.95. URL http://www.oreilly.com/catalog/9781565920620.

Liu:2003:DBC

[1071] Cricket Liu. DNS and BIND cookbook: Solutions and examples for system administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00410-9. xvi + 222 pp. LCCN TK5105.8835 .L58

2003. US$34.95. URL http://safari.oreilly.com/0596004109;
http://www.oreilly.com/catalog/9780596004101;   http://www.
oreilly.com/catalog/dnsbindckbk.

Liu:2006:DB

[1072] Cricket Liu and Paul Albitz. DNS and BIND. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-10057-4. xxi + 616 pp. LCCN TK5105.875.I57; TK5105.875.I57 A43 2006; TK5105.875.I57 A43

2006eb. URL http://www.oreilly.com/catalog/9780596100575.

Lloyd:2006:BYO

[1073] Ian Lloyd. Build your own web site the right way using HTML and CSS. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2006. ISBN 0-9752402-9-3. xxvi + 488 pp. LCCN TK5105.888 .L56x 2006; TK5105.888 .L56

2006. URL http://www.oreilly.com/catalog/9780975240298.

Lockhart:2004:NSH

[1074] Andrew Lockhart. Network security hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00643-8. xv + 298 pp. LCCN TK5105.59 .L63 2004. URL http:

//www.oreilly.com/catalog/9780596006433.

Lockhart:2007:NSH

[1075] Andrew Lockhart. Network security hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52763-2. 455 pp.

LCCN TK5105.59 .L63 2004. URL http://www.oreilly.com/catalog/
9780596527631.

Lomasky:1998:OS

[1076] Brian Lomasky and David C. Kreines. Oracle Scripts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-438-X. xvi + 185 pp. LCCN QA76.73.S67 L66 1998; QA76.9.D3 L648 1998. US$29.95.

URL http://www.oreilly.com/catalog/9781565924383;   http://
www.oreilly.com/catalog/oraclescrp. Covers Oracle8.

Lomax:1997:LV

[1077] Paul Lomax. Learning VBScript. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1997. ISBN 1-56592-247-6. xvi + 598 pp. LCCN QA76.73.V27 L 1997; QA76.73.V27 L65 1997. US$39.95.

URL http://www.oreilly.com/catalog/9781565922471;   http://
www.oreilly.com/catalog/vbscript. Includes CD-ROM.

Lomax:1998:VVN

[1078] Paul Lomax. VB & VBA in a Nutshell: The Language. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-358-8. xiv + 633 pp. LCCN QA76.73.B3 L647 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565923584;   http://
www.oreilly.com/catalog/vbanut.

Lomax:2003:VN

[1079] Paul Lomax, Matt Childs, and Ronald Petrusha. VBScript in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00488-5. xvii + 491 pp. LCCN QA76.76.V27 L66 2003. URL http://

www.oreilly.com/catalog/9780596004880.

Long:2008:CDR

[1080] Ben Long. The Canon Digital Rebel XSi companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52086-7. xiv + 260 + 24 pp. LCCN TR267 .L66 2008.

Long:2008:CEDa

[1081] Ben Long. The Canon EOS Digital Rebel XS/1000D companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-15452-6. 288 (est.) pp. LCCN TR267 .L66 2008.

Long:2008:CEDb

[1082] Ben Long. The Canon EOS Digital Rebel XSi/450D companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52086-7. xiv + 260 + 24 pp. LCCN TR267 .L66 2008.

Long:2019:RCP

[1083] J. D. Long and Paul Teetor. R Cookbook: Proven Recipes for Data Analysis, Statistics, and Graphics. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2019. ISBN 1-4920-4068-1 (paperback), 1-4920-4067-3 (e-book). xvii + 579 pp. LCCN QA276.45.R3 L66 2019.

Lott:2003:AC

[1084] Joey Lott. ActionScript cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00490-7. xxiii + 869 pp. LCCN TR897.7 .L68 2003.

URL http://www.oreilly.com/catalog/9780596004903.

Lott:2006:AC

[1085] Joey Lott, Darron Schall, and Keith Peters. ActionScript 3.0 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52695-4. xxvii + 556 pp. LCCN TR897.7; TR897.7 .L682eb. URL http://www.oreilly.

com/catalog/9780596526955.

Lott:2006:FC

[1086] Joey Lott, Jeffrey Bardzell, et al. Flash 8 cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10240-2. xxi + 510 pp. LCCN TR897.7; TR897.7 .L68 2006eb. URL http://www.oreilly.com/

catalog/9780596102401.

Lott:2008:PF

[1087] Joey Lott and Chafic Kazoun. Programming Flex 3. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51621-5. xx + 636 pp. LCCN QA76.625 .L68 2008.

Loudon:1999:MAC

[1088] Kyle Loudon. Mastering Algorithms with C. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-453-3. xvii + 540 pp. LCCN QA76.73.C15 L68 1999; QA76.9.A43 L68 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565924536;   http://
www.oreilly.com/catalog/masteralgoc.

Loudon:2003:CPR

[1089] Kyle Loudon. C++ pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00496-6. viii + 130 pp. LCCN QA76.73.C153 L68

2003. URL http://www.oreilly.com/catalog/9780596004965.

Loukides:1990:SPT

[1090] Mike Loukides. System Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1990. ISBN 0-937175-60-9. 336 pp. LCCN QA76.76.O63 L66 1992. US$29.95. URL http://www.oreilly.com/

catalog/9780937175606;   http://www.oreilly.com/catalog/spt.

Loukides:1990:UFP

[1091] Mike Loukides. UNIX for FORTRAN Programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1990. ISBN 0-937175-51-X. 264 pp. LCCN QA76.76.O63 L67 1990. US$24.95. URL http://www.oreilly.com/

catalog/9780937175514.

Loukides:1997:PGS

[1092] Michael Kosta Loukides and Andrew Oram. Programming with GNU Software. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-112-7. xiv + 244 pp. LCCN QA76.76.O63

L65 1997. US$39.95. URL http://www.ora.com/catalog/prognu/;
http://www.oreilly.com/catalog/9781565921122;   http://www.
oreilly.com/catalog/prognu. Includes CD-ROM.

Loukides:2008:MDG

[1093] Mike Loukides, editor. Maven: the definitive guide: [everything you need to know from ideation to deployment]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51733-5. xiv + 452 pp. LCCN QA76.76.D47 M38

2008. URL http://swbplus.bsz-bw.de/bsz278243878cov.htm.

Love:2007:LSP

[1094] Robert Love. Linux system programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00958-5 (paperback). xvii + 368 pp. LCCN QA76.76.O63 L6735 2007; QA76.76.O63 L69 2007. URL http://www.

oreilly.com/catalog/9780596009588.

Lowe-Norris:2000:WAD

[1095] Alistair G. Lowe-Norris. Windows 2000 Active Directory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-638-2. xix + 621 pp. LCCN QA76.76.O63 L69 2000 Bar. US$39.95.

URL http://www.oreilly.com/catalog/9781565926387;   http://
www.oreilly.com/catalog/win2000ads.

Lowe:2005:HNM

[1096] Scott Lowe. Home networking: the missing manual. The missing manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00558-X (paperback). xvi + 245 pp. LCCN TK5105.75 .L69 2005. URL

http://www.oreilly.com/catalog/9780596005580.

Lowy:2001:CNC

[1097] Juval Löwy. COM and .NET Component Services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2001. ISBN 0-596-00103-7. xx + 362 pp. LCCN QA76.9.A73 L69 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596001032;   http://
www.oreilly.com/catalog/comdotnetsvs.

Lowy:2003:PNC

[1098] Juval Lowy. Programming .NET components. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00347-1. xvii + 459 pp. LCCN QA76.76.M52

L69 2003. URL http://www.oreilly.com/catalog/9780596003470.

Lowy:2005:PNC

[1099] Juval Löwy. Programming .NET components. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00762-0. xviii + 624 pp. LCCN QA76.625 .L69 2005; QA76.625; QA76.76.C66 L69

2005; QA76.625 .L69 2005eb. URL http://www.oreilly.com/catalog/
9780596007621.

Lowy:2007:PWS

[1100] Juval Löwy. Programming WCF services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52699-7. xxii + 610 pp. LCCN QA76.76.A65 L6885 2007; QA76.76.A65 L6885 2007eb; QA76.76.A65; TK5105.88813; TK5105.88813eb. URL http://www.oreilly.com/

catalog/9780596526993.

Lowy:2008:PWS

[1101] Juval Löwy. Programming WCF services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-15721-5. 750 (est.) pp. LCCN QA76.76.D47 L69 2008.

Lowy:2010:PWS

[1102] Juval Löwy. Programming WCF services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 1-4493-8247-9, 0-596-80548-9. ???? pp. LCCN ????

Loy:2002:JS

[1103] Marc Loy, Robert Eckstein, David Wood, James E. lliott, and Brian Cole. Java Swing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 J3855 2003.

US$54.95. URL http://www.oreilly.com/catalog/jswing2.

Loy:2003:JS

[1104] Marc Loy, Robert Eckstein, David Wood, James Elliott, and Brian Cole, editors. Java Swing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 J3855 2003.

US$54.95. URL http://www.oreilly.com/catalog/9780596004088;
 http://www.oreilly.com/catalog/jswing2.

Loy:2006:CRA

[1105] Marc Loy and Tom Mornini. Capistrano and the Rails application lifecycle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52962-7. LCCN QA76.64 .L69 2006eb; QA76.64. URL http://www.oreilly.

com/catalog/9780596529628.

Lucas:2003:AOU

[1106] Michael Lucas. Absolute OpenBSD: Unix for the practical paranoid. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-99-9. xviii + 489 pp. LCCN QA76.76.O63 L835 2003eb. URL http://www.oreilly.

com/catalog/9781886411999.

Lucas:2004:CRD

[1107] Michael Lucas. Cisco routers for the desperate: router management, the easy way. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-049-6. xii + 129 pp. LCCN TK5105.543 .L83 2004eb; TK5105.543

.L83 2004. URL http://www.oreilly.com/catalog/9781593270490.

Lucas:2018:AF

[1108] Michael W. Lucas. Absolute FreeBSD. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2018. ISBN 1-59327-892-6. LCCN ????

Lunde:1993:UJI

[1109] Ken Lunde. Understanding Japanese information processing = [Nihongo jōhō shori]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-043-0. xxxii + 435 pp. LCCN PL524.5 .L86 1993. US$29.95.

URL http://www.oreilly.com/catalog/9781565920439. See [2157].

Lunde:1999:CIP

[1110] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-224-7. xxiv + 1101 pp. LCCN PL1074.5 .L85 1999.

US$64.95. URL http://www.oreilly.com/catalog/9781565922242;
 http://www.oreilly.com/catalog/cjkvinfo/.

Lundh:2001:PSL

[1111] Fredrik Lundh. Python Standard Library: Annotated reference for Python 2.0. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 0-596-00096-0. xvi + 281 pp. LCCN QA76.73.P98 L86 2001.

US$29.95. URL http://www.oreilly.com/catalog/9780596000967;
 http://www.oreilly.com/catalog/pythonsl.

Lutz:1996:PP

[1112] Mark Lutz. Programming Python. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-197-6. xxii + 880 pp. LCCN QA76.73.P98 L88 1996. US$44.95.

URL http://www.oreilly.com/catalog/9781565921979;   http://
www.oreilly.com/catalog/python.

Lutz:1998:PPR

[1113] Mark Lutz. Python Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-500-9. iii + 75 pp. LCCN QA76.73.P98 L882 1998.

US$6.95. URL http://www.oreilly.com/catalog/9781565925007;
http://www.oreilly.com/catalog/pythonpr/.

Lutz:1999:LP

[1114] Mark Lutz and David Ascher. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-464-9. xvi + 366 pp. LCCN QA76.73.P98

L8798 1999. US$29.95. URL http://www.oreilly.com/catalog/
9781565924642;   http://www.oreilly.com/catalog/lpython/.

Lutz:2001:PPO

[1115] Mark Lutz. Programming Python: Object-Oriented Scripting. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, March 2001. ISBN 0-596-00085-5. xxxvii + 1255 pp. LCCN QA76.73.P98 L88 2001. US$54.95.

URL http://www.oreilly.com/catalog/9780596000851;   http://
www.oreilly.com/catalog/python2. Includes CD-ROM.

Lutz:2002:PPR

[1116] Mark Lutz. Python Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00189-4. iv + 124 pp. LCCN QA76.73.P98 L89 2002. US$11.95. URL http://safari.oreilly.com/

0596001894;   http://www.oreilly.com/catalog/9780596001896;
 http://www.oreilly.com/catalog/pythonpr2.

Lutz:2004:LP

[1117] Mark Lutz and David Ascher. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00281-5. xxvi + 591 pp. LCCN QA76.73.P98 L877 2004. URL http://www.oreilly.com/

catalog/9780596002817.

Lutz:2005:PPR

[1118] Mark Lutz. Python pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005.

ISBN 0-596-00940-2 (paperback). ix + 148 pp. LCCN QA76.73.P98 L89

2005. URL http://www.oreilly.com/catalog/9780596009403.

Lutz:2006:PP

[1119] Mark Lutz. Programming Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00925-9. xlii + 1552 pp. LCCN QA76.73.P98 L88 2006eb; QA76.73.P98 L88 2006; QA76.73.P98. URL

http://www.oreilly.com/catalog/9780596009250.

Lutz:2007:EPM

[1120] Mark Lutz, David Ascher, and Dinu C.. Gherman. Einführung in Python: [moderne OO-Programmierung; behandelt Python 2.5]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 3-89721-488-1. xxvii + 624

pp. LCCN ???? EUR 39.90. URL http://www.gbv.de/dms/ilmenau/
toc/527924601.PDF.

Lutz:2008:LP

[1121] Mark Lutz. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2008. ISBN 0-596-51398-4, 0-596-51398-4. xliv + 700 pp. LCCN QA76.73.P98 L877 2008; QA76.73.P98 L877 2008eb; QA76.73.P98 L8798 2008; QA76.73.P98. URL http://www.oreilly.

com/catalog/9780596513986.

Lutz:2009:LPa

[1122] Mark Lutz. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2009. ISBN 0-596-15806-8 (paperback). xlix + 1160 pp. LCCN QA76.73.P98 L877 2009.

Lutz:2009:LPb

[1123] Mark Lutz. Learning Python. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2009. ISBN 0-596-80539-X. ???? pp. LCCN QA76.73.P98 L877 2009.

Lutz:2013:LP

[1124] Mark Lutz. Learning Python. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fifth edition, 2013. ISBN 1-4493-5573-0. l + 1540 pp. LCCN QA76.73.P98 L877 2013. URL http:

//proquestcombo.safaribooksonline.com/9781449355722.

Mabe:2005:BH

[1125] Dave Mabe. Blackberry hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10115-5. xviii + 307 pp. LCCN QA76.8.B53

M33 2005eb; QA76.8.B53. URL http://www.oreilly.com/catalog/
9780596101152.

MacCarty:2002:LDG

[1126] Bill MacCarty. Learning Debian GNU Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-705-2. ???? pp. LCCN QA76.76.O63 M372 1999. US$34.95. Includes CD-ROM.

MacDonald:2005:BVB

[1127] Matthew MacDonald. The Book of Visual Basic 2005: .NET Insight for classic VB Developers. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-074-7. xviii + 474 pp. LCCN QA76.73.B3 M282 2005eb.

URL http://www.oreilly.com/catalog/9781593270742.

MacDonald:2005:CWS

[1128] Matthew MacDonald. Creating Web sites: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00842-2. ix + 548 pp. LCCN TK5105.8888 .M33 2006eb; TK5105.8888 .M33 2006; TK5105.8888. URL

http://www.oreilly.com/catalog/9780596008420.

MacDonald:2005:EMM

[1129] Matthew MacDonald. Excel: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00664-0. xix + 770 pp. LCCN HF5548.4.M523 M33 2005; HF5548.4.M523; HF5548.4.M523

M33 2005eb. URL http://www.oreilly.com/catalog/9780596006648.

MacDonald:2005:VBD

[1130] Matthew MacDonald. Visual Basic 2005: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00726-4. xvi + 243 pp. LCCN QA76.73.B3 M2863 2005. URL http://www.oreilly.

com/catalog/9780596007263.

MacDonald:2006:AMM

[1131] Matthew MacDonald. Access 2007: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52760-8. xv + 734 pp. LCCN HF5548.4.M525; HF5548.4.M525eb. URL http://www.

oreilly.com/catalog/9780596527600.

MacDonald:2006:ESM

[1132] Matthew MacDonald and Matthew Excel MacDonald. Excel for starters: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10154-6. xiii + 380 pp. LCCN HF5548.4.M523 M185 2006; HF5548.4.M523 M185

2006eb; HF5548.4.M523. URL http://www.oreilly.com/catalog/
9780596101541.

MacDonald:2007:ASM

[1133] Matthew MacDonald. Access 2007 for starters: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52833-7. xiv + 377 pp. LCCN HF5548.4.M525; HF5548.4.M525eb. URL http://www.

oreilly.com/catalog/9780596528331.

MacDonald:2007:EMM

[1134] Matthew MacDonald. Excel 2007: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52759-4. xxi + 831 pp. LCCN HF5548.4.M523eb; HF5548.4.M523. URL http://www.

oreilly.com/catalog/9780596527594.

MacDonald:2007:ESM

[1135] Matthew MacDonald. Excel 2007 for starters: the missing manual: your best friend for answers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52832-9. xiii + 332 pp. LCCN HF5548.4.M523eb; HF5548.4.M523.

URL http://www.oreilly.com/catalog/9780596528324.

MacHaffie:2005:PTH

[1136] Scott MacHaffie. Palm and Treo hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10054-X. 256 pp. LCCN QA76.5 .M18745 2005eb;

QA76.5. URL http://www.oreilly.com/catalog/9780596100544.

Macrakis:1996:UWT

[1137] Stavros Macrakis, editor. Understanding Web Technology. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-202-6. 400 pp. LCCN ????

US$24.95. URL http://www.oreilly.com/catalog/9781565922020.

MacVittie:2006:XN

[1138] Lori A. MacVittie. XAML in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52673-3. xv + 284 pp. LCCN QA76.76.D47; QA76.76.D47 M33 2006eb; QA76.76.D47 M33 2006. URL

http://www.oreilly.com/catalog/9780596526733.

Mahapatra:2000:OPP

[1139] Tushar Mahapatra and Sanjay Mishra. Oracle Parallel Processing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-701-X. xvii + 268 pp. LCCN QA76.58 .M34 2000. US$32.95.

URL http://www.oreilly.com/catalog/9781565927018;   http://
www.oreilly.com/catalog/oraclepp.

Mahemoff:2006:ADP

[1140] Michael Mahemoff. Ajax Design Patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10180-5. xv + 635 pp. LCCN TK5105.8885.A52

M24 2006. EUR 38.00. URL http://www.oreilly.com/catalog/
9780596101800.

Mahmoud:2002:LWJ

[1141] Qusay H. Mahmoud. Learning Wireless Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00243-2. xiii + 245 pp. LCCN QA76.73.J38 M323 2002; TK5103.2 .M33

2002. US$34.95. URL http://safari.oreilly.com/0596002432;
http://www.oreilly.com/catalog/9780596002435;   http://www.
oreilly.com/catalog/wirelessjava. Help for new J2ME developers.

Covers Motorola and Palm hardware.

Malamud:1993:ENC

[1142] Carl Malamud. European Networking: Carl Malamud interviews Glenn Kowack and Rob Blokzijl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-999-3. LCCN ???? US$9.95. URL http://www.

oreilly.com/catalog/9781565929999. Audio tape: 1 hour.

Malamud:1993:NIO

[1143] Carl Malamud. Networked Information and Online Libraries: Carl Malamud interviews Peter Deutsch and Cliff Lynch. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-998-5. LCCN ???? US$9.95.

URL http://www.oreilly.com/catalog/9781565929982. Audio tape:

1 hour.

Malamud:1993:SNC

[1144] Carl Malamud. Security and Networks: Carl Malamud interviews Jeff Schiller and John Romkey. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-997-7. LCCN ???? US$9.95. URL http://www.

oreilly.com/catalog/9781565929975. Audio tape: 1 hour.

Malamud:1994:FIP

[1145] Carl Malamud. The Future of the Internet Protocol: Carl Malamud interviews Steve Deering, Bob Braden, Christian Huitema, Bob Hinden, Peter Ford, Steve Casner, Bernhard Stockman and Noel Chiappa. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1994. ISBN 1-56592-996-

9. LCCN ???? US$24.95. URL http://www.oreilly.com/catalog/

9781565929968. Audio tape: 4 hours.

Malamud:1994:GNO

[1146] Carl Malamud. Global Network Operations: Carl Malamud interviews Brian Carpenter, Bernhard Stockman, Mike O’Dell and Geoff Huston. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1994. ISBN 1-56592-993-

4. LCCN ???? US$16.95. URL http://www.oreilly.com/catalog/

9781565929937. Audio tape: 2 hours.

Malamud:1994:MIN

[1147] Carl Malamud. Mobile IP Networking: Carl Malamud interviews Phil Karn and Jun Murai. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1994. ISBN 1-56592-994-2. LCCN ???? US$9.95. URL http://www.oreilly.

com/catalog/9781565929944. Audio tape: 1 hour.

Malat:1996:PMY

[1148] Randy Malat. Passport Mexico: Your Pocket Guide to Mexican Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-885073-30-5. 96 pp. LCCN HF5389.M35 1996. US$6.95. URL

http://www.oreilly.com/catalog/9781885073303.

Malhotra:2002:IR

[1149] Ravi Malhotra. IP Routing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00275-0. xv + 219 pp. LCCN TK5105.585 .M35

2002. US$34.95. URL http://safari.oreilly.com/0596002750;
http://www.oreilly.com/catalog/9780596002756;   http://www.
oreilly.com/catalog/iprouting.

Mangano:2002:XC

[1150] Sal Mangano. XSLT cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002.

ISBN 0-596-00372-2. xiv + 654 pp. LCCN QA76.73.X58 M36 2003. URL

http://www.oreilly.com/catalog/9780596003722.

Mangano:2006:XC

[1151] Sal Mangano. XSLT cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00974-7 (paperback). xix + 751 pp. LCCN QA76.73.X58; QA76.73.X58 M36 2005eb; QA76.73.X58 M36 2005. URL

http://www.oreilly.com/catalog/9780596009748.

Mangano:2010:MC

[1152] Sal Mangano. Mathematica Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-52099-9. 782 (est.) pp. LCCN ????

Manske:1999:FTF

[1153] Laura Manske, editor. Family Travel: The Farther You Go, the Closer You Get. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-33-3. 368 pp. LCCN G151.F35 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211330;   http://
www.oreilly.com/catalog/family.

Mantaro:2005:FMM

[1154] Jessica Mantaro. FrontPage 2003: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00950-X. xx + 411 pp. LCCN TK5105.888.M53; TK5105.888.M53 M358 2005; TK5105.888.M535 M358 2005; TK5105.888.M53 M358 2005eb. URL

http://www.oreilly.com/catalog/9780596009502.

Margulies:2000:UJT

[1155] Edwin Margulies, editor. Understanding Java Telephony. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-199-00626-2. ???? pp. LCCN ????

URL http://www.oreilly.com/catalog/.

Marinacci:2005:SHT

[1156] Joshua Marinacci and Chris Adamson. Swing Hacks: Tips and Tools for Killer GUIs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00907-0. xix + 519 pp. LCCN QA76.73.J38 M37 2005. US$9.95, CAN$41.95, UK£20.95. URL

http://www.oreilly.com/catalog/9780596009076.

Marino:2004:GBF

[1157] Paul Marino. 3D game-based filmmaking: the art of machinima. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2004. ISBN 1-932111-85-9. xxvi + 470 pp. LCCN T385 .M27 2004. URL http://www.oreilly.

com/catalog/9781932111859.

Markel:2003:SAN

[1158] Scott Markel and Darryl León. Sequence analysis in a nutshell: a guide to tools and databases. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00494-X. xiv + 286 pp. LCCN QH441.2 .M37 2003. URL http:/

/www.oreilly.com/catalog/9780596004941.

Marrs:2006:JWP

[1159] Tom Marrs and Scott Davis. JBoss at work: a practical guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00734-5. xvi + 287 pp. LCCN TK5105.8885.J43; TK5105.8885.J43 M37 2006eb. URL http://www.

oreilly.com/catalog/9780596007348.

Marschke:2008:JER

[1160] Doug Marschke and Harry Reynolds. JUNOS enterprise routing. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51442-5. xxv + 783 pp. LCCN TK5105.543 .M367 2008.

Marshall:2006:WSR

[1161] Kevin Marshall. Web services on rails. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52796-9. 32 pp. LCCN TK5105.88813;

TK5105.88813 .M37 2006eb. URL http://www.oreilly.com/catalog/
9780596527969.

Martelli:2002:PC

[1162] Alex Martelli and David Ascher, editors. Python Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00167-3. xxix + 574 pp. LCCN QA76.73.P98 P983 2002; QA76.73.P98 P95

2002. US$39.95. URL http://safari.oreilly.com/0596001673;
http://www.oreilly.com/catalog/9780596001674;   http://www.
oreilly.com/catalog/pythoncook.

Martelli:2003:PN

[1163] Alex Martelli. Python in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00188-6. xv + 636 pp. LCCN QA76.73.P98 M37 2003.

URL http://www.oreilly.com/catalog/9780596001889.

Martelli:2005:PC

[1164] Alex Martelli, Anna Martelli Ravenscroft, and David Ascher, editors. Python cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00797-3. xxxiii + 807 pp. LCCN QA76.73.P98 P983 2005. URL http://www.

oreilly.com/catalog/9780596007973.

Martelli:2006:PN

[1165] Alex Martelli. Python in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10046-9. xiv + 695 pp. LCCN QA76.73.P98 M37 2006eb; QA76.73.P98 M37 2006; QA76.73.P98. URL

http://www.oreilly.com/catalog/9780596100469.

Martelli:2006:PPE

[1166] Alex Martelli, Anna Martelli Ravensroft, and David Ascher. Python par l’exemple. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 2-84177-379-5. xxv + 525 pp. LCCN ???? French translation by Éric Jacoboni.

Martelli:2007:PC

[1167] Alex Martelli. Python en concentré. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 2-84177-452-X. xvi + 802 pp. LCCN ???? French translation by Éric Jacoboni and Yann Serra.

Mason:1990:LY

[1168] Tony Mason and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-49-8. xviii + 216 pp. LCCN QA76.76.O63

M37 1990. URL http://www.oreilly.com/catalog/9780937175491.

Mason:2006:PVC

[1169] Mike Mason. Pragmatic Version Control using Subversion, volume 1 of Pragmatic starter kit. Pragmatic, Lewisville, TX, USA, second edition, 2006. ISBN 0-9776166-5-7 (paperback). xiii + 231 pp. LCCN

QA76.76.D47 M3653 2006. URL http://www.loc.gov/catdir/toc/
fy0702/2006281498.html;   http://www.oreilly.com/catalog/9780977616657.

Massol:2005:MDN

[1170] Vincent Massol, Timothy O’Brien, and Michael Kosta Loukides. Maven: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00750-7. xxv + 191 pp. LCCN QA76.76.D47 M37 2005. URL

http://www.oreilly.com/catalog/9780596007508.

Matloff:2008:ADG

[1171] Norman S. Matloff and P. J. Salzman. The art of debugging with GDB and DDD and Eclipse. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 1-59327-174-3. 200 (est.) pp. LCCN QA76.9.D43 M38 2008.

Matsumoto:2002:RND

[1172] Yukihiro Matsumoto. Ruby in a Nutshell: a Desktop Quick Reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00214-9. xi + 204 pp. LCCN QA76.64 M385 2002; QA76.73.R82 M37

2002. US$24.95. URL http://safari.oreilly.com/0596002149;
http://www.oreilly.com/catalog/9780596002145;   http://www.

oreilly.com/catalog/ruby. Translated from Japanese by David L. Reynolds, Jr.

Matzan:2007:FCC

[1173] Jem Matzan. The FreeBSD 6.2 crash course. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51016-0. LCCN QA76.754.T53; QA76.754.T53

M38 2007eb. URL http://www.oreilly.com/catalog/9780596510169.

Matzan:2007:OCC

[1174] Jem Matzan. The OpenBSD 4.0 crash course. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51015-2. LCCN QA76.76.O63; QA76.76.O63

M38 2007eb. URL http://www.oreilly.com/catalog/9780596510152.

Mauro:2001:ES

[1175] Douglas R. Mauro and Kevin J. (Kevin James) Schmidt. Essential SNMP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2001. ISBN 0-596-00020-0. xiv + 313 pp. LCCN TK5105.583 .M38 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596000202;   http://
www.oreilly.com/catalog/esnmp.

Mauro:2005:ES

[1176] Douglas R. Mauro and Kevin J. (Kevin James) Schmidt. Essential SNMP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00840-6. xv + 442 pp. LCCN TK5105.583; TK5105.583 .M38 2005eb; TK5105.583 .M38 2005. URL http://www.oreilly.com/

catalog/9780596008406.

Mayer:1997:HTL

[1177] Musa Mayer. Holding Tight, Letting Go: living with metastatic breast cancer. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-254-9. xxix + 478 pp. LCCN RC280.B8 M353 1997. US$19.95. URL http://www.

ora.com/catalog/holding/;   http://www.oreilly.com/catalog/
9781565922549;   http://www.oreilly.com/catalog/holding.

Mayer:1998:ABC

[1178] Musa Mayer. Advanced Breast Cancer: a Guide to Living with Metastatic Disease. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1998. ISBN 1-56592-522-X. xxix + 503 pp. LCCN RC280.B8

M353 1998. US$19.95. URL http://www.oreilly.com/catalog/
9781565925229;   http://www.oreilly.com/catalog/advbc.

McCallum:2007:MRB

[1179] Ethan McCallum. Managing RPM-based systems with Kickstart and Yum. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51382-8. LCCN QA76.76.O63; QA76.76.O63 M34 2007eb. URL http://www.

oreilly.com/catalog/9780596513825.

McCarty:1999:LDG

[1180] Bill McCarty. Learning Debian GNU/Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-705-2. xiii + 343 pp. LCCN QA76.76.O63

M3758 1999. US$34.95. URL http://www.oreilly.com/catalog/
9781565927056;   http://www.oreilly.com/catalog/debian. In-

cludes CD-ROM.

McCarty:1999:LRL

[1181] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-627-7. xiii + 378 pp. LCCN QA76.76.O63

M379 1999. US$34.95. URL http://www.oreilly.com/catalog/
9781565926271;   http://www.oreilly.com/catalog/redhat.

McCarty:2002:LRL

[1182] Bill McCarty. Learning Red Hat Linux: a Guide to Red Hat Linux for New Users. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00071-5. xx + 346 pp. LCCN QA76.76.O63 M376

2002. US$34.95. URL http://safari.oreilly.com/0596000715;
http://www.oreilly.com/catalog/9780596000714;   http://www.
oreilly.com/catalog/redhat2.

McCarty:2003:LRL

[1183] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00469-9. xvi + 319 pp. LCCN

QA76.76.O63 M376 2003. URL http://www.oreilly.com/catalog/
9780596004699.

McCarty:2004:LRE

[1184] Bill McCarty. Learning Red Hat Enterprise Linux and Fedora. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2004. ISBN 0-596-00589-X. xvii + 326 pp. LCCN QA76.76.O63 M376 2004. URL http://www.oreilly.

com/catalog/9780596005894.

McCarty:2005:SNO

[1185] Bill McCarty. SELINUX: NSA’s open source Security Enhanced Linux. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00716-7. xiii + 238 pp. LCCN TK5105.59

.M37 2005. URL http://www.oreilly.com/catalog/9780596007164.

McCauley:1995:TTS

[1186] Lucy McCauley, editor. Travelers’ Tales Spain. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1995. ISBN 1-885211-07-4. xviii + 452 pp. LCCN DP43.2 .S63 1995. US$17.95.

URL http://www.ora.com/gnn/bus/ora/item/spain.html;   http:
//www.oreilly.com/catalog/9781885211071;   http://www.oreilly.
com/catalog/spain.

McCauley:1997:TTW

[1187] Lucy McCauley, editor. Traveler’s Tales: Women in the Wild. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-21-X. 350 (est.) pp.

LCCN ???? US$17.95. URL http://www.ora.com/catalog/wild/;
http://www.oreilly.com/catalog/9781885211217;   http://www.
oreilly.com/catalog/wild.

McClanahan:2012:GSP

[1188] Nicholas McClanahan and Jeff Ledger. Getting started with the Propeller: learn to use the powerful eight-core microcontroller. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1650-6 (paperback). 1 v. pp. LCCN ????

McClelland:2004:AIC

[1189] Deke McClelland. Adobe InDesign CS one-on-one. Deke Press/O’Reilly, Sebastopol, CA, USA, 2004. ISBN 0-596-00736-1 (paperback). xxii + 473 pp. LCCN Z253.532.A34 M435 2004. URL http://www.oreilly.

com/catalog/9780596007362.

McClelland:2004:APC

[1190] Deke McClelland. Adobe Photoshop CS one-on-one. Deke Press, Sebastopol, CA, USA, 2004. ISBN 0-596-00618-7. xxii + 465 pp.

LCCN T385 .M37777 2004. URL http://www.oreilly.com/catalog/
9780596006181.

McClelland:2005:APC

[1191] Deke McClelland. Adobe Photoshop CS2 one-on-one. Deke Press, Sebastopol, CA, USA, 2005. ISBN 0-596-10096-5. xxvi + 475 pp.

LCCN T385 .M37777 2005. URL http://www.oreilly.com/catalog/
9780596100964.

McClelland:2005:APE

[1192] Deke McClelland and Galen Fott. Adobe Photoshop Elements 4 oneon-one. Deke Press, Sebastopol, CA, USA, 2005. ISBN 0-596-10098-1 (paperback). xxii + 407 pp. LCCN TR267.5.A33 M44 2005.

McClelland:2005:PEW

[1193] Deke McClelland. Photoshop Elements 3: for Windows one-on-one. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00844-9. xxii + 393 pp. LCCN TR267 .M33 2005. URL http://www.oreilly.com/

catalog/9780596008444.

McClelland:2007:APC

[1194] Deke McClelland. Adobe Photoshop CS3: one-on-one. Deke Press/O’Reilly, Sebastopol, CA, USA, 2007. ISBN 0-596-52975-9. xxviii + 515 pp. LCCN T385 .M37776 2007eb; T385. URL http://www.

oreilly.com/catalog/9780596529758.

McClelland:2008:AIC

[1195] Deke McClelland. Adobe InDesign CS3 one-on-one. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52976-7. 560 pp. LCCN Z286.D47 K93 2008eb; Z286.D47. URL http://www.oreilly.com/

catalog/9780596529765.

McClelland:2009:APCb

[1196] Deke McClelland. Adobe Photoshop CS4 one-on-one. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15610-3, 0-596-52189-8. 513 (est.) pp. LCCN TR267 .M33 2008.

McConnell:2001:BCG

[1197] Brian McConnell. Beyond Contact: a Guide to SETI and Communicating with Alien Civilizations. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2001. ISBN 0-596-00037-5. vi + 417 pp. LCCN QB54 .M23 2001.

US$24.95. URL http://www.oreilly.com/catalog/9780596000370;
 http://www.oreilly.com/catalog/alien.

McCool:2007:RGO

[1198] Matthew McCool. Reaching a global online audience. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51373-9. LCCN HF5415.1265

M34 2007eb; HF5415.1265. URL http://www.oreilly.com/catalog/
9780596513733.

McDaniel:2002:TPR

[1199] Jim McDaniel and Patrick McGrath. TOAD Pocket Reference for Oracle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00337-4. 128 (est.)

pp. LCCN ???? US$12.95. URL http://www.oreilly.com/catalog/
9780596003371;   http://www.oreilly.com/catalog/tdpr.

McDaniel:2005:TPR

[1200] Jim McDaniel, Jeff Smith, Patrick McGrath, and Bert Scalzo. TOAD pocket reference for Oracle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00971-2 (paperback). vi + 122 pp. LCCN QA76.9.D3 M395 2005. URL http://www.oreilly.com/

catalog/9780596009717.

McFarland:2001:DMM

[1201] David Sawyer McFarland. Dreamweaver 4: The Missing Manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, July 2001. ISBN 0-596-00097-9. xii + 516 pp. LCCN TK5105.8885.D74 M39 2001.

US$24.95. URL http://www.oreilly.com/catalog/9780596000974;
 http://www.oreilly.com/catalog/dreamweavermm.

McFarland:2002:DMM

[1202] David Sawyer McFarland. Dreamweaver MX: The Missing Manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2002. ISBN 0-596-00349-8. xii + 779 pp. LCCN TK5105.8885.D74 M3923 2002. US$34.95. URL

http://www.oreilly.com/catalog/9780596003494;   http://www.
oreilly.com/catalog/drmweavmxtmm.

McFarland:2004:DM

[1203] David Sawyer McFarland. Dreamweaver MX 2004. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, third edition, 2004. ISBN 0-596-00631-4. xii + 836 pp. LCCN TK5105.8885.D74

M3923 2004. URL http://www.oreilly.com/catalog/9780596006310.

McFarland:2006:CMM

[1204] David Sawyer McFarland. CSS: the missing manual. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-52687-3. xv + 476 pp. LCCN TK5105.888; TK5105.888 .M44 2006eb. URL http://www.oreilly.

com/catalog/9780596526870.

McFarland:2006:DMM

[1205] David Sawyer McFarland and David Pogue. Dreamweaver 8: the missing manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10056-6. xiv + 936 pp. LCCN TK5105.8885.D74 M3923 2006eb; TK5105.8885.D74. URL http://www.

oreilly.com/catalog/9780596100568.

McFarland:2007:DC

[1206] David Sawyer McFarland. Dreamweaver CS3. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2007. ISBN 0-596-51043-8, 0-596-51043-8. xxii + 995 pp. LCCN TK5105.8885.D74 M3922 2007eb; TK5105.8885.D74 M3922 2007; TK5105.8885.D74. URL

http://www.oreilly.com/catalog/9780596510435.

McFarland:2008:JMM

[1207] David Sawyer McFarland. JavaScript: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2008. ISBN 0-596-51589-8. xiv + 528 pp. LCCN QA76.73.J39

M388 2008; **See. URL http://proquest.safaribooksonline.com/
9780596515898.

McFarland:2009:DC

[1208] David Sawyer McFarland. Dreamweaver CS4. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2009. ISBN 0-596-52292-4 (paperback). xxii + 1064 pp. LCCN TK5105.8885.D74 M39225 2009.

McFarland:2010:DC

[1209] David Sawyer McFarland. Dreamweaver CS5. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2010. ISBN 1-4493-8181-2 (paperback). xviii + 1074 pp. LCCN TK5105.8885.D74 M34 2010.

McFarlane:2005:FHT

[1210] Nigel McFarlane. Firefox hacks: tips and tools for next-generation web browsing. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00928-3. xvii + 377 pp. LCCN TK5105.882 .M34 2005. URL http://www.oreilly.com/

catalog/9780596009281.

McGavren:2019:HFG

[1211] Jay McGavren. Head First Go: a brain-friendly guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4919-6955-5. xxxiv + 520 pp. LCCN QA76.73.G63 M34 2019.

McGuire:2007:GSP

[1212] Paul McGuire. Getting started with pyparsing. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51423-9. LCCN QA76.76.A65 M33 2007eb; QA76.76.A65. URL http://www.oreilly.

com/catalog/9780596514235.

McGuire:2012:BFM

[1213] Hugh McGuire and Brian (Brian Francis) O’Leary, editors. Book: a futurist’s manifesto: essays from the bleeding edge of publishing. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2012. ISBN 1-4493-0560-1 (paperback). viii + 307 pp. LCCN Z278 .B66 2012. Â£18.99Â£18.99.

McIntosh:2003:MXN

[1214] Jason McIntosh, Chuck Toporek, and Chris Stone. Mac OS X in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00370-6. xxii + 801 pp. LCCN QA76.76.O63 M38 2003; QA76.76.O63 M39873 2003.

URL http://www.oreilly.com/catalog/9780596003708.

McKinney:2012:PDA

[1215] Wes McKinney. Python for Data Analysis. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1979-3 (paperback). 400 pp. LCCN ???? EUR 32.00; UK £30.99.

McLaughlin:2000:JX

[1216] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2000. ISBN 0-596-00016-2. xvi + 479 pp. LCCN QA76.73.J38 M39 2000. US$39.95.

URL http://www.oreilly.com/catalog/9780596000165;   http://
www.oreilly.com/catalog/javaxml.

McLaughlin:2001:JX

[1217] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00197-5 (paperback). xv + 509 pp. LCCN QA76.73.J38 M39 2001. Also available via the Internet.

McLaughlin:2001:JXE

[1218] Brett McLaughlin. Java and XML: Solutions to Real-World Problems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, September 2001. ISBN 0-596-00197-5. xv + 509 pp. LCCN QA76.73.J38 .M39 2001.

US$39.95. URL http://www.oreilly.com/catalog/9780596001971;
 http://www.oreilly.com/catalog/javaxml2.

McLaughlin:2002:BJE

[1219] Brett McLaughlin. Building Java enterprise applications: vol. 1: architecture. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00123-1. xiii + 302 pp. LCCN QA76.73.J38 M38 2002; QA76.73.J38

M393 2002. URL http://www.oreilly.com/catalog/9780596001230.

Designing with EJBs, databases, and directory servers.

McLaughlin:2002:JXD

[1220] Brett McLaughlin. Java and XML data binding. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00278-5. xi + 200 pp. LCCN QA76.73.J38 M39 2002. URL http://www.oreilly.com/

catalog/9780596002787.

McLaughlin:2004:JTD

[1221] Brett McLaughlin and David Flanagan. Java 1.5 Tiger: a developer’s notebook. The developer’s notebook series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00738-8. xix + 177 pp. LCCN QA76.73.J38 M3897 2004. URL

http://www.oreilly.com/catalog/9780596007386.

McLaughlin:2005:HTH

[1222] Brett McLaughlin. Home theater hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00704-3. xix + 381 pp. LCCN TK7881.3 .M35 2005. URL http:

//www.oreilly.com/catalog/9780596007041.

McLaughlin:2006:HRA

[1223] Brett McLaughlin. Head rush Ajax. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10225-9. xxxi + 413 pp. LCCN QA76.73.J39

M35 2006eb; QA76.73.J39. URL http://www.oreilly.com/catalog/
9780596102258.

McLaughlin:2006:JX

[1224] Brett McLaughlin. Java and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-10149-X. xii + 465 pp. LCCN QA76.73.J38 M39 2006eb; QA76.73.J38; QA76.73.J38 M39 2006. URL

http://www.oreilly.com/catalog/9780596101497.

McLaughlin:2007:HFO

[1225] Brett McLaughlin, Gary Pollice, and David West. Head first objectoriented analysis and design. Head first series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00867-8. xxxiv + 600 pp. LCCN

QA76.9.O35eb; QA76.9.O35. URL http://www.oreilly.com/catalog/
9780596008673.

McLaughlin:2007:JX

[1226] Brett McLaughlin. Java and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-10149-X (paperback). 465 pp. LCCN QA76.73.J38 M39 2007. URL http://www.loc.gov/

catdir/enhancements/fy0715/2007273941-d.html;   http://www.
loc.gov/catdir/toc/fy0710/2007273941.html.

McNab:2004:NSA

[1227] Chris (Chris R.) McNab. Network security assessment. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00611-X. xxi + 371 pp. LCCN QA76.9.A25 M4355

2004. URL http://www.oreilly.com/catalog/9780596006112.

McNab:2007:NSA

[1228] Chris (Chris R.) McNab. Network security assessment. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51030-6. LCCN

TK5105.59. URL http://www.oreilly.com/catalog/9780596510305.

McShaffry:2005:GCC

[1229] Mike McShaffry. Game coding complete. Paraglyph Press, Inc., Scottsdale, AZ, USA, second edition, 2005. ISBN 1-932111-91-3. xxix + 906 pp. LCCN QA76.76.C672 M35 2005. URL http://www.oreilly.com/

catalog/9781932111910.

McWilliams:2005:SKR

[1230] Brian (Brian S.) McWilliams. Spam kings: the real story behind the highrolling hucksters pushing porn, pills and @*#?% enlargements. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00732-9. xv + 333 + 16 pp. LCCN HE7551; TK5105.743 .M38 2005eb; TK5105.743 .M38 2005 .M38 2005; TK5105.743; HE7551 .M38 2005. URL http://www.oreilly.com/

catalog/9780596007324.

Mecklenburg:2005:MPG

[1231] Robert Mecklenburg. Managing Projects with GNU Make. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00610-1. xviii + 280 pp. LCCN QA76.76.U84 O73 2005. US$29.95. URL http://www.

oreilly.com/catalog/9780596006105.

Mednieks:2012:PA

[1232] Zigurd R. Mednieks et al. Programming Android: [Java programming for the new generation of mobile devices]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2012. ISBN 1-4493-1664-6 (paperback). xvii + 542 pp. LCCN ???? EUR 37.00; UK £34.50.

Meggitt:1997:WNU

[1233] Ashley J. Meggitt and Timothy D. Ritchey. Windows NT User Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-301-4. xi + 204 pp. LCCN QA76.76.O63 M4338 1997. US$29.95. URL http://www.

ora.com/catalog/winuser/;   http://www.oreilly.com/catalog/
9781565923010;   http://www.oreilly.com/catalog/winuser.

Meyer:1997:JVM

[1234] Jon Meyer and Troy Downing. Java Virtual Machine. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1997. ISBN 1-56592-194-1. xxiv + 426 pp. LCCN QA76.73.J38 M49 1997.

US$32.95.  URL http://www.ora.com/www/item/javavm.html;
http://www.oreilly.com/catalog/9781565921948;   http://www.
oreilly.com/catalog/javavm.

Meyer:19xx:JVM

[1235] Jon Meyer and Troy Downing. The Java Virtual Machine. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-63-X. LCCN ???? 4,500 yen. Japanese translation.

Meyer:2000:CSS

[1236] Eric A. Meyer. Cascading Style Sheets: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-622-6. xiv + 453 pp. LCCN QA76.76.H94 M47 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565926226;   http://
www.oreilly.com/catalog/css.

Meyer:2001:CPR

[1237] Eric A. Meyer. CSS Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 0-596-00120-7. 91 pp. LCCN QA76.76.H94 M472 2001.

US$9.95. URL http://www.oreilly.com/catalog/9780596001209;
http://www.oreilly.com/catalog/csspr.

Meyer:2004:CPR

[1238] Eric A. Meyer. CSS pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00777-9. vi + 128 pp. LCCN TK5105.888 .M472 2004. URL

http://www.oreilly.com/catalog/9780596007775.

Meyer:2004:CSS

[1239] Eric A. Meyer. Cascading Style Sheets: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00525-3. xviii + 507 pp. LCCN QA76.76.H94 M47 2004. URL http://www.oreilly.

com/catalog/9780596005252.

Meyer:2005:SHH

[1240] Gordon Meyer. Smart home hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00722-1. xxi + 376 pp. LCCN TK7881.25; TK7881.25 .M49 2005; TK7881.25 .M49 2005eb. URL http://www.oreilly.com/

catalog/9780596007225.

Meyer:2006:CSS

[1241] Eric A. Meyer. Cascading Style Sheets: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-52733-0. 535 (est.) pp. LCCN QA76.76.H94 M47 2004. URL http://www.oreilly.

com/catalog/9780596527334.

Meyer:2008:CPR

[1242] Eric A. Meyer. CSS pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2008. ISBN 0-596-51505-7 (paperback). 163 pp. LCCN TK5105.888; TK5105.888 .M472 2008xeb. URL http://www.oreilly.

com/catalog/9780596515058.

Milburn:2004:DPE

[1243] Ken Milburn. Digital photography: expert techniques. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00547-4. xiv + 467 pp. LCCN TR267 .M57 2004. URL http://www.oreilly.com/

catalog/9780596005474.

Milburn:2006:DL

[1244] Ken Milburn. From darkroom to Lightroom. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52977-5. LCCN TR267.5.L5; TR267.5.L5 M55 2006eb. URL http://www.oreilly.com/

catalog/9780596529772.

Milburn:2006:DPE

[1245] Ken Milburn. Digital photography: expert techniques. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52690-3. xiv + 387 pp. LCCN TR267; TR267 .M57 2006eb. URL http://www.oreilly.

com/catalog/9780596526900.

Milburn:2007:SAD

[1246] Ken Milburn. Sony Alpha DSLR A100: a better manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52925-2. LCCN TR256; TR256 .M55

2007eb. URL http://www.oreilly.com/catalog/9780596529253.

Miles:2005:AC

[1247] Russ Miles. AspectJ cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00654-3. xix + 331 pp. LCCN QA76.64 .M54 2005 .M55 2004; QA76.64 .M55 2004; QA76.64 .M54 2005; QA76.64 .M54 2005eb; QA76.64. URL http:

//www.oreilly.com/catalog/9780596006549.

Miles:2006:LU

[1248] Russ (Russell) Miles and Kim Hamilton. Learning UML 2.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00982-8. xiii + 269 pp. LCCN QA76.76.D47; QA76.76.D47 M55 2006eb. URL http://www.oreilly.

com/catalog/9780596009823.

Millsap:2003:OOP

[1249] Cary V. Millsap and Jeff Holt. Optimizing Oracle performance. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00527-X. xxv + 388 pp. LCCN QA76.9.D3 M577 2003. URL http://www.oreilly.com/

catalog/9780596005276.

Milstein:2004:G

[1250] Sarah Milstein and Rael Dornfest. Google. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2004. ISBN 0-596-00613-6. xii + 299 pp. LCCN ZA4234.G64 M55 2004. URL http://

www.oreilly.com/catalog/9780596006136.

Milstein:2006:G

[1251] Sarah Milstein, J. D. Biersdorfer, and Matthew MacDonald. Google. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2006. ISBN 0-596-10019-1 (paperback). xv + 446 pp. LCCN ZA4234.G64 M55 2006. URL http://www.loc.gov/

catdir/enhancements/fy0715/2006280379-d.html;   http://www.
oreilly.com/catalog/9780596100193.

Milstein:2006:GMM

[1252] Sarah Milstein, J. D. Biersdorfer, and Matthew MacDonald. Google: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10019-1. 446 pp. LCCN TK5105.885.G66; TK5105.885.G66 M63 2006eb; TK5105.885.G66 M63 2006. URL http:

//www.oreilly.com/catalog/9780596100193.

Minar:2001:PPH

[1253] Nelson Minar, Marc Hedlund, Clay Shirky, Tim O’Reilly & Dan Bricklin, David Anderson, Jeremie Miller, Adam Langley, Gene Kan, Alan Brown, Marc Waldman, Lorrie Cranor, Aviel Rubin, Roger Dingledine, Michael Freedman, David Molnar, Rael Dornfest, Dan Brickley, Theodore Hong, Richard Lethin, Jon Udell, Nimisha Asthagiri, Walter Tuvell, and Brandon Wiley. Peer-to-Peer: Harnessing the Power of Disruptive Technologies. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2001. ISBN 0-596-00110-X. xv + 432 pp. LCCN TK5105.7 .P43 2001; TK5105.7 .P44 2001.

US$29.95. URL http://www.oreilly.com/catalog/9780596001100;
 http://www.oreilly.com/catalog/peertopeer. Edited by Andrew

Oram.

Miner:2012:MDP

[1254] Donald Miner and Adam Shook. MapReduce Design Patterns: Building Effective Algorithms and Analytics for Hadoop and Other Systems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2717-6, 1-4493-5855-1. 256 pp. LCCN ???? EUR 41.00.

Mishra:2002:MOS

[1255] Sanjay Mishra and Alan Beaulieu. Mastering Oracle SQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00129-0. xiv + 321 pp. LCCN QA76.73.S67 M57 2002. US$34.95. URL http://safari.oreilly.com/

0596001290;   http://www.oreilly.com/catalog/9780596001292;
 http://www.oreilly.com/catalog/mastorasql.

Mishra:2004:MOS

[1256] Sanjay Mishra and Alan Beaulieu. Mastering Oracle SQL. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00632-2. xvii + 472 pp. LCCN

QA76.9.D3 M5787 2004. URL http://www.oreilly.com/catalog/
9780596006327.

Mishra:2004:OUP

[1257] Sanjay Mishra. Oracle utilities: pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00899-6. viii + 117 pp. LCCN QA76.9.D3 M57 2004. URL

http://www.oreilly.com/catalog/9780596008994.

Mitchell:1996:PSA

[1258] Charles Mitchell, editor. Passport South Africa: Your Pocket Guide to South African Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885073-19-4. 96 pp. LCCN HD58.7.M566 1997.

US$6.95. URL http://www.oreilly.com/catalog/9781885073198.

Mitchell:1997:WII

[1259] Stan Mitchell. Windows 95 IFS: Inside the Windows 95 Installable File System Manager. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-200-X. xv + 360 pp. LCCN QA76.76.O63 M59 1997.

US$32.95. URL http://www.oreilly.com/catalog/9781565922006;
 http://www.oreilly.com/catalog/winfile.

Mitchell:1998:PRY

[1260] Charles Mitchell, editor. Passport Russia: Your Pocket Guide to Russian Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-885073-32-1. 96 pp. LCCN HD58.7 .M564 1998. URL

http://www.oreilly.com/catalog/9781885073327.

Mitchell:2000:DAS

[1261] Scott Mitchell. Designing Active Server Pages: Scott Mitchell’s Guide to Writing Reusable Code. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 0-596-00044-8. xii + 348 pp. LCCN TK5105.8885.A26 M58 2000. US$29.95. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0715/00062331-d.html;
 http://catdir.loc.gov/catdir/enhancements/fy0912/00062331-
b.html;   http://catdir.loc.gov/catdir/enhancements/fy1008/
00062331-t.html;   http://www.oreilly.com/catalog/9780596000448;
 http://www.oreilly.com/catalog/designasp.

Mitchell:2005:WMI

[1262] Tyler Mitchell. Web mapping illustrated. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00865-1 (paperback). xvi + 349 pp. LCCN GA139 .M58 2005.

URL http://www.oreilly.com/catalog/9780596008659.

Mitchell:2006:FAA

[1263] Phil Mitchell and Evan Callahan. Fixing Access annoyances: how to fix the most annoying things about your favorite database. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00852-X. xix + 357 pp. LCCN QA76.9.D3 M58 2006eb; QA76.9.D3. URL http://www.oreilly.com/

catalog/9780596008529.

Mitchell:2015:WSP

[1264] Ryan Mitchell. Web scraping with Python: collecting data from the modern web. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-1029-1 (paperback), 1-4919-1028-3. xiii + 238 pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.com/9781491910283.

Mojica:2001:CPV

[1265] Jose Mojica. COM+ Programming with Visual Basic: Developing COM+ Servers with COM, COM+, and .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 1-56592-840-7. xiv + 348 pp. LCCN QA76.9.A73

M65 2001. US$34.95. URL http://www.oreilly.com/catalog/
9781565928404;   http://www.oreilly.com/catalog/compluspvb.

Mojica:2002:CVN

[1266] Jose Mojica. C# & VB.NET Conversion Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00319-6. iv + 139 pp. LCCN QA76.73.C154 M65 2002. US$14.95. URL

http://safari.oreilly.com/0596003196;   http://www.oreilly.
com/catalog/9780596003197;   http://www.oreilly.com/catalog/
csharpvbpr.

Molinaro:2006:SC

[1267] Anthony Molinaro. SQL cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00976-3. xxix + 595 pp. LCCN QA76.73.S67 M64

2006. URL http://www.oreilly.com/catalog/9780596009762.

Molkentin:2007:BQA

[1268] Daniel Molkentin. The book of Qt 4: the art of building Qt applications. Open Source Press, Munich, Germany, 2007. ISBN 1-59327-147-6. 440 pp. LCCN QA76.9.U83 M6213 2007. URL http://www.loc.gov/

catdir/enhancements/fy0715/2007013181-d.html;   http://www.
oreilly.com/catalog/9781593271473.

Moncur:1998:MCE

[1269] Michael G. Moncur. MCSE: The Core Exams in a Nutshell: a desktop quick reference Microsoft Certified Systems Engineer. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-376-6. xii + 401 pp. LCCN QA76.3 .M644 1998; QA76.76.O63 M66 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565923768;   http://
www.oreilly.com/catalog/mcsecore.

Moncur:1998:MEN

[1270] Michael G. Moncur. MCSE: The Electives in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1998. ISBN 1-56592-482-7. xii + 357 pp. LCCN QA76.3 .M6443 1998; QA76.3 .M66 1998.

US$24.95. URL http://www.oreilly.com/catalog/9781565924826;
 http://www.oreilly.com/catalog/mcselect.

Moncur:1999:NEF

[1271] Michael Moncur. Networking Essentials Flashcards: MCSE Core Exam 70-058. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-568-8. 280

pp. LCCN ???? US$29.95. URL http://www.oreilly.com/catalog/
9781565925687;   http://www.oreilly.com/catalog/netessfc/.

Moncur:1999:TIW

[1272] Michael Moncur. TCP/IP Windows NT 4.0 Flashcards: MCSE Elective Exam #70-059. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-583-1. 240 pp. LCCN QA76.76.O63 M597 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565925830;   http://
www.oreilly.com/catalog/tcpipfc/.

Moncur:1999:WNS

[1273] Michael Moncur. Windows NT Server 4.0 in the Enterprise Flashcards: MCSE Core Exam #70-068. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-584-X. 260 pp. LCCN ???? US$29.95.

URL http://www.oreilly.com/catalog/9781565925847;   http://
www.oreilly.com/catalog/enterprisefc.

Moncur:2000:MCE

[1274] Michael G. Moncur. MCSE: the core exams in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2000. ISBN 1-56592-721-4. xii + 472 pp. LCCN QA76.3 .M644

2000. URL http://www.oreilly.com/catalog/9781565927216.

Moncur:2001:MNW

[1275] Michael G. Moncur and Paul Murphy. MCSE in a Nutshell: The Windows 2000 Exams. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2001. ISBN 0-596-00030-8. xvi + 459 pp. LCCN QA76.3 .M643 2001. US$29.95.

URL http://www.oreilly.com/catalog/9780596000301;   http://
www.oreilly.com/catalog/mcseian.

Monson-Haefel:1999:EJ

[1276] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-605-6. xvi + 317 pp. LCCN QA76.76.J38

M66 1999. US$32.95. URL http://www.oreilly.com/catalog/
9781565926059;   http://www.oreilly.com/catalog/entjbeans.

Monson-Haefel:2000:EJ

[1277] Richard Monson-Haefel. Enterprise JavaBeans. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, March 2000. ISBN 1-56592-869-5. xvii + 472 pp. LCCN QA76.73.J38 M65 2000. US$39.95.

URL http://www.oreilly.com/catalog/9781565928695;   http://
www.oreilly.com/catalog/entjbeans2.

Monson-Haefel:2001:EJ

[1278] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2001. ISBN 0-596-00226-2. xix + 567 pp. LCCN QA76.73.J38 M65 2001. URL http://www.oreilly.com/

catalog/9780596002268. Covers EJB 2.0.

Monson-Haefel:2001:JMS

[1279] Richard Monson-Haefel and David A. Chappell. Java Message Service. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00068-5. xvi + 220 pp. LCCN QA76.73.J38 M655 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596000684;   http://
www.oreilly.com/catalog/javmesser.

Monson-Haefel:2004:EJ

[1280] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2004. ISBN 0-596-00530-X. xxi + 763 pp. LCCN QA76.73.J38 M65 2004. URL http://www.oreilly.com/

catalog/9780596005306.

Monson-Haefel:2006:EJ

[1281] Richard Monson-Haefel and Bill Burke. Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-00978-X.

760 (est.) pp. LCCN ???? URL http://www.oreilly.com/catalog/
9780596009786;   http://www.oreilly.com/catalog/entjbeans5/.

Moock:2001:ADG

[1282] Colin Moock. ActionScript: The Definitive Guide: Mastering Flash Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 1-56592-852-0. xxxii + 696 pp. LCCN TR897.7 .M55 2001; TR897.7 .M65 2001.

US$39.95. URL http://www.oreilly.com/catalog/9781565928527;
 http://www.oreilly.com/catalog/actscript.

Moock:2003:AFMa

[1283] Colin Moock. ActionScript for Flash MX: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00396-X. xxxiv + 1051 pp. LCCN TR897.7 .M65 2003. US$54.95.

URL http://www.oreilly.com/catalog/9780596003968;   http://
www.oreilly.com/catalog/actscript2.

Moock:2003:AFMb

[1284] Colin Moock. ActionScript for Flash MX: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00514-8. v + 138 pp. LCCN TR897.7 .M6497 2003. URL http://www.oreilly.com/

catalog/9780596005146.

Moock:2004:EA

[1285] Colin Moock. Essential ActionScript 2.0. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00652-7. xxiv + 502 pp. LCCN TR897.7 .M6523 2004. URL

http://www.oreilly.com/catalog/9780596006525.

Moock:2007:EA

[1286] Colin Moock. Essential ActionScript 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52694-6 (paperback). xxxi + 911 pp. LCCN TR897.7; TR897.7 .M6524 2007eb; TR897.7 .M6524 2007. URL http:/

/www.oreilly.com/catalog/9780596526948.

Morrison:2008:HFJ

[1287] Michael Morrison. Head first JavaScript. Head first series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-52774-8. xxxiv + 615 pp. LCCN QA76.73.J39eb; QA76.73.J39. URL http://www.oreilly.com/

catalog/9780596527747.

Morville:2005:AF

[1288] Peter Morville. Ambient findability. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00765-5 (paperback). vii–xiv + 188 pp. LCCN

QA76.9.D26 M673 2005. URL http://www.oreilly.com/catalog/
9780596007652;   http://www.oreilly.com/catalog/ambient/.

Morville:2007:IAW

[1289] Peter Morville, Louis Rosenfeld, and Louis Information architecture for the World Wide Web Rosenfeld. Information architecture for the World Wide Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2007. ISBN 0-596-52734-9 (paperback). xix + 504 pp. LCCN TK5105.888 .R67

2007. URL http://www.oreilly.com/catalog/9780596527341.

Mott:2001:LCa

[1290] Troy Mott, editor. Learning Carbon. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 0-596-00161-4. xiv + 341 pp. LCCN QA76.9.U83 L428 2001; QA76.9.U83 L43 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596001612;   http://
www.oreilly.com/catalog/learncarbon.

Mott:2001:LCb

[1291] Troy Mott. Learning Cocoa. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 0-596-00160-6. xiii + 366 pp. LCCN QA76.64 .L434 2001; QA76.64 .L437 2001. US$34.95. URL

http://www.oreilly.com/catalog/9780596001605;   http://www.
oreilly.com/catalog/learncocoa.

Mountjoy:2004:WDG

[1292] Jon Mountjoy and Avinash Chugh. WebLogic: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00432-X. xvii + 828 pp. LCCN TK5105.8885.B4 M68 2004. URL http://www.oreilly.

com/catalog/9780596004323.

Muck:2003:FRD

[1293] Tom Muck. Flash remoting: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00401-X. xxvi + 612 pp. LCCN TK5105.888

.M835 2003. URL http://www.oreilly.com/catalog/9780596004019.

Muench:2000:BOX

[1294] Steve Muench. Building Oracle XML Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 1-56592-691-9. xvi + 792 pp. LCCN QA76.76.H94 M75 2000. US$44.95.

URL http://www.oreilly.com/catalog/9781565926912;   http://
www.oreilly.com/catalog/orxmlapp. Includes CD-ROM.

Mui:1990:PBR

[1295] Linda Mui. Pick BASIC: a Reference Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel:

+1 617 354 5800, 1990. ISBN 0-937175-42-0. xxvii + 307 pp. LCCN QA76.73.P225 M85 1990. US$39.95. URL http://www.oreilly.com/

catalog/9780937175422.

Mui:1992:XWS

[1296] Linda Mui and Eric Pearce. X Window System Administrator’s Guide for X11 Release 4 and Release 5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 0-937175-83-8 (paperback), 1-56592-052-X (with CD ROM). xxiv + 346 pp. LCCN QA76.76.W56 D44 v.8 1992. US$29.95 (without CD ROM), US$59.95 (with CD ROM).

URL http://www.oreilly.com/catalog/9780937175835;   http://
www.oreilly.com/catalog/v8. Contains CD ROM with X11R4 and

X11R5 source code, plus compiled versions for Sun 3 SunOS 4.1.1, Sun 4 SunOS 4.1.1, DECstation ULTRIX 4.2, and IBM RS/6000 AIX 3.2.

Mui:1994:XUT

[1297] Linda Mui and Valerie Quercia. X User Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1994. ISBN 1-56592-019-8. 856 pp. LCCN QA76.76.W56 M85 1994. US$49.95 (CD-ROM included). URL http:/

/www.oreilly.com/catalog/9781565920194.

Mui:1995:WYC

[1298] Linda Mui. When You Can’t Find Your UNIX System Administrator. What you need to know. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1995. ISBN 1-56592-104-6. xiii + 139 pp. LCCN QA76.76.O63 M845 1995.

US$17.95. URL http://www.oreilly.com/catalog/9781565921047;
 http://www.oreilly.com/catalog/usa.

Muller:2006:SVP

[1299] Al Muller, Andy Jones, David E. Williams, Stephen Beaver, David A. Payne, Jeremy Pries, and David E. Hart. Scripting VMware Power Tools: Automating Virtual Infrastructure Administration. Syngress Publishing, Inc., Rockland, MA, USA, 2006. ISBN 1-59749-059-8. xviii + 398 pp. LCCN QA76.76.O63 S385 2006. URL http://www.oreilly.

com/catalog/9781597490597.

Mullet:2000:MI

[1300] Dianna Mullet and Kevin Mullet. Managing IMAP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 0-596-00012-X. xv + 390 pp. LCCN TK5105.73 .M84 2000. US$34.95.

URL http://www.oreilly.com/catalog/9780596000127;   http://
www.oreilly.com/catalog/mimap.

Murphy:2005:INA

[1301] Niall Richard Murphy and David Malone. IPv6 network administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00934-8 (paperback). xxvii + 275 pp. LCCN TK5105.585 .M87 2005. URL http://

www.oreilly.com/catalog/9780596009342.

Murray:1994:EGF

[1302] James D. Murray and William vanRyper. Encyclopedia of Graphics File Formats. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1994. ISBN 1-56592-058-9. xxxii + 894 pp. LCCN T385 .M87 1994. US$59.95. URL http://www.

oreilly.com/catalog/9781565920583.

Murray:1996:EGF

[1303] James D. Murray and William vanRyper. Encyclopedia of Graphics File Formats: The Complete Reference on CD-ROM with Links to Internet Resources. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, May 1996. ISBN 1-56592-161-5. xxxvi + 1116 pp. LCCN T385 .M87 1996.

US$59.95. URL http://www.oreilly.com/catalog/9781565921610;
 http://www.oreilly.com/catalog/gffcd.

Murray:1998:WNE

[1304] James D. Murray. Windows NT Event Logging. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-514-9. xvii + 295 pp. LCCN QA76.76.O63 M8664 1998; QA76.76.O63 M87 1998. US$32.95.

URL http://www.oreilly.com/catalog/9781565925144;   http://
www.oreilly.com/catalog/winlog.

Murray:1998:WNS

[1305] James D. Murray. Windows NT SNMP: Simple Network Management Protocol. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-338-3. xvi + 446 pp. LCCN TK5105.583 .M87 1998. US$34.95. URL http://www.

ora.com/catalog/winsnmp/;   http://www.oreilly.com/catalog/
9781565923386;   http://www.oreilly.com/catalog/winsnmp.

Murty:2008:PAW

[1306] James Murty. Programming Amazon Web Services — S3, EC2, SQS, FPS, and SimpleDB. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51581-2. xvi + 581 pp. LCCN QA76.625 .M88 2008.

Musciano:1996:HDG

[1307] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-175-5. xx + 385 pp. LCCN QA76.76.H94 M87 199. US$29.95, CDN$42.95. URL

http://www.oreilly.com/catalog/9781565921757.

Musciano:1997:HDG

[1308] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-235-2. xvii + 531 pp. LCCN QA76.76.H94

M88 1997. US$32.95. URL http://www.ora.com/catalog/html2/;
http://www.oreilly.com/catalog/9781565922358;   http://www.
oreilly.com/catalog/html2.

Musciano:1998:HDG

[1309] Chuck Musciano and Bill Kennedy. HTML: The Definitive Guide. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1998. ISBN 1-56592-492-4. xviii + 587 pp. LCCN QA 76.76 H94 M88 1998.

US$32.95. URL http://www.oreilly.com/catalog/9781565924925;
 http://www.oreilly.com/catalog/html3.

Musciano:2000:HXD

[1310] Chuck Musciano and Bill Kennedy. HTML and XHTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2000. ISBN 0-596-00026-X. xvi + 655 pp. LCCN QA76.76.H94 M875 2000 Bar.

US$34.95. URL http://www.oreilly.com/catalog/9780596000264;
 http://www.oreilly.com/catalog/html4.

Musciano:2002:HXD

[1311] Chuck Musciano and Bill Kennedy. HTML & XHTML: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2002. ISBN 0-596-00382-X. xxi + 645 pp. LCCN QA76.76.H94 M875

2002. US$39.95. URL http://safari.oreilly.com/059600382X;
http://www.oreilly.com/catalog/9780596003821;   http://www.
oreilly.com/catalog/html5.

Musciano:2007:HXD

[1312] Chuck Musciano and Bill Kennedy. HTML and XHTML: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, sixth edition, 2007. ISBN 0-596-52732-2 (paperback). xxi + 654 pp. LCCN QA76.76.H94; QA76.76.H94 M875 2007; QA76.76.H94 M88 2006; QA76.76.H94 M875 2006eb. URL

http://www.oreilly.com/catalog/9780596527327.

Musumeci:2002:SPT

[1313] Gian-Paolo D. Musumeci and Mike Loukides. System Performance Tuning. Nutshell handbooks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00284-X. xiv + 334 pp. LCCN QA76.76.O63 L66

2002. US$39.95. URL http://safari.oreilly.com/059600284X;
http://www.oreilly.com/catalog/9780596002848;   http://www.
oreilly.com/catalog/spt2.

Mutton:2004:IH

[1314] Paul Mutton. IRC hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00687-X. xix + 409 pp. LCCN TK5105.886

.M87 2004. URL http://www.oreilly.com/catalog/9780596006877.

Myer:2005:NNX

[1315] Thomas Myer. No nonsense XML Web development with PHP. Site-Point Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9752402-0-X (paperback). xiv + 354 pp. LCCN QA76.76.H94 M94 2005. URL

http://www.oreilly.com/catalog/9780975240205.

Myer:2007:LGW

[1316] Thomas Myer. Lead generation on the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51021-7. LCCN HF5415.1265

.M94 2007eb; HF5415.1265. URL http://www.oreilly.com/catalog/
9780596510213.

Naftalin:2006:JGC

[1317] Maurice Naftalin and Philip Wadler. Java Generics and Collections. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52775-6. 294 (est.) pp. LCCN ???? US$34.99.

Naftalin:2007:JGC

[1318] Maurice Naftalin and Philip Wadler. Java Generics and Collections. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52775-6. xvii + 273 pp. LCCN QA76.73.J38 N325 2007. US$34.99. URL http://www.loc.

gov/catdir/enhancements/fy0715/2006287224-d.html;   http://
www.loc.gov/catdir/toc/fy0705/2006287224.html;   http://www.
oreilly.com/catalog/9780596527754.

Nagar:1997:WNF

[1319] Rajeev Nagar. Windows NT File System Internals: a Developer’s Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-249-2. xviii + 774 pp. LCCN QA76.76.O63 N34 1997.

US$49.95. URL http://www.ora.com/catalog/wininternals/;
http://www.oreilly.com/catalog/9781565922495;   http://www.
oreilly.com/catalog/wininternals.

Nanda:2005:OPS

[1320] Arup Nanda and Steven Feuerstein. Oracle PL/SQL for DBAs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00587-3. xxi + 429 pp. LCCN QA76.73.P258; QA76.73.P258eb. URL http://www.oreilly.

com/catalog/9780596005870.

Negrino:2003:MXU

[1321] Tom Negrino and Dori Smith. Mac OS X unwired: a guide for home, office, and the road. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00508-3. xiv + 207 pp. LCCN QA76.76.O63 N415 2004. URL

http://www.oreilly.com/catalog/9780596005085.

Neil:2012:MDP

[1322] Theresa Neil. Mobile design pattern gallery: [UI patterns for mobile applications]. Safari Tech Books Online. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2012. ISBN 1-4493-1432-5 (paperback), 1-4493-1893-2, 1-4493-3644-2. xii + 261 pp. LCCN QA76.76.A65 N45 2012. EUR 24.50; UK £21.50. URL http://

proquest.safaribooksonline.com/9781449318901;   http://www.
gbv.de/dms/tib-ub-hannover/684155222.pdf;   http://www.gbv.
eblib.com/patron/FullRecord.aspx?p=871556.

Neilson:2000:LDA

[1323] Greg Neilson. Lotus Domino administration in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-717-6. xii + 356 pp. LCCN HF5548.4.L673 N44 2000. US$24.95.

URL http://www.oreilly.com/catalog/9781565927179;   http://
www.oreilly.com/catalog/domino.

Neuburg:1998:FDG

[1324] Matt Neuburg. Frontier: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1998. ISBN 1-56592-383-9. xxi + 591 pp. LCCN QA76.9.D3 N483 1998. US$34.95.

URL http://www.oreilly.com/catalog/9781565923836;   http://
www.oreilly.com/catalog/frontier.

Neuburg:1999:RDG

[1325] Matt Neuburg. REALbasic: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-657-9. xxiv + 660 pp. LCCN QA76.73.B3 N47 1999; QA76.8.M3 N48 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565926578;   http://
www.oreilly.com/catalog/realbasic.

Neuburg:2001:RDG

[1326] Matt Neuburg. REALbasic: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00177-0. xxvi + 724 pp. LCCN QA76.8.M3 N48 2001 Stacks SEMSTAX-

NBKS. US$39.95. URL http://safari.oreilly.com/0596001770;
http://www.oreilly.com/catalog/9780596001773;   http://www.
oreilly.com/catalog/realbasic2.

Neuburg:2004:ADG

[1327] Matt Neuburg. AppleScript: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00557-1. xix + 453 pp. LCCN QA76.73.A67

N48 2004. URL http://www.oreilly.com/catalog/9780596005573.

Neuburg:2006:ADG

[1328] Matt Neuburg. AppleScript: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10211-9. 504 pp. LCCN QA76.73.A67 N48 2006; QA76.73.A67 N48 2006eb; QA76.73.A67. URL

http://www.oreilly.com/catalog/9780596102111.

Neumann:2000:GPR

[1329] Sven Neumann. Gimp pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-731-1. iii + 97 pp. LCCN T385 .N46713 2000;

T385 .N47 2000. US$9.95. URL http://www.oreilly.com/catalog/

9781565927315. Translated by Sven Riedel.

Newham:1995:LBS

[1330] Cameron Newham and Bill Rosenblatt. Learning the bash Shell. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1995. ISBN 1-56592-147-X. xvi + 292 pp. LCCN QA76.76.O63 N458 1995. US$27.95.

URL http://www.ora.com/gnn/bus/ora/item/bash.html;   http:/
/www.oreilly.com/catalog/9781565921474.

Newham:1998:LBS

[1331] Cameron Newham and Bill Rosenblatt. Learning the bash Shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second (revised and updated) edition, 1998. ISBN 1-56592-347-2. xvi + 318 pp. LCCN QA76.76.O63 N458 1998. US$29.95.

URL http://www.oreilly.com/catalog/9781565923478;   http://
www.oreilly.com/catalog/bash2.

Newham:2005:LBS

[1332] Cameron Newham and Bill Rosenblatt. Learning the bash shell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00965-8 (paperback). xvi + 333 pp. LCCN QA76.76.O63 N458 2005. UK£24.95. URL http://www.

oreilly.com/catalog/9780596009656.

Newman:2015:BMD

[1333] Samuel Newman. Building microservices: designing fine-grained systems. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-5035-8, 1-4919-5033-1 (e-book). xviii + 259 pp. LCCN QA76.9.A73 N49 2015. URL http://proquest.

safaribooksonline.com/?fpi=9781491950340.

Nichols:1996:PP

[1334] Bradford Nichols, Bick Buttlar, and Jackie Proulx Farrell. Pthreads Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-115-1. xvi + 267 pp. LCCN QA76.642.N53 1996. US$32.95.

URL http://www.oreilly.com/catalog/9781565921153;   http://
www.oreilly.com/catalog/pthread.

Nichols:1998:PP

[1335] Bradford Nichols, Dick Buttlar, and Jacqueline Proulx Farrell. Pthreads programming. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-115-1. xvi + 267 pp. LCCN QA76.642 .N53 1998.

Niederst:1996:DWG

[1336] Jennifer Niederst and Edie Freedman. Designing for the Web: Getting Started in a New Medium. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1996. ISBN 1-56592-165-8. xiv + 165 pp. LCCN TK5105.888.N54 1996.

US$24.95. URL http://www.oreilly.com/catalog/9781565921658;
 http://www.oreilly.com/catalog/wdesign.

Niederst:1998:WDN

[1337] Jennifer Niederst. Web Design in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-515-7. xviii + 560 pp. LCCN TK5105.888

.N542 1999. US$24.95. URL http://www.oreilly.com/catalog/wdnut.

Niederst:1999:WDN

[1338] Jennifer Niederst. Web design in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-515-7. xviii + 560 pp. LCCN TK5105.888 .N54 1999; TK5105.888 .N542 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565925151;   http://
www.oreilly.com/catalog/wdnut.

Niederst:2000:HPR

[1339] Jennifer Niederst. HTML Pocket Reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-579-3. iii + 92 pp. LCCN QA76.76.H94 N52 2000; QA76.76 .N54 2000. US$9.95.

URL http://www.oreilly.com/catalog/9781565925793;   http://
www.oreilly.com/catalog/htmlpr.

Niederst:2001:LWD

[1340] Jennifer Niederst. Learning Web Design: a Beginner’s Guide to HTML, Graphics, and Beyond. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2001. ISBN 0-596-00036-7. xii + 388 pp. LCCN TK5105.888 .N489 2001; TK5105.8883 .N545 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596000363;   http://
www.oreilly.com/catalog/learnweb.

Niederst:2001:WDN

[1341] Jennifer Niederst. Web Design in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00196-7. xix + 618 pp. LCCN TK5105.888 .N542

2001. US$29.95. URL http://safari.oreilly.com/0596001967;
http://www.oreilly.com/catalog/9780596001964;   http://www.
oreilly.com/catalog/wdnut2.

Niederst:2002:HPR

[1342] Jennifer Niederst. HTML Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00296-3. v + 95 pp. LCCN QA76.76.H94 N53 2002. US$12.95. URL http://safari.oreilly.com/

0596002963;   http://www.oreilly.com/catalog/9780596002961;
 http://www.oreilly.com/catalog/htmlpr2.

Niederst:2002:WDC

[1343] Jennifer Niederst. The Web design CD bookshelf: six bestselling books on CD-ROM, 2002. ISBN 0-596-00271-8. URL http://www.oreilly.

com/catalog/9780596002718.

NiederstRobbins:2006:HXP

[1344] Jennifer Niederst Robbins. HTML and XHTML Pocket Reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2006. ISBN 0-596-52727-6. vi + 97 pp. LCCN QA76.76.H94 N52 2006. EUR 9.90, US$13.09. URL http://

www.oreilly.com/catalog/9780596527273.

NiederstRobbins:2006:WDN

[1345] Jennifer Niederst Robbins. Web design in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00987-9. xxvii + 796 pp. LCCN TK5105.888 .N54 2006eb; TK5105.888 .N54 2006; TK5105.888.

URL http://www.oreilly.com/catalog/9780596009878.

NiederstRobbins:2007:LWD

[1346] Jennifer Niederst Robbins. Learning Web design: a beginner’s guide to (X)HTML, style sheets, and Web graphics. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2007. ISBN 0-596-52752-7. xiv + 464 pp. LCCN TK5105.888;

TK5105.888 .N489 2007. URL http://www.oreilly.com/catalog/
9780596527525.

Niemetzky:2005:PRP

[1347] Gerard Niemetzky. Photo retouching with Photoshop: a designer’s notebook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00860-0. 96 pp. LCCN T385 .N45

2005. URL http://www.oreilly.com/catalog/9780596008604.

Niemeyer:1996:EJ

[1348] Patrick Niemeyer and Joshua Peck. Exploring Java. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-184-4. xv + 407 pp. LCCN QA76.73.J38 N54 1996. US$24.95.

URL http://www.ora.com/www/item/expjava.html;   http://www.
oreilly.com/catalog/9781565921849.

Niemeyer:1996:JEI

[1349] Patrick Niemeyer and Joshua Peck. Java — Expedition ins Programmierreich. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1996. ISBN 3-930673-52-5 (??invalid ISBN??). ???? pp. LCCN ???? 69 DM. URL

http://www.ora.de/. Includes CD-ROM.

Niemeyer:1996:JPP

[1350] Patrick Niemeyer and Joshua Peck. Java par La Pratique. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1996. ISBN 2-84177-022-2. xxi + 413 pp. LCCN ???? 220 FF. URL http://www.editions-

oreilly.fr/international/france/prog/expjava.html;   http:/

/www.ora.de/. French translation of [1348]. Includes CD-ROM.

Niemeyer:1997:EJa

[1351] Pat Niemeyer and Josh Peck. Exploring Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 4-900900-17-6. ???? pp. LCCN ???? 4300

yen. URL http://www.ora.com/international/catalog/japanese/
inet/exjava.html;   http://www.ora.com/international/japan/
;   http://www.oreilly.co.jp/BOOK/exjava.htm. Japanese trans-

lation.

Niemeyer:1997:EJb

[1352] Patrick Niemeyer and Joshua Peck. Exploring Java. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-271-9. xviii + 594 pp. LCCN QA76.73.J38 N54 1997.

US$32.95. URL http://www.ora.com/catalog/expjava2/;   http:/
/www.oreilly.com/catalog/9781565922716.

Niemeyer:2000:LJ

[1353] Patrick Niemeyer and Jonathan Knudsen. Learning Java. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-718-4. xvi + 706 pp. LCCN QA76.73.J38 N545 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565927186;   http://
www.oreilly.com/catalog/learnjava. Includes CD-ROM.

Niemeyer:2002:LJ

[1354] Pat Niemeyer and Jonathan Knudsen. Learning Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00285-8. xvii + 807 pp. LCCN QA76.73.J38 N545

2002. US$44.95. URL http://safari.oreilly.com/0596002858;
http://www.oreilly.com/catalog/9780596002855;   http://www.
oreilly.com/catalog/learnjava2.

Niemeyer:2005:LJ

[1355] Patrick Niemeyer and Jonathan Knudsen. Learning Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00873-2. xx + 954 pp. LCCN QA76.73.J38 N545 2005. URL http://www.oreilly.

com/catalog/9780596008734.

Noble:2008:FC

[1356] Joshua J. Noble and Todd Anderson. Flex 3 cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52985-6. xxiv + 673 pp. LCCN QA76.625 .N633 2008.

Nolan:1994:CBP

[1357] James L. Nolan, Karla C. Shippey, J. D., et al. China Business: The Portable Encyclopedia for Doing Business with China. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1994. ISBN 0-9631864-3-4. 418 pp. LCCN HC427.92.C4636 1995. US$24.95. URL http://www.oreilly.

com/catalog/9780963186430.

Nolan:1994:HKB

[1358] James L. Nolan, Karla C. Shippey, J. D., et al. Hong Kong Business: The Portable Encyclopedia for Doing Business with Hong Kong. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1994. ISBN 0-9631864-7-7. 306 pp. LCCN HC470.3.H6655 1994. US$24.95. URL http://www.oreilly.

com/catalog/9780963186478.

Nolan:1994:JBP

[1359] James L. Nolan, Karla C. Shippey, J. D., et al. Japan Business: The Portable Encyclopedia for Doing Business with Japan. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1994. ISBN 0-9631864-2-6. 374 pp. LCCN HC462.95.J355 1994. US$24.95. URL http://www.oreilly.

com/catalog/9780963186423.

Nolan:1994:MBP

[1360] James L. Nolan, Karla C. Shippey, J. D., et al. Mexico Business: The Portable Encyclopedia for Doing Business with Mexico. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1994. ISBN 0-9631864-0-X. 488 pp. LCCN HC135.M54534 1994. US$24.95. URL http://www.oreilly.

com/catalog/9780963186409.

Nolan:1994:SBP

[1361] James L. Nolan, Karla C. Shippey, J. D., et al. Singapore Business: The Portable Encyclopedia for Doing Business with Singapore. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1994. ISBN 0-9631864-6-9. 309 pp. LCCN HC445.8.S543 1994. US$24.95. URL http://www.oreilly.com/

catalog/9780963186461.

Nolan:1994:TBP

[1362] James L. Nolan, Karla C. Shippey, J. D., et al. Taiwan Business: The Portable Encyclopedia for Doing Business with Taiwan. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1994. ISBN 0-9631864-5-0. 310 pp. LCCN HC430.5.T3375 1994. US$24.95. URL http://www.oreilly.

com/catalog/9780963186454.

Nolan:1995:ABP

[1363] James L. Nolan, Karla C. Shippey, J. D., et al. Argentina Business: The Portable Encyclopedia for Doing Business with Argentina. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1995. ISBN 1-885073-04-6. 372 pp. LCCN HC175.A844 1996. US$24.95. URL http://www.oreilly.com/

catalog/9781885073044.

Nolan:1995:PBP

[1364] James L. Nolan, Karla C. Shippey, J. D., et al. Philippines Business: The Portable Encyclopedia for Doing Business with Philippines. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1995. ISBN 1-885073-08-9. 342 pp. LCCN HC455.P5273 1996. US$24.95. URL http://www.oreilly.com/

catalog/9781885073082.

Nolan:1995:UBP

[1365] James L. Nolan, Karla C. Shippey, J. D., et al. USA Business: The Portable Encyclopedia for Doing Business with the United States. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1995. ISBN 1-885073-01-1. 504 pp. LCCN HC106.82.U83 1995. US$24.95. URL http://www.oreilly.

com/catalog/9781885073013.

Nolan:1996:ABP

[1366] James L. Nolan, Karla C. Shippey, J. D., et al. Australia Business: The Portable Encyclopedia for Doing Business with Australia. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-885073-03-8. 328 pp. LCCN HC605.A7736 1996. US$24.95. URL http://www.oreilly.com/

catalog/9781885073037.

Norberg:2000:SWN

[1367] Stefan Norberg. Securing Windows NT/2000 servers for the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-768-0. xiv + 199 pp. LCCN QA76.76.O63 N65 2001.

Norberg:2001:SWN

[1368] Stefan Norberg. Securing Windows NT/2000 Servers for the Internet: a Checklist for System Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-768-0. xiv + 199 pp. LCCN QA76.76.O63 N65 2001.

US$29.95. URL http://www.oreilly.com/catalog/9781565927681;
 http://www.oreilly.com/catalog/securwinserv.

Northrup:2004:HHP

[1369] Anthony Northrup and Eric Faulkner. Home hacking projects for geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00405-2 (paperback). xiv + 319 pp.

LCCN TK7881.25 .N67 2004. URL http://www.oreilly.com/catalog/
9780596004057.

Notomi:2004:IFB

[1370] Yasukuni Notomi. iPod fan book: go everywhere with iPod. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, revised and updated edition, 2004. ISBN 0-596-00776-0 (paperback). 96 pp. LCCN ML74.4.I48 N68 2004. URL

http://www.oreilly.com/catalog/9780596007768.

Novas:1996:PSY

[1371] Himilce Novas. Passport Spain: Your Pocket Guide to Spanish Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-885073-35-6. 96 pp. LCCN HD58.7.S559 1996. US$6.95. URL

http://www.oreilly.com/catalog/9781885073358.

Nyberg:2002:WSW

[1372] Greg Nyberg. WebLogic 6.1 Server Workbook for Enterprise JavaBeans. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00417-6. 256 (est.) pp. LCCN ???? US$24.95.

URL http://www.oreilly.com/catalog/9780596004170;   http://
www.oreilly.com/catalog/weblogicworkbook.

Nye:1988:XPM

[1373] Adrian Nye. Xlib Programming Manual for Version 11, volume 1. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1988. ISBN 0-937175-26-9, 0-937175-89-7. xxxiii + 615 pp. LCCN QA76.76.W56 D44 v.1 1988. URL http:

//www.oreilly.com/catalog/9780937175262.

Nye:1988:XRM

[1374] Adrian Nye. Xlib Reference Manual for Version 11, volume 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1988. ISBN 0-937175-27-7. xiv + 701 pp. LCCN

QA76.76.W56 D44 v.2 1988. URL http://www.oreilly.com/catalog/
9780937175279.

Nye:1990:XPM

[1375] Adrian Nye. Xlib Programming Manual: for Version 11 of the X Window System, volume 1 of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 1990. ISBN 0-937175-11-0 (6th printing). xxxi + 635 pp. LCCN QA76.76.W56 N93

1990. URL http://www.oreilly.com/catalog/9780937175118.

Nye:1990:XPR

[1376] Adrian Nye. X Protocol Reference Manual for Version 11 of the X Window System, volume 0 of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1990. ISBN 0-937175-50-1. xv + 483 pp. LCCN QA76.76.W56 X215 1990. URL

http://www.oreilly.com/catalog/9780937175507.

Nye:1990:XRM

[1377] Adrian Nye. Xlib Reference Manual: for Version 11 of the X Window System, volume 1. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 1990. ISBN 0-937175-12-9. xxiii + 765 pp. LCCN QA76.76.W56 X52

1990. URL http://www.oreilly.com/catalog/9780937175125.

Nye:1990:XTIa

[1378] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1990. ISBN 0-937175-34-X, 0-937175-56-0. xxxi + 543 pp. LCCN QA76.9.W56

N94 1990. URL http://www.oreilly.com/catalog/9780937175347.

Nye:1990:XTIb

[1379] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Reference Manual: OSF/Motif 1.1 Edition for X11, Release 4, volume 4. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, December 1990. ISBN 0-937175-62-5 (vol. 4), 0-937175-66-8 (set). xxxi + 632 pp. LCCN QA76.9.W56N94

1990. URL http://www.oreilly.com/catalog/9780937175620.

Nye:1990:XTIc

[1380] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Reference Manual — For X11 Release 4, volume 5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1990. ISBN 0-937175-57-9 (vol. 5), 0-937175-58-7 (vol. 4 and 5 set). xii + 764 pp. LCCN QA76.76.W56X225 1990. URL http:/

/www.oreilly.com/catalog/9780937175576.

Nye:1991:XRI

[1381] Adrian Nye, editor. The X Resource: Issue 0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1991. ISBN 0-937175-79-X. 253 pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9780937175798.

Nye:1992:XPM

[1382] Adrian Nye. Xlib Programming Manual, volume 1. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, July 1992. ISBN 1-56592-002-3. 824 pp. LCCN QA76.76.W56N93 1990. US$34.95. URL http://www.oreilly.com/

catalog/9781565920026;   http://www.oreilly.com/catalog/v1.

Nye:1992:XPR

[1383] Adrian Nye. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1992. ISBN 1-56592-008-2. LCCN

QA76.76.W56 X215 1990. URL http://www.oreilly.com/catalog/
9781565920088.

Nye:1992:XRIb

[1384] Adrian Nye, editor. The X Resource: Issue 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1992. ISBN 0-937175-97-8. 190 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9780937175972.

Nye:1992:XRIc

[1385] Adrian Nye, editor. The X Resource: Issue 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1992. ISBN 0-937175-98-6. 220 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9780937175989.

Nye:1992:XRId

[1386] Adrian Nye, editor. The X Resource: Issue 4. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1992. ISBN 0-937175-99-4. 276 pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9780937175996.

Nye:1992:XRM

[1387] Adrian Nye. Xlib Reference Manual, volume 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, June 1992. ISBN 1-56592-006-6. xvii + 1122 pp. LCCN QA76.76.W56 X55 1992. US$34.95.

URL http://www.oreilly.com/catalog/9781565920064;   http://
www.oreilly.com/catalog/v2.

Nye:1992:XTI

[1388] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual: Motif Edition, volume 4M. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, August 1992. ISBN 1-56592-013-9. xxxviii + 674 pp. LCCN QA 76.76.W56 N928 1992; QA 76.76 W56 N94 1992. US$34.95.

URL http://www.oreilly.com/catalog/9781565920132;   http://
www.oreilly.com/catalog/v4m.

Nye:1993:XRIb

[1389] Adrian Nye, editor. The X Resource: Issue 6. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1993. ISBN 1-56592-021-X. 234 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565920217.

Nye:1993:XRIc

[1390] Adrian Nye, editor. The X Resource: Issue 7. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1993. ISBN 1-56592-022-8. 150 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565920224.

Nye:1993:XRId

[1391] Adrian Nye, editor. The X Resource: Issue 8. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1993. ISBN 1-56592-023-6. 176 pp. LCCN ???? US$14.95.

URL http://www.oreilly.com/catalog/9781565920231.

Nye:1993:XTI

[1392] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, April 1993. ISBN 1-56592-003-1. xxxvi + 567 pp. LCCN QA76.76.W56 N945 1993.

URL http://www.oreilly.com/catalog/9781565920033.

Nye:1994:XRIb

[1393] Adrian Nye, editor. The X Resource: Issue 10. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1994. ISBN 1-56592-067-8. 212 pp. LCCN ????

US$14.95. URL http://www.oreilly.com/catalog/9781565920675.

Nye:1995:PSR

[1394] Adrian Nye and David Flanagan, editors. Programmer’s Supplement for Release 6 of the X Window System, version 11. The definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, R6 update edition, September 1995. ISBN 1-56592-089-9. xxi + 430 pp. LCCN QA76.76.W56 F55 1995. US$29.95. URL http://www.oreilly.com/

catalog/9781565920897;   http://www.oreilly.com/catalog/r6.

Nye:1995:VXP

[1395] Adrian Nye, editor. Volume 0: X Protocol Reference Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, February 1995. ISBN 1-56592-083-X. xix + 434 pp. LCCN ???? US$34.95. URL http://www.oreilly.

com/catalog/9781565920835.

Nystrom:2007:SID

[1396] Marty Nystrom. SQL injection defenses. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52964-3. LCCN QA76.73.S67

N97 2007eb; QA76.73.S67. URL http://www.oreilly.com/catalog/
9780596529642.

Oakleycomputerscientist:2005:MIM

[1397] Andy Oakley, computer scientist. Monad: introducing the MSH command shell and language. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10009-4. 206 pp. LCCN QA76.76.O63; QA76.76.O63 O23 2005eb.

URL http://www.oreilly.com/catalog/9780596100094.

Oaks:1997:JT

[1398] Scott Oaks and Henry Wong. Java Threads. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1997. ISBN 1-56592-216-6. xiii + 252 pp. LCCN QA76.73.J38 O25 1997. US$29.95.

URL http://www.oreilly.com/catalog/9781565922167;   http://
www.oreilly.com/catalog/jthreads.

Oaks:1998:JS

[1399] Scott Oaks. Java Security. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-403-7. xv + 454 pp. LCCN QA76.73.J38

O249 1998. US$32.95. URL http://www.oreilly.com/catalog/
9781565924031;   http://www.oreilly.com/catalog/javasec.

Oaks:1999:JT

[1400] Scott Oaks and Henry Wong. Java Threads. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-418-5. xiii + 319 pp. LCCN QA76.73.J38 O25 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565924185;   http://
www.oreilly.com/catalog/jthreads2/.

Oaks:19xx:JT

[1401] Scott Oaks and Henry Wong. Java Threads. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 19xx. ISBN 4-900900-54-0. ???? pp. LCCN ???? 3,400 yen. URL

http://www.oreilly.co.jp/BOOK/jthread.htm. Japanese translation.

Oaks:2000:JDQ

[1402] Scott Oaks and Henry Wong. Jini: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 2000. ISBN 1-56592-759-1. xiii + 400 pp. LCCN QA76.9.C55 O25 2000. US$29.95.

URL http://www.oreilly.com/catalog/9781565927599;   http://
www.oreilly.com/catalog/jininut. es.

Oaks:2001:JS

[1403] Scott Oaks. Java Security. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, May 2001. ISBN 0-596-00157-6. xvi + 599 pp. LCCN QA76.73.J38 O247 2001. US$44.95.

URL http://www.oreilly.com/catalog/9780596001575;   http://
www.oreilly.com/catalog/javasec2.

Oaks:2002:JN

[1404] Scott Oaks, Bernard Traversat, and Li Gong. JXTA in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00236-X. xii + 401 pp. LCCN TK5105.525 .O255 2002. US$34.95.

URL http://www.oreilly.com/catalog/9780596002367;   http://
www.oreilly.com/catalog/jxtaian.

Oaks:2004:JT

[1405] Scott Oaks and Henry Wong. Java threads. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2004. ISBN 0-596-00782-5. xvi + 340 pp. LCCN QA76.73.J38 O25 2004; QA76.73.J38 O25 2004eb; QA76.73.J38. URL

http://www.oreilly.com/catalog/9780596007829.

Oaks:2014:JPD

[1406] Scott Oaks. Java Performance: The Definitive Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-5845-4 (paperback), 1-4493-6354-7 (e-book). xiv + 408 pp. LCCN QA76.73.J38. URL http://proquest.

safaribooksonline.com/9781449363512.

OBrien:2005:JCC

[1407] Timothy M. O’Brien. Jakarta commons cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00706-X. xix + 377 pp. LCCN QA76.76.A65 O246 2005.

URL http://www.oreilly.com/catalog/9780596007065.

Odewahn:1999:OWA

[1408] Andrew Odewahn. Oracle Web Applications: PL/SQL Developer’s Introduction. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-687-0. xi + 241 pp. LCCN QA76.625 .O34 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565926875;   http://
www.oreilly.com/catalog/oracleweb.

Ogasawara:2006:WIM

[1409] Todd Ogasawara. Windows for Intel Macs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52840-X. LCCN QA76.76.O63. URL http://

www.oreilly.com/catalog/9780596528409.

Olsen:2007:AJ

[1410] Steven Douglas Olsen. Ajax on Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10187-2. xiv + 211 pp. LCCN QA76.73.J38

O58 2007eb; QA76.73.J38. URL http://www.oreilly.com/catalog/
9780596101879.

OLuanaigh:2006:GDC

[1411] Patrick O’Luanaigh. Game design complete. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2006. ISBN 1-933097-00-0. xxi + 430 pp. LCCN

QA76.76.C672 O48 2006. URL http://www.oreilly.com/catalog/
9781933097008.

OQuinn:1997:PN

[1412] Donnie O’Quinn and Matt LeClair. Photoshop in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-313-8. 650 pp. LCCN

???? US$19.95. URL http://www.ora.com/catalog/photonut/;
http://www.oreilly.com/catalog/9781565923133;   http://www.
oreilly.com/catalog/photonut.

OQuinn:1998:QND

[1413] Donnie O’Quinn. QuarkXPress in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1998. ISBN 1-56592-399-5. xviii + 526 pp. LCCN Z253.532.Q37 O65 1998; Z253.532.Q37 O78 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565923997;   http://
www.oreilly.com/catalog/qxpnut.

OQuinn:1999:PND

[1414] Donnie O’Quinn. Photoshop in a Nutshell: a Desktop Quick Reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-565-3. xxvi + 632 pp. LCCN T385 .O68 1999.

US$24.95. URL http://www.oreilly.com/catalog/9781565925656;
 http://www.oreilly.com/catalog/photonut2/. Covers Adobe Pho-

toshop version 5.0.

OR:2001:PIO

[1415] O’Reilly Research. 2001 P2P Industry Overview: An O’Reilly Research Report. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN ???? 200 (est.) pp.

LCCN ???? US$1,295.00. URL http://www.oreilly.com/catalog/
p2presearch.

ORA:1996:WTW

[1416] O’Reilly & Associates, Inc. WebBoard (TM): Web Conferencing System Software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1996. ISBN 1-56592-181-X. 98 pp. LCCN ???? US$149. URL http://www.oreilly.com/

catalog/9781565921818. Includes 3 diskettes.

ORA:1996:WWW

[1417] O’Reilly & Associates, Inc. World Wide Web Journal: Volume 1, Issue 2: Key Specifications of the World Wide Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Spring 1996. ISBN 1-56592-190-9. 356 pp. LCCN ????

US$24.95. URL http://www.oreilly.com/catalog/9781565921900.

Oram:1991:MPM

[1418] Andrew Oram and Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1991. ISBN 0-937175-90-0. xiv + 136 pp. LCCN QA76.76.O63 T35 199.

URL http://www.oreilly.com/catalog/9780937175903;   http://
www.oreilly.com/catalog/make2.

Orebaugh:2005:SC

[1419] Angela Orebaugh, Simon Biles, and Jacob Babbin. Snort cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00791-4 (paperback). xiii + 270 pp. LCCN TK5105.59 .O74 2005. URL http://www.oreilly.com/

catalog/9780596007911.

OReilly:1986:UNS

[1420] Tim O’Reilly, Jean Diaz, et al. UNIX in a Nutshell: System V Edition: a Desktop Quick Reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1986. ISBN 0-937175-19-6 (paperback). 260 (est.) pp. LCCN QA76.76.O63 U5453 1990.

OReilly:1988:XWS

[1421] Tim O’Reilly, Valerie Quercia, and Linda Lamb. X Window System User’s Guide for Version 11, volume 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1988. ISBN 0-937175-29-3. xviii + 344 pp. LCCN QA76.76.W56 D44

v.3 1988. URL http://www.oreilly.com/catalog/9780937175293.

OReilly:1990:XTI

[1422] Tim O’Reilly. X Toolkit Intrinsics Reference Manual, volume 5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1990. ISBN 0-937175-35-8. xii + 543 pp. LCCN

QA76.76.W56 D44 v.5 1990. URL http://www.oreilly.com/catalog/
9780937175354.

OReilly:1991:GO

[1423] The Staff of O’Reilly & Associates. Guide to OSF/1. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1991. ISBN 0-937175-78-1. 304 pp. LCCN QA76.755.G85 1991. US$21.95. URL http://www.oreilly.com/

catalog/9780937175781.

OReilly:1993:COB

[1424] O’Reilly staff. Catalog of O’Reilly Books, 1993. URL http://www.

oreilly.com/catalog/. You get can it (book.catalog.Z) by anony-

mous ftp from ftp.ora.com where you can also find source codes for examples in many books in the Nutshell and X series … Or read the catalog on the O’Reilly gopher server (telnet gopher.ora.com, login:

gopher or gopher gopher.ora.com (if a gopher client is installed) . . .).

And to get a hard copy of their catalog-plus-magazine called ora.com,

mail letters@ora.com.

OReilly:1993:TTT

[1425] James O’Reilly and Larry Habegger, editors. Travelers’ Tales Thailand. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1993. ISBN 1-885211-05-8. 405 pp. LCCN ???? US$17.95.

URL http://www.oreilly.com/catalog/9781885211057;   http://
www.oreilly.com/catalog/thai.

OReilly:1994:TTM

[1426] James O’Reilly and Larry Habegger, editors. Travelers’ Tales Mexico. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1994. ISBN 1-885211-00-7. xvi + 405 pp. LCCN F1209 .T721 1994. US$17.95.

URL http://www.oreilly.com/catalog/9781885211002;   http://
www.oreilly.com/catalog/mexico.

OReilly:1995:TTF

[1427] James O’Reilly, Larry Habegger, and Sean O’Reilly, editors. Travelers’ Tales France. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1995. ISBN 1-885211-02-3. xvi + 414 pp. LCCN DC16 .T73 1995. US$17.95.

URL http://www.oreilly.com/catalog/9781885211026;   http://
www.oreilly.com/catalog/france.

OReilly:1995:TTI

[1428] James O’Reilly and Larry Habegger, editors. Travelers’ Tales India. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1995. ISBN 1-885211-01-5. xvi + 462 pp. LCCN DS406 .T73 1995. US$17.95.

URL http://www.oreilly.com/catalog/9781885211019;   http://
www.oreilly.com/catalog/india.

OReilly:1995:WWW

[1429] O’Reilly & Associates, Inc. and Web Consortium (W3C), editors. World Wide Web Journal: The Fourth International WWW Conference Proceedings, pages xii + 735, 1995. ISBN 1-56592-169-0. ISSN 1085-2301.

US$39.95. URL http://www.ora.com/gnn/bus/ora/item/wj1.html;
 http://www.oreilly.com/catalog/9781565921696. O’Reilly & Asso-

ciates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800. The World Wide Web Journal is a quarterly publication that provides timely, in-depth coverage of the issues, techniques, and research developments in the World Wide Web. The December issue contains the Conference Proceeding papers that were chosen for the 4th International World Wide Web conference in Boston, MA.

OReilly:1995:XCC

[1430] The Staff of O’Reilly & Associates. The X Companion CD for R6. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-084-8. 80

pp. LCCN ???? US$29.95. URL http://www.oreilly.com/catalog/
9781565920842.

OReilly:1996:RLCa

[1431] O’Reilly & Associates, Inc. and Red Hat Software. Running Linux Companion CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-171-2. 100 pp. LCCN ????

US$24.95. URL http://www.ora.com/gnn/bus/ora/item/runuxcd.
html;   http://www.oreilly.com/catalog/9781565921719.

OReilly:1996:RLCb

[1432] O’Reilly & Associates, Inc. and Red Hat Software. Running Linux Companion CD-ROM: RedHat Linux Release 3.03. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, August 1996. ISBN 1-56592-212-3. 140

pp. LCCN ???? US$24.95. URL http://www.oreilly.com/catalog/
9781565922129.

OReilly:1996:TTH

[1433] James O’Reilly, Larry Habegger, and Sean O’Reilly, editors. Travelers’ Tales Hong Kong. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885211-03-1. xvi + 409 pp. LCCN DS796.H73 T77 1996. US$17.95.

URL http://www.ora.com/gnn/bus/ora/item/hong.html;   http:/
/www.oreilly.com/catalog/9781885211033;   http://www.oreilly.
com/catalog/hong.

OReilly:1996:TTS

[1434] James O’Reilly, Larry Habegger, and Sean O’Reilly, editors. Travelers’ Tales San Francisco. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-885211-08-2. xvi + 409 pp. LCCN F869.S33 T73 1996. US$17.95.

URL http://www.oreilly.com/catalog/9781885211088;   http://
www.oreilly.com/catalog/sf.

OReilly:1996:W

[1435] O’Reilly & Associates, Inc., Susan Peck, and Stephen Arrants. Web-Site(TM) 1.1. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-173-9. 494 pp. LCCN ???? US$249.00. URL http://www.oreilly.com/ catalog/9781565921733. Includes 4 diskettes and WebSite T-shirt.

OReilly:1996:WPT

[1436] O’Reilly & Associates, Inc. and Susan Peck. WebSite Professional (TM). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1996. ISBN 1-56592-174-7.

LCCN ???? US$499.00. URL http://www.oreilly.com/catalog/

9781565921740. Includes CD-ROM and 3 books.

OReilly:1997:ST

[1437] O’Reilly & Associates, Inc. Statisphere(TM). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-233-6. 135 pp. LCCN ????

US$399. URL http://www.ora.com/catalog/statisphere/;   http:
//www.oreilly.com/catalog/9781565922334. Includes diskettes.

OReilly:1997:TTP

[1438] James O’Reilly & Larry Habegger and Sean O’Reilly. Travelers’ Tales Paris. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-10-4. xx + 395 pp. LCCN DC707 .T73 1997. US$17.95.

URL http://www.ora.com/catalog/paris/noframes.html;   http:
//www.oreilly.com/catalog/9781885211101;   http://www.oreilly.
com/catalog/paris.

OReilly:1997:TTR

[1439] Sean O’Reilly, James O’Reilly, and Tim O’Reilly, editors. Travelers’ Tales: The Road Within. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-19-8. xix + 435 pp. LCCN G180

.R62 1997. US$17.95. URL http://www.ora.com/catalog/road/;
http://www.oreilly.com/catalog/9781885211194;   http://www.
oreilly.com/catalog/road.

OReilly:1997:WDL

[1440] O’Reilly & Associates, Inc. Web Developer’s Library on the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-363-4. ???? pp.

LCCN ???? US$59.95. URL http://www.ora.com/catalog/webrlw/
;   http://www.oreilly.com/catalog/9781565923638;   http://
www.oreilly.com/catalog/webrlw.

OReilly:1997:WN

[1441] O’Reilly & Associates, Inc. WebMaster in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, deluxe edition, 1997. ISBN 1-56592-305-7. 356 (est.) pp. LCCN ???? US$69.95. URL http://www.ora.com/catalog/

webdeluxe/;   http://www.oreilly.com/catalog/9781565923058;
 http://www.oreilly.com/catalog/webdeluxe.

OReilly:1998:CWN

[1442] O’Reilly & Associates, Inc. The Complete Windows NT and UNIX System Administration Pack. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN ???? ???? pp. LCCN ???? US$149.00. URL http://www.

oreilly.com/catalog/hp12pack.

OReilly:1999:CWN

[1443] O’Reilly & Associates, Inc. Complete Windows NT & UNIX System Administration Pack. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ????

pp. LCCN ???? US$149.00. URL http://www.oreilly.com/catalog/
hp12pack.

OReilly:1999:DTS

[1444] James O’Reilly & Larry Habegger and Sean O’Reilly. Danger! True Stories of Trouble and Survival. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-32-5. 375 pp. LCCN GB5018.D25 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211323;   http://
www.oreilly.com/catalog/danger/.

OReilly:1999:GCT

[1445] Sean O’Reilly & James O’Reilly & Larry Habegger. Grand Canyon: True Stories of Life Below the Rim. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-34-1. 296 pp. LCCN F788.G756 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211347;   http://
www.oreilly.com/catalog/canyon.

OReilly:1999:NCB

[1446] O’Reilly & Associates, Inc. The Networking CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-523-8. 456 pp. LCCN ????

US$79.95. URL http://www.oreilly.com/catalog/9781565925236;
 http://www.oreilly.com/catalog/netcd/. Six O’Reilly networking

books on CD-ROM.

OReilly:1999:TPT

[1447] Sean O’Reilly, Larry Habegger, and James O’Reilly, editors. Testosterone Planet: True Stories from a Man’s World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-43-0. 328 pp. LCCN G525.T39 1999.

US$17.95. URL http://www.oreilly.com/catalog/9781885211439;
 http://www.oreilly.com/catalog/tplanet.

OReilly:1999:WND

[1448] Tim O’Reilly & Troy Mott and Walter J. Glenn. Windows 98 in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-486-X. xxii + 618 pp. LCCN QA76.76.O63 O72 1999; QA76.76.O63 O735 1999. US$24.95. URL http:

//www.oreilly.com/catalog/9781565924864.

OReilly:1999:WWC

[1449] O’Reilly & Associates, Inc. WebBoard 4.0: Web Conferencing Software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-602-1. ????

pp. LCCN ???? US$1,199. URL http://www.oreilly.com/catalog/
9781565926028;   http://www.oreilly.com/catalog/wboard4.

OReilly:2000:OPS

[1450] O’Reilly & Associates, Inc. The Oracle PL/SQL CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2000. ISBN 1-56592-849-0. 285 pp.

LCCN ???? US$89.95. URL http://www.oreilly.com/catalog/
9781565928497;   http://www.oreilly.com/catalog/oraclecdbs.

Includes CD-ROM.

OReilly:2002:NCB

[1451] O’Reilly & Associates, Inc. Networking CD Bookshelf, Version 2.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00334-X. 752 (est.)

pp. LCCN ???? US$119.95. URL http://www.oreilly.com/catalog/
9780596003340;   http://www.oreilly.com/catalog/netcd2.

OReilly:2002:PCB

[1452] O’Reilly & Associates, Inc. Perl CD Bookshelf, Version 3.0: 7 Bestselling Books on CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00389-7. 800 (est.) pp. LCCN ???? US$119.95.

URL http://www.oreilly.com/catalog/9780596003890;   http://
www.oreilly.com/catalog/perlcdbs3. Includes a Perl in a Nutshell,

2nd Edition.

OReilly:2002:XCB

[1453] O’Reilly and Inc. Associates. XML CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00335-8. 600 (est.) pp. LCCN ????

US$119.95. URL http://www.oreilly.com/catalog/xmlcdbksf.

Oreilly:MUU88

[1454] Tim O’Reilly and Grace Todino. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1988. ISBN 0-937175-09-9. xvi + 256 pp. LCCN QA76.8.U65 O64 1988. URL http://www.oreilly.com/

catalog/9780937175095.

Orsini:2007:RC

[1455] Rob Orsini. Rails cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52731-4. xviii + 514 pp. LCCN QA76.73.R83 O77 2007; QA76.73.R83; QA76.73.R83 O77 2007eb. URL http://www.oreilly.

com/catalog/9780596527310.

Orwant:1999:MAP

[1456] Jon Orwant, Jarkko Hietaniemi, and John Macdonald. Mastering algorithms with Perl: Practical programming through computer science. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-398-7. xvii + 684 pp. LCCN QA76.73.P22 O788 1999; QA76.9.A43 .O78 1999.

US$34.95. URL http://www.oreilly.com/catalog/9781565923980;
 http://www.oreilly.com/catalog/maperl.

Orwant:2002:CSP

[1457] Jon Orwant, editor. Computer Science & Perl Programming: Best of The Perl Journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00310-2. xvii + 737 pp. LCCN QA76.73.P22 C383 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596003104;   http://
www.oreilly.com/catalog/tpj1.

Orwant:2003:GDP

[1458] Jon Orwant, editor. Games, diversions, and Perl culture: best of the Perl Journal. The best of the Perl journal series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00312-9. xiii + 569 pp. LCCN QA76.73.P22

G36 2003. URL http://www.oreilly.com/catalog/9780596003128.

Orwant:2003:WGP

[1459] Jon Orwant, editor. Web, graphics, and Perl/TK: best of the Perl Journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00311-0. xiii + 432 pp. LCCN T385 .W394 2003. URL http://www.oreilly.com/

catalog/9780596003111.

Oster:2000:MIM

[1460] Nancy Oster, Lucy Thomas, and Darol Joseff. Making informed medical decisions: where to look and how to use what you find. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-459-2. xvii + 364 pp. LCCN RA776.5 .O684 2000. URL http://www.oreilly.com/

catalog/9781565924598.

OSullivan:2009:RWH

[1461] Bryan O’Sullivan, Don Stewart, and John Goerzen. Real world Haskell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15533-6, 0-596-51498-0. xxxviii + 670 pp. LCCN QA76.73.H37 O88 2008.

Oualline:1991:PCP

[1462] Steve Oualline. Practical C Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-65-X. xxii + 396 pp. LCCN QA76.73.C15 O84 1991. US$24.95. URL http://www.oreilly.com/ catalog/9780937175651. Yet another good C book describing how to create programs that are easy to read, maintain and debug.

Oualline:1993:PCP

[1463] Steve Oualline. Practical C Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, January 1993. ISBN 1-56592-035-X. xxii + 396 pp. LCCN QA76.73.C15O93 1993. US$24.95. URL http://www.oreilly.

com/catalog/9781565920354.

Oualline:1995:PCP

[1464] Steve Oualline. Practical C++ Programming. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-139-9. xxiv + 557 pp. LCCN QA76.73.C153 O833 1995. US$32.95.

URL http://www.oreilly.com/catalog/9781565921399;   http://
www.oreilly.com/catalog/cplus.

Oualline:1997:PCP

[1465] Steve Oualline. Practical C Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1997. ISBN 1-56592-306-5. xxiii + 428 pp. LCCN QA76.73.C15 O84 1997. US$32.95. URL

http://www.ora.com/catalog/pcp3/;   http://www.oreilly.com/
catalog/9781565923065;   http://www.oreilly.com/catalog/pcp3.

Oualline:2003:HPC

[1466] Steve Oualline. How not to program in C++: 111 broken programs and 3 working ones, or why does 2+2 = 5986 ? No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-95-6. 265 pp. LCCN QA76.73.C153 O832

2003eb. URL http://www.oreilly.com/catalog/9781886411951.

Oualline:2003:PCP

[1467] Steve Oualline. Practical C++ programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00258-0. xxi + 549 pp. LCCN

QA76.73.C153 O833 2003. URL http://www.oreilly.com/catalog/
9780596002589.

Oualline:2006:WCP

[1468] Steve Oualline. Wicked cool Perl scripts: useful Perl scripts that solve difficult problems. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-062-3 (paperback). xvii + 312 pp. LCCN QA76.73.P22 Q523 2006eb; QA76.73.P22 Q523 2006. URL http://www.oreilly.

com/catalog/9781593270629.

Oxer:2006:UH

[1469] Jonathan Oxer, Kyle Rankin, and Bill Childers. Ubuntu Hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52720-9. xix + 426 pp. LCCN QA76.76.O63 2006. EUR 29.00. URL http://www.oreilly.com/

catalog/9780596527204.

Pachev:2006:UMI

[1470] Alexander Pachev. Understanding MySQL internals. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00957-7. ???? pp. LCCN QA76.73.S67; Internet; QA76.73.S67 .P33eb. URL http://www.

oreilly.com/catalog/9780596009571.

Palaia:2005:DMO

[1471] Wayne Palaia and Christina Palaia. Degunking Microsoft Office. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-932111-95-6. xxvi + 385 pp. LCCN HF5548.4.M525 P35 2005. URL http://www.

oreilly.com/catalog/9781932111958.

Palmer:2005:ASM

[1472] Scott D. Palmer and Kate Chase. Access for starters: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00665-9 (paperback). xiv + 379 pp. LCCN HF5548.4.M525 C43

2006eb; HF5548.4.M525. URL http://www.oreilly.com/catalog/
9780596006655.

Parker:2005:SYE

[1473] Jim Parker. Start your engines: developing driving and racing games. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-01-9. xxii + 434 pp. LCCN GV1469.17.A98 P37 2005. URL http://www.

oreilly.com/catalog/9781933097015.

Parks:2006:MAK

[1474] Bob Parks. Makers: all kinds of people making amazing things in garages, basements, and backyards. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10188-0 (hardcover). 183 pp. LCCN T39 .P265 2006. URL http://www.

oreilly.com/catalog/9780596101886.

Parr:2007:DAR

[1475] Terence Parr. The definitive ANTLR reference: building domain-specific languages. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 0-9787392-5-6. xx + 361 pp. LCCN QA76.7 .P37

2007. URL http://www.oreilly.com/catalog/9780978739256.

Parr:2010:LIP

[1476] Terence (Terence John) Parr. Language implementation patterns: create your own domain-specific and general programming languages. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-45-X (paperback). xix + 374 pp. LCCN QA76.7 .P375 2010.

Patrick:2006:VBC

[1477] Tim Patrick and John Clark Craig. Visual Basic 2005 cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10177-5. xxiv + 713 pp. LCCN QA76.73.B3;

QA76.73.B3 P37 2006eb. URL http://www.oreilly.com/catalog/
9780596101770.

Patrick:2008:PVB

[1478] Tim Patrick. Programming Visual Basic 2008: build .NET 3.5 applications with Microsoft’s popular RAD Tool for business. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51843-9. xx + 757 pp. LCCN QA76.73.B3 P38 2008.

Patwardhan:2002:PND

[1479] Nate Patwardhan, Ellen Siever, and Stephen Spainhour. Perl in a nutshell: a Desktop Quick Reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00241-6. xx + 737 pp. LCCN QA76.73.P22 S554 2002. US$39.95. URL http://safari.oreilly.com/

0596002416;   http://www.oreilly.com/catalog/9780596002411;
 http://www.oreilly.com/catalog/perlnut2.

Paul:2005:DVH

[1480] Joshua Paul. Digital video hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00946-1 (paperback). xix + 404 pp. LCCN TK6680.5 .P38 2005. URL http://

www.oreilly.com/catalog/9780596009465.

Pawson:2002:XFM

[1481] Dave Pawson. XSL-FO: Making XML Look Good in Print. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00355-2. xv + 263 pp. LCCN QA76.76.H94 P38 2002. US$34.95. URL http://safari.oreilly.com/

0596003552;   http://www.oreilly.com/catalog/9780596003555;
 http://www.oreilly.com/catalog/xslfo.

Pearce:1997:WNN

[1482] Eric Pearce. Windows NT in a Nutshell: a Desktop Quick Reference for System Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-251-4. 348 pp. LCCN QA76.76.W56 P43

1997. US$24.95. URL http://www.ora.com/catalog/winnut/;
http://www.oreilly.com/catalog/9781565922518;   http://www.
oreilly.com/catalog/winnut.

Peck:1995:W

[1483] Susan Peck and Linda Mui. WebSite. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1995. ISBN 1-56592-143-7. 342 pp. LCCN ???? US$499. URL http://

www.oreilly.com/catalog/9781565921436. Includes two diskettes.

Peck:1996:BYO

[1484] Susan B. Peck and Stephen Arrants. Building Your Own WebSite: Everything You Need to Reach Your Audience on the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1996. ISBN 1-56592-232-8. xiv + 500 pp. LCCN TK5105.888 .P4 1996. US$59.00.

URL http://www.oreilly.com/catalog/9781565922327;   http://
www.oreilly.com/catalog/buildweb. Includes CD-ROM.

Peck:1997:BYOa

[1485] Susan B. Peck and Beverly Murray Scherf. Building Your Own Web Conferences. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-279-4. xii + 246 pp. LCCN TK5105.888 .P43

1997. US$59.95. URL http://www.ora.com/catalog/buildcon/;
http://www.oreilly.com/catalog/9781565922792;   http://www.
oreilly.com/catalog/buildcon.

Peck:1997:BYOb

[1486] Susan B. Peck and Beverly Murray Scherf. Building your own Web conferences. Build your own. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-339-1. xii + 246 pp. LCCN TK5105.888 .P43 1997. URL http:

//www.oreilly.com/catalog/9781565923393. Special edition: docu-

mentation for electronic software distribution only.

Peck:1998:W

[1487] Susan B. Peck, Jay York, and Chris Duke. WebBoard 3.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-56592-429-0. 358 (est.)

pp. LCCN ???? US$699.00. URL http://www.oreilly.com/catalog/
9781565924291;   http://www.oreilly.com/catalog/wboard3.

Pedro:2007:PSL

[1488] Bruno Pedro and Vitor Rodrigues. PHP and Smarty on large scale Web development. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51379-8. LCCN TK5105.888; TK5105.888 .P43 2007eb. URL http://www.

oreilly.com/catalog/9780596513795.

Pedroni:2002:JE

[1489] Samuele Pedroni and Noel Rappin. Jython essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00247-5. xx + 277 pp. LCCN QA76.73.J38 P43 2002. US$24.95.

URL http://www.oreilly.com/catalog/9780596002473;   http://
www.oreilly.com/catalog/jythoness.

Peek:1992:MXE

[1490] Jerry Peek. MH and xmh: E-mail for Users and Programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, September 1992. ISBN 1-56592-027-9. 728 pp. LCCN HE6239.E54 P44 1992. US$29.95. URL

http://www.oreilly.com/catalog/9781565920279.

Peek:1993:UPT

[1491] Jerry Peek, Mike Loukides, and Tim O’Reilly & others. UNIX Power Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1993. ISBN 0-679-79073-X (Random House). xlii + 1119 pp. LCCN QA76.76.O63 P44 1993.

US$59.95. URL http://www.oreilly.com/catalog/9780679790730.

Peek:1994:UPT

[1492] Jerry D. Peek and Tim O’Reilly & Michael Kosta Loukides. UNIX power tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1994. ISBN 1-56592-260-3. xliii + 1073 pp. LCCN QA76.76.O63 P44 1993. URL

http://www.oreilly.com/catalog/9781565922600.

Peek:1995:MXE

[1493] Jerry D. Peek, Bill Wohler, and Brent B. Welch. MH and xmh: e-mail for users and programmers. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, April 1995. ISBN 1-56592-093-7. xlii + 738 pp. LCCN HE6239.E54 P44 1995; HE7551 .P44 1995. US$34.95. URL

http://www.oreilly.com/catalog/9781565920934.

Peek:1997:UPT

[1494] Jerry Peek, Tim O’Reilly & Mike Loukides, Linda Mui, et al. UNIX Power Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-260-3. xliii + 1073 pp. LCCN QA76.76.O63

P44 1997. US$59.95. URL http://www.ora.com/catalog/upt2/;
http://www.oreilly.com/catalog/9781565922600;   http://www.
oreilly.com/catalog/upt2.

Peek:1998:LUO

[1495] Jerry D. Peek, Grace Todino, and John Strang. Learning the UNIX operating system. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 1998. ISBN 1-56592-390-1. xi + 92 pp. LCCN QA76.76.O63 T62 1998. US$10.95.

URL http://www.oreilly.com/catalog/9781565923904;   http://
www.oreilly.com/catalog/lunix4. Additions and revisions by Jerry

Peek.

Peek:2001:LUO

[1496] Jerry D. Peek, Grace Todino, and John Strang. Learning the UNIX operating system. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2001. ISBN 0-596-00261-0. xiii + 157 pp. LCCN QA76.76.O63 T62 2002. URL

http://www.oreilly.com/catalog/9780596002619.

Peikari:2004:SW

[1497] Cyrus Peikari and Anton Chuvakin. Security warrior. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00545-8. xvii + 531 pp. LCCN

TK5105.59 .P44985 2004. URL http://www.oreilly.com/catalog/
9780596005450.

Perry:2001:AND

[1498] Bruce W. Perry. AppleScript in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 1-56592-841-5. xxiii + 499 pp. LCCN QA76.73.A67 P47 2001.

US$29.95. URL http://www.oreilly.com/catalog/9781565928411;
 http://www.oreilly.com/catalog/aplscptian.

Perry:2002:JME

[1499] J. Steven (James Steven) Perry. Java Management Extensions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00245-9. x + 300 pp. LCCN QA76.73.J38 P472 2002. US$34.95. URL http://safari.oreilly.com/

0596002459;   http://www.oreilly.com/catalog/9780596002459;
 http://www.oreilly.com/catalog/javamngext.

Perry:2004:JSJ

[1500] Bruce W. Perry. Java servlet and JSP cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00572-5. xix + 723 pp. LCCN QA76.73.J38

P469 2004. URL http://www.oreilly.com/catalog/9780596005726.

Perry:2006:AH

[1501] Bruce W. Perry. Ajax hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10169-4. xxi + 414 pp. LCCN QA76.73.J39

P47 2006eb; QA76.73.J39. URL http://www.oreilly.com/catalog/
9780596101695.

Perry:2006:GWT

[1502] Bruce W. Perry. Google Web toolkit for Ajax. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-51022-5. LCCN TK5105.8885.A52 P47

2006eb; TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596510220.

Persse:2006:PIE

[1503] James R. Persse. Process improvement essentials. Theory in practice (Sebastopol, Calif.). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10217-8. xiv + 334 pp. LCCN QA76.76.Q35 P473 2006eb; QA76.76.Q35 P47 2006; QA76.76.Q35; QA76.76.Q35 P473 2006. URL http://www.

oreilly.com/catalog/9780596102173.

Peterson:2005:WSM

[1504] Eric T. Peterson. Web site measurement hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00988-7. 352 pp. LCCN TK5105.888 .P48 2005; TK5105.888; TK5105.888 .P48 2005eb. URL http://www.

oreilly.com/catalog/9780596009885.

Petreley:2005:LDH

[1505] Nick Petreley and Jono Bacon. Linux desktop hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00911-9. xxi + 318 pp. LCCN QA76.76.O63 P528645

2005. URL http://www.oreilly.com/catalog/9780596009113.

Petrich:1997:NIN

[1506] Dean Petrich and David Flanagan. Netscape IFC in a nutshell: a desktop quick reference for Java programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-343-X. xiii + 354 pp. LCCN QA76.73.J38 P38 1997; QA76.73.J38 P477 1997. US$19.95.

URL http://www.ora.com/catalog/ifcnut/author.html;   http:/
/www.oreilly.com/catalog/9781565923430. Covers IFC 1.1.

Petrov:2019:DID

[1507] Alex Petrov. Database Internals: a Deep Dive Into How Distributed Data Systems Work. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4920-4034-7 (paperback), 1-4920-4030-4 (e-book), 1-4920-4031-2 (e-book),. xix + 349 pp. LCCN

QA76.9.D3 P48 2019. URL http://proquest.safaribooksonline.
com/?fpi=9781492040330.

Petrusha:1996:IWR

[1508] Ron Petrusha. Inside the Windows 95 Registry: a Guide for Programmers, System Administrators, and Users. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-56592-170-4. 594 pp. LCCN ???? US$32.95.

URL http://www.oreilly.com/catalog/9781565921702;   http://
www.oreilly.com/catalog/winreg.

Picht:2009:XKI

[1509] Hans-Joachim Picht and Sebastian Ott. Xen Kochbuch: intelligente Virtualisierungslösungen mit XEN 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 3-89721-729-5. xix + 461 pp. LCCN ???? EUR 44.90. URL

http://www.gbv.de/dms/ilmenau/toc/572426836.PDF.

Pilato:2008:VCS

[1510] C. Michael Pilato, Ben Collins-Sussman, and Brian W. Fitzpatrick. Version control with Subversion. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-51033-0 (paperback). xxiv + 404 pp. LCCN

QA76.6 .C6274 2008. URL http://svnbook.red-bean.com/.

Pilgrim:2005:GH

[1511] Mark Pilgrim. Greasemonkey hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10165-1. 352 pp. LCCN TK5105.882eb; TK5105.882.

URL http://www.oreilly.com/catalog/9780596101657.

Pilgrim:2010:HR

[1512] Mark Pilgrim. HTML5: up and running. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-80602-7. 250 pp. LCCN ????

Pilone:2003:UPR

[1513] Dan Pilone. UML pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00497-4. vi + 81 pp. LCCN QA76.9.O35 P55 2003.

URL http://www.oreilly.com/catalog/9780596004972.

Pilone:2005:UN

[1514] Dan Pilone and Neil Pitman. UML 2.0 in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00795-7 (paperback). xv + 216 pp. LCCN QA76.9.O35

P54 2003. URL http://www.oreilly.com/catalog/9780596007959.

Pilone:2007:HFS

[1515] Dan Pilone. Head first software development. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52735-7 (paperback). xxxvi + 459 pp. LCCN QA76.76.D47; QA76.76.D47eb. URL http://www.oreilly.com/

catalog/9780596527358.

Pine:2005:LP

[1516] Chris Pine. Learn to program. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-4-2 (paperback). xiv + 149 pp. LCCN QA76.64

.P528 2005. URL http://www.oreilly.com/catalog/9780976694045.

Plate:2005:DYP

[1517] Shannon Plate. Degunking your personal finances. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-02-7. xxviii + 334

pp. LCCN HG179 .P53 2005. URL http://www.oreilly.com/catalog/
9781933097022.

Platzer:2013:BKG

[1518] Jörg Platzer. Bitcoin — kurz and gut. (German) [Bitcoin — short and good]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2013. ISBN 3-95561-650-9 (print), 3-95561-653-3. 204 pp. LCCN HG1710 .P53 2014.

Plumlee:2007:TDA

[1519] Philip C. Plumlee. Test Driven Ajax (on Rails). O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51065-9. LCCN TK5105.8885.A52 P56

2007eb; TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596510657.

Pogue:1998:PUG

[1520] David Pogue. PalmPilot: The Ultimate Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1998. ISBN 1-56592-420-7. xviii + 489 pp. LCCN QA

76.5 P64 1998. US$29.95. URL http://www.oreilly.com/catalog/
9781565924208;   http://www.oreilly.com/catalog/palmpilot.

Pogue:1999:PUG

[1521] David Pogue. PalmPilot: The Ultimate Guide: Mastering Palm Organizers from Pilot 1000 to Palm VII. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-600-5. xxiii + 597 pp. LCCN QA76.8.P138

.P64 1999. US$29.95. URL http://www.oreilly.com/catalog/
9781565926004;   http://www.oreilly.com/catalog/palmpilot2.

Pogue:2000:IMM

[1522] David Pogue. iMovie: The Missing Manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, May 2000. ISBN 1-56592-859-8. xii + 373 pp. LCCN TR899 .P64 2000. US$19.95.

URL http://www.oreilly.com/catalog/9781565928596;   http://
www.oreilly.com/catalog/imoviemm.

Pogue:2000:MMM

[1523] David Pogue. Mac OS 9: The Missing Manual. The missing manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, March 2000. ISBN 1-56592-857-1. x + 461 pp. LCCN QA76.76.O63 P63 2000; QA76.76.O63

P634 2000. US$19.95. URL http://www.oreilly.com/catalog/
9781565928572;   http://www.oreilly.com/catalog/macosmm.

Pogue:2000:WMM

[1524] David Pogue. Windows Me: The Missing Manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, September 2000. ISBN 0-596-00009-X. x + 413 pp. LCCN QA76.76.O63 P635 2000; QA76.76.O63 .P64 2000.

US$19.95. URL http://www.oreilly.com/catalog/9780596000097;
 http://www.oreilly.com/catalog/winmemm.

Pogue:2001:IMM

[1525] David Pogue. iMovie 2: The Missing Manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, January 2001. ISBN 0-596-00104-5. xii + 402 pp. LCCN TR896 .P64 2001; TR896

.P642 2001. US$19.95. URL http://www.oreilly.com/catalog/
9780596001049;   http://www.oreilly.com/catalog/imoviemm2.

Pogue:2002:IMM

[1526] David Pogue, Joseph Schorr, and Derrick Story. iPhoto: The Missing Manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2002. ISBN 0-596-00365-X. xii + 288 pp. LCCN

TR267 .P64 2002. US$24.95. URL http://www.oreilly.com/catalog/
9780596003654;   http://www.oreilly.com/catalog/iphototmm.

Pogue:2002:MX

[1527] David Pogue. Mac OS X. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2002. ISBN 0-596-00450-8. xii + 712 pp. LCCN QA76.76.O63 P634 2002.

Pogue:2002:MXM

[1528] David Pogue. Mac OS X: The Missing Manual. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2002. ISBN 0-596-00450-8. xii + 712 pp. LCCN QA76.76.O63

P6344 2002. US$29.95. URL http://www.oreilly.com/catalog/
9780596004507;   http://www.oreilly.com/catalog/macosxmm2.

Pogue:2002:PPG

[1529] David Pogue and Nancy Kotary. Palm Pocket Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00426-5. 128 (est.) pp. LCCN ????

US$12.95. URL http://www.oreilly.com/catalog/9780596004262;
 http://www.oreilly.com/catalog/palmpg.

Pogue:2002:WXH

[1530] David Pogue. Windows XP Home Edition: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00260-2. xii + 571 pp. LCCN QA76.76.O63 P645 2002. US$24.95.

URL http://www.oreilly.com/catalog/9780596002602;   http://
www.oreilly.com/catalog/windowsxptmm.

Pogue:2003:I

[1531] David Pogue, Joseph Schorr, and Derrick Story. iPhoto 2. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00506-7. xii + 336 pp. LCCN TR267 .P64 2003. URL

http://www.oreilly.com/catalog/9780596005061.

Pogue:2003:IIM

[1532] David Pogue. IMovie 3 and iDVD: the missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00507-5. xii + 450 pp. LCCN TR896 .P64 2003. URL http://www.

oreilly.com/catalog/9780596005078.

Pogue:2003:MXP

[1533] David Pogue. Mac OS X Panther edition: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00615-2. xii + 763 pp. LCCN QA76.76.O63 P63443

2003. URL http://www.oreilly.com/catalog/9780596006150.

Pogue:2003:SMM

[1534] David Pogue. Switching to the Mac: The Missing Manual. Missing manual series. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2003. ISBN 0-596-00452-4. xii + 434 pp. LCCN QA76.76.O63

P63445 2003. US$19.95. URL http://www.oreilly.com/catalog/
9780596004521;   http://www.oreilly.com/catalog/switchmacmm.

Pogue:2003:WXP

[1535] David Pogue, Craig Zacker, and L. J. Zacker. Windows XP Pro: the missing manual. The missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00348-X. xii + 658 pp. LCCN QA76.76.O63 P636

2003. URL http://www.oreilly.com/catalog/9780596003487.

Pogue:2004:I

[1536] David Pogue and Derrick Story. iPhoto 4. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2004. ISBN 0-596-00692-6. xii + 361 pp. LCCN TR267.5.I64 P64 2004. URL http://www.

oreilly.com/catalog/9780596006921.

Pogue:2004:IIM

[1537] David Pogue. IMovie 4 and iDVD: the missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2004. ISBN 0-596-00693-4 (paperback). xii + 484 pp. LCCN TR896 .P64 2004. URL

http://www.oreilly.com/catalog/9780596006938.

Pogue:2004:WXH

[1538] David Pogue. Windows XP Home Edition: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, second edition, 2004. ISBN 0-596-00897-X. 612 pp. LCCN QA76.76.O63

P6355 2004. URL http://www.oreilly.com/catalog/9780596008970.

Pogue:2004:WXP

[1539] David Pogue, Craig Zacker, and L. J. Zacker. Windows XP Pro: the missing manual. The missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004.

ISBN 0-596-00898-8. 689 pp. LCCN QA76.76.O63 P636 2003. URL

http://www.oreilly.com/catalog/9780596008987.

Pogue:2005:GMM

[1540] David Pogue. GarageBand 2: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10035-3 (paperback). x + 256

pp. LCCN MT56 .P64eb. URL http://www.oreilly.com/catalog/
9780596100353.

Pogue:2005:IHI

[1541] David Pogue. IMovie HD and iDVD 5: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-10033-7 (paperback). xii + 516 pp. LCCN TR899

.P64 2005. URL http://www.oreilly.com/catalog/9780596100339.

Pogue:2005:IMMa

[1542] David Pogue. iLife ’05: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10036-1 (paperback). xvi + 724 pp.

LCCN QA76.575 .P64 2005. URL http://www.oreilly.com/catalog/
9780596100360;   http://www.oreilly.com/catalog/ilife5tmm/.

Pogue:2005:IMMb

[1543] David Pogue and Derrick Story. iPhoto 5: the missing manual. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10034-5. xii + 388 pp.

LCCN TR267 .P64 2005. URL http://www.oreilly.com/catalog/
9780596100346.

Pogue:2005:MXT

[1544] David Pogue. Mac OS X, Tiger edition: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00941-0 (paperback). xvi + 847 pp. LCCN QA76.76.O63 P634 2005. URL http:/

/www.oreilly.com/catalog/9780596009410.

Pogue:2005:SMM

[1545] David Pogue. Switching to the Mac: the missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, tiger edition, 2005. ISBN 0-596-00660-8. 450 pp. LCCN QA76.76.O63; QA76.76.O63eb. URL http://www.oreilly.com/

catalog/9780596006600.

Pogue:2006:DP

[1546] David Pogue, Barbara (Barbara K.) Brundage, and Andy Rathbone. Digital photography. The missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00841-4 (paperback). xv + 403 pp. LCCN

TR267 .G47 2006. US$20.99. URL http://www.oreilly.com/catalog/
9780596008413.

Pogue:2006:IIM

[1547] David Pogue. IMovie 6 and iDVD: the missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-52726-8 (paperback). xii + 499 pp. LCCN TR896 .P646 2006eb;

TR896 .P646 2006; TR896. URL http://www.oreilly.com/catalog/
9780596527266.

Pogue:2006:IMMa

[1548] David Pogue and Derrick Story. iPhoto 6: the missing manual. The missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-52725-X. xii + 395 pp. LCCN TR267; TR267 .P64 2006; TR267 .P64 2006eb. URL http://

www.oreilly.com/catalog/9780596527259.

Pogue:2006:IMMb

[1549] David Pogue and J. D. Biersdorfer. The Internet: the missing manual. The missing manual series. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-52742-X. xv + 442 pp. LCCN TK5105.875.I57; TK5105.875.I57 P584 2006eb; TK5105.875.I57 P584

2006. URL http://www.oreilly.com/catalog/9780596527426.

Pogue:2006:WXS

[1550] David Pogue. Windows XP for starters: exactly what you need to get started. The missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10155-4. xv + 385 pp. LCCN QA76.76.O63; QA76.76.O63 P6353 2006eb; QA76.76.O63 P6353 2006.

URL http://www.oreilly.com/catalog/9780596101558.

Pogue:2007:IIM

[1551] David Pogue. iMovie 08 and iDVD: the missing manual. The missing manual series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51619-3. 560 (est.) pp. LCCN TR899 .P642 2007; TR899; TR899 .P642 2007eb. URL http://

www.oreilly.com/catalog/9780596516192.

Pogue:2007:IMM

[1552] David Pogue. iPhone: the missing manual. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51374-7 (paperback), 0-596-51374-7. x + 292 pp. LCCN TK6570.M6 P64 2007eb; TK6570.M6; TK6570.M6 P64 2007.

URL http://www.oreilly.com/catalog/9780596513740.

Pogue:2007:MXL

[1553] David Pogue. Mac OS X Leopard edition: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2007. ISBN 0-596-52952-X. xvi + 893 pp. LCCN QA76.76.O63

P634428 2007. URL http://www.loc.gov/catdir/enhancements/
fy0715/2004558625-d.html;   http://www.oreilly.com/catalog/
9780596529529.

Pogue:2007:WV

[1554] David Pogue. Windows Vista. The missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52827-2. xvi + 827 pp. LCCN QA76.76.O63; QA76.76.O63 P63525 2007eb. URL http://www.

oreilly.com/catalog/9780596528270.

Pogue:2007:WVS

[1555] David Pogue. Windows Vista for starters: the missing manual: exactly what you need to get started. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52826-4. xiv + 408 pp. LCCN QA76.76.O63;

QA76.76.O63 P643 2007eb. URL http://www.oreilly.com/catalog/
9780596528263.

Pogue:2012:WMM

[1556] David Pogue. Windows 8: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1403-1. 900 pp. LCCN ???? EUR 32.00.

Popp:2015:DSA

[1557] Stefan Popp and Ralf Peters. Durchstarten mit Swift: [Apples neue Programmiersprache Schritt für Schritt verständlich und gut erklärt]. (German) [Getting started with Swift: [Apple’s new programming language step by step understandable and well-explained]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 3-95561-884-6. xiii + 249 pp. LCCN ???? URL

http://d-nb.info/105779788X/04.

Potter:2003:S

[1558] Bruce Potter and Bob Fleck. 802.11 Security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00290-4. xiii + 176 pp. LCCN TK5105.78 .P68 2003. US$34.95, CDN$54.95, UK£24.95.

URL http://www.oreilly.com/catalog/80211security/;   http:/
/www.oreilly.com/catalog/9780596002909.

Powers:1993:LYJ

[1559] Paul Powers and Deborah Russell. Love Your Job!: Loving the Job You Have… Finding a Job You Love. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1993. ISBN 1-56592-036-8. xi + 198 pp. LCCN HF5382.7 .P69 1993.

US$12.95. URL http://www.oreilly.com/catalog/9781565920361.

Powers:1999:DAC

[1560] Shelley Powers. Developing ASP Components. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-446-0. xvii + 490 pp. LCCN TK5105.8885.A26 P69 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565924468;   http://
www.oreilly.com/catalog/devaspcom/.

Powers:2001:DAC

[1561] Shelley Powers. Developing ASP Components. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, March 2001. ISBN 1-56592-750-8. xvii + 811 pp. LCCN TK5105.8885.A26 P69 2001.

URL http://www.oreilly.com/catalog/9781565927506;   http://
www.oreilly.com/catalog/devaspcom2.

Powers:2002:UPT

[1562] Shelley Powers, Jerry Peek, Tim O’Reilly, Mike Loukides, et al. Unix Power Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00330-7. xxxviii + 1116 pp. LCCN QA76.76.O63 P44 2003.

US$69.95. URL http://www.oreilly.com/catalog/upt3.

Powers:2003:PR

[1563] Shelley Powers. Practical RDF. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00263-7. xv + 331 pp. LCCN QA76.76.H94 P49485 2003.

URL http://www.oreilly.com/catalog/9780596002633.

Powers:2003:UPT

[1564] Shelley Powers, Jerry Peek, Tim O’Reilly, Mike Loukides, et al. Unix Power Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00330-7. xxxviii + 1116 pp. LCCN QA76.76.O63 P44 2003.

US$69.95. URL http://www.oreilly.com/catalog/upt3.

Powers:2005:WSF

[1565] Shelley Powers. What are syndication feeds. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-52697-0. 24 pp. LCCN HD30.37 P69 2005eb;

HD30.37. URL http://www.oreilly.com/catalog/9780596526979.

Powers:2007:AA

[1566] Shelley Powers. Adding Ajax. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52936-8 (paperback). xvi + 382 pp. LCCN TK5105.8885.A52eb; TK5105.8885.A52. URL http://www.oreilly.

com/catalog/9780596529369.

Powers:2007:LJ

[1567] Shelley Powers. Learning JavaScript. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52746-2. xiv + 335 pp. LCCN QA76.73.J39 P7 2007eb; QA76.73.J39.

URL http://www.oreilly.com/catalog/9780596527464.

Powers:2008:PW

[1568] Shelley Powers. Painting the Web. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51509-X. xvi + 638 pp. LCCN T385 .P6769 200.

Preston:1999:UBR

[1569] W. Curtis Preston. UNIX Backup & Recovery. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-642-0. xxiii + 707 pp. LCCN QA76.76.O63 P744 1999; TK5105.8885.A26 P69 2001. US$36.95.

URL http://www.oreilly.com/catalog/9781565926424;   http://
www.oreilly.com/catalog/unixbr.

Preston:2002:USN

[1570] W. Curtis Preston. Using SANs and NAS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00153-3. xiv + 205 pp. LCCN TK5105.86 .P74

2002. US$29.95. URL http://safari.oreilly.com/0596001533;
http://www.oreilly.com/catalog/9780596001537;   http://www.
oreilly.com/catalog/sansnas.

Preston:2007:BR

[1571] W. Curtis Preston. Backup and recovery. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10246-1 (paperback). xxviii + 729 pp. LCCN QA76.9.B32 P74 2007. URL http://

www.loc.gov/catdir/enhancements/fy0715/2007273944-d.html;
http://www.oreilly.com/catalog/9780596102463;   http://www.
oreilly.com/catalog/9780596102463/index.html.

Pribyl:2002:LOP

[1572] Bill Pribyl and Steven Feuerstein. Learning Oracle PL/SQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00180-0. xix + 401 pp. LCCN QA76.73.P258 P75 2002. US$39.95. URL

http://safari.oreilly.com/0596001800;   http://www.oreilly.
com/catalog/9780596001803;   http://www.oreilly.com/catalog/
learnoracle.

Price:2001:JPO

[1573] Jason Price. Java Programming with Oracle SQLJ. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2001. ISBN 0-596-00087-1. xiii + 381 pp. LCCN QA76.73.J38 P74 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596000875;   http://
www.oreilly.com/catalog/orasqlj.

Pring:2005:PFE

[1574] Roger Pring. Photoshop filter effects encyclopedia: the hands-on desktop reference for digital photographers. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10021-3. 192 pp. LCCN TR590.5; TR590.5 .P75 2005eb; TR590.5 .P75 2005. URL http://www.

oreilly.com/catalog/9780596100216.

Prinz:2002:CPR

[1575] Peter Prinz and Ulla Kirch-Prinz. C Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00436-2. viii + 134 pp. LCCN QA76.73.C15 P74 2002. US$14.95.

URL http://www.oreilly.com/catalog/9780596004361;   http://
www.oreilly.com/catalog/cpr. Translated by Tony Crawford.

Prinz:2005:CN

[1576] Peter Prinz and Tony Crawford. C in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00697-7. 400 (est.) pp. LCCN QA76.73.C15; QA76.73.C15 P75 2005eb; Internet. URL http://www.oreilly.com/

catalog/9780596006976.

Prosser:2006:FIW

[1577] Susan Prosser. FileMaker 8.5: integrating the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52823-X. LCCN QA76.9.D3; QA76.9.D3

P76 2006eb. URL http://www.oreilly.com/catalog/9780596528232.

Pruett:2006:AWS

[1578] Mark Pruett. Ajax and Web services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52853-1. LCCN TK5105.8885.A52 P78

2006eb; TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596528539.

Pruett:2007:YP

[1579] Mark Pruett. Yahoo! Pipes. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51453-0. LCCN QA76.76.A63 P78 2007eb; QA76.76.A63.

URL http://www.oreilly.com/catalog/9780596514532.

Prusa:2012:GSR

[1580] Josef Prusa. Getting started with the RepRap: 3D printing on your desktop. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1042-7 (paperback). 1 v. pp. LCCN ????

Pugh:2005:PEA

[1581] Ken Pugh. Prefactoring: Extreme Abstraction, Extreme Separation, Extreme Readability. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00874-0. xvi + 220 pp. LCCN QA76.76.D47 P84 2005. US$29.95, CDN$41.95. URL

http://www.oreilly.com/catalog/9780596008741;   http://www.
oreilly.com/catalog/prefactoring/.

Pugh:2006:IOD

[1582] Kenneth Pugh. Interface-oriented design. Pragmatic Bookshelf, Raleigh, NC, USA, 2006. ISBN 0-9766940-5-0 (paperback). xv + 215 pp. LCCN

QA76.76.A63 P84 2006. URL http://www.oreilly.com/catalog/
9780976694052.

Purdy:2000:CPR

[1583] Gregor N. Purdy. CVS Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 0-596-00003-0. iii + 75 pp. LCCN QA76.6 .P87 2000; QA76.76.D47 P863 2000. US$9.95.

URL http://www.oreilly.com/catalog/9780596000035;   http://
www.oreilly.com/catalog/cvspr.

Purdy:2003:CPR

[1584] Gregor N. Purdy. CVS pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00567-9. vi + 84 pp. LCCN QA76.76.D47

P863 2003. URL http://www.oreilly.com/catalog/9780596005672.

Purdy:2004:LIP

[1585] Gregor N. Purdy. Linux iptables: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00569-5. iii + 91 pp. LCCN QA76.76.O63

P873 2004. URL http://www.oreilly.com/catalog/9780596005696.

Qualline:2002:PCP

[1586] Steve Qualline. Practical C++ Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00419-2. 549 (est.) pp. LCCN QA76.73.C153 O833 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596004194;   http://
www.oreilly.com/catalog/cplus2.

Quedenbaum:2015:MVD

[1587] Martin Quedenbaum. MAGIX Video deluxe. O’Reillys basics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 3-95561-686-X. 448 pp. LCCN ????

URL http://d-nb.info/1058570366/04.

Quercia:1989:XWS

[1588] Valerie Quercia and Tim O’Reilly. X Window System user’s guide: for Version 11 of the X Window System, volume 3 of The definitive guides to the X Window System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 1989. ISBN 0-937175-36-6. xxi + 546 pp. LCCN

QA76.76.W56 D44 v.3 1989. URL http://www.oreilly.com/catalog/
9780937175361.

Quercia:1990:XWS

[1589] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide, volume 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1990. ISBN 0-937175-14-5. xxvi + 723 pp. LCCN QA76.76.W56 Q83 1990. URL

http://proquestcombo.safaribooksonline.com/9780937175149;
http://www.oreilly.com/catalog/9780937175149.

Quercia:1991:XWS

[1590] Valerie Quercia and Tom O’Reilly. X Window System User’s Guide: OSF/Motif edition. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-61-7. xxvi + 709 pp. LCCN QA76.76.W56 Q4 1991. URL

http://www.oreilly.com/catalog/9780937175613.

Quercia:1993:XWSa

[1591] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide: Standard Edition, volume 3. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, May 1993. ISBN 1-56592-014-7. xxx + 835 pp. LCCN QA76.76.W56D43 1993. US$34.95. URL http://www.oreilly.com/

catalog/9781565920149.

Quercia:1993:XWSb

[1592] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide: OSF/Motif 1.2 Edition, volume 3M. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, May 1993. ISBN 1-56592-015-5. xxx + 835 pp. LCCN QA76.76 W56Q4 1993.

Quercia:1993:XWSc

[1593] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide: OSF/Motif 1.2 Edition, volume 3M. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, May 1993. ISBN 1-56592-015-5. xxx + 835 pp. LCCN QA76.76 W56Q4 1993.

Quercia:1997:IND

[1594] Valerie Quercia, Linda Mui, Clairemarie Fisher O’Leary, and Ed Krol. Internet in a Nutshell: a Desktop Quick Reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-323-5. xvi + 431 pp. LCCN TK5105.875.I57 Q47

1997. US$24.95. URL http://www.ora.com/catalog/internut/;
http://www.oreilly.com/catalog/9781565923232;   http://www.
oreilly.com/catalog/internut.

Radermacher:2000:NP

[1595] Todd Radermacher and Matthew Gast. Network Printing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 0-596-00038-3. xv + 285 pp. LCCN TK7887.7 .R33 2000. US$34.95.

URL http://www.oreilly.com/catalog/9780596000387;   http://
www.oreilly.com/catalog/netprint.

Radin:1994:BSS

[1596] David Radin. Building a Successful Software Business. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1994. ISBN 1-56592-064-3. xix + 371 pp. LCCN QA76.76.D47 R328 1994. US$19.95. URL http://www.oreilly.com/

catalog/9781565920644.

Raines:1998:TTP

[1597] Paul Raines. Tcl/Tk Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-498-3. iv + 90 pp. LCCN QA76.73.T44 R35 1998.

US$7.95. URL http://www.oreilly.com/catalog/9781565924987;
http://www.oreilly.com/catalog/tclpr/.

Raines:1999:TTN

[1598] Paul Raines and Jeff Tranter. Tcl/TK in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-433-9. xiii + 440 pp. LCCN QA76.73.T44 R33 1999; QA76.73.T44

R35 1999. US$24.95. URL http://www.oreilly.com/catalog/
9781565924338;   http://www.oreilly.com/catalog/tclnut.

Ramalho:2015:FPC

[1599] Luciano Ramalho. Fluent Python: clear, concise, and effective programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-4600-8. xxiv + 743 pp. LCCN AA76.73.P98. URL http://proquest.

safaribooksonline.com/?fpi=9781491946237.

Ramalho:2018:PVG

[1600] Luciano Ramalho. Python vs. Go. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2018. ISBN 1-4920-3317-0, 1-4920-3318-9. ???? pp. LCCN QA76.73.P98.

Rampersad:2007:MYM

[1601] Tara Rampersad. Making your mark in Second Life: business, land, and money. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51417-4. LCCN GV1469.25.S425; GV1469.25.S425 R36 2007eb. URL http://

www.oreilly.com/catalog/9780596514174.

Randal:2003:PE

[1602] Allison Randal, Dan Sugalski, and Leopold Tötsch. Perl 6 essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00499-0. ix + 195 pp. LCCN QA76.73.P22 R36 2003. URL http://www.oreilly.com/

catalog/9780596004996.

Randal:2004:PPE

[1603] Allison Randal, Dan Sugalski, and Leopold Tötsch. Perl 6 and Parrot essentials. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00737-X. xi + 279 pp. LCCN QA76.73.P22 R36 2004. URL http://www.oreilly.

com/catalog/9780596007379.

Rankin:2004:KHI

[1604] Kyle Rankin. Knoppix hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00787-6 (paperback). xx + 314 pp. LCCN

QA76.76.O63 R368 2004. URL http://www.oreilly.com/catalog/
9780596007874.

Rankin:2005:KPR

[1605] Kyle Rankin. Knoppix pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10075-2. vii + 84 pp. LCCN QA76.76.O63 R36 2005.

URL http://www.oreilly.com/catalog/9780596100759.

Rankin:2006:LMH

[1606] Kyle Rankin. Linux multimedia hacks: tips and tools for taming images, audio and video. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10076-0. xviii + 310 pp. LCCN QA76.575; QA76.575 .R345 2006eb.

URL http://www.oreilly.com/catalog/9780596100766.

Rankin:2007:KH

[1607] Kyle Rankin. Knoppix hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51493-X (paperback). xxv + 391 pp. LCCN QA76.76.O63; QA76.76.O63 R36 2007eb. URL http://www.oreilly.

com/catalog/9780596514938.

Rao:2019:NLP

[1608] Delip Rao and Brian McMahan. Natural Language Processing with Py-Torch: Build Intelligent Language Applications Using Deep Learning. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4919-7818-X, 1-4919-7820-1 (e-book), 1-4919-7822-8, 1-4919-7823-6 (paperback). xiii + 238 pp. LCCN QA76.9.N38 R36 2019.

Rash:2007:LFA

[1609] Michael Rash. Linux firewalls: attack detection and response with iptables, psad, and fwsnort. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-141-7. xix + 308 pp. LCCN QA76.9.A25 R36

2007. URL http://www.oreilly.com/catalog/9781593271411.

Rathbone:2006:PMM

[1610] Andy Rathbone. PCs: the missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2006. ISBN 0-596-10093-0. xi + 583 pp. LCCN QA76.5 .R36 2006eb; QA76.5. URL http://www.oreilly.

com/catalog/9780596100933.

Rauschmayer:2014:SJD

[1611] Axel Rauschmayer. Speaking JavaScript: [an in-depth guide for programmers]. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-6503-5. xvii + 437 pp. LCCN QA76.73.J39 R38 2014.

Ravin:1996:UMU

[1612] Ed Ravin, Tim O’Reilly & Dale Dougherty, and Grace Todino. Using & Managing UUCP. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1996. ISBN 1-56592-153-4. xix + 401 pp. LCCN QA76.76.O63 R3913 1996.

US$29.95. URL http://www.ora.com/catalog/umuucp/;   http://
www.oreilly.com/catalog/9781565921535.

Ray:2001:LXC

[1613] Erik T. Ray. Learning XML: Creating Self-Describing Data. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00046-4. xii + 354 pp. LCCN QA76.76.H94 R3945 2001. US$34.95. URL http://www.oreilly.com/

catalog/9780596000462.

Ray:2002:PWS

[1614] Randy J. Ray and Pavel Kulchenko. Programming Web Services with Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00206-8. xiii + 470 pp. LCCN TK5105.888 .R39 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596002060;   http://
www.oreilly.com/catalog/pwebserperl.

Ray:2002:PX

[1615] Erik T. Ray and Jason McIntosh. Perl and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00205-X. xii + 202 pp. LCCN QA76.73.P22

R39 2002. URL http://www.oreilly.com/catalog/9780596002053.

Ray:2003:LX

[1616] Erik T. Ray. Learning XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00420-6. xiv + 400 pp. LCCN QA76.76.H94

R3945 2003. URL http://www.oreilly.com/catalog/9780596004200.

Raymond:1999:CBM

[1617] Eric S. Raymond. The Cathedral & the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-724-9. xi + 268 pp. LCCN QA76.76.O63 R396 1999. US$19.95. URL http://www.oreilly.com/

catalog/9781565927247;   http://www.oreilly.com/catalog/cb.

Raymond:1999:CBP

[1618] Eric S. Raymond. The Cathedral and the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1999. ISBN 0-596-00108-8. 208 pp. LCCN QA76.76.O63 R396 1999. US$16.95.

URL http://www.oreilly.com/catalog/9780596001087;   http://
www.oreilly.com/catalog/cathbazpaper.

Raymond:2001:CBM

[1619] Eric S. Raymond. The cathedral and the bazaar: musings on Linux and open source by an accidental revolutionary. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, revised edition, 2001. ISBN 0-596-00131-2. xiv + 241 pp. LCCN QA76.76.O63 R397 2001. URL http://www.oreilly.com/

catalog/9780596001315.

Raymond:2006:PQR

[1620] Scott Raymond and Sergio Pereira. Prototype quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-51018-7. LCCN QA76.73.J38

R39 2006eb; QA76.73.J38. URL http://www.oreilly.com/catalog/
9780596510183.

Raymond:2007:AR

[1621] Scott Raymond. Ajax on Rails. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52744-6. xii + 336 pp. LCCN TK5105.8885.A52eb; TK5105.8885.A52. URL http://www.oreilly.

com/catalog/9780596527440.

Redmond:2012:SDS

[1622] Eric Redmond and Jim R. Wilson. Seven databases in seven weeks: a guide to modern databases and the NoSQL movement. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2012. ISBN 1-934356-92-1 (paperback). xiii + 333 pp. LCCN QA76.9.D32 .R43 2012. Edited by Jacquelyn Carter.

Reed:2009:LX

[1623] Aaron Reed. Learning XNA 3.0. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15490-9. 512 (est.) pp. LCCN QA76.76.C672 R44 2008.

Reese:1997:DPJ

[1624] George Reese. Database Programming with JDBC and Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-270-0. 224 pp. LCCN QA76.625 .R44 1997. US$29.95. URL

http://www.ora.com/catalog/javadata/;   http://www.oreilly.
com/catalog/9781565922709;   http://www.oreilly.com/catalog/
javadata.

Reese:1998:JJ

[1625] George Reese. JDBC et Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1998. ISBN 2-84177-042-7. LCCN ????

220 FF. URL http://www.editions-oreilly.fr/international/
france/prog/java-jdbc.html.

Reese:2000:DPJ

[1626] George Reese. Database Programming with JDBC and Java. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2000. ISBN 1-56592-616-1. xvii + 328 pp. LCCN QA76.9.D26 R44 2000. US$34.95.

URL http://www.oreilly.com/catalog/9781565926165;   http://
www.oreilly.com/catalog/jdbc2.

Reese:2002:MUM

[1627] George Reese, Randy Jay Yarger, Tim King, and Hugh E. Williams. Managing & Using MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00211-4. 425 pp. LCCN QA76.73.S67 R43

2002. US$39.95. URL http://safari.oreilly.com/0596002114;
http://www.oreilly.com/catalog/9780596002114;   http://www.
oreilly.com/catalog/msql2.

Reese:2003:JDB

[1628] George Reese. Java database best practices. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00522-9. xvi + 267 pp. LCCN QA76.73 .J38

R445 2003. URL http://www.oreilly.com/catalog/9780596005221.

Reese:2003:MPR

[1629] George Reese. MySQL pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00446-X. vi + 87 pp. LCCN QA76.73.S67 R43

2003. URL http://www.oreilly.com/catalog/9780596004460.

Reese:2007:MPR

[1630] George Reese. MySQL pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51426-3 (paperback). vi + 126 pp. LCCN QA76.73.S67 R43 2007. URL http://www.oreilly.com/

catalog/9780596514266.

Reinders:2007:ITB

[1631] James Reinders. Intel Threading Building Blocks: Outfitting C++ for Multi-core Processor Parallelism. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51480-8. xxv + 303 pp. LCCN QA76.58 .R446 2007.

URL http://www.oreilly.com/catalog/9780596514808.

Retford:2007:HBS

[1632] Brian Retford. How to build an SMS service. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51513-8. LCCN QA76.59; QA76.59 .R48

2007eb. URL http://www.oreilly.com/catalog/9780596515133.

RevealNet:2001:ODC

[1633] RevealNet. Oracle DBA checklists: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00122-3. iii + 76 pp. LCCN QA76.9.D3 O725 2001. US$9.95.

URL http://www.oreilly.com/catalog/9780596001223;   http://
www.oreilly.com/catalog/ordbacheck.

Rhodes:1999:PPD

[1634] Neil Rhodes and Julie McKeehan. Palm Programming: The Developer’s Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-525-4. xxi + 457 pp. LCCN QA76.5 .R56 1999; QA76.8.P138 R556 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565925250;   http://
www.oreilly.com/catalog/palmprog. Includes CD-ROM.

Rhodes:2002:PPD

[1635] Neil Rhodes and Julie McKeehan. Palm OS programming: the developer’s guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 1-56592-856-3. xvii + 681 pp. LCCN QA76.8.P138 R555

2002. US$39.95. URL http://safari.oreilly.com/1565928563;

http://www.oreilly.com/catalog/9781565928565;   http://www.
oreilly.com/catalog/palmprog2.

Rice:2017:HCY

[1636] Liz Rice. How to Containerize Your Go Code. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-8230-6. ???? pp. LCCN ????

Richards:2005:JDN

[1637] Norman Richards and Sam Griffith. JBoss: a developer’s notebook. The developer’s notebook series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10007-8. xix + 149 pp. LCCN TK5105.8885.J43 R53 2005. URL http://www.

oreilly.com/catalog/9780596100070.

Richards:2009:JMS

[1638] Mark Richards, Richard Monson-Haefel, and David Chappell. Java Message Service. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-52204-5. 328 (est.) pp. LCCN ???? US$39.99.

Richardson:2005:SIP

[1639] Jared R. Richardson and William A. Gwaltney. Ship it!: a practical guide to successful software projects. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-4-7. xiii + 198 pp. LCCN QA76.76.D47 R53 2005; Qa76.76.D47 R53 2005. URL http://www.oreilly.com/

catalog/9780974514048.

Richardson:2007:RWS

[1640] Leonard Richardson and Sam Ruby. RESTful Web services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52926-0 (paperback). xxiv + 419 pp. LCCN TK5105.88813; TK5105.88813eb. URL http://www.

oreilly.com/catalog/9780596529260.

Richardson:2012:BBC

[1641] Matt Richardson. Beginning BeagleBone: Creating Linux-Powered Electronics Projects. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-4537-9. 75 pp. LCCN ???? EUR 8.50.

Richardson:2014:BEL

[1642] Matt Richardson. BeagleBone für Einsteiger: [Linux-basierte Elektronik-Projekte mit Python und JavaScript]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 3-95561-409-3. xii + 134 pp. LCCN ????

Richardson:2015:RPE

[1643] Matt Richardson and Shawn Wallace. Raspberry Pi für Einsteiger. (German) [Raspberry Pi for beginners]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2015. ISBN 3-95561-992-3. 208 pp. LCCN ????

Riordan:2008:HFA

[1644] Rebecca Riordan. Head first Ajax. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51578-2. xxx + 485 pp. LCCN TK5105.8885.A52 R56 2008.

Ristic:2005:AS

[1645] Ivan Ristic. Apache security. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00724-8. xxii + 396 pp. LCCN TK5105.8885.A63 R57 2005. URL http://www.

oreilly.com/catalog/9780596007249.

Rizzo:2005:MAH

[1646] John Rizzo. Mac annoyances: how to fix the most annoying things about your Mac. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00723-X. xiv + 156 pp. LCCN QA76.8.M3; QA76.8.M3 R588 2004eb; QA76.8.M3 R588 2004. URL http://www.oreilly.com/

catalog/9780596007232.

Robbins:1999:UNS

[1647] Arnold Robbins. UNIX in a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http://www.oreilly.com/

catalog/unixnut3.

Robbins:1999:VEK

[1648] Arnold Robbins. Vi-Editor: kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 3-89721-213-7 (paperback). 62 pp. LCCN ???? DM 12.80, S 93.00.

Robbins:1999:VEP

[1649] Arnold Robbins. vi editor: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-497-5. v + 66 pp. LCCN QA76.76.T49 R617 1999; QA76.76.T49 R63 1999. US$6.95.

URL http://www.oreilly.com/catalog/9781565924970;   http://
www.oreilly.com/catalog/vipr/.

Robbins:2000:HUC

[1650] Arnold Robbins. HP-UX: a companion to Unix in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-760-5. ???? pp. LCCN

QA76.76.O63 R565 2000. URL http://www.oreilly.com/catalog/
9781565927605.

Robbins:2000:SAP

[1651] Arnold Robbins. sed and awk Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-729-X. iii + 45 pp. LCCN QA76.76.U84 R63 2000. US$9.95.

URL http://www.oreilly.com/catalog/9781565927292;   http://
www.oreilly.com/catalog/sedawkrepr.

Robbins:2001:EAP

[1652] Arnold Robbins. Effective AWK Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2001. ISBN 0-596-00070-7 (paperback), 0-596-80537-3 (e-book). xxiv + 421 pp. LCCN QA76.73.A95 R63 2001.

US$39.95. URL http://www.oreilly.com/catalog/awkprog3/.

Robbins:2002:SAK

[1653] Arnold Robbins. Sed und awk. Kurz und gut. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 3-89721-246-3. 62 pp. LCCN ????

Robbins:2002:SAP

[1654] Arnold Robbins. sed & awk Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00352-8. vi + 46 pp. LCCN QA76.76.U84 R62 2002. US$12.95. URL http://safari.oreilly.com/

0596003528;   http://www.oreilly.com/catalog/9780596003524;
 http://www.oreilly.com/catalog/sedawkrepr2.

Robbins:2003:UCB

[1655] Arnold Robbins, Bill Rosenblatt, Dale Dougherty, Jerry Peek, Linda Lamb, and Brian Jepson, editors. The UNIX CD bookshelf: 7 bestselling books on CD-ROM: Unix power tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00392-7. ???? pp. LCCN ????

Robbins:2005:CSS

[1656] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63

R633 2005. URL http://www.oreilly.com/catalog/shellsrptg/.

Also available in Chinese [1664], French [1658], German [1661], Japanese [?], and Polish [?] translations.

Robbins:2005:GPG

[1657] Arnold Robbins. GDB Pocket Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10027-2. v + 69 pp. LCCN QA76.9.D43 R63 2005. US$9.95, CAN$13.95.

URL http://www.oreilly.com/catalog/9780596100278.

Robbins:2005:ISS

[1658] Arnold Robbins and Nelson H. F. Beebe. Introduction aux Scripts Shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 2-84177-375-2. xxii +

558 pp. LCCN ???? URL http://www.silicon.fr/getarticle.asp?

id=14015. French translation of [1656] by Eric Jacoboni.

Robbins:2005:UN

[1659] Arnold Robbins. Unix in a Nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-10029-9. xviii + 885 pp. LCCN QA76.76.O63 R63 2005. US$34.95, CAN$48.95, UK£24.95. URL http://www.oreilly.com/

catalog/unixnut4/index.html.

Robbins:2006:BQR

[1660] Arnold Robbins. Bash quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52776-4. LCCN QA76.76.O63; QA76.76.O63 R63

2006eb. URL http://www.oreilly.com/catalog/9780596527761.

Robbins:2006:KSP

[1661] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung: [automatisieren Sie Ihre Unix/Linux-Tasks]. (German) [Classic Shell Scripting: [automating your Unix/Linux jobs]]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 3-89721-441-5. xxiii + 572 pp. LCCN QA76.76.O63 R563 2005. EUR 44.00; EUR 45.65 (AT). URL http://www.gbv.

de/dms/hebis-darmstadt/toc/17645067X.pdf. German translation of

[1656] by Kathrin Lichtenberg.

Robbins:2006:UN

[1662] Arnold Robbins. Unix in a Nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2006. ISBN 0-596-10029-9. xviii + 885 pp. LCCN QA76.76.O63 R566 2006. US$34.95, CAN$48.95, UK£24.95. URL

http://www.oreilly.com/catalog/9780596100292;   http://www.
oreilly.com/catalog/unixnut4/index.html.

Robbins:2008:LVV

[1663] Arnold Robbins, Elbert Hannah, and Linda Lamb. Learning the vi and Vim Editors. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, seventh edition, 2008. ISBN 0-596-52983-X. 492 (est.) pp. LCCN QA76.76.O63 L355 2008.

URL http://www.oreilly.com/catalog/9780596529833.

Robbins:2008:SJB

[1664] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi NaN = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 7-111-25504-6. vi + 494 pp. LCCN ???? 79.00 renminbi. Simplified Chinese translation of [1656].

Robbins:2009:UD

[1665] Jeff Robbins et al., editors. Using Drupal. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15793-2. 496 pp. LCCN TK5105.8885.D78 U85 2008.

Robichaux:1998:MWN

[1666] Paul E. Robichaux. Managing the Windows NT Registry. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-56592-378-2. xii + 360 pp. LCCN QA76.76.O63 R6225 1998; QA76.76.O63 R625 1998. US$34.95, CDN$49.95. URL

http://www.oreilly.com/catalog/9781565923782;   http://www.
oreilly.com/catalog/manwinreg;   http://www.oreilly.com/catalog/
manwinreg/noframes.html.

Robichaux:1999:MME

[1667] Paul E. Robichaux. Managing Microsoft Exchange Server. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-545-9. xiv + 701 pp. LCCN QA76.9.C55 R627 1999; QA76.9.C55 R63 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565925458;   http://
www.oreilly.com/catalog/managexsvr.

Robichaux:2000:MWR

[1668] Paul E. Robichaux. Managing the Windows 2000 Registry. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-943-

8. xiii + 541 pp. LCCN QA76.76.O63 R6226 2000. US$39.95.

URL http://www.oreilly.com/catalog/9781565929432;   http://
www.oreilly.com/catalog/mwin2reg.

Robichaux:2005:ESC

[1669] Paul E. Robichaux, Missy Koslosky, and Devin L. Ganger. Exchange Server cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00717-5 (paperback). xix + 437 pp. LCCN QA76.9.C55 R62 2005. URL http://www.oreilly.

com/catalog/9780596007171.

Roelofs:1999:PDG

[1670] Greg Roelofs. PNG: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-542-4. xix + 327 + 4 pp. LCCN T385 .R588 1999; T385 .R58 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565925427;   http://
www.oreilly.com/catalog/pngdefg.

Roerden:1997:NLW

[1671] Laura Parker Roerden. Net Lessons: Web-Based Projects for Your Classroom. Songline Studios, Inc. and O’Reilly & Associates, Inc., Sebastopol, CA, 1997. ISBN 1-56592-291-3. xxi + 284 pp. LCCN LB1044.87 .R64

1997. US$24.95. URL http://www.ora.com/catalog/netlessons/;
http://www.oreilly.com/catalog/9781565922914;   http://www.
oreilly.com/catalog/netlessons. Foreword by Bonnie Bracey.

Roff:2001:AAD

[1672] Jason T. Roff. ADO: ActiveX Data Objects. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 1-56592-415-0. xiii + 601 pp. LCCN QA76.625 .R64 2001. US$44.95. URL http://www.oreilly.com/

catalog/9781565924154;   http://www.oreilly.com/catalog/ado.

Rolland:1999:LPP

[1673] Christian Rolland. LaTEX par la pratique. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 2-84177-073-7. xx + 559 pp. LCCN ????

Rolsky:2002:EPH

[1674] Dave Rolsky and Ken Williams. Embedding Perl in HTML with Mason. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00225-4. xviii + 297 pp. LCCN QA76.73.P22 R65 2002. US$34.95. URL http://www.

oreilly.com/catalog/perlhtmlmason.

Rolsky:2003:EPH

[1675] Dave Rolsky and Ken Williams. Embedding Perl in HTML with Mason. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00225-4. xviii + 297 pp. LCCN QA76.73.P22 R65 2003. US$34.95.

URL http://www.oreilly.com/catalog/9780596002251;   http://
www.oreilly.com/catalog/perlhtmlmason.

Roman:1997:ADD

[1676] Steven Roman. Access database design and programming. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-297-2. xvi + 251 pp. LCCN QA76.9.D3 R657

1997. US$24.95. URL http://www.ora.com/catalog/accessdata/;
http://www.oreilly.com/catalog/9781565922976;   http://www.
oreilly.com/catalog/accessdata.

Roman:1998:DVB

[1677] Steven Roman. Developing Visual Basic Add-Ins: The VB IDE Extensibility Model. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-527-0. xii + 171 pp. LCCN QA76.73.B3 R65 1999. US$27.95.

URL http://www.oreilly.com/catalog/9781565925274;   http://
www.oreilly.com/catalog/devvbad.

Roman:1998:LWP

[1678] Steven Roman. Learning Word Programming: Creating Word Macros and Beyond. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-524-6. xviii + 388 pp. LCCN QA76.73.B3 R655 1998; QA76.73.B3 R66 1998.

US$27.95. URL http://www.oreilly.com/catalog/9781565925243;
 http://www.oreilly.com/catalog/lrnwdpr.

Roman:1999:ADD

[1679] Steve Roman. Access Database Design & Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 1999. ISBN 1-56592-626-9. xx + 409 pp. LCCN QA76.9.D3 R65 1999. US$27.95.

URL http://www.oreilly.com/catalog/9781565926264;   http://
www.oreilly.com/catalog/accessdata2.

Roman:1999:WEM

[1680] Steven Roman. Writing Excel Macros: Automating Excel to Work for You. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-587-4. xix + 529 pp. LCCN QA76.6 .R647 1999; QA76.6 .R65 1999. US$27.95.

URL http://www.oreilly.com/catalog/9781565925878;   http://
www.oreilly.com/catalog/exlmacro.

Roman:1999:WWM

[1681] Steven Roman. Writing Word Macros: An Introduction to Programming Word using VBA. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, October 1999. ISBN 1-56592-725-7. xix + 388 pp. LCCN QA76.6 .R68 1999.

US$29.95. URL http://www.oreilly.com/catalog/9781565927254;
 http://www.oreilly.com/catalog/writewordmacro.

Roman:2000:WAP

[1682] Steven Roman. Win32 API programming with Visual Basic: Taking advantage of Windows system services. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-631-5. xix + 511 pp. LCCN QA76.76.O63 R66 2000; QA76.66 .R66 2000. US$39.95.

URL http://www.oreilly.com/catalog/9781565926318;   http://
www.oreilly.com/catalog/win32api.

Roman:2001:VNL

[1683] Steven Roman, Ronald Petrusha, and Paul Lomax. VB.NET language in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00092-8. xvi + 636 pp. LCCN QA76.73.B3 R657 2001.

URL http://www.oreilly.com/catalog/9780596000929.

Roman:2002:ADD

[1684] Steven Roman. Access Database Design & Programming. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00273-4. xix + 425 pp. LCCN QA76.9.D3 R657

2002. US$34.95. URL http://safari.oreilly.com/0596002734;
http://www.oreilly.com/catalog/9780596002732;   http://www.
oreilly.com/catalog/accessdata3.

Roman:2002:VNLa

[1685] Steven Roman, Ronald Petrusha, and Paul Lomax. VB.NET Language in a Nutshell: a Desktop Quick Reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00308-0. xvii + 662 pp. LCCN QA76.73.B3 R657

2002. US$44.95. URL http://safari.oreilly.com/0596003080;
http://www.oreilly.com/catalog/9780596003081;   http://www.
oreilly.com/catalog/vbdotnetnut2.

Roman:2002:WEM

[1686] Steven Roman. Writing Excel Macros with VBA. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00359-5. xiii + 549 pp. LCCN QA76.6 .R65 2002. US$34.95. URL http://safari.oreilly.com/

0596003595;   http://www.oreilly.com/catalog/9780596003593;
 http://www.oreilly.com/catalog/exlmacro2.

Roman:2003:VNL

[1687] Steven Roman, Ronald Petrusha, and Paul Lomax. VB.NET language pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00428-1. vii + 142 pp. LCCN QA76.73.B3 R6573 2003. US$14.95.

URL http://www.oreilly.com/catalog/9780596004286;   http://
www.oreilly.com/catalog/vbnetlpr.

Rose:2002:BDG

[1688] Marshall T. Rose. BEEP: The Definitive Guide: Developing New Applications for the Internet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00244-0. xv + 204 pp. LCCN TK5105.55 .R66 2002. US$34.95.

URL http://www.oreilly.com/catalog/9780596002442;   http://
www.oreilly.com/catalog/beep.

Rosenberry:1992:UD

[1689] Ward Rosenberry, David Kenney, and Gerry Fisher. Understanding DCE. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1992. ISBN 1-56592-005-8. xxv + 233 pp. LCCN QA76.9.D5 R67 1992. US$24.95.

URL http://www.oreilly.com/catalog/9781565920057;   http://
www.oreilly.com/catalog/udce.

Rosenberry:1993:DAA

[1690] Ward Rosenberry and Jim Teague. Distributing Applications Across DCE and Windows NT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1993. ISBN 1-56592-047-3. xxvi + 274 pp. LCCN QA76.9.D5 R65 1993; QA76.9.D5 R66 1993. US$24.95. URL http://www.oreilly.com/

catalog/9781565920477.

Rosenblatt:1993:LKS

[1691] Bill Rosenblatt. Learning the Korn Shell. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1993. ISBN 1-56592-054-6. xxii + 338 pp. LCCN QA76.73.K67 R68 1993. US$29.95.

URL http://www.oreilly.com/catalog/9781565920545;   http://
www.oreilly.com/catalog/korn.

Rosenblatt:1994:LKS

[1692] Bill Rosenblatt. Learning the Korn Shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1994. ISBN 1-56592-054-6. xxii + 336 pp. LCCN QA76.73.K67

R68 1994. US$29.95. URL http://www.oreilly.com/catalog/korn.

Rosenblatt:2002:LKS

[1693] Bill Rosenblatt and Arnold Robbins. Learning the Korn Shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00195-9 (paperback), 1-4493-7127-2 (e-book), 1-4493-7128-0 (e-book). xviii + 412 pp. LCCN QA76.73.K67 R68 2002. US$34.95.

Rosenfeld:1998:IAW

[1694] Louis Rosenfeld and Peter Morville. Information Architecture for the World Wide Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-282-4. xix + 202 pp. LCCN TK5105.888 .R67 1998; TK5105.888 .R673

1998. US$24.95. URL http://www.ora.com/catalog/infotecture/
;   http://www.oreilly.com/catalog/9781565922822;   http://
www.oreilly.com/catalog/infotecture.

Rosenfeld:2002:IAW

[1695] Louis Rosenfeld and Peter Morville. Information Architecture for the World Wide Web: Designing Large-Scale Web Sites. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00035-9. xxi + 461 pp. LCCN TK5105.888 .R67

2002. US$39.95. URL http://safari.oreilly.com/0596000359;
http://www.oreilly.com/catalog/9780596000356;   http://www.
oreilly.com/catalog/infotecture2.

Rosenthal:1996:PIY

[1696] Donna Rosenthal. Passport Israel: Your Pocket Guide to Israeli Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-885073-22-4. 96 pp. LCCN HD58.7.R67 1996. US$6.95.

URL http://www.oreilly.com/catalog/9781885073228.

Rosenthol:2012:CCR

[1697] Leonard Rosenthol. Creating and Consuming Rich PDFs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-2791-5. 200 pp. LCCN ???? EUR 24.50.

Ross:2003:BWF

[1698] John Ross. The book of Wi-Fi: install, configure, and use 802.11b wireless networking. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-45-X (paperback). xvii + 260 pp. LCCN TK5105.78 .R67 2003; TK5105.78 .R67 2003eb. URL http://www.oreilly.com/

catalog/9781886411456.

Ross:2006:NDB

[1699] John Ross. It’s never done that before: a guide to troubleshooting Windows XP. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-139-5. xix + 263 pp. LCCN QA76.76.O63 R682 2006eb. URL

http://www.oreilly.com/catalog/9781593271398.

Rothman:2005:BCD

[1700] Johanna Rothman and Esther Derby. Behind closed doors: secrets of great management. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-2-6. xii + 167 pp. LCCN ????

Rothman:2008:MXU

[1701] Ernest E. (Ernest Eric) Rothman, Brian Jepson, and Rich Rosen. Mac OS X for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2008. ISBN 0-596-52062-X. [380] pp. LCCN QA76.76.O63 R68 2008.

Rubini:1998:LDD

[1702] Alessandro Rubini. Linux Device Drivers. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-292-1. xviii + 421 pp. LCCN QA76.76.D49 R92 1998; QA87.76.O63 R82

1998. US$29.95. URL http://www.ora.com/catalog/linuxdrive/;
http://www.oreilly.com/catalog/9781565922921;   http://www.
oreilly.com/catalog/linuxdrive.

Rubini:2001:LDD

[1703] Alessandro Rubini and Jonathan Corbet. Linux Device Drivers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 2001. ISBN 0-596-00008-1. xix + 564 pp. LCCN QA76.76.D49 R92

2001. US$39.95. URL http://safari.oreilly.com/0596000081;
http://www.oreilly.com/catalog/9780596000080;   http://www.
oreilly.com/catalog/linuxdrive2.

Russell:1991:CSB

[1704] Deborah Russell and G. T. Gangemi, Sr. Computer Security Basics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1991. ISBN 0-937175-71-4. 464 pp. LCCN QA76.9.A25 R8 1991. US$29.95.

URL http://www.oreilly.com/catalog/9780937175712;   http://
www.oreilly.com/catalog/csb.

Russell:2007:GRD

[1705] Matthew Russell. Get up and running with Dojo. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51705-X. LCCN QA76.64 .R87 2007eb; QA76.64. URL http://www.oreilly.

com/catalog/9780596517052.

Russell:2011:MSW

[1706] Matthew A. Russell. Mining the Social Web: Finding Needles in the Social Haystack. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8834-5. xx + 332 pp. LCCN ???? URL http://paperc.de/16024-

mining-the-social-web-9781449304164.

Russell:2014:MSW

[1707] Matthew A. Russell. Mining the social web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2014. ISBN 1-4493-6761-5 (paperback), 1-4493-7045-4 (e-book), 1-4493-6821-2 (e-book), 1-4493-6822-0 (e-book). xxiv + 421 pp. LCCN QA76.9.D343 R87 2013.

Sadun:2004:MMX

[1708] Erica Sadun. Modding Mac OS X: extreme makeovers for your Mac. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00709-4. xxi + 276 pp. LCCN QA76.76.O63; QA76.76.O63 S33 2004eb; QA76.76.O63

S33 2004. URL http://www.oreilly.com/catalog/9780596007096.

Sadun:2007:EGW

[1709] Erica Sadun. Eight great ways to get the most from your Zune. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52991-0. LCCN ML74.4.Z86;

ML74.4.Z86 S26 2007eb. URL http://www.oreilly.com/catalog/
9780596529918.

Saffer:2009:DGI

[1710] Dan Saffer. Designing gestural interfaces. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15675-8, 0-596-51839-0. 272 (est.) pp. LCCN QA76.9 .U83 2008.

Sample:2006:PHT

[1711] C. K. (Clinton Kennedy) Sample. PSP hacks: tips and tools for your mobile gaming and entertainment handheld. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10143-0. xix + 298 pp. LCCN QA76.76.C672eb; QA76.76.C672. URL http://www.oreilly.

com/catalog/9780596101435.

Sanders:2006:SMT

[1712] Chris Sanders. Saving money and time with Virtual Server. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52801-9. LCCN QA76.76.O63;

QA76.76.O63 S26 2006eb. URL http://www.oreilly.com/catalog/
9780596528010.

Sanders:2007:ADP

[1713] William B. Sanders and Chandima Cumaranatunge. ActionScript 3.0 design patterns: object-oriented programming techniques. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52846-9 (paperback). xx + 509 pp. LCCN QA76.73.A137 S26 2007; QA76.73.A137; QA76.73.A137 S26

2007eb. URL http://www.oreilly.com/catalog/9780596528461.

Sanders:2007:APO

[1714] Bill Sanders. ActionScript 3.0 programming: overview, getting started, and examples of new concepts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52923-6. LCCN TR897.7; TR897.7 .S36 2007eb. URL

http://www.oreilly.com/catalog/9780596529239.

Sanders:2007:PPA

[1715] Chris Sanders. Practical packet analysis: using Wireshark to solve realworld network problems. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-149-2. xx + 164 pp. LCCN TK5105.55 .S265 2007.

URL http://www.loc.gov/catdir/toc/ecip0714/2007013453.html;
 http://www.oreilly.com/catalog/9781593271497.

Sanders:2008:LFM

[1716] William B. Sanders. Learning Flash Media Server 3. Adobe developer library. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51590-1. 260 (est.) pp. LCCN TK5105.386 .S36 2008.

Sanderson:1993:S

[1717] David W. Sanderson and Dale Dougherty. Smileys. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1993. ISBN 1-56592-041-4. 93 pp. LCCN QA76.9.K48 S26 1993. US$5.95. URL

http://www.oreilly.com/catalog/9781565920415;   http://www.
oreilly.com/catalog/smileys.

Sanderson:2013:PGA

[1718] Dan Sanderson. Programming Google App Engine. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2013. ISBN 1-4493-9826-X (paperback). xxiii + 509 pp. LCCN ???? UK£38.50.

Sanderson:2015:PGA

[1719] Dan Sanderson. Programming Google App Engine with Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0025-3, 1-4919-0024-5, 1-4919-0368-6 (e-book), 1-4919-0367-8 (e-book). xxi + 439 pp. LCCN TK5105.88813.

Satir:1995:CCL

[1720] Gregory Satir and Doug Brown. C++: The Core Language. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1995. ISBN 1-56592-116-X. xix + 207 pp. LCCN QA76.73.C153 S278 1995. US$24.95.

URL http://www.ora.com/gnn/bus/ora/item/core.html;   http:/
/www.oreilly.com/catalog/9781565921160;   http://www.oreilly.
com/catalog/core.

Sato:2002:JSL

[1721] Jin Sato and Arnie Rusoff. Jin Sato’s Lego Mindstorms: the master’s technique. No Starch Press, San Francisco, CA, USA, 2002. ISBN 1-886411-56-5 (paperback). xix + 361 pp. LCCN TJ211.15 .S2813 2002.

URL http://www.oreilly.com/catalog/9781886411562.

Sauer:2014:WBB

[1722] Moritz “mo.” Sauer. Das WordPress-Buch: [behandelt WordPress 4]. (German) [The WordPress book: [including WordPress 4]]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2014. ISBN 3-95561-

860-9. 288 pp. LCCN ???? URL http://bvbr.bib-bvb.de:8991/
F?func=service\%26doc\_library=BVB01\%26local\_base=BVB01\%
26doc\_number=027654943\%26line\_number=0001\%26func\_code=
DB\_RECORDS\%26service\_type=MEDIA;   http://swbplus.bsz-bw.
de/bsz420324720cov.htm.

Savikas:2005:WH

[1723] Andrew Savikas. Word hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00493-1 (paperback). xxii + 372 pp. LCCN Z52.5.M52 S28 2004. URL http://

www.oreilly.com/catalog/9780596004934.

Sayers:1996:WMD

[1724] Charles Sayers. Web Makeovers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1996. ISBN 1-56592-196-8. 180 pp. LCCN ???? US$21.95. URL

http://www.oreilly.com/catalog/9781565921962.

Scheifler:1989:XPM

[1725] Robert W. Scheifler. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1989. ISBN 0-937175-40-4. xv + 398

pp. LCCN ???? US$30.00. URL http://www.oreilly.com/catalog/
9780937175408.

Scheifler:1989:XPR

[1726] Robert W. Scheifler. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1989. ISBN 0-937175-40-4. xv + 398 pp. LCCN QA76.76.W56 X5 1989. US$30.00.

Schmalz:2005:IEA

[1727] Michael Schmalz. Integrating Excel and Access. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00973-9. 504 pp. LCCN QA76.9.D3. URL

http://www.oreilly.com/catalog/9780596009731.

Schmitt:1997:SSD

[1728] Bob Schmitt. Shockwave studio: designing multimedia for the Web. Web review studio series. Songline Studios, Inc. and O’Reilly & Associates, Inc., Sebastopol, CA, March 1997. ISBN 1-56592-231-X. xii + 183 pp. LCCN QA76.76.I59 S36 1997. URL http://www.ora.com/catalog/

shock/;   http://www.oreilly.com/catalog/9781565922310;   http:
//www.oreilly.com/catalog/shock.

Schmitt:2004:CC

[1729] Christopher Schmitt. CSS Cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00576-8. xvi + 252 pp. LCCN TK5105.888 .S3524 2004. URL http://

www.oreilly.com/catalog/9780596005764.

Schmitt:2006:CC

[1730] Christopher Schmitt. CSS cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52741-1. xx + 516 pp. LCCN Internet.

URL http://www.oreilly.com/catalog/9780596527419.

Schmitt:2006:RC

[1731] Christopher Schmitt. Releasing CSS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52935-X. LCCN TK5105.888 S36 2006eb; TK5105.888.

URL http://www.oreilly.com/catalog/9780596529352.

Schmitt:2010:CC

[1732] Christopher Schmitt. CSS cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2010. ISBN 0-596-15593-X (paperback). xxiv + 702 pp. LCCN TK5105.888 .S3524 2010. Foreword by Dan Cederholm.

Schrenk:2007:WSS

[1733] Michael Schrenk. Webbots, spiders, and screen scrapers: a guide to developing Internet agents with PHP/CURL. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-120-4. xviii + 306 pp.

LCCN TK5105.884 .S37 2007. URL http://www.loc.gov/catdir/toc/
ecip0619/2006026680.html;   http://www.oreilly.com/catalog/
9781593271206.

Schroder:2005:LC

[1734] Carla Schroder. Linux cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00640-3 (paperback). xxiii + 553 pp. LCCN QA76.76.O63; QA76.76.O63 S377 2004eb; QA76.76.O63 S377 2004. URL http://www.

oreilly.com/catalog/9780596006402.

Schroder:2007:LNC

[1735] Carla Schroder. Linux networking cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10248-8 (paperback). xxiii + 612 pp.

LCCN TK5105.5 .S384 2007. URL http://www.oreilly.com/catalog/
9780596102487.

Schroder:2008:LNC

[1736] Carla Schroder. Linux networking cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-10248-8. xxiii + 612 pp. LCCN QA76.76.O63 S377 2008.

Schulman:1998:WND

[1737] Andrew Schulman, Troy Mott, and Tim O’Reilly. Windows 95 in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 1998. ISBN 1-56592-316-2. xxii + 503 pp. LCCN QA76.76.O63 O734 1998; QA76.76.W56 O74 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565923164;   http://
www.oreilly.com/catalog/win95nut.

Schutt:2013:DDS

[1738] Rachel Schutt and Cathy O’Neil. Doing Data Science. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2013. ISBN 1-4493-5865-9 (paperback). xxiv + 375 pp. LCCN QA76.9.D343 S27 2013.

Schwartz:1993:LP

[1739] Randal L. Schwartz. Learning Perl. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, November 1993. ISBN 1-56592-042-2. xxv + 246 pp. LCCN QA76.73.P22 S37 1993; QA76.73.P225 S39 1993. US$24.95.

URL http://www.oreilly.com/catalog/9781565920422. Foreword by

Larry Wall.

Schwartz:1996:NHC

[1740] Edward A. Schwartz. NetActivism: How Citizens Use the Internet. Songline guides. Songline Studios, Inc., Sebastopol, CA, Fall 1996. ISBN 1-56592-160-7. xi + 212 pp. LCCN JK1764 .S38 1996.

US$24.95. URL http://www.oreilly.com/catalog/9781565921603.

Includes CD-ROM. Same ISBN as [167].

Schwartz:1997:LP

[1741] Randal L. Schwartz and Tom Christiansen. Learning Perl. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-284-0. xxix + 269 pp. LCCN QA76.73.P22 S37 1997; QA76.73.P22 S39 1997. US$32.95. URL http://www.ora.com/catalog/

lperl2/;   http://www.oreilly.com/catalog/9781565922846;   http:
//www.oreilly.com/catalog/lperl2. Foreword by Larry Wall.

Schwartz:1997:LPW

[1742] Randal L. Schwartz, Erik W. Olson, and Tom Christiansen. Learning Perl on Win32 Systems. The Perl series; A Nutshell handbook.

O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-324-3. xxii + 282 pp. LCCN QA76.73.P22 S38 1997. US$29.95. URL

http://www.ora.com/catalog/lperlwin/;   http://www.oreilly.
com/catalog/9781565923249;   http://www.oreilly.com/catalog/
lperlwin.

Schwartz:1998:MML

[1743] Alan Schwartz. Managing Mailing Lists: Majordomo, LISTSERV, Listproc, and SmartList. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1998. ISBN 1-56592-259-X. xiv + 282 pp. LCCN HE7551 .S39 1998; ZA4480 .S39 1998. US$29.95.

URL http://www.oreilly.com/catalog/9781565922594;   http:/
/www.oreilly.com/catalog/mailing;   http://www.oreilly.com/
catalog/mailing/noframes.html.

Schwartz:1998:SSS

[1744] Alan Schwartz and Simson Garfinkel. Stopping Spam: Stamping Out Unwanted Email and News Postings. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-388-X. x + 191 pp. LCCN HE7551 .S393 1998; TK5105.73 .S38 1998. US$19.95.

URL http://www.oreilly.com/catalog/9781565923881;   http://
www.oreilly.com/catalog/spam.

Schwartz:2001:LPE

[1745] Randal L. Schwartz and Tom Phoenix. Learning Perl: Making Easy Things Easy and Hard Things Possible. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, July 2001. ISBN 0-596-00132-0. xvi + 316 pp. LCCN QA76.73.P22 S37 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596001322;   http://
www.oreilly.com/catalog/lperl3.

Schwartz:2003:LPO

[1746] Randal L. Schwartz, Tom Phoenix, and Linda Mui. Learning Perl objects, references, and modules. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00478-8. xv + 205 pp. LCCN QA76.73.P22 S379 2003.

Schwartz:2004:S

[1747] Alan Schwartz. SpamAssassin. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00707-8. xi + 207 pp. LCCN TK5105.73 .S388 2004; TK5105.73 .S388 2004eb; TK5105.73. URL http://www.oreilly.com/

catalog/9780596007072.

Schwartz:2005:LP

[1748] Randal L. Schwartz, Tom Phoenix, and Brian D. Foy. Learning Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2005. ISBN 0-596-10105-8. xvii + 283 pp. LCCN QA76.73.P33 S34 2005. URL http://

www.oreilly.com/catalog/9780596101053.

Schwartz:2006:IP

[1749] Randal L. Schwartz, Brian D. Foy, and Tom Phoenix. Intermediate Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10206-2. xix + 256 pp. LCCN QA76.73.P22 S368 2006; QA76.73.P22; QA76.73.P22 S348

2006eb. URL http://www.oreilly.com/catalog/9780596102067.

Schwartz:2008:HPM

[1750] Baron Schwartz, Peter Zaitsev, Vadim Tkachenko, Jeremy D. Zawodny, Arjen Lentz, and Derek J. Balling. High performance MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 0-596-10171-6. xxii + 684 pp. LCCN ????

Schwartz:2008:LP

[1751] Randal L. Schwartz, Tom Phoenix, and Brian D. Foy. Learning Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2008. ISBN 0-596-52010-7. xviii + 328 pp. LCCN QA76.73.P22 S37 2008.

Scott:1998:VPN

[1752] Charlie Scott, Paul Wolfe, and Mike Erwin. Virtual Private Networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1998. ISBN 1-56592-319-7. xi + 177 pp. LCCN TK5105.875.I57 S36 1998; TK5105.875.I57 S395 1998.

US$29.95 (Can. US$42.95). URL http://www.oreilly.com/catalog/
9781565923195;   http://www.oreilly.com/catalog/vpn; http:/
/www.oreilly.com/catalog/vpn/.

Scott:1999:VPN

[1753] Charlie Scott, Paul Wolfe, and Mike Erwin. Virtual Private Networks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-529-7. xiv + 211 pp. LCCN TK5105.875.I57 S395 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565925298;   http://
www.oreilly.com/catalog/vpn2.

Sedayao:2001:CIA

[1754] Jeff Sedayao. Cisco IOS Access Lists. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 1-56592-385-5. x + 260 pp. LCCN TK5105.59

.S444 2001. US$39.95. URL http://www.oreilly.com/catalog/
9781565923850;   http://www.oreilly.com/catalog/cisrtlist.

Seefeld:2005:RB

[1755] Kim Seefeld and Ernst Linder. R for Bioinformatics. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00544-X. 320 (est.) pp. LCCN ???? URL http://

www.oreilly.com/catalog/9780596005443.

Segaran:2007:PCI

[1756] Toby Segaran. Programming collective intelligence: building Smart Web 2.0 applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52932-5, 0-596-55068-5. xxi + 334 pp. LCCN T58.5 .S43 2007. URL

http://www.oreilly.com/catalog/9780596529321.

Seibold:2008:BBA

[1757] Chris Seibold. Big book of Apple hacks. Hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52982-1. xi + 627 pp. LCCN QA76.76.O63 S45 2008.

Seiwert:2014:ZOZ

[1758] Lothar Seiwert, Holger Wöltje, and Christian Obermayr. Zeitmanagement mit Outlook: die Zeit im Griff mit Microsoft Outlook 2003-2013; Strategien, Tipps und Techniken. (German) [Time management with Outlook: controlling time with Microsoft Outlook 2003–2013; Strategies, tips and techniques]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, tenth edition, 2014. ISBN 3-95561-974-5. 256 pp. LCCN ???? URL http://d-nb.info/

1059839326/04.

Sells:2005:PWP

[1759] Chris Sells and Ian Griffiths. Programming Windows Presentation Foundation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10113-9. xv + 430 + 8 pp. LCCN QA76.76.A65 S45 2005eb; QA76.76.A65. URL

http://www.oreilly.com/catalog/9780596101138.

Serim:1996:NWT

[1760] Ferdi Serim and Melissa Koch. NetLearning: Why Teachers Use the Internet. Songline guides. Songline Studios, Inc. and O’Reilly & Associates, Inc., Sebastopol, CA, 1996. ISBN 1-56592-201-8. xii + 292 pp. LCCN LB1044.87 .S47 1996. US$24.95. URL

http://www.oreilly.com/catalog/9781565922013;   http://www.
oreilly.com/catalog/netlearn. Includes CD-ROM.

Setterberg:1998:TTA

[1761] Fred Setterberg. Travelers’ Tales America. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-885211-28-7. 400 pp. LCCN ???? US$19.95.

URL http://www.oreilly.com/catalog/9781885211286;   http://
www.oreilly.com/catalog/america.

Shafer:2006:HUD

[1762] Dan Shafer and Rachel Andrew. HTML utopia: designing without tables using CSS. SitePoint Pty. Ltd., Collingwood, VIC, Australia, second edition, 2006. ISBN 0-9752402-7-7 (paperback). xvi + 497 pp. LCCN TK5105.888 .S43 2006. URL http://www.oreilly.com/

catalog/9780975240274.

Shakespeare:1998:YND

[1763] Norman Shakespeare. Year 2000 in a Nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-421-5. xiii + 314 pp. LCCN QA76.76.S64 S53 1998. US$24.95.

URL http://www.oreilly.com/catalog/9781565924215;   http://
www.oreilly.com/catalog/y2knut.

Shapiro:1997:NHT

[1764] Michael Shapiro. NetTravel: how travelers use the Internet. Songline guides. Songline Studios, Inc. and O’Reilly & Associates, Inc., Sebastopol, CA, 1997. ISBN 1-56592-172-0. xxviii + 284 pp. LCCN G153.4 .S42 1997; G155.A1 S47 1997. US$24.95. URL

http://www.oreilly.com/catalog/9781565921726;   http://www.
oreilly.com/catalog/nettravel. Foreword by Arthur Frommer.

Shelbourne:2005:PPE

[1765] Tim Shelbourne. Photoshop photo effects cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10022-1. 176 pp. LCCN T385; T385

.S53 2005eb. URL http://www.oreilly.com/catalog/9780596100223.

Shelbourne:2007:PCP

[1766] Tim Shelbourne. Photoshop CS3 photo effects cookbook. O’Reilly digital studio. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51504-9. 176 pp. LCCN

T385. URL http://www.oreilly.com/catalog/9780596515041.

Sheppard:2006:SH

[1767] Andrew Sheppard. Skype hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10189-9. xx + 320 pp. LCCN TK5105.8865eb; TK5105.8865.

URL http://www.oreilly.com/catalog/9780596101893.

Shiflett:2006:EPS

[1768] Chris Shiflett. Essential PHP security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00656-X. xii + 109 pp. LCCN QA76.73.P224; QA76.73.P224 S522 2006eb; QA76.73.P224 S52 2006. URL http://www.

oreilly.com/catalog/9780596006563.

Shirazi:2000:JPT

[1769] Jack Shirazi. Java Performance Tuning. Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 0-596-00015-4. xiv + 426 pp. LCCN QA76.73.J38 S47 2000. US$34.95.

URL http://www.oreilly.com/catalog/9780596000158;   http://
www.oreilly.com/catalog/javapt.

Shirazi:2003:JPT

[1770] Jack Shirazi. Java performance tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00377-3. xv + 570 pp. LCCN QA76.73.J38

S47 2003. URL http://www.oreilly.com/catalog/9780596003777.

Shirky:2002:PWS

[1771] Clay Shirky. Planning for Web Services: Obstacles and Opportunities: An O’Reilly Research Report. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00364-1. 150 (est.) pp. LCCN ???? US$195.00.

URL http://www.oreilly.com/catalog/9780596003647;   http://
www.oreilly.com/catalog/wsrep.

Shirley:1992:GWD

[1772] John Shirley. Guide to Writing DCE Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1992. ISBN 1-56592-004-X. xxi + 251 pp. LCCN QA76.9.D5 S5 1992; QA76.9.D5 S54 1992. US$29.95. URL http://www.

oreilly.com/catalog/9781565920040.

Shirley:1994:GWD

[1773] John Shirley, Wei Hu, and David Magid. Guide to Writing DCE Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, May 1994. ISBN 1-56592-045-7. xxx + 429 pp. LCCN QA76.9.D5 S54 1994. US$29.95.

URL http://www.oreilly.com/catalog/9781565920453;   http://
www.oreilly.com/catalog/gwdce.

Shirley:1995:MRP

[1774] John Shirley and Ward Rosenberry. Microsoft RPC Programming Guide. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, March 1995. ISBN 1-56592-070-8. xix + 232 pp. LCCN QA76.9.D5 S544 1995.

US$24.95. URL http://www.oreilly.com/catalog/9781565920705.

Shnier:1996:DPH

[1775] Mitchell Shnier. Dictionary of PC Hardware and Data Communications Terms. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1996. ISBN 1-56592-158-5. xiv + 516 pp. LCCN TK7887.5 .S53 1996. US$19.95 (US$28.95 Can.).

URL http://www.oreilly.com/catalog/9781565921580;   http://
www.oreilly.com/catalog/dcterms.

Shore:2008:AAD

[1776] James Shore and Shane Warden. The art of agile development. Theory in practice. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52767-5. xviii + 409 pp. LCCN QA76.76.D47; QA76.76.D47eb. URL http://www.oreilly.com/

catalog/9780596527679.

Shupe:2006:FPL

[1777] Rich Shupe and Robert Hoekman. Flash 8: projects for learning animation and interactivity. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10223-2. xvi + 340 pp. LCCN TR897.7; TR897.7 .S48 2006; TR897.7 .S48 2006eb. URL http://www.oreilly.

com/catalog/9780596102234.

Shupe:2007:LAD

[1778] Rich Shupe. Learning ActionScript 3.0: a beginner‘s guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52787-X (paperback). xviii + 363 pp. LCCN TR897.7; TR897.7eb. URL http://www.oreilly.com/

catalog/9780596527877.

Sierra:2003:HFE

[1779] Kathy Sierra and Bert Bates. Head first EJB. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00571-7. xxx + 700 pp. LCCN QA76.3

.S468 2003. URL http://www.oreilly.com/catalog/9780596005719.

Sierra:2003:HFJ

[1780] Kathy Sierra and Bert Bates. Head first Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00465-6. xxx + 619 pp. LCCN QA76.73.J38

S535 2003. URL http://www.oreilly.com/catalog/9780596004651.

Sierra:2005:HFJ

[1781] Kathy Sierra and Bert Bates. Head first Java. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00920-8 (paperback). xxxii + 688 pp. LCCN QA76.73.J38 S535 2005. URL http://www.oreilly.com/

catalog/9780596009205.

Siever:1998:PN

[1782] Ellen Siever, Stephen Spainhour, and Nathan Patwardhan. Perl in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-286-7. xviii + 654 pp. LCCN QA76.73.P22 S53 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565922860;   http://
www.oreilly.com/catalog/perlnut/.

Siever:1999:LND

[1783] Ellen Siever and Jessica Perry Hekman. Linux in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-585-8. xiv + 612 pp. LCCN QA76.76.O63 .S4 1999; QA76.76.O63 H453 1999. US$24.95. URL http://www.oreilly.

com/catalog/9781565925854.

Siever:1999:PCB

[1784] Ellen Siever, Tom Christiansen, Randal L. Schwartz, Larry Wall, and Sriram Srinivasan. The Perl CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-462-2. LCCN QA76.73.P22 P47 1999. US$59.95.

URL http://www.oreilly.com/catalog/9781565924628;   http://
www.oreilly.com/catalog/perlcdbs/. Six O’Reilly Perl books on a

CD-ROM.

Siever:2000:LND

[1785] Ellen Siever et al., editors. Linux in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2000. ISBN 0-596-00025-1. xv + 797 pp. LCCN QA76.76.O63 L5459 2000; QA76.76.O63 L54595 2000. URL http://www.oreilly.

com/catalog/9780596000257.

Siever:2001:LWS

[1786] Ellen Siever, Matt Welsh, Ben Laurie, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. Linux Web Server CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00208-4. 812 (est.) pp. LCCN QA76.76.O63 L56 2001. US$79.95.

URL http://www.oreilly.com/catalog/9780596002084;   http://
www.oreilly.com/catalog/linuxwebcdbs.

Siever:2001:PCB

[1787] Ellen Perl Siever, Larry Wall, Sriram Srinivasan, Tom Christiansen, and David N. Blank-Edelman, editors. The Perl CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 2.0. edition, 2001. ISBN 0-596-00164-9. LCCN QA76.73.P22. URL

http://www.oreilly.com/catalog/9780596001643;   http://www.
oreilly.com/catalog/perlcdbs2. Includes CD-ROM.

Siever:2003:LN

[1788] Ellen Siever, Stephen Figgins, and Aaron Weber. Linux in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2003. ISBN 0-596-00482-6. xiv + 928 pp. LCCN QA76.76.O63 S5582 2003. URL

http://www.oreilly.com/catalog/9780596004828.

Siever:2005:LN

[1789] Ellen Siever. Linux in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 928 pp. LCCN QA76.76.O63; QA76.76.O63 H453 2005eb; QA76.76.O63 H453 2005; QA76.76.O63 L5459 2003. URL http://www.oreilly.com/

catalog/9780596009304.

Sigglekow:2005:JSC

[1790] Bill Sigglekow. Jakarta Struts cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00771-X. xvi + 507 pp. LCCN QA76.73.J38 S445 2005. URL

http://www.oreilly.com/catalog/9780596007713.

Simmons:2004:HJ

[1791] Robert Simmons, Jr. Hardcore Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00568-7. xvii + 324 pp. LCCN QA76.73.J38 S553

2004. URL http://www.oreilly.com/catalog/9780596005689.

Simmons:2004:HJS

[1792] Robert Simmons, Jr. Hardcore Java: Secrets of the Java masters. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00568-7. xvii + 324 pp. LCCN QA76.73.J38 S553 2004.

Simon:2000:ENP

[1793] Jinjer L. Simon. Excel 2000 in a Nutshell: a power user’s quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-714-1. xii + 591 pp. LCCN HF5548.4.M523 S56 2000.

URL http://www.oreilly.com/catalog/9781565927148;   http://
www.oreilly.com/catalog/excel2000ian.

Simpson:2002:XXL

[1794] John E. Simpson. XPath and XPointer: Locating Content in XML Documents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00291-2. x + 196 pp. LCCN QA76.76.H94 S578

2002. US$24.95. URL http://safari.oreilly.com/0596002912;
http://www.oreilly.com/catalog/9780596002916;   http://www.
oreilly.com/catalog/xpathpointer.

Skinner:2007:UA

[1795] Jesse Skinner. Unobtrusive Ajax. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51024-1. LCCN TK5105.8885.A52.

URL http://www.oreilly.com/catalog/9780596510244.

Sklar:2003:PC

[1796] David Sklar and Adam Trachtenberg. PHP Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 1-56592-681-1. xxi + 608 pp. LCCN QA76.73.P224 S56 2003. US$39.95.

URL http://www.oreilly.com/catalog/9781565926813;   http://
www.oreilly.com/catalog/phpckbk.

Sklar:2004:LP

[1797] David Sklar. Learning PHP 5. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00560-1. xviii + 350 pp. LCCN QA76.73.P224 S55 2004.

URL http://www.oreilly.com/catalog/9780596005603.

Sklar:2006:PC

[1798] David Sklar and Adam Trachtenberg. PHP cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10101-5. xxiv + 784 pp. LCCN QA76.73.P224 S56 2006; QA76.73.P224; QA76.73.P224 S55

2006eb. URL http://www.oreilly.com/catalog/9780596101015.

Sloan:2001:NTT

[1799] Joseph D. (Joseph Donald) Sloan. Network Troubleshooting Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00186-X. xv + 346 pp. LCCN TK5105.5 .S557 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596001865;   http://
www.oreilly.com/catalog/nettroubletools.

Sloan:2005:HPL

[1800] Joseph D. (Joseph Donald) Sloan. High performance Linux clusters with OSCAR, Rocks, openMosix, and MPI. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00570-9. xv + 350 pp. LCCN QA76.58; QA76.58 .S56 2005eb; QA76.58 .S56 2005; QA76.58 .S58 2005; QA76.58 .S595 2005.

URL http://www.oreilly.com/catalog/9780596005702.

Smart:2008:JPT

[1801] John Ferguson Smart. Java power tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52793-4. xxxv + 871 pp. LCCN QA76.73.J38 S63 2008.

Smith:2005:AFT

[1802] Jared Smith, Jim Van Meggelen, and Leif Madsen. Asterisk: The Future of Telephony. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00962-3. 380 (est.) pp. LCCN TK5105.8865 .V26 2005. US$39.95, CDN$55.95, UK£28.50. URL

http://www.oreilly.com/catalog/9780596009625.

Smith:2005:DL

[1803] Roderick Smith and Jeff Duntemann. Degunking Linux. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-04-3. xx + 332 pp. LCCN QA76.76.O63 S58 2005. URL http://www.oreilly.com/

catalog/9781933097046.

Smith:2005:LWW

[1804] Roderick W. Smith. Linux in a Windows world. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005.

ISBN 0-596-00758-2. xiv + 478 pp. LCCN QA76.76.O63 S548 2005.

URL http://www.oreilly.com/catalog/9780596007584.

Smith:2007:DLP

[1805] Jennifer Smith. Dynamic learning Photoshop CS3: with digital classroom video tutorials. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51061-6. xiii + 349 pp. LCCN T385 .S62283 2007eb; T385 .S62283 2007; T385. URL http://

www.oreilly.com/catalog/9780596510619.

Smith:2009:PF

[1806] Chris Smith. Programming F#. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-80208-0. ???? pp. LCCN ????

Smith:2010:PF

[1807] Chris Smith. Programming F#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-15364-3. xx + 383 pp. LCCN ????

Smith:2015:C

[1808] Kurt W. Smith. Cython. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0155-1. 240 pp. LCCN ???? URL http://proquest.

safaribooksonline.com/9781491901731.

Snell:2002:PWS

[1809] James Snell, Doug Tidwell, and Pavel Kulchenko. Programming Web Services with SOAP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00095-2. xiii + 244 pp. LCCN QA76.625 .S64 2002

Bar. US$34.95. URL http://safari.oreilly.com/0596000952;
http://www.oreilly.com/catalog/9780596000950;   http://www.
oreilly.com/catalog/progwebsoap.

Sofield:2004:PH

[1810] Shannon Sofield, Dave Nielsen, and Dave Burchell. PayPal hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00751-5. xxiv + 343 pp. LCCN HG1710 .S64 2004. URL http://www.oreilly.

com/catalog/9780596007515.

Souders:2007:HPW

[1811] Steve Souders. High performance Web sites: essential knowledge for frontend engineers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52930-9 (paperback). xix + 146 pp. LCCN QA76.625; QA76.625 .S68

2007; QA76.625 .S68 2007eb. URL http://www.oreilly.com/catalog/
9780596529307.

Southwick:2012:JNW

[1812] Peter Southwick. Juniper Networks Warrior: a Guide to the Rise of Juniper Networks Implementations. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1663-8. 400 pp. LCCN ???? EUR 49.00.

Spainhour:1996:MIE

[1813] Stephen Spainhour. The Microsoft Internet Explorer 2.0 Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-192-5. 230

pp. LCCN ???? US$17.95. URL http://www.oreilly.com/catalog/
9781565921924.

Spainhour:1996:WND

[1814] Stephen Spainhour and Valerie Quercia. WebMaster in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1996. ISBN 1-56592-229-8. xv + 356 pp. LCCN TK5105.888 .S63 1996; TK5105.888 .S68 1996. US$19.95.

URL http://www.oreilly.com/catalog/9781565922297;   http://
www.oreilly.com/catalog/webmaster.

Spainhour:1999:WN

[1815] Stephen Spainhour and Robert Eckstein. Webmaster in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-325-1. 523 pp. LCCN TK5105.888 .S68 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565923256;   http://
www.oreilly.com/catalog/webmaster2.

Spainhour:2003:WN

[1816] Stephen Spainhour and Robert Eckstein. Webmaster in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00357-9. xiv + 561 pp. LCCN TK5105.888 .S68 2003. URL http://

www.oreilly.com/catalog/9780596003579.

Spainhour:2003:WPC

[1817] Stephen Spainhour, Robert Eckstein, Danny Goodman, Hugh E. Williams, David Lane, Rasmus Lerdorf, Kevin Tatroe, David Sklar, and Adam Trachtenberg, editors. The Web programming CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 1.0. edition, 2003. ISBN 0-596-00510-5. LCCN TK5105.888. URL http://www.oreilly.com/

catalog/9780596005108.

Spector:2000:BLC

[1818] David H. M. Spector. Building Linux clusters. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-625-0. xviii + 332 pp. LCCN QA76.76.O63

S6647 2000. URL http://www.oreilly.com/catalog/9781565926257.

Spencer:1998:MU

[1819] Henry Spencer and David Lawrence. Managing Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1998. ISBN 1-56592-198-4. xiii + 492 pp. LCCN TK5105.875.U83 S64

1998. US$32.95. URL http://www.ora.com/catalog/musenet/;
http://www.oreilly.com/catalog/9781565921986;   http://www.
oreilly.com/catalog/musenet;   http://www.oreilly.com/catalog/
musenet/noframes.html.

Spetic:2002:TSC

[1820] Ales Spetic and Jonathan Gennick. Transact-SQL Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-756-7. xviii + 282 pp. LCCN QA76.73.S67 S64 2002. US$34.95. URL http://safari.oreilly.com/

1565927567;   http://www.oreilly.com/catalog/9781565927568;
 http://www.oreilly.com/catalog/transqlcook.

Spinellis:2009:BA

[1821] Diomidis Spinellis and Georgios Gousios, editors. Beautiful architecture. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15578-6. xix + 404 pp. LCCN

QA76.754 .B43 2009. URL http://proquest.safaribooksonline.
com/9780596155780.

Spry:1994:IBC

[1822] SPRY, Inc. Internet In a Box: The Complete Internet Solution. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1994. ISBN ???? ???? pp. LCCN ????

US$149.00. URL http://www.oreilly.com/catalog/.

Spry:1996:IBC

[1823] SPRY, Inc. Internet In a Box (TM), Version 2.0: The Complete Internet Solution. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1996. ISBN ???? ???? pp. LCCN ???? US$149.00. URL http://www.oreilly. com/catalog/. Includes two diskettes and a copy of The Whole Internet Users Guide & Catalog.

Spurgeon:2000:EDG

[1824] Charles (Charles E.) Spurgeon. Ethernet: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2000. ISBN 1-56592-660-9. 500 pp. LCCN TK5105.8.E83 S67 2000. US$44.95.

URL http://www.oreilly.com/catalog/9781565926608;   http://
www.oreilly.com/catalog/enettdg.

Srinivasan:1997:APP

[1825] Sriram Srinivasan. Advanced Perl Programming. Perl series; A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-220-4. xxiii + 404 pp. LCCN QA76.73.P22 S655

1997. US$34.95. URL http://www.ora.com/catalog/advperl/;
http://www.oreilly.com/catalog/9781565922204;   http://www.
oreilly.com/catalog/advperl.

Stafford:2005:MH

[1826] Tom Stafford and Matt Webb. Mind hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00779-5. xxvii + 363 pp. LCCN QP376 .S7615 2005; QP376 .S83

2005; QP376 .S7615 2005eb. URL http://www.oreilly.com/catalog/
9780596007799.

Stanek:2006:MCR

[1827] William R. Stanek. MCSE Core required exams in a nutshell: a desktop quick reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2006. ISBN 0-596-10228-3. xiv + 720 pp. LCCN QA76.3; QA76.3 .S73 2006eb. URL http://www.

oreilly.com/catalog/9780596102289.

Stanek:2007:WVD

[1828] William R. Stanek and Paul Marquardt. Windows Vista: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52800-0. xvii + 922 pp. LCCN QA76.76.O63; QA76.76.O63 S74 2007eb. URL http:

//www.oreilly.com/catalog/9780596528003.

Stark:2010:BIA

[1829] Jonathan Stark. Building iPhone apps with HTML, CSS, and JavaScript: Making App Store apps without Objective-C or Cocoa. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 1-4493-8023-9, 0-596-80578-0. xv + 166 pp. LCCN ????

Stay:2008:FE

[1830] Jesse Stay. FBML essentials. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51918-4. 120 (est.) pp. LCCN QA76.76.A65 S73 2008.

Stefanov:2010:JP

[1831] Stoyan Stefanov. JavaScript patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-9911-8 (paperback), 0-596-80675-2 (paperback). xvi + 216 pp. LCCN QA76.73.J39 S734 2010.

Stein:1997:SLA

[1832] Lincoln Stein, Clint Wong, Ron Petrusha, Shishir Gundavaram, et al. Scripting Languages: Automating the Web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-265-4. 230 pp. LCCN ????

US$29.95. URL http://www.ora.com/catalog/wj6/;   http://www.
oreilly.com/catalog/9781565922655.

Stein:1999:WAM

[1833] Lincoln D. Stein and Doug MacEachern. Writing Apache Modules with Perl and C: The Apache API and mod perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-567-X. xix + 724 pp. LCCN TK5105.8885.A63.S74 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565925670;   http://
www.oreilly.com/catalog/wrapmod/.

Steinberg:2004:MXP

[1834] Gene Steinberg. Mac OS X 10.3 Panther. Little black book. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2004. ISBN 1-932111-86-7. xxv + 548 pp. LCCN QA76.76.O63 S75 2004. URL http://www.oreilly.com/

catalog/9781932111866.

Steinberg:2005:ZCN

[1835] Daniel H. (Daniel Howard) Steinberg and Stuart Cheshire. Zero configuration networking: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10100-7. 256 pp. LCCN TK5105.5eb; TK5105.5. URL

http://www.oreilly.com/catalog/9780596101008.

Stellman:2005:ASP

[1836] Andrew Stellman and Jennifer Greene. Applied software project management. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00948-8. 256 pp. LCCN QA76.758 .S74 2005; QA76.758; QA76.758 .S74 2005eb. URL

http://www.oreilly.com/catalog/9780596009489.

Stellman:2008:HFC

[1837] Andrew Stellman and Jennifer Greene. Head first C#. O’Reilly’s Head first series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51482-4 (paperback). xl + 738 pp. LCCN QA76.73.C154; QA76.73.C154eb. URL http://www.

oreilly.com/catalog/9780596514822.

Stellman:2010:HFC

[1838] Andrew Stellman and Jennifer Greene. Head first C#. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2010. ISBN 1-4493-8034-4 (paperback). ???? pp. LCCN ???? US$35.50.

Stephens:2006:CC

[1839] D. Ryan Stephens et al. C++ cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00761-2 (paperback). xvi + 573 pp. LCCN

QA76.73.C153 S84 2006. URL http://www.oreilly.com/catalog/
9780596007614.

Sterling:1996:TTF

[1840] Richard Sterling. Travelers’ Tales Food: a Taste of the Road. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-885211-09-0. xxiv + 436 pp. LCCN TX357 .T73 1996. US$17.95.

URL http://www.ora.com/catalog/food/noframes.html;   http:/
/www.oreilly.com/catalog/9781885211095;   http://www.oreilly.
com/catalog/food.

Sterling:1997:TTG

[1841] Richard Sterling. Travelers’ Tales: Gutsy Gourmet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-22-8. 100 (est.) pp. LCCN

???? US$7.95. URL http://www.ora.com/catalog/gourmet/;
http://www.oreilly.com/catalog/9781885211224;   http://www.
oreilly.com/catalog/diner.

Sterling:1999:AF

[1842] Richard Sterling, editor. The Adventure of Food: True Stories of Eating Everything. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-37-6. 375 (est.) pp. LCCN TX357.A38 1999. US$17.95.

URL http://www.oreilly.com/catalog/9781885211378;   http://
www.oreilly.com/catalog/advenfood.

Stern:1991:MNN

[1843] Hal Stern. Managing NFS and NIS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-75-7. xxiv + 410 pp. LCCN TK5105.5 .S74 1991.

URL http://www.oreilly.com/catalog/9780937175750;   http://
www.oreilly.com/catalog/nfs.

Stern:2001:MNN

[1844] Hal Stern, Mike Eisler, and Ricardo Labiaga. Managing NFS and NIS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 2001. ISBN 1-56592-510-6. xviii + 490 pp. LCCN TK5105.574 .S74 2001. US$39.95.

URL http://www.oreilly.com/catalog/9781565925106;   http://
www.oreilly.com/catalog/nfs2.

Stern:2005:IIH

[1845] Hadley Stern. iPod and iTunes hacks: Tips and tools for ripping, mixing, and burning. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00778-7 (paperback). xxi + 433 pp. LCCN ML74.4.I48 S74 2005. URL http://www.oreilly.

com/catalog/9780596007782.

Steward:2004:PH

[1846] Sid Steward. PDF Hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00655-1 (paperback). xvi + 278 pp. LCCN QA76.76.F48 S74 2004. URL http://www.oreilly.com/

catalog/9780596006556.

Stigliani:1995:CUS

[1847] Joan Stigliani. The Computer User’s Survival Guide: Staying Healthy in a High Tech World. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 1995. ISBN 1-56592-030-9. 276 pp. LCCN RC965.V53 S75 1995. US$21.95.

URL http://www.ora.com/gnn/bus/ora/item/cusg.html;   http:/
/www.oreilly.com/catalog/9781565920309.

StLaurent:2001:PWS

[1848] Simon St. Laurent, Joe Johnston, and Edd Dumbill. Programming Web Services with XML-RPC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 0-596-00119-3. xvii + 213 pp. LCCN QA76.76.H94 O7 2001; QA76.76.H94 S717 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596001193;   http://
www.oreilly.com/catalog/progxmlrpc.

StLaurent:2004:UOS

[1849] Andrew M. St. Laurent. Understanding Open Source and Free Software Licensing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00581-4. xii + 193 pp. LCCN QA76.76.S46 S7 2004; K1443.C6 S35 2004.

US$24.95. URL http://www.oreilly.com/catalog/9780596005818;
 http://www.oreilly.com/catalog/osfreesoft/book/.

StLaurent:2005:XPR

[1850] Simon St. Laurent, Michael Fitzgerald, and Robert Eckstein. XML pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-10050-7 (paperback). iv + 171 pp. LCCN QA76.76.H94 E25 2005. URL http://

www.oreilly.com/catalog/9780596100506.

StLaurent:2009:LR

[1851] Simon St Laurent and Edd Dumbill. Learning Rails. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15494-1. 448 (est.) pp. LCCN QA76.73.R83 S25 2008.

StLaurent:2010:LRL

[1852] Simon St Laurent and Edd Dumbill. Learning Rails: live editions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 1-4493-8325-4, 0-596-51877-3. 417 (est.) pp. LCCN QA76.76.A65 S8L 2010.

Stokes:2007:IMI

[1853] Jon Stokes. Inside the machine: an illustrated introduction to microprocessors and computer architecture. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-104-2. xx + 292 pp. LCCN TK7895.M5

S76 2007. URL http://www.loc.gov/catdir/enhancements/fy0715/
2005037262-d.html;   http://www.loc.gov/catdir/toc/ecip065/
2005037262.html;   http://www.oreilly.com/catalog/9781593271046.

Stolarz:2005:CPH

[1854] Damien Stolarz. Car PC hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00871-6 (paperback). xviii + 361 pp. LCCN TL272.53 .S76 2005. URL http://

www.oreilly.com/catalog/9780596008710.

Story:2002:DPP

[1855] Derrick Story. Digital Photography Pocket Guide. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00454-0. xiii + 114 pp. LCCN TR267 .S78 2003. US$14.95.

URL http://www.oreilly.com/catalog/9780596004545;   http://
www.oreilly.com/catalog/digphotopg.

Story:2003:DVP

[1856] Derrick Story. Digital video pocket guide. O’Reilly Digital Studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00523-7. xii + 113 pp. LCCN TR882.3 .S76 2003. URL http://www.oreilly.com/

catalog/9780596005238.

Story:2004:DPH

[1857] Derrick Story. Digital photography hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00666-7 (paperback). xx + 309 pp. LCCN TR267 .S7797 2004.

URL http://www.oreilly.com/catalog/9780596006662.

Story:2004:DPP

[1858] Derrick Story. Digital photography pocket guide. O’Reilly digital studio. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00627-6. xiii + 113

pp. LCCN TR267 .S78 2004. URL http://www.oreilly.com/catalog/
9780596006273.

Story:2005:DPP

[1859] Derrick Story. Digital Photography Pocket Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-10015-9. xvi + 155 pp. LCCN TR267 .S78 2005. US$14.95, CDN$20.95, UK£9.95.

URL http://www.oreilly.com/catalog/9780596100155;   http://
www.oreilly.com/catalog/digphotopg3/.

Story:2005:IFB

[1860] Derrick Story. iBook fan book: smart and beautiful to boot. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00861-9. xvi + 104 pp. LCCN QA76.8.I3132

S76 2004. URL http://www.oreilly.com/catalog/9780596008611.

Story:2008:DPC

[1861] Derrick Story. The digital photography companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51766-1. xvi + 214 pp. LCCN TR267 .S7795 2008.

Story:2008:PCC

[1862] Derrick Story. The Photoshop CS4 companion for photographers. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-15587-5, 0-596-52193-6. xiii + 190 pp. LCCN TR267.5.A33 S76 2008.

Strang:1986:PC

[1863] John Strang. Programming with curses. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1986. ISBN 0-937175-02-1. 76 pp. LCCN QA76.73.C15S78 1991.

US$12.95. URL http://www.oreilly.com/catalog/9780937175026;
 http://www.oreilly.com/catalog/curses.

Strang:1986:PTC

[1864] John Strang. Programming with curses. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1986. ISBN 0-937175-02-1. 71 pp.

Strang:1986:RWT

[1865] John Strang. Reading and Writing TERMCAP Entries. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1986. ISBN ???? 73 pp. LCCN ???? URL

http://www.oreilly.com/catalog/.

Strang:1988:TT

[1866] John Strang, Linda Mui, and Tim O’Reilly. termcap & terminfo. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, April 1988. ISBN 0-937175-22-6. xv + 253 pp. LCCN QA76.76.O63 S765 1990; QA76.8.U65

S79 1988. US$21.95. URL http://www.oreilly.com/catalog/
9780937175224;   http://www.oreilly.com/catalog/term.

Strother:1993:S

[1867] Lesley Strother. smileys. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1993. ISBN 1-56592-041-4. 595 pp. A collection of 650 “smileys”. While not an Internet book per se, smileys are certainly used enough on the Internet to warrant an entry here.:-).

Stubblebine:2003:REP

[1868] Tony Stubblebine. Regular expression pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00415-X. vi + 93 pp. LCCN QA76.9.T48

S78 2003. URL http://www.oreilly.com/catalog/9780596004156.

Stubblebine:2004:SHD

[1869] Tony Stubblebine and Junko Mishima. Seiki hyogen desukutoppu rifarensu: regular expressions for Perl, C, PHP, Python, Java, and .NET. Orairi Japan, Tokyo, Japan, 2004. ISBN 4-87311-170-6. vi + 96 pp. LCCN ????

Stubblebine:2007:REP

[1870] Tony Stubblebine. Regular expression pocket reference: Regular expressions for Perl, Ruby, PHP, Python, C, Java, and .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51427-1 (paperback). vii + 117 pp. LCCN QA76.9.T48 S78 2007. URL http://

proquest.safaribooksonline.com/9780596514273;   http://www.
loc.gov/catdir/toc/fy0802/2007281074.html.

Stubblebine:2008:RAK

[1871] Toni Stubblebine, Peter Klicman, and Lars Schulten. Reguläre Ausdrücke — kurz and gut [für Perl, Ruby, PHP, C#, Python, Java and .NET]. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2008. ISBN 3-89721-535-7. 131 pp. LCCN ???? EUR 9.90.

Stutz:2003:SSC

[1872] David Stutz, Ted Neward, and Geoff Shilling. Shared Source CLI Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00351-X. xviii + 357 pp. LCCN QA76.625 .S76 2003. US$34.95.

URL http://www.oreilly.com/catalog/9780596003517;   http://
www.oreilly.com/catalog/sscliess.

Stutz:2004:LCT

[1873] Michael Stutz. The Linux cookbook: tips and techniques for everyday use. No Starch Press, San Francisco, CA, USA, second edition, 2004. ISBN 1-59327-031-3. xxxiv + 788 pp. LCCN QA76.76.O63 S788 2004eb; QA76.76.O63 S788 2004. URL http://www.oreilly.com/

catalog/9781593270315.

Subramaniam:2005:NGW

[1874] Venkat Subramaniam. .NET gotchas: 75 ways to improve your C# and VB.NET programs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00909-7. xviii + 372 pp. LCCN QA76.625 .S93 2005. URL http:

//www.oreilly.com/catalog/9780596009090.

Suda:2006:UM

[1875] Brian Suda. Using microformats. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52821-3. LCCN QA76.76.H94. URL http://www.

oreilly.com/catalog/9780596528218.

Sun:1999:UMP

[1876] Andrew Sun. Using and Managing PPP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-321-9. xii + 430 pp. LCCN TK5105.582

.S86 1999. US$32.95. URL http://www.ora.com/catalog/umppp/;
http://www.oreilly.com/catalog/9781565923218.

Swan:1998:SSW

[1877] Sheila Swan and Peter Laufer. Safety and Security for Women Who Travel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-885211-29-5. xii + 147 pp. LCCN G151 .S93 1998. US$12.95.

URL http://www.oreilly.com/catalog/9781885211293;   http://
www.oreilly.com/catalog/security.

Swan:2015:BBNa

[1878] Melanie Swan. Bitcoin: A Blueprint for a New World Currency. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-2049-1. 120 pp. LCCN ????

URL http://deposit.d-nb.de/cgi-bin/dokserv?id=5125084&prov=
M&dok_var=1&dok_ext=htm.

Swan:2015:BBNb

[1879] Melanie Swan. Blockchain: blueprint for a new economy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-2049-1, 1-4919-2045-9 (e-book), 1-4919-2047-5 (e-book). xviii + 130 pp. LCCN HG1710 .S93 2015. URL http://

proquest.safaribooksonline.com/?fpi=9781491920480.

Swarr:2007:F

[1880] Robert Swarr. Facelets. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52924-4. LCCN QA76.76.D47 S93 2007eb; QA76.76.D47. URL

http://www.oreilly.com/catalog/9780596529246.

Swinford:2006:FPA

[1881] Echo Swinford. Fixing PowerPoint annoyances: how to fix the most annoying things about your favorite presentation program. Annoyances series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10004-3. xviii + 305 pp. LCCN T385 .S85 2006; T385 .S88 2006; T385 .S85 2006eb; T385. URL http://www.

oreilly.com/catalog/9780596100049.

Sydik:2007:DAW

[1882] Jeremy J. Sydik. Design accessible web sites: thirty-six keys to creating content for all audiences and platforms. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 1-934356-02-6 (paperback). xii + 318 pp. LCCN TK5105.888 .S955 2007. URL http://

www.oreilly.com/catalog/.

Sydow:1999:POS

[1883] Dan Parks Sydow. Programming the Be Operating System: Writing Programs for the Be Operating System. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-467-3. x + 390 pp. LCCN QA76.76.O63 S93 1999; QA76.76.O63 S94345 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565924673;   http://
www.oreilly.com/catalog/beosprog.

Sylwester:1997:MFS

[1884] Peter Sylwester. Macromedia Flash Studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-352-9. 250 pp. LCCN ????

US$34.95. URL http://www.ora.com/catalog/flash/;   http://
www.oreilly.com/catalog/9781565923522.

Syroid:2000:ONP

[1885] Tom Syroid and Bo Leuf. Outlook 2000 in a nutshell: a power user’s quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-704-4. xvi + 642 pp. LCCN HD69.T54 S97 2000.

US$29.95. URL http://www.oreilly.com/catalog/9781565927049;
 http://www.oreilly.com/catalog/out2000ian.

Tahaghoghi:2007:LM

[1886] Seyed M. M. Tahaghoghi and Hugh E. Williams. Learning MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00864-3 (paperback). xvii + 598 pp. LCCN QA76.7.S67; QA76.7.S67 T34 2007;

QA76.7.S67 T34 2007eb. URL http://www.loc.gov/catdir/toc/
fy0706/2007272020.html;   http://www.oreilly.com/catalog/9780596008642.

Talbot:2007:DSB

[1887] Steve Talbot. Devices of the soul: battling for our selves in an age of machines. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52680-6. xv + 281 pp. LCCN T14.5 .T33 2007; T14.5 .T33 2007eb; T14.5; HM851 .T35 2007.

URL http://www.oreilly.com/catalog/9780596526801.

Talbott:1986:MPM

[1888] Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1986. ISBN ???? LCCN ???? URL http://www.oreilly.

com/catalog/.

Talbott:1988:MPM

[1889] Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1988. ISBN 0-937175-04-8. 77 pp. LCCN QA76.8.U65 T34 1988. URL http://www.oreilly.com/

catalog/9780937175040.

Talbott:1992:PRM

[1890] Steve Talbott, editor. PEXlib Reference Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1992. ISBN 1-56592-029-5. xxv + 551 pp. LCCN QA76.76.W56 P4 1992; QA76.76.W56 P4757 1992. US$34.95. URL

http://www.oreilly.com/catalog/9781565920293.

Talbott:1995:FDC

[1891] Stephen L. Talbott. The Future Does Not Compute: Transcending the Machines in Our Midst. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1995. ISBN 1-56592-085-6. xix + 481 pp. LCCN QA76.9.C66 T26 1995; QA76.9.C66 T33 1995. US$22.95.

URL http://www.oreilly.com/catalog/9781565920859;   http://
www.oreilly.com/catalog/future.

Tapp:2006:PCM

[1892] Eddie Tapp and Rick Lucas. Practical color management: Eddie Tapp on digital photography. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52768-3. 150 pp. LCCN TR510 .T37 2006eb; TR510. URL http:

//www.oreilly.com/catalog/9780596527686.

Tapp:2006:PWS

[1893] Eddie Tapp. Photoshop workflow setups: Eddie Tapp on digital photography. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2006. ISBN 0-596-10168-6. 197 pp. LCCN TR267; TR267 .T37 2006eb; TR267 .T37 2006. URL http://www.oreilly.

com/catalog/9780596101688.

Tate:2004:BFL

[1894] Bruce Tate and Justin Gehtland. Better, Faster, Lighter Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00676-4. 243 (est.) pp. LCCN QA76.73 T38 2004. US$34.95, CDN$50.95, UK£24.95.

URL http://www.oreilly.com/catalog/9780596006761;   http://
www.oreilly.com/catalog/bfljava/.

Tate:2005:BJ

[1895] Bruce Tate. Beyond Java. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10094-9. xii + 185 pp. LCCN QA76.73.J38 T38 2005; QA76.73.J38 T38 2005eb; QA76.73.J38. URL http://www.oreilly.

com/catalog/9780596100940.

Tate:2005:SDN

[1896] Bruce Tate, Justin Gehtland, and Michael Kosta Loukides. Spring: a developer’s notebook. A developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00910-0. xv + 184 pp. LCCN QA76.73.S67

T38 2005. URL http://www.oreilly.com/catalog/9780596009106.

Tate:2006:RRR

[1897] Bruce Tate and Curt Hibbs. Ruby on rails: up and running. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10132-5. xi + 167 pp. LCCN QA76.73.R82 T38 2006eb; QA76.73.R82 T38 2006; QA76.73.R82. URL

http://www.oreilly.com/catalog/9780596101329.

Tate:2010:SLS

[1898] Bruce Tate. Seven languages in seven weeks: a pragmatic guide to learning programming languages. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-59-X (paperback). xvii + 317 pp. LCCN QA76.7 .T38 2010.

Taylor:2001:KPS

[1899] Don Taylor, Jim Mischel, and Tim Gentry. Kylix power solutions. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-932111-29-8. xviii + 331 pp. LCCN QA76.76.A65 T415 2001eb. URL http://www.oreilly.

com/catalog/9781932111293.

Taylor:2002:LUM

[1900] Dave Taylor, Jerry D. Peek, Grace Todino, and John Strang. Learning Unix for Mac OS X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00342-0. xiv + 139 pp. LCCN QA76.76.O63 T388

2002. US$19.95. URL http://safari.oreilly.com/0596003420;
http://www.oreilly.com/catalog/9780596003425;   http://www.
oreilly.com/catalog/lunixmacosx.

Taylor:2003:LUM

[1901] Dave Taylor and Brian Jepson. Learning Unix for Mac OS X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00470-2. xiv + 141 pp. LCCN QA76.76.O63 T388 2003. URL http://www.oreilly.

com/catalog/9780596004705. Covers Mac OS X Jaguar.

Taylor:2004:LUM

[1902] Dave Taylor and Brian Jepson. Learning UNIX for Mac OS X Panther. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2004. ISBN 0-596-00617-9. xiii + 168 pp. LCCN QA76.76.O63 T3884 2004. URL http://www.oreilly.com/

catalog/9780596006174.

Taylor:2004:WCS

[1903] Dave Taylor. Wicked cool shell scripts. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-012-7. xx + 365 pp. LCCN

QA76.76.O63 T3895 2004eb. URL http://www.oreilly.com/catalog/
9781593270124.

Taylor:2005:LUM

[1904] Dave Taylor. Learning UNIX for Mac OS X Tiger. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-00915-1. xviii + 260 pp. LCCN QA76.76.O63

T388 2005. URL http://www.oreilly.com/catalog/9780596009151.

Team:1998:ATO

[1905] The Be Development Team. Be Advanced Topics: Official documentation for the BeOS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1998. ISBN 1-56592-396-0. xi + 365 pp. LCCN QA76.76.O63 B4 1998. US$39.95.

URL http://www.oreilly.com/catalog/beadv. Includes BeSpecific 5

CD-ROM.

Teetor:2011:RC

[1906] Paul Teetor. R cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 0-596-80915-8. xviii + 413

pp.  LCCN QA276.45.R3 T44 2011.  URL http://www.loc.

gov/catdir/enhancements/fy1113/2011282768-b.html;   http://
www.loc.gov/catdir/enhancements/fy1113/2011282768-d.html;
http://www.loc.gov/catdir/enhancements/fy1113/2011282768-t.
html.

Teilhet:2001:SHV

[1907] Stephen Teilhet. Subclassing and Hooking with Visual Basic. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 0-596-00118-5. xvii + 685 pp. LCCN QA76.73.B3 T4 2001. US$49.95.

URL http://www.oreilly.com/catalog/9780596001186;   http://
www.oreilly.com/catalog/subhookvb.

Teilhet:2004:CC

[1908] Stephen Teilhet and Jay Hilyard. C# cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00339-0. xxii + 831 pp. LCCN

QA76.73.C154 T43 2004. URL http://www.oreilly.com/catalog/
9780596003395.

Thai:1999:LD

[1909] Thuan L. Thai. Learning DCOM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-581-5. xx + 479 pp. LCCN QA76.9.A73 T48 1999; QA76.9.D5 T476 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565925816;   http://
www.oreilly.com/catalog/ldcom/.

Thai:2001:NFE

[1910] Thuan L. Thai and Hoang Q. Lam. .NET Framework Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 0-596-00165-7. xiv + 304 pp. LCCN QA76.625 .T47 2001. US$29.95.

URL http://www.oreilly.com/catalog/9780596001650;   http://
www.oreilly.com/catalog/dotnetfrmess.

Thai:2002:NFE

[1911] Thuan L. Thai and Hoang Lam. .NET Framework Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00302-1. xii + 307 pp. LCCN QA76.625 .T47

2002. US$29.95. URL http://safari.oreilly.com/0596003021;
http://www.oreilly.com/catalog/9780596003029;   http://www.
oreilly.com/catalog/dotnetfrmess2.

Thai:2003:NFE

[1912] Thuan L. Thai and Hoang Q. Lam. .NET framework essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00505-9. xiv + 363 pp. LCCN QA76.625 .T47 2003. URL http://www.oreilly.com/

catalog/9780596005054.

Thau:2006:BJP

[1913] Thau. The book of JavaScript: a practical guide to interactive Web pages. No Starch Press, San Francisco, CA, USA, second edition, 2006. ISBN 1-59327-106-9. xxx + 490 pp. LCCN QA76.73.J39 T37 2006eb; QA76.73.J39 T37 2007. URL http://www.oreilly.com/

catalog/9781593271060.

Theriault:1998:OS

[1914] Marlene Theriault and William Heney. Oracle Security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-450-9. xviii + 425 pp. LCCN QA76.9.A25 T54 1998; QA76.9.D314 T453 1998. US$34.95.

URL http://www.oreilly.com/catalog/9781565924505;   http://
www.oreilly.com/catalog/orasec.

Thomas:2004:PRP

[1915] David Thomas, Chad Fowler, and Andrew Hunt. Programming Ruby: the pragmatic programmers’ guide. Pragmatic Bookshelf, Raleigh, NC, USA, second edition, 2004. ISBN 0-9745140-5-5 (paperback). xxxii + 830 pp. LCCN QA76.64 .T494 2004; QA76.64 .T494 2005; QA76.64 .T48 2004. URL http://www.loc.gov/

catdir/enhancements/fy0715/2005299026-d.html;   http://www.
oreilly.com/catalog/9780974514055.

Thomas:2004:PVC

[1916] David Thomas and Andrew Hunt. Pragmatic version control: using CVS, volume 1 of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-0-4. xiv + 161 pp. LCCN QA76.76.D47

T46 2004. URL http://www.oreilly.com/catalog/9780974514000.

Thomas:2005:AWD

[1917] David Thomas, David Heinemeier Hansson, and Leon Breedt. Agile Web development with rails: a Pragmatic guide. The facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-0-X. x + 558 pp. LCCN TK5105.888 .T46 2005.

Thomas:2007:AWD

[1918] David Thomas, David Heinemeier Hansson, Leon Breedt, et al. Agile Web development with Rails. The facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, second edition, 2007. ISBN 0-9776166-3-0. xvi + 719 pp. LCCN TK5105.888 .T477 2007; TK5105.888 .T46 2006b.

URL http://www.oreilly.com/catalog/9780977616633.

Thompson:1997:WNS

[1919] Robert Bruce Thompson. Windows NT Server 4.0 for NetWare Administrators. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-280-8. xxii + 732 pp. LCCN QA76.76.O63 T4847 1997. US$39.95. URL http://www.

ora.com/catalog/netware/;   http://www.oreilly.com/catalog/
9781565922808;   http://www.oreilly.com/catalog/netware.

Thompson:2000:PHN

[1920] Robert Bruce Thompson and Barbara Fritchman Thompson. PC hardware in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-599-8. xxi + 501 pp. LCCN TK7887.5 .T48 2000. US$29.95. URL http://www.oreilly.

com/catalog/9781565925991.

Thompson:2002:PHN

[1921] Robert Bruce Thompson and Barbara Fritchman Thompson. PC Hardware in a Nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00353-6. xxiii + 780 pp. LCCN TK7887.5 .T48 2002. US$39.95. URL http://safari.oreilly.com/

0596003536;   http://www.oreilly.com/catalog/9780596003531;
 http://www.oreilly.com/catalog/pchardnut2.

Thompson:2003:PHN

[1922] Robert Bruce Thompson and Barbara Fritchman Thompson. PC hardware in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00513-X. xxii + 850 pp. LCCN TK7887.5 .T48 2003. URL

http://www.oreilly.com/catalog/9780596005139.

Thompson:2004:BPP

[1923] Robert Bruce Thompson and Barbara Fritchman Thompson. Building the perfect PC. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00663-2. xvi + 332 pp.

LCCN QA76.5 .T56 2004. URL http://www.oreilly.com/catalog/
9780596006631.

Thompson:2005:AH

[1924] Robert Bruce Thompson and Barbara Fritchman Thompson. Astronomy hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10060-4 (paperback). xx + 388 pp. LCCN QB64 .T46 2005. URL http://www.oreilly.com/

catalog/9780596100605.

Thompson:2005:PHB

[1925] Robert Bruce Thompson and Barbara Fritchman Thompson. PC hardware buyer’s guide: choosing the perfect components. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00938-0. xiv + 127 pp. LCCN TK7887.5 .T478 2005.

URL http://www.oreilly.com/catalog/9780596009380.

Thompson:2006:BPP

[1926] Robert Bruce Thompson and Barbara Fritchman Thompson. Building the perfect PC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52686-5. xviii + 402 pp. LCCN TK9969; TK9969 .T55

2006eb. URL http://www.oreilly.com/catalog/9780596526863.

Thompson:2006:RUY

[1927] Robert Bruce Thompson and Barbara Fritchman Thompson. Repairing and upgrading your PC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00866-X. xvi + 429 pp. LCCN QA76.5; QA76.5 .T46 2006eb. URL

http://www.oreilly.com/catalog/9780596008666.

Thota:2006:PMN

[1928] Chandu Thota. Programming MapPoint in .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00906-2. xvi + 354 pp. LCCN GA102.4.E4; GA102.4.E4 T46 2006; GA102.4.E4 T46 2006eb. URL http://www.

oreilly.com/catalog/9780596009069.

Tidwell:2001:X

[1929] Doug Tidwell. XSLT. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00053-7. xvi + 460 pp. LCCN QA76.73.X58 T53 2001. US$39.95.

URL http://www.oreilly.com/catalog/9780596000530;   http://
www.oreilly.com/catalog/xslt.

Tidwell:2006:DI

[1930] Jenifer Tidwell. Designing interfaces. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00803-1. xx + 331 pp. LCCN QA76.9.U83; QA76.9.U83 T43 2006eb; QA76.9.U83 T53 2006; QA76.9.U83 T55 2006.

URL http://www.oreilly.com/catalog/9780596008031.

Tidwell:2007:XMX

[1931] Doug Tidwell. XSLT: mastering XML transformations. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52721-7 (paperback). xviii + 965 pp. LCCN QA76.73.X58 T53 2008. US$31.99.

Tilly:2002:ADG

[1932] Jesse E. Tilly and Eric M. Burke. Ant: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00184-3. xviii + 269 pp. LCCN QA76.73.J38 T57 2002. US$34.95. URL http://safari.oreilly.com/

0596001843;   http://www.oreilly.com/catalog/9780596001841;
 http://www.oreilly.com/catalog/anttdg.

Tisdall:2001:BPB

[1933] James D. Tisdall. Beginning Perl for bioinformatics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00080-4. xiii + 368 pp. LCCN QA76.73.P22 T57 2001 Copy 5 on Re-

serve. US$39.95. URL http://safari.oreilly.com/0596000804;
http://www.oreilly.com/catalog/9780596000806;   http://www.
oreilly.com/catalog/begperlbio.

Tisdall:2003:MPB

[1934] James D. Tisdall. Mastering Perl for bioinformatics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00307-2. xvi + 377 pp. LCCN QA76.73.P22 T59 2003. URL http://www.oreilly.com/

catalog/9780596003074.

Titmus:2021:CNG

[1935] Matthew Titmus. Cloud Native Go: building reliable services in unreliable environments. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2021. ISBN 1-4920-7633-3. ???? pp. LCCN ????

Todino:1985:UUU

[1936] Grace Todino. Using UUCP and USENET. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1985. ISBN (none). 90 pp. LCCN QA76.8.U655

T63 1985. URL http://www.oreilly.com/catalog/.

Todino:1986:UUU

[1937] Grace Todino and Dale Dougherty. Using UUCP and Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1986. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA 76.76 O63 T63 1991. US$21.95. URL http://www.oreilly.

com/catalog/9780937175101.

Todino:1987:UUUa

[1938] Grace Todino. Using UUCP and Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 1987. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA 76.76 O63 T63 1991.

Todino:1987:UUUb

[1939] Grace Todino and Tim O’Reilly & Dale Dougherty. Using UUCP and USENET. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 1987. ISBN (none). xi + 172 pp. LCCN QA76.6 .T62 1987.

URL http://www.oreilly.com/catalog/.

Todino:1991:UUU

[1940] Grace Todino and Dale Dougherty. Using UUCP and Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1991. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA76.76 O63 T63 1991. US$21.95. Well written on how to use uucp and Netnews.

Todino:1992:MUU

[1941] Grace Todino and Tim O’Reilly. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, tenth edition, January 1992. ISBN 0-937175-93-5. 368 pp. LCCN QA76.76.O63 O73 1992. US$27.95. URL http://www.

oreilly.com/catalog/9780937175934.

Todino:1993:LUO

[1942] Grace Todino, John Strang, and Jerry Peek. Learning the UNIX Operating System. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, August 1993. ISBN 1-56592-060-0. xv + 92 pp. LCCN QA76.76.O63 T62 1994. US$9.95. URL http://www.

loc.gov/catdir/enhancements/fy0915/94242752-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0915/94242752-d.html;
http://www.loc.gov/catdir/enhancements/fy1113/94242752-t.html.

Todino:1994:LUO

[1943] Grace Todino, John Strang, and Jerry D. Peek. Learning the UNIX Operating System. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1994. ISBN 1-56592-060-0. xv + 92 pp. LCCN QA76.76.O63 T62 1994. US$9.95. URL http://www.oreilly.com/

catalog/9781565920606. Minor corrections.

Todino:1998:LUO

[1944] Grace Todino, Jerry D. Peek, and John Strang. Learning the UNIX operating system. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 1998. ISBN 1-56592-390-1. xi + 92 pp. LCCN QA76.76.O63 T62 1998. Additions and revisions by Jerry Peek.

Toledo:2000:ONC

[1945] Hugo Toledo and Jonathan Gennick. Oracle Net8 Configuration and Troubleshooting. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 2000. ISBN 1-56592-753-2. 408 pp. US$39.95. URL http://www.oreilly.

com/catalog/net8trouble.

Toledo:2001:ONC

[1946] Hugo Toledo and Jonathan Gennick. Oracle Net8 Configuration and Troubleshooting. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-753-2. xvii + 387 pp. LCCN QA76.76.O63 T64 2001. US$39.95.

URL http://www.oreilly.com/catalog/9781565927537;   http://
www.oreilly.com/catalog/net8trouble.

Topley:2002:JND

[1947] Kim Topley. J2ME in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00253-X. xv + 450 pp. LCCN QA76.73.J38 T67 2002; QA76.73.J38 T69 2002.

URL http://www.oreilly.com/catalog/9780596002534.

Topley:2003:JWS

[1948] Kim Topley. Java Web services in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00399-4. xviii + 642 pp. LCCN TK5105.88813

.T67 2003. URL http://www.oreilly.com/catalog/9780596003999.

Toporek:1999:HGP

[1949] Chuck Toporek and Kellie Robinson. Hydrocephalus: a Guide for Patients, Families, and Friends. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-410-X. xxiv + 355 pp. LCCN

RC391 .T66 1999. US$19.95. URL http://www.oreilly.com/catalog/
9781565924109;   http://www.oreilly.com/catalog/hydro.

Toporek:2002:MXPa

[1950] Chuck Toporek. Mac OS X Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00346-3. vii + 115 pp. LCCN QA76.6.O63 P634A 2002; QA76.76.O63 T67

2002. US$12.95. URL http://safari.oreilly.com/0596003463;
http://www.oreilly.com/catalog/9780596003463;   http://www.
oreilly.com/catalog/macosxpr.

Toporek:2003:MXPb

[1951] Chuck Toporek. Mac OS X Pocket Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00458-3. vii + 141 pp. LCCN QA76.76.O63 T66 2002. US$14.95.

URL http://www.oreilly.com/catalog/9780596004583;   http://
www.oreilly.com/catalog/macosxpg2.

Toporek:2004:IM

[1952] Chuck Toporek. Inside .Mac. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00501-6. xxii + 360 pp. LCCN QA76.8.M3 T67 2004. URL

http://www.oreilly.com/catalog/9780596005016.

Toporek:2004:MXPa

[1953] Chuck Toporek, Chris Stone, and Jason McIntosh. Mac OS X Panther in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00606-3. xxv + 1025 pp. LCCN QA76.76.O63 M3873 2004. URL http://www.

oreilly.com/catalog/9780596006068.

Toporek:2004:MXPb

[1954] Chuck Toporek. Mac OS X Panther pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2004. ISBN 0-596-00616-0. ix + 158 pp. LCCN QA76.76.O63 T668

2004. URL Errata:;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.oreilly.com/catalog/9780596006167;
 http://www.oreilly.com/catalog/macpantherpg/errata/.

Toporek:2005:MXT

[1955] Chuck Toporek. Mac OS X Tiger: pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-00914-3 (paperback). xi + 223 pp. LCCN

QA76.76.O63 T668 2005. URL http://www.oreilly.com/catalog/
9780596009144.

Toporek:2006:MXT

[1956] Chuck Toporek, Chris Stone, Jason McIntosh, et al. Mac OS X Tiger in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00943-7 (paperback). xv + 498 pp. LCCN QA76.76.O63 M3873 2006.

URL http://www.oreilly.com/catalog/9780596009434.

Toporek:2006:RBC

[1957] Chuck Toporek. Running Boot Camp. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52772-1. LCCN QA76.76.O63. URL http://www.

oreilly.com/catalog/9780596527723.

Toporek:2007:MXL

[1958] Chuck Toporek. Mac OS X Leopard pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52981-3. xi + 211 pp. LCCN QA76.76.O63.

URL http://www.oreilly.com/catalog/9780596529819.

Torres:2005:SPD

[1959] Jesse M. Torres and Peter Sideris. Surviving PC disasters, mishaps, and blunders. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-932111-98-0. xxx + 408 pp. LCCN QA76.9.F34 T67 2005. URL http:

//www.oreilly.com/catalog/9781932111989.

Torres:2006:WAS

[1960] Jesse M. Torres. Windows admin scripting little black book. Paraglyph Press, Inc., Scottsdale, AZ, USA, third edition, 2006. ISBN 1-933097-10-8. xxxvii + 466 pp. LCCN QA76.76.O63 T6775 2006. URL http://

www.oreilly.com/catalog/9781933097107.

Tow:2003:ST

[1961] Dan Tow. SQL tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00573-3. xix + 314 pp. LCCN QA76.73.S67 T69 2003. URL http:

//www.oreilly.com/catalog/9780596005733.

Trachtenberg:2004:UP

[1962] Adam Trachtenberg. Upgrading to PHP 5. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00636-5. xviii + 327 pp. LCCN QA76.73.P224 T73 2004. URL

http://www.oreilly.com/catalog/9780596006365.

Tranter:1995:MLD

[1963] Jeff Tranter. Multimedia on Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-212-3. 400 pp. LCCN ???? US$29.95.

URL http://www.oreilly.com/catalog/9781565922129.

Tranter:1996:LMG

[1964] Jeff Tranter. Linux Multimedia Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-219-0. xx + 363 pp. LCCN QA76.575 .T73 1996.

US$32.95. URL http://www.oreilly.com/catalog/9781565922198;
 http://www.oreilly.com/catalog/multilinux.

Tranter:1997:SIS

[1965] Jeff Tranter and Eric Dumas. Son et images sous Linux. O’Reilly, Paris, France, 1997. ISBN 2-84177-014-1. 380 pp. LCCN ????

Troy:2010:VC

[1966] Ryan Troy and Matthew Helmke. VMware cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2010. ISBN 0-596-15725-8 (paperback). xv + 280 pp. LCCN QA76.9.V5 T76 2010.

Truelove:2001:PNO

[1967] Kelly Truelove, Clay Shirky, Lucas Gonze, and Rael Dornfest. 2001 P2P Networking Overview: The Emergent P2P Platform of Presence, Identity, and Edge Resources. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00185-1. 289 (est.) pp. LCCN ???? US$495.00.

URL http://www.oreilly.com/catalog/9780596001858;   http://
www.oreilly.com/catalog/p2presearch.

Ts:2003:SPR

[1968] Jay Ts, Robert Eckstein, David Collier-Brown, and Robert Eckstein. Samba pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00546-6. vi + 133 pp. LCCN QA76.76.O63 E266

2003. URL http://www.oreilly.com/catalog/9780596005467.

Ts:2003:USF

[1969] Jay Ts, Robert Eckstein, and David Collier-Brown. Using Samba: a File & Print Server for Linux, Unix & Mac OS X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00256-4. xiii + 539 pp. LCCN QA76.9.C55 E38 2003. US$39.95, EUR 44.00, SFr 72.90.

URL http://www.oreilly.com/catalog/9780596002565;   http://
www.oreilly.com/catalog/samba2/.

Tulloch:1999:MES

[1970] Mitch Tulloch. Microsoft Exchange Server in a nutshell: a desktop quick reference. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-601-3. xi + 390 pp. LCCN QA76.9.C55 T85 1999.

US$24.95. URL http://www.oreilly.com/catalog/9781565926011;
 http://www.oreilly.com/catalog/exservnut/.

Tulloch:2001:WAN

[1971] Mitch Tulloch. Windows 2000 administration in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 2001. ISBN 1-56592-713-3. x + 786 pp. LCCN QA76.76.O63 T85 2001.

US$39.95. URL http://www.oreilly.com/catalog/9781565927131;
 http://www.oreilly.com/catalog/win2000nut.

Tulloch:2003:WSN

[1972] Mitch Tulloch. Windows Server 2003 in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00404-4. xi + 647 pp. LCCN

QA76.76.O63 T8515 2003. URL http://www.oreilly.com/catalog/
9780596004040.

Tulloch:2004:WSH

[1973] Mitch Tulloch. Windows server hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00647-0. xxvi + 357 pp. LCCN

QA76.76.O63 T8517 2004. URL http://www.oreilly.com/catalog/
9780596006471.

Turnbull:2006:YLW

[1974] Giles Turnbull. Your life in Web apps. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52806-X. LCCN QA76.76.D47. URL http://www.

oreilly.com/catalog/9780596528065.

Turner:2006:IN

[1975] Andrew Turner. Introduction to Neogeography. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52995-3. LCCN GA139 .T87 2006eb;

GA139. URL http://www.oreilly.com/catalog/9780596529956.

Tyler:2006:FL

[1976] Chris Tyler. Fedora Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52682-2. xiii + 639 pp. LCCN QA76.76.O63; QA76.76.O63

T9 2006eb. URL http://www.oreilly.com/catalog/9780596526825.

Tyler:2008:XPT

[1977] Chris Tyler. X power tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-10195-3. xiii + 254 pp. LCCN QA76.76.W56. URL http:/

/www.oreilly.com/catalog/9780596101954.

Tynan:2005:CPA

[1978] Dan Tynan. Computer privacy annoyances: how to avoid the most annoying invasions of your personal and online privacy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00775-2 (paperback). xii + 177 pp. LCCN QA76.9.A25

T96 2005. URL http://www.oreilly.com/catalog/9780596007751.

Udell:1999:PIG

[1979] Jon G. Udell. Practical Internet Groupware. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-537-8. xxiv + 497 pp. LCCN QA76.573.I58 U34 1999; QA76.625 .U34 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565925373;   http://
www.oreilly.com/catalog/pracintgr.

Vadala:2002:MRL

[1980] Derek Vadala. Managing RAID on Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-730-3. xi + 245 pp. LCCN QA76.76.O63

V32 2002. US$39.95. URL http://www.oreilly.com/catalog/
9781565927308;   http://www.oreilly.com/catalog/mraidlinux.

vanBeijnum:2002:BBR

[1981] Iljitsch van Beijnum. BGP: Building Reliable Networks with the Border Gateway Protocol. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00254-8. xiv + 272 pp. LCCN TK5105.555 .B45 2002. US$39.95.

URL http://www.oreilly.com/catalog/9780596002541;   http://
www.oreilly.com/catalog/bgp.

Vanderplas:2016:PDS

[1982] Jake Vanderplas. Python Data Science Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2016. ISBN 1-4919-1205-7. ???? pp. LCCN ????

VanderVeer:2006:FMM

[1983] Emily A. Vander Veer. Flash 8: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10137-6. xi + 446 pp. LCCN TR897.7 .V36 2006; TR897.7; TR897.7 .V36 2006eb. URL

http://www.oreilly.com/catalog/9780596101374.

VanderVeer:2007:FCM

[1984] Emily A. Vander Veer and Chris Grover. Flash CS3: the missing manual. Missing manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51044-6 (paperback). xii + 513 pp. LCCN TR897.7; TR897.7 .V384 2007eb; TR897.7 .V384 2007. URL http://www.oreilly.com/

catalog/9780596510442.

VanderVeer:2007:PMM

[1985] Emily A. Vander Veer. PowerPoint 2007: the missing manual. Missing manual. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52738-1. xv + 469 pp. LCCN T385 .V36 2007eb; T385. URL http://www.oreilly.com/

catalog/9780596527389.

VanderVeer:2007:PSM

[1986] Emily A. Vander Veer. PowerPoint 2007 for starters: the missing manual. The Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2007. ISBN 0-596-52831-0. xii + 305 pp. LCCN T385 .V335 2007eb; T385 .V335 2007; T385. URL http://www.oreilly.com/

catalog/9780596528317.

VanderVeer:2008:FMM

[1987] Emily A. Vander Veer. Facebook: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2008. ISBN 0-596-51769-6. xviii + 268 pp. LCCN TK5105.888 .V363 2008. URL

http://www.oreilly.com/catalog/9780596517694.

vanderVlist:2002:XS

[1988] Eric Van der Vlist. XML Schema. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00252-1. xiii + 379 pp. LCCN QA76.76.H94 V55 2002

Stacks. URL http://www.oreilly.com/catalog/9780596002527.

VanderVlist:2004:RN

[1989] Eric Van der Vlist. RELAX NG. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00421-4. xviii + 486 pp. LCCN QA76.76.H94 V3597

2004. URL http://www.oreilly.com/catalog/9780596004217.

VanderVlist:2007:S

[1990] Eric Van der Vlist. Schematron. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52771-3. LCCN QA76.76.H94 V36 2007eb; QA76.76.H94.

URL http://www.oreilly.com/catalog/9780596527716.

VanMeggelen:2006:ATO

[1991] Jim Van Meggelen, Jared Smith, and Leif Madsen. Asterisk: la téléphonie Open Source. O’Reilly, Paris, France, édition française edition, 2006. ISBN 2-84177-394-9. xvi + 414 pp. LCCN ????

VanMeggelen:2007:AFT

[1992] Jim Van Meggelen, Leif Madsen, and Jared Smith. Asterisk: the future of telephony. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51048-9. xxv + 574 pp. LCCN TK5105.8865 .M43 2007; TK5105.8865; TK5105.8865 .M43 2007eb. URL http://www.oreilly.

com/catalog/9780596510480.

vanMeggelen:2011:ADG

[1993] Jim van Meggelen, Leif Madsen, and Russell Bryant. Asterisk: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2011. ISBN 0-596-51734-3. ???? pp. LCCN ????

VanRaalte:1991:XRM

[1994] Thomas Van Raalte. XView Reference Manual, volume 7B. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, September 1991. ISBN 0-937175-88-9. 311 pp. LCCN QA76.76.W56 X85 1991. US$24.95. URL http:

//www.oreilly.com/catalog/9780937175880. Latest update: August

1993.

VanWyk:2001:IR

[1995] Kenneth R. Van Wyk and Richard Forno. Incident Response. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00130-4. xvii + 214 pp. LCCN T58.64 .V365 2001. US$34.95.

URL http://www.oreilly.com/catalog/9780596001308;   http://
www.oreilly.com/catalog/incidentres.

Vesperman:2003:EC

[1996] Jennifer Vesperman. Essential CVS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00459-1. xiv + 320 pp. LCCN QA76.76.D47 V46 2003.

URL http://www.oreilly.com/catalog/9780596004590.

Vesperman:2006:EC

[1997] Jennifer Vesperman. Essential CVS. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-52703-9. xviii + 407 pp. LCCN QA76.76.D47 V47 2007. URL http://www.oreilly.com/

catalog/9780596527037.

Viega:2002:NSO

[1998] John Viega, Matt Messier, and Pravir Chandra. Network Security with OpenSSL: Cryptography for Secure Communications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00270-X. xiv + 367 pp. LCCN TK5105.59 .V54 2002. US$39.95. URL http://safari.oreilly.com/

059600270X;   http://www.oreilly.com/catalog/9780596002701;
 http://www.oreilly.com/catalog/openssl.

Viega:2003:SPC

[1999] John Viega and Matt Messier. Secure programming cookbook for C and C++: Recipes for cryptography, authentication, networking, input validation and more. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00394-3. xxv + 762 pp. LCCN QA76.73.C15 V53 2003. URL http://

www.oreilly.com/catalog/9780596003944.

Vincent:2005:RE

[2000] Jesse Vincent, Robert Spier, Dave Rolsky, Darren Chamberlain, and Richard Foley. RT essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00668-3. xiii + 200 pp. LCCN HD30.2 .V453 2005. URL

http://www.oreilly.com/catalog/9780596006686.

VonHagen:2006:LSH

[2001] William Von Hagen and Brian K. (Brian Kenneth) Jones. Linux server hacks. Vol. 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10082-5. xx + 456 pp. LCCN QA76.76.O63; QA76.76.O63eb. URL

http://www.oreilly.com/catalog/9780596100827.

Vromans:1996:PDR

[2002] Johan Vromans. Perl 5 Desktop Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1996. ISBN 1-56592-187-9. v + 39 pp. LCCN QA76.73

.P33 V76 1996. US$6.95. URL http://www.oreilly.com/catalog/
9781565921870;   http://www.oreilly.com/catalog/perlqr.

Vromans:1998:PPR

[2003] Johan Vromans. Perl 5 Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1998. ISBN 1-56592-495-9. 67 pp. LCCN QA76.73.P22 V76 1998; QA76.73.P22 V765 1998. US$6.95.

URL http://www.oreilly.com/catalog/9781565924956;   http://
www.oreilly.com/catalog/perlpr2.

Vromans:2000:PPR

[2004] Johan Vromans. Perl 5 Pocket Reference: Programming Tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, May 2000. ISBN 0-596-00032-4. vi + 90 pp. LCCN QA76.73.P22 V765 2000. US$9.95.

URL http://www.oreilly.com/catalog/9780596000325;   http://
www.oreilly.com/catalog/perlpr3.

Vromans:2002:PPR

[2005] Johan Vromans. Perl Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2002. ISBN 0-596-00374-9. iv + 91 pp. LCCN QA76.73.P22 V7652 2002. US$12.95.

URL http://www.oreilly.com/catalog/9780596003746;   http://
www.oreilly.com/catalog/perlpr4.

Vromans:2003:PKG

[2006] Johan Vromans. Perl: kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2003. ISBN 3-89721-247-1 (paperback). 108 pp. LCCN ???? EUR 8.00. URL http://www.

gbv.de/dms/hbz/toc/ht013606323.

Walker:2004:XFB

[2007] Mark H. Walker. Xbox fan book: rock the box. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00884-8 (paperback). xiv + 114 pp.

LCCN TK6681 .W35 2004. URL http://www.oreilly.com/catalog/
9780596008840.

Walker:2005:OMM

[2008] Mark H. Walker and Franklin N. Tessler. Office 2004 for Macintosh: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00820-1. xii + 736 pp. LCCN

HF5548.4.M525 W35 2005. URL http://www.oreilly.com/catalog/
9780596008208.

Walker:2011:GSN

[2009] Chris Walker. Getting started with Netduino: [programming open source electronics with .NET]. Make: Projects. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2012. ISBN 1-4493-3119-X, 1-4493-0245-9 (paperback). xii + 128 pp. LCCN QA76.73.B3 .W65 2012. EUR 19.00; UK £15.50. URL http://www.

gbv.de/dms/tib-ub-hannover/671810693.pdf;   http://www.gbv.
eblib.com/patron/FullRecord.aspx?p=862733.

Wall:1991:PP

[2010] Larry Wall and Randal L. Schwartz. Programming perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1991. ISBN 0-937175-64-1. 482 pp. LCCN QA76.73.P43 W35 1990, QA76.73.P226 W34 1992. US$29.95.

URL http://www.oreilly.com/catalog/9780937175644.

Wall:1992:PP

[2011] Larry Wall and Randal L. Schwartz. Programming Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1992. ISBN 0-937175-64-1. xxv + 454 pp. LCCN QA76.73.P43 W35 1990. US$29.95. The authoritative guide to perl— the programming language for any serious UNIX users.

Wall:1996:PP

[2012] Larry Wall, Randal L. Schwartz, and Tom Christiansen. Programming Perl. A Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1996. ISBN 1-56592-149-6. xxi + 645 pp. LCCN QA76.73.P22 W35 1996; QA76.73.P43 W35

1996. US$39.95. URL http://www.ora.com/catalog/pperl2/;
http://www.oreilly.com/catalog/9781565921498;   http://www.
oreilly.com/catalog/pperl2.

Wall:1997:PRK

[2013] Larry Wall, Clay Irving, Nate Patwardhan, Ellen Siever, and Brian Jepson. The Perl Resource Kit — UNIX Edition. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-370-7. 1700 (est.) pp. LCCN

???? US$149.95. URL http://www.ora.com/catalog/prkunix/;
http://www.oreilly.com/catalog/9781565923706;   http://www.
oreilly.com/catalog/prkunix.

Wall:2000:PP

[2014] Larry Wall, Tom Christiansen, and Jon Orwant. Programming Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, July 2000. ISBN 0-596-00027-8. xxxiii + 1067 pp. LCCN QA76.73.P22 W35 2000. US$49.95.

URL http://www.oreilly.com/catalog/9780596000271;   http://
www.oreilly.com/catalog/pperl3.

Wallace:1999:PWG

[2015] Shawn P. Wallace. Programming Web Graphics with Perl and GNU Software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-478-9. xiv + 454 pp. LCCN QA76.73.P22 W34 1999; T385 .W3523 1999. US$29.95.

URL http://www.oreilly.com/catalog/9781565924789;   http://
www.oreilly.com/catalog/prowg/.

Wallace:2002:PGP

[2016] Shawn P. Wallace. Perl graphics programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00219-X. xiii + 462 pp. LCCN T385 .W3522

2002. URL http://www.oreilly.com/catalog/9780596002190.

Wallingford:2005:SV

[2017] Ted Wallingford. Switching to VoIP. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00868-6. xxi + 477 pp. LCCN TK5105.8865 .W35

2005. URL http://www.oreilly.com/catalog/9780596008680.

Wallingford:2006:VH

[2018] Ted Wallingford. VoIP hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10133-3. xviii + 306 pp. LCCN TK5105.8865;

TK5105.8865 W35 2005eb. URL http://www.oreilly.com/catalog/
9780596101336.

Walmsley:2006:X

[2019] Priscilla Walmsley. XQuery. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00634-9. xv + 491 pp. LCCN QA76.7; QA76.7 .W35 2006eb.

URL http://www.oreilly.com/catalog/9780596006341.

Walsh:1994:MTW

[2020] Norman Walsh. Making TEX Work. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1994. ISBN 1-56592-051-1. xxxvi + 483 pp. LCCN TK5105.875.I57 M36 1994; Z253.4.T47 W34 1994. US$29.95.

URL ftp://ftp.ora.com/pub/products/catalogs/book.catalog;
http://www.oreilly.com/catalog/9781565920514. Written specifi-

cally for database maintainers and system administrators, this Unixbased technical guide covers installing, setting up, and running Internet applications such as gopher holes, FTP and telnet sites, mailing lists, WAIS, and World Wide Web pages, and other sites, as well as keeping them secure.

Walsh:1999:DDG

[2021] Norman Walsh and Leonard Muellner. DocBook: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-580-7. xii + 635 pp. LCCN QA76.76.H94 W35 1999. US$36.95.

URL http://www.oreilly.com/catalog/9781565925809;   http://
www.oreilly.com/catalog/docbook. Includes CD-ROM.

Walsh:1999:LPT

[2022] Nancy Walsh. Learning Perl/Tk: Graphical User Interfaces with Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-314-6. xvi + 357 pp. LCCN QA76.73.P22 W36 1999. US$32.95.

URL http://www.oreilly.com/catalog/9781565923140;   http://
www.oreilly.com/catalog/lperltk/.

Walsh:2003:RL

[2023] Matt Walsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer, editors. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2003. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63

R855 2003. US$44.95. URL http://www.oreilly.com/catalog/
9780596002725;   http://www.oreilly.com/catalog/runux4.

Waltz:1999:PDD

[2024] Mitzi Waltz. Pervasive Developmental Disorders: Finding a Diagnosis and Getting Help. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-530-0. xiii + 567 pp. LCCN RJ506.D47 W34 1999; WS 350 W241p 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565925304;   http://
www.oreilly.com/catalog/pdd/.

Waltz:2000:BDG

[2025] Mitzi Waltz. Bipolar disorders: a guide to helping children and adolescents. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-656-0. xvi + 442 pp. LCCN RJ506.D4 W35 2000. URL http://www.oreilly.com/

catalog/9781565926561.

Waltz:2000:OCD

[2026] Mitzi Waltz. Obsessive-compulsive disorder: help for children and adolescents. Patient-centered guides. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-758-3. xvii + 384 pp. LCCN RJ506.O25 W35 2000. URL

http://www.oreilly.com/catalog/9781565927582.

Wampler:2011:FPJ

[2027] Dean Wampler. Functional programming for Java developers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-1265-9, 1-4493-1103-2. xi + 72 pp. LCCN QA76.62 .W36 2011. URL http://proquest.

safaribooksonline.com/9781449312657.

Wang:2003:SCB

[2028] Wally Wang and Wally Steal this computer book Wang. Steal this computer book 3: what they won’t tell you about the Internet. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-59327-000-3. xx + 359 pp.

LCCN HV6773 .W35 2003eb. URL http://www.oreilly.com/catalog/
9781593270001.

Wang:2004:SFS

[2029] Wally Wang. Steal this file sharing book: what they won’t tell you about file sharing. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-050-X. 259 pp. LCCN TK5105.525 .W36 2004; TK5105.525 .W36

2004eb. URL http://www.oreilly.com/catalog/9781593270506.

Wang:2005:VBE

[2030] Wally Wang. Visual Basic 2005 express: now playing. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-059-3 (cd-rom). xxi + 452 pp. LCCN QA76.73.B3 W367435 2005eb. URL http://www.oreilly.

com/catalog/9781593270599.

Wang:2006:BNC

[2031] Wally Wang. The book of Nero 7: CD and DVD burning made easy. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-110-7. xvi + 250 pp. LCCN TK7895.C39 W363 2006eb. URL http://www.oreilly.

com/catalog/9781593271107.

Wang:2006:SCB

[2032] Wally Wang. Steal this computer book 4.0: what they won’t tell you about the Internet. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-105-0. 363 pp. LCCN TK5105.875.I57 W3727 2006eb.

URL http://www.oreilly.com/catalog/9781593271053.

Ward:2002:BVC

[2033] Brian Ward. The book of VMware: the complete guide to VMware workstation. No Starch Press, San Francisco, CA, USA, 2002. ISBN 1-886411-72-7 (paperback). xv + 249 pp. LCCN QA76.76.O63 W3653 2002eb.

US$39.95. URL http://www.oreilly.com/catalog/9781886411722.

Ward:2004:HLW

[2034] Brian Ward. How Linux works: what every super-user should know. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-035-6. xvi + 347 + 4 pp. LCCN QA76.76.O63 W3654 2004eb; QA76.76.O63 W3654

2004. URL http://www.oreilly.com/catalog/9781593270353.

Warkus:2004:OGD

[2035] Matthias Warkus. The official GNOME 2 developer’s guide. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-030-5. 497 pp. LCCN QA76.9.U83 W375 2004; QA76.9.U83 W375 2004eb. URL http:

//www.oreilly.com/catalog/9781593270308.

Watters:2002:SAG

[2036] Paul A. Watters. Solaris 8 Administrator’s Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00073-1. xv + 282 pp. LCCN QA76.76.O63 W37 2002; QA76.76.O63 W3918

2002. US$39.95. URL http://safari.oreilly.com/0596000731;
http://www.oreilly.com/catalog/9780596000738;   http://www.
oreilly.com/catalog/solaris8.

Weaver:2008:SFF

[2037] Matthew Weaver and Duane Wessels. Small form factor PCs. Make: projects. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52076-X. xv + 275 pp. LCCN TK9969 .W43 2008.

Webb:2004:EPD

[2038] Jeff Webb. Excel 2003 programming: a developer’s notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00767-1. xv + 294 pp. LCCN HF5548.4.M523 W42 2004eb; HF5548.4.M523; HF5548.4.M523 W42 2004; HF5548.4.M523 W43 2004. URL http://www.oreilly.com/

catalog/9780596007676.

Webb:2005:ES

[2039] Jeff Webb. Essential SharePoint. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00880-5. xv + 310 pp. LCCN TK5105.875.I6 W43

2005. URL http://www.oreilly.com/catalog/9780596008802.

Webb:2005:SOP

[2040] Jeff Webb. SharePoint office pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10112-0. vi + 77 pp. LCCN TK5105.875.I6 W432 2005. URL

http://www.oreilly.com/catalog/9780596101121.

Webb:2006:PEV

[2041] Jeff Webb and Steve Saunders. Programming Excel with VBA and .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00766-3. xxi + 1090 pp. LCCN HF5548.4.M523 W43 2006; HF5548.4.M523 W43

2006eb; HF5548.4.M523. URL http://www.oreilly.com/catalog/
9780596007669.

Webb:2007:ES

[2042] Jeff Webb. Essential SharePoint 2007. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51407-7 (paperback). xv + 428 pp. LCCN TK5105.875.I6; TK5105.875.I6 W43 2007; TK5105.875.I6 W43

2007eb. URL http://www.oreilly.com/catalog/9780596514075.

Weber:2004:OOW

[2043] Jean Hollis Weber. OpenOffice.org Writer: the free alternative to Microsoft Word. O’Reilly Community Press, Sebastopol, CA, USA, 2004. ISBN 0-596-00826-0. xvii + 213 pp. LCCN HF5548.4.O57 W43 2004.

URL http://www.oreilly.com/catalog/9780596008260.

Weeks:2004:LU

[2044] Roger Weeks, Edd Dumbill, and Brian Jepson. Linux unwired. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00583-0. xii + 297 pp. LCCN QA76.76.O63

W433 2004. URL http://www.oreilly.com/catalog/9780596005832.

Weissinger:1999:AND

[2045] A. Keyton Weissinger. ASP in a Nutshell: a Desktop Quick Reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-490-8. xv + 407 pp. LCCN Tk5105.8885.A26 W45 1999. US$24.95.

URL http://www.oreilly.com/catalog/9781565924901;   http://
www.oreilly.com/catalog/aspnut/.

Weissinger:2000:AND

[2046] A. Keyton Weissinger. ASP in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, July 2000. ISBN 1-56592-843-1. xv + 473 pp. LCCN TK5105.8885.A26 W45 2000.

US$29.95. URL http://www.oreilly.com/catalog/9781565928435;
 http://www.oreilly.com/catalog/aspnut2.

Weissinger:2002:DJC

[2047] A. Keyton Weissinger. Developing JSP custom tag libraries. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00155-X. ???? pp. LCCN

QA76.73.J38 W46 2002. URL http://www.oreilly.com/catalog/
9780596001551.

Wells:2007:SAA

[2048] Christopher Wells. Securing Ajax applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52931-7 (paperback). xiii + 233 pp.

LCCN TK5105.8885.A52. URL http://www.oreilly.com/catalog/
9780596529314.

Welsh:1995:RL

[2049] Matt Welsh and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1995. ISBN 1-56592-100-3. xxii + 576 pp. LCCN QA76.76.O63 W465 1995. US$24.95. URL http://www.oreilly.com/

catalog/9781565921009.

Welsh:1996:RL

[2050] Matt Welsh and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, Summer 1996. ISBN 1-56592-151-8. xviii + 630 pp. LCCN QA76.76.O63 W465 1996. US$29.95.

URL http://www.oreilly.com/catalog/9781565921511;   http://
www.oreilly.com/catalog/runux2.

Welsh:1997:SL

[2051] Matt Welsh, Lar Kaufman, René Cougnenc, Manuel Makarevitch, and Nat Makarevitch. Le Système Linux. O’Reilly, Paris, France, 2e rev. et augm. edition, 1997. ISBN 2-84177-033-8. xix + 653 pp. LCCN ????

Welsh:1999:RL

[2052] Matt Welsh, Matthias Kalle Dalheimer, and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1999. ISBN 1-56592-469-X. xix + 730 pp. LCCN QA76.76.O63 W465 1999. US$34.95.

URL http://www.oreilly.com/catalog/9781565924697;   http://
www.oreilly.com/catalog/runux2.

Welsh:2001:LWS

[2053] Matt Welsh, Ben Laurie, Ellen Siever, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00208-4. LCCN Disk 2057 Protect. CD-ROM with Running Linux, 3rd ed. – Linux in a nutshell, 3rd ed. – CGI programming with Perl, 2nd ed. – Apache: the definitive guide, 2nd ed. – MySQL and mSQL – Programming the Perl DBI.

Welsh:2002:RL

[2054] Matt Welsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2002. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63 R855 2003. US$44.95. URL http://www.oreilly.com/

catalog/runux4.

Wenz:2006:PA

[2055] Christian Wenz. Programming Atlas. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52672-5 (paperback). xix + 382 pp. LCCN QA76.73.A83 W46 2006; TK5105.888 .W46 2006. URL http://www.

oreilly.com/catalog/9780596526726.

Wenz:2007:PAN

[2056] Christian Wenz. Programming ASP.NET AJAX. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51424-7. xviii + 454 pp.

LCCN TK5105.8885.A26. URL http://www.oreilly.com/catalog/
9780596514242.

Wenz:2008:ESD

[2057] Christian Wenz. Essential Silverlight 2 up-to-date. O’Reilly shortcuts. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51998-2. 212 (est.) pp. LCCN QA76.575 .W46 2008.

Wessels:2001:WC

[2058] Duane Wessels. Web Caching. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, June 2001. ISBN 1-56592-536-X. xv + 300 pp. LCCN TK7895.M4 W45 2001.

US$39.95. URL http://www.oreilly.com/catalog/9781565925366;
 http://www.oreilly.com/catalog/webcaching.

Wessels:2004:SDG

[2059] Duane Wessels. Squid: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00162-2. xx + 441 pp. LCCN TK7895.M4 W4497

2004. URL http://www.oreilly.com/catalog/9780596001629.

Wheaton:2004:DBF

[2060] Wil Wheaton. Dancing barefoot: five short but true stories about life in the so-called space age. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00674-8. 115 pp.

LCCN ???? URL http://www.oreilly.com/catalog/9780596006747.

Wheaton:2004:JGU

[2061] Wil Wheaton. Just a geek: unflinchingly honest tales of the search for life, love, and fulfillment beyond the Starship Enterprise. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00768-X (paperback). xxvi + 267 pp. LCCN

PN2287.W4568 A3 2004. URL http://www.oreilly.com/catalog/
9780596007683.

White:2012:BAW

[2062] John Myles White. Bandit Algorithms for Website Optimization. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-4133-0. 80 pp. LCCN ???? EUR 16.00.

White:2012:HDG

[2063] Tom (Tom E.) White. Hadoop: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2012. ISBN 1-4493-1152-0 (paperback). xxiii + 657 pp. LCCN ???? EUR 41.00; UK £35.50. URL http://www.

gbv.de/dms/weimar/toc/688614469\_toc.pdf.

Whittle:1997:CBP

[2064] Janet Whittle, Joe Reif, et al. Canada Business: The Portable Encyclopedia for Doing Business with Canada. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885073-13-5. 304 pp. LCCN HC115.C1984 1996.

US$24.95. URL http://www.oreilly.com/catalog/9781885073136.

Wickham:2017:RDS

[2065] Hadley Wickham and Garrett Grolemund. R for data science: import, tidy, transform, visualize, and model data. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-1039-9 (paperback), 1-4919-1033-X (e-book), 1-4919-1034-8 (ebook), 1-4919-1036-4 (e-book). LCCN QA76 .W53 2017. URL http://

ebookcentral.proquest.com/lib/imperial/detail.action?docID=
4770093.

Widenius:2002:MRM

[2066] Michael “Monty” Widenius and David Axmark. MySQL Reference Manual: Documentation from the Source. O’Reilly Community Press, Sebastopol, CA, USA, 2002. ISBN 0-596-00265-3. x + 802 pp. LCCN QA76.73.S67 W54 2002. US$39.95. URL http://

press.oreilly.com/commpress.html;   http://www.oreilly.com/
catalog/9780596002657;   http://www.oreilly.com/catalog/mysqlref.

Widzer:1999:PPP

[2067] Joel L. Widzer. The Penny Pincher’s Passport to Luxury Travel: The Art of Cultivating Preferred Customer Status. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-885211-31-7. 253 pp. LCCN G153.4.W334 1999.

US$12.95. URL http://www.oreilly.com/catalog/9781885211316;
 http://www.oreilly.com/catalog/penny/.

Wildermuth:2006:PSS

[2068] Shawn Wildermuth and Richard Blewett. Programming SQL Server 2005. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00479-6. xi + 571 pp. LCCN QA76.73.S67 W55 2006eb; QA76.73.S67

W55 2006; QA76.73.S67. URL http://www.oreilly.com/catalog/
9780596004798.

Wildermuth:2007:GSS

[2069] Shawn Wildermuth. Getting started with Silverlight. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51068-3. LCCN TR897.7; TR897.7 .W55

2007eb. URL http://www.oreilly.com/catalog/9780596510688.

Williams:2001:WDA

[2070] Hugh E. Williams and David Lane. Web Database Applications with PHP & MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00041-3. 563 pp. LCCN QA76.73.P224 W55 2001

Stacks. US$44.95. URL http://safari.oreilly.com/0596000413;
http://www.oreilly.com/catalog/9780596000417;   http://www.
oreilly.com/catalog/webdbapps.

Williams:2002:FFR

[2071] Sam Williams. Free as in Freedom: Richard Stallman’s Crusade for Free Software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00287-4. xii + 225 pp. LCCN QA76.2.S73 W55 2002; QA76.76.C73 W55 2002.

US$22.95. URL http://www.oreilly.com/catalog/9780596002879;
 http://www.oreilly.com/catalog/freedom.

Williams:2002:WDA

[2072] Hugh E. Williams and David (David John) Lane. Web database applications with PHP and MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00041-3. xvi + 563 pp. LCCN QA76.73.P224 W55

2002. US$44.95. URL http://safari.oreilly.com/0596000413;
http://www.oreilly.com/catalog/9780596000417;   http://www.
oreilly.com/catalog/webdbapps.

Williams:2004:WDA

[2073] Hugh E. Williams and David (David John) Lane. Web database applications with PHP and MySQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00543-1. xviii + 796 pp. LCCN QA76.73.P224

W55 2004. URL http://www.oreilly.com/catalog/9780596005436.

Williamson:2002:DND

[2074] Heather Williamson and Bruce A. Epstein. Dreamweaver in a Nutshell: a Desktop Quick Reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00239-4. xxvi + 442 pp. LCCN TK5105.8885.D74 W55

2002. US$29.95. URL http://safari.oreilly.com/0596002394;
http://www.oreilly.com/catalog/9780596002398;   http://www.
oreilly.com/catalog/dreamweavernut.

Wilson:2005:DCS

[2075] Greg Wilson. Data crunching: solve everyday problems using Java, Python and more. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-7-1. viii + 193 pp. LCCN QA63 .W55 2005; QA63 .W56 2005. URL http://www.oreilly.com/

catalog/9780974514079.

Wilson:2012:BNA

[2076] Mike Wilson. Building Node Applications with MongoDB and Backbone. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-3739-2. 150 pp. LCCN ???? EUR 16.00.

Windley:2005:DI

[2077] Phillip J. Windley. Digital identity. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00878-3. xviii + 234 pp. LCCN TK5105.59 .W45 2005.

URL http://www.oreilly.com/catalog/9780596008789.

Wingerd:2005:PP

[2078] Laura Wingerd. Practical Perforce. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10185-6. 256 pp. LCCN QA76.76.D47; QA76.76.D47eb. URL http://www.oreilly.com/

catalog/9780596101855.

Winton:2001:PNP

[2079] Greg Winton. Palm OS Network Programming: Writing Connected Applications for the Palm. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2001. ISBN 0-596-00005-7. xv + 382 pp. LCCN QA76.76.O63 W578 2001. US$39.95. URL

http://www.oreilly.com/catalog/9780596000059;   http://www.
oreilly.com/catalog/palmosnetpro.

Wise:1996:PTY

[2080] Naomi Wise. Passport Thailand: Your Pocket Guide to Thai Business, Customs and Etiquette. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-885073-26-7. 96 pp. LCCN HD58.7.W57 1996. US$6.95. URL

http://www.oreilly.com/catalog/9781885073266.

Wist:2001:CS

[2081] Brad Wist and Max Ramsay. Commerce Server 2000. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00147-9. ???? pp. LCCN QA76.9.C55

W575 2001. URL http://www.oreilly.com/catalog/9780596001476.

Wong:1997:WCP

[2082] Clinton Wong. Web Client Programming with Perl. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Summer 1997. ISBN 1-56592-214-X. xii + 213 pp. LCCN QA76.625 .W66 1997. US$29.95.

URL http://www.oreilly.com/catalog/9781565922143;   http://
www.oreilly.com/catalog/webclient.

Wong:2000:HPR

[2083] Clinton Wong. HTTP Pocket Reference: Hypertext Transfer Protocol. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-862-8. iii + 75 pp. LCCN QA76.76.H94 W66 2000. US$9.95.

URL http://www.oreilly.com/catalog/9781565928626;   http://
www.oreilly.com/catalog/httppr.

Wood:1999:PIE

[2084] David Wood. Programming Internet Email. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-479-7. xiii + 362 pp. LCCN TK5105.73 .W66 1999. US$34.95. URL

http://www.oreilly.com/catalog/9781565924796;   http://www.
oreilly.com/catalog/progintemail.

Woodman:2006:HBR

[2085] Mark Woodman. How to build an RSS 2.0 feed. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52938-4. LCCN TK5105.888;

TK5105.888 .W66 2006eb. URL http://www.oreilly.com/catalog/
9780596529383.

Woodring:1997:WPT

[2086] Mike Woodring and Aaron Cohen. WebSite Professional (TM) V2.0. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-327-8. 700

(est.) pp. LCCN ???? US$799. URL http://www.ora.com/catalog/
webpro2/;   http://www.oreilly.com/catalog/9781565923270;
http://www.oreilly.com/catalog/webpro2.

Woods:2003:ESA

[2087] Dan Woods. Enterprise services architecture. AN O’Reilly field guide to Enterprise software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00551-2. xvi + 205 pp. LCCN TK5105.88813 .W66 2003. URL

http://www.oreilly.com/catalog/9780596005511.

Woods:2003:PCA

[2088] Dan Woods. Packaged composite applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00552-0. xiv + 187 pp. LCCN HF5548.375

.W66 2003. URL http://www.oreilly.com/catalog/9780596005528.

Woods:2005:OSE

[2089] Dan Woods and Gautam Guliani. Open source for the enterprise: managing risks, reaping rewards. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10119-8. xiii + 217 pp. LCCN T58.5 .W66 2005. URL

http://www.oreilly.com/catalog/9780596101190.

Woods:2006:ESD

[2090] Dan Woods and Thomas Mattern. Enterprise SOA: designing IT for business innovation. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10238-0. xv + 433 pp. LCCN QA76.9.A73; QA76.9.A73 W66

2006eb. URL http://www.oreilly.com/catalog/9780596102388.

Worsley:2002:PP

[2091] John C. Worsley and Joshua D. Drake. Practical PostgreSQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-846-6. xiv + 619 pp. LCCN QA76.9.D3 W67 2002; QA76.9.D3 W72 2002. US$44.95.

URL http://www.oreilly.com/catalog/9781565928466;   http://
www.oreilly.com/catalog/ppostgresql. CD-ROM contains LXP ver-

sion 0.8.0 and PostgreSQL version 7.1.3.

Worsley:2002:PPP

[2092] John C. Worsley and Joshua D. Drake. PostgreSQL par la pratique. O’Reilly, Paris, France, 2002. ISBN 2-84177-211-X. xiv + 614 pp. LCCN ????

Worsley:2002:PPPb

[2093] John C. Worsley and Joshua D. Drake. PostgreSQL par la pratique. O’Reilly, Paris, France, 2002. ISBN 2-84177-211-X. xiv + 614 pp. LCCN ????

Wright:2006:DGB

[2094] Shauna Wright. Don’t get burned on eBay. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10178-3. xii + 160 pp. LCCN HF5478 .W75 2006eb; HF5478. URL http://www.oreilly.com/

catalog/9780596101787.

WWW:1997:AHS

[2095] World Wide Web Consortium. Advancing HTML: style and substance, volume 2(1) of World Wide Web journal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1997. ISBN 1-56592-264-6. ISSN 1085-2298. vii + 246 pp.

LCCN ???? URL http://www.oreilly.com/catalog/9781565922648.

Wylie:1997:TTL

[2096] Judith Babcock Wylie, editor. Travelers’ Tales: Love & Romance. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-885211-18-X. 350 (est.) pp. LCCN ???? US$17.95. URL http://www.ora.com/catalog/

romance/;   http://www.oreilly.com/catalog/9781885211187;
http://www.oreilly.com/catalog/romance.

Yaghmour:2003:BEL

[2097] Karim Yaghmour. Building Embedded Linux Systems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00222-X. xxii + 391 pp. LCCN QA76.76.O63 Y35 2003. US$44.95, CDN$69.95, UK£31.95.

URL http://www.oreilly.com/catalog/9780596002220;   http://
www.oreilly.com/catalog/belinuxsys/.

Yaghmour:2008:BEL

[2098] Karim Yaghmour et al., editors. Building embedded Linux systems. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2008. ISBN 0-596-52968-6. 439 (est.) pp. LCCN QA76.76.O63 Y35 2008.

Yaghmour:2011:EA

[2099] Karim Yaghmour. Embedded Android. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2011. ISBN 1-4493-0829-5. ???? pp. LCCN

QA76.76.A65 Y345 2011. URL http://proquest.safaribooksonline.
com/?fpi=9781449327958.

Yaghmour:2012:EAP

[2100] Karim Yaghmour. Embedded Android Porting, Extending, and Customizing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-0829-5. ???? pp. LCCN ????

Yank:2004:BYO

[2101] Kevin Yank. Build your own database driven website using PHP and MySQL. SitePoint Pty. Ltd., Collingwood, VIC, Australia, third edition, 2004. ISBN 0-9752402-1-8. xiv + 359 pp. LCCN TK5105.888 .Y36

2004; TK5105.888 .Y25 2004. URL http://www.oreilly.com/catalog/
9780975240212.

Yarger:1999:MM

[2102] Randy Jay Yarger, George Reese, and Tim King. MySQL and mSQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-434-7. xv + 487 pp. LCCN QA76.73.S67 Y37 1999. US$34.95. URL http://www.

oreilly.com/catalog/9781565924345.

Yoder:2015:BCS

[2103] Mark A. Yoder and Jason Kridner. BeagleBone Cookbook: Software and Hardware Problems and Solutions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0539-5. ???? pp. LCCN ????

Youens-Clark:2022:CLR

[2104] Ken Youens-Clark. Command-line Rust: a Project-Based Primer for Writing Rust CLIs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2022. ISBN 1-0981-0943-0 (paperback). xviii + 377 pp. LCCN QA76.73.R87 Y68 2022.

Yuan:2005:NSH

[2105] Michael Juntao Yuan. Nokia smartphone hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00961-5. xxi + 382 pp. LCCN TK6570.M6; TK6570.M6 Y83 2005eb; TK6570.M6 Y83 2005. URL http://www.

oreilly.com/catalog/9780596009618.

Zacker:1999:BWN

[2106] Craig Zacker. Building Windows 98 networks: design, implementation, and use. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-548-3. ???? pp. LCCN QA76.76.O63 Z34 1999. URL http://www.oreilly.com/

catalog/9781565925489.

Zacker:1999:ZAW

[2107] Craig Zacker. Zero Administration for Windows. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-508-4. xiii + 544 pp. LCCN QA76.76.C69

Z33 1999. US$34.95. URL http://www.oreilly.com/catalog/
9781565925083;   http://www.oreilly.com/catalog/zeroadm.

Zakas:2010:HPJ

[2108] Nicholas C. Zakas. High performance JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2010. ISBN 0-596-80279-x, 1-4493-8230-4. xviii + 209 pp. LCCN ????

Zalewski:2005:SWF

[2109] Michal Zalewski. Silence on the wire: a field guide to passive reconnaissance and indirect attacks. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-046-1. xxv + 281 pp. LCCN TK5105.59 .Z35 2005; TK5105.59 .Z35 2005eb. URL http://www.oreilly.com/

catalog/9781593270469.

Zarella:2011:FMB

[2110] Dan Zarella and Alison Zarella. The FaceBook marketing book. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8848-5. vii + 272 pp. LCCN ????

Zawodny:2004:HPM

[2111] Jeremy D. Zawodny and Derek J. Balling. High performance MySQL: optimization, backups, replication, and load balancing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00306-4. xvi + 276 pp. LCCN QA76.73.S67 Z39 2004. URL http://www.oreilly.com/

catalog/9780596003067.

Zdziarski:2005:ESB

[2112] Jonathan A. Zdziarski. Ending spam: Bayesian content filtering and the art of statistical language classification. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-052-6. xx + 287 pp. LCCN TK5105.743 .Z35 2005; TK5105.743 .Z35 2005eb. URL http://www.oreilly.com/

catalog/9781593270520.

Zdziarski:2008:IOA

[2113] Jonathan A. Zdziarski. iPhone open application development. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51855-2 (paperback). 267 (est.) pp. LCCN TK6570.M6 Z397 2008.

Zdziarski:2009:IOA

[2114] Jonathan A. Zdziarski. iPhone open application development. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2009. ISBN 0-596-15534-4. xii + 253 pp. LCCN TK6570.M6 .Z397 2008.

Zukowski:1997:JAR

[2115] John Zukowski. Java AWT Reference. The Java series. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-240-9. xxv + 1045 pp. LCCN QA76.73.J38 Z841 1997; QA76.73.J38 Z85 1997. US$39.95.

URL http://www.oreilly.com/catalog/9781565922402;   http://
www.oreilly.com/catalog/javawt.

Zwicky:2000:BIF

[2116] Elizabeth D. Zwicky, Simon Cooper, and D. Brent Chapman. Building Internet Firewalls: Internet and Web security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 2000. ISBN 1-56592-871-7. xxi + 869 pp. LCCN TK5105.59 .Z85 2000. US$44.95.

URL http://www.oreilly.com/catalog/9781565928718;   http://
www.oreilly.com/catalog/fire2.

Adams:2007:ASC

[2117] Cameron Adams et al., editors. The art and science of CSS. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2007. ISBN 0-9758419-7-1 (paperback). x + 213 pp. LCCN QA76.76.H94 A25 2007. URL http://

www.oreilly.com/catalog/9780975841976.

Brunner:2008:SC

[2118] Stefan Brunner et al., editors. ScreenOS cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51003-9. xxi + 814 pp. LCCN TK5105.59. URL http://

www.oreilly.com/catalog/9780596510039.

Collandre:2005:PCA

[2119] Patrick Collandre et al., editors. Photomontages créatifs avec Photoshop. Creating photomontages with Photoshop: a designer’s notebook. Designer notebooks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00858-9. 96 pp. LCCN T385 .P4713 2005; T385 .P48513 2005; T385 .P4613

2005. URL http://www.oreilly.com/catalog/9780596008581.

Dabbs:2005:IPC

[2120] Alistair Dabbs et al., editors. InDesign production cookbook: 169 easy-to-follow recipes for graphic designers and desktop publishers. O’Reilly digital studio. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10048-5. 192 pp. LCCN Z286.D47 I53 2005eb; Z286.D47; Z286.D47 I53 2005.

URL http://www.oreilly.com/catalog/9780596100483.

Ferguson:1995:XRIa

[2121] Paula Ferguson, editor. The X Resource: Issue 13: Proceedings of the 9th Annual X Technical Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1995. ISBN 1-56592-121-6. LCCN ???? US$14.95. URL http:/

/www.oreilly.com/catalog/9781565921214.

Ferguson:1996:XRIb

[2122] Paula Ferguson, editor. The X Resource: Issue 17: Proceedings of the 10th Annual X Technical Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1996. ISBN 1-56592-178-X. LCCN ???? US$14.95. URL http:

//www.oreilly.com/catalog/9781565921788.

Ford:2006:NFJ

[2123] Neal Ford, editor. No fluff, just stuff anthology: the 2006 edition. Pragmatic Bookshelf, Raleigh, NC, USA, 2006. ISBN 0-9776166-6-5 (paperback). xi + 240 pp. LCCN QA76.76.D47 N646 2006. URL http://www.

oreilly.com/catalog/9780977616664.

Freeman:2004:HFD

[2124] Eric Freeman, Elisabeth Freeman, Kathy Sierra, and Bert Bates, editors. Head first design patterns. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00712-4, 0-596-00712-4. xxxvi + 638 pp. LCCN QA76.76.D47 H427 2004eb; QA76.76.D47 F74 2004; QA76.76.D47;

QA76.76.D47 H427 2004. URL http://www.oreilly.com/catalog/
9780596007126.

Goelker:2007:GPI

[2125] Klaus Goelker, editor. GIMP 2 for photographers: image editing with open source software. Rocky Nook, Santa Barbara, CA, USA, 2007. ISBN 1-933952-03-2. ix + 186 pp. LCCN TR267.5.G5 .G64 2007. URL

http://www.oreilly.com/catalog/9781933952031.

Merholz:2008:SCC

[2126] Peter Merholz, editor. Subject to change: creating great products and services for an uncertain world. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51683-5. viii + 178 pp. LCCN TS170 .S83 2008. URL http://www.oreilly.com/

catalog/9780596516833.

Nye:1992:XRIa

[2127] Adrian Nye, editor. The X Resource: Issue 1: Proceedings of the 6th Annual X Technical Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1992. ISBN 0-937175-96-X. LCCN ???? US$14.95. URL http://www.

oreilly.com/catalog/9780937175965.

Nye:1993:XRIa

[2128] Adrian Nye, editor. The X Resource: Issue 5: Proceedings of the 7th Annual X Technical Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1993. ISBN 1-56592-020-1. LCCN ???? US$14.95. URL http://www.

oreilly.com/catalog/9781565920200.

Nye:1994:XRIa

[2129] Adrian Nye, editor. The X Resource: Issue 9: Proceedings of the 8th Annual X Technical Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Winter 1994. ISBN 1-56592-066-X. LCCN ???? US$14.95. URL http://www.

oreilly.com/catalog/9781565920668.

Oram:2007:BC

[2130] Andrew Oram and Greg Wilson, editors. Beautiful Code. Theory in practice. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51004-7 (paperback). xxi + 593 pp. LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http://

proquest.safaribooksonline.com/9780596510046;   http://www.
oreilly.com/catalog/9780596510046.

OReilly:1997:HCIa

[2131] O’Reilly & Associates, Inc., editor. The Harvard Conference on the Internet and Society. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 0-674-45931-8. LCCN ???? US$32.95. URL http://www.ora.com/catalog/

harvard/;   http://www.oreilly.com/catalog/9780674459311;
http://www.oreilly.com/catalog/harvard. Conference held in Cam-

bridge, MA, USA May 28–31, 1996.

OReilly:1997:HCIb

[2132] O’Reilly & Associates, Inc., editor. The Harvard Conference on the Internet and Society. Harvard University Press, Cambridge, MA, 1997. ISBN 1-56592-277-8 (hardcover + CD-ROM), 0-674-45932-6 (paperback + CD-ROM), 0-674-45931-8 (paperback). LCCN TK5105.875.I57 H37

1997 Desk. URL http://www.oreilly.com/catalog/9781565922778.

Conference held in Cambridge, MA, USA, May 28–31, 1996.

Patrick:2006:VBN

[2133] Tim Patrick et al., editors. Visual Basic 2005 in a nutshell. In a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2006. ISBN 0-596-10152-X (paperback). xvii + 746 pp. LCCN QA76.73.B3; QA76.73.B3eb; QA76.73.B3 V56 2006;

QA76.73.B3 V5694 2006. URL http://www.oreilly.com/catalog/
9780596101527.

Pritchard:2006:LLC

[2134] Steven Pritchard et al., editors. LPI Linux certification in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00528-8. xviii + 961 pp. LCCN QA76.76.O63; QA76.76.O63 L65 2006eb.

URL http://www.oreilly.com/catalog/9780596005283.

Pyles:2007:SDG

[2135] James Pyles et al., editors. Sharepoint 2007: the definitive guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52958-9. LCCN TK5105.875.I6.

URL http://www.oreilly.com/catalog/9780596529581.

Siever:2005:LND

[2136] Ellen Siever et al., editors. Linux in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 925 pp. LCCN QA76.76.O63 L56 2005; QA76.76.O63 L5459 2005. URL http://www.oreilly.com/

catalog/9780596009304.

Smith:2007:LAD

[2137] Bob Smith et al., editors. Linux appliance design: a hands-on guide to building Linux appliances. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-140-9. xxi + 356 pp. LCCN QA76.76.O63 L545115

2007. URL http://www.oreilly.com/catalog/9781593271404.

Abelson:2011:AI

[2138] Hal Abelson and David Wolber. App Inventor. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9748-4. ???? pp. LCCN ????

Aimonetti:2011:MDG

[2139] Matt Aimonetti. MacRuby: The Definitive Guide: Ruby and Cocoa on OS X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8037-9. 350 (est.) pp. LCCN ????

Berry:2011:UD

[2140] Addison Berry and Angela Byron. Using Drupal. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2011. ISBN 1-4493-9052-8. ???? pp. LCCN ????

Biersdorfer:2011:IMM

[2141] J. D. Biersdorfer. iPad: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-0173-8. ???? pp. LCCN ????

Brantley:2011:BT

[2142] Peter Brantley, James Bridle, Bob Stein, Sara Lloyd, Tim O’Reilly, Liza Daly, Joseph J. Esposito, Kate Eltham, Adam Hodgkin, Kassia Krozser, Bill McCoy, Andrew Savikas, and Mac Slocum. Best of TOC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 0-596-80212-9. ???? pp. LCCN ????

Cecco:2011:SJG

[2143] Raffaele Cecco. Supercharged JavaScript Graphics. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9363-2. 400 (est.) pp. LCCN ????

Donahue:2011:NW

[2144] Gary A. Donahue. Network Warrior. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2011. ISBN 1-4493-8786-1. ???? pp. LCCN ????

Fulton:2011:HC

[2145] Jeff Fulton and Steve Fulton. HTML5 Canvas. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9390-X. ???? pp. LCCN ????

Gralla:2011:DXM

[2146] Preston Gralla. Droid X: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9386-1. 384 (est.) pp. LCCN ????

Gralla:2011:GTM

[2147] Preston Gralla. Galaxy Tab 2: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9685-2. 320 (est.) pp. LCCN ????

Grover:2011:TMM

[2148] Chris Grover. Title 207: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9825-1. 816 (est.) pp. LCCN ????

Hughes-Croucher:2011:NRS

[2149] Tom Hughes-Croucher. Node: Up and Running: Scalable Server-Side Code with JavaScript. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9858-8. ???? pp. LCCN ????

Karvinen:2011:ABG

[2150] Kimmo Karvinen and Tero Karvinen. Arduino Bots and Gadgets. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8971-6. ???? pp. LCCN ????

MacDonald:2011:CWM

[2151] Matthew MacDonald. Creating a Website: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2011. ISBN 1-4493-0172-X. ???? pp. LCCN ????

McFarland:2011:TMM

[2152] David Sawyer McFarland. Title 206: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9797-2. ???? pp. LCCN ????

Pfister:2011:GSI

[2153] Cuno Pfister. Getting Started with the Internet of Things: Connecting Sensors and Microcontrollers to the Cloud. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9357-8. 200 (est.) pp. LCCN ????

Simon:2011:HFA

[2154] Jonathan Simon. Head First Android Development. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9330-6. ???? pp. LCCN ????

Southwick:2011:JER

[2155] Peter Southwick, Harry Reynolds, and Doug Marschke. Junos Enterprise Routing: a Practical Guide to Junos Routing and Certification. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9863-4. 700 (est.) pp. LCCN ????

VanderVeer:2011:FMM

[2156] Emily A. Vander Veer. Facebook: The Missing Manual. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-9741-7. 271 (est.) pp. LCCN ????

Reischauer:1988:JTC

[2157] Edwin O. Reischauer. The Japanese Today: Change and Continuity. Belknap Press of Harvard University Press, Cambridge, MA and London, UK, 1988. ISBN 0-674-47181-4. viii + 426 pp. LCCN DS806.R35 1988. US$39.95.

Unger:1987:FGF

[2158] J. Marshall Unger. The Fifth Generation Fallacy—Why Japan is Betting its Future on Artificial Intelligence. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-504939-X. x + 230 pp. LCCN QA76.85 .U541 1987. US$19.95. See [2157].